A Complete Bibliography of Publications in IEEE Transactions on Parallel and Distributed Systems:

2020–2029

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

18 December 2024 Version 2.22

Title word cross-reference

2 [DA20]. 3 [DA20, GHG^{+}20, HJEV^{+}21, OMD^{+}21]. ^{+} [HZB^{+}24]. ^{+}+ [LLR^{+}24, YWL^{+}24]. ^{2} [YLW^{+}22]. B^{+} [ZZN^{+}24]. f(r, k + 1)/k [HLH22]. K [YZL^{+}20, WSX^{+}23, ZLCW23]. LU [GLL22]. N [AMW^{+}21, ZLCW23, vDIDB23]. Q [ZRXF23].

-Ary [ZLCW23]. -Body [AMW^{+}21]. -Cubes [ZLCW23]. -Diagnosis [HLH22]. -Means [YZL^{+}20]. -Networks [ZRXF23]. -Party [vDIDB23]. -Shape [WSX^{+}23]. -Tree [ZZN^{+}24].

2-Disjoint [KK24]. 2.0 [ASH^{+}22, BTL^{+}22]. 2D [KK23, TPV20]. 2S [YLW^{+}22].

3 [CYZ^{+}23, TLGA^{+}22]. 3-D [CYZ^{+}23]. 3D [HCA24, LLT^{+}23, QQD^{+}24, WPG^{+}24, ZGZ^{+}24]. 3DLF [CYZ^{+}23]. 3PIP [SJLN20].

50-Photon [LGC^{+}22].

6G [WXH^{+}24].

Abnormal [ZLZ^{+}23]. Abstraction

[ALAK20, GXC^{+}23]. Accelerate [GDS^{+}23]. Accelerated [CLZ^{+}22a, CMLH20, KS23, LYL^{+}20a, MDC24, SSH21, ZMS^{+}22]. Accelerating [ACP^{+}22, CYZ^{+}23, CLG^{+}21, FFQ^{+}22, FLW^{+}23, GWG^{+}22, GZY21, GWLZ21, HLW^{+}21a, HW22, JXX^{+}23, JLL^{+}22, JLY^{+}23b, LLY^{+}20, LS21, LH22, LLD22, LCCZ20b, LLP^{+}23, SZS^{+}23, SHC^{+}22, WLY22, WPZ22, WGN^{+}23, WSX^{+}23, WFY^{+}24, WCT21, WHLM21, WDZ^{+}23, XLZD24, XHY^{+}24, ZKP20, ZLW^{+}23]. Acceleration [AMW^{+}21, CZP^{+}23, DLY^{+}24, GLX^{+}22, GJCC21, KY22, LHQ^{+}20, LLK24, LJW^{+}23, LWZ^{+}24, MKKP22, SLY^{+}23, WLH^{+}23, WGW^{+}23, YCA^{+}20, YFB^{+}23, YZSX23, YZL24, ZPL^{+}22, ZGZ^{+}24, ZLC^{+}22, ZJGD21]. Accelerator [AMW^{+}21, CCYC21, JMF22, JLK^{+}20, LLK22, LZWL22, MKJ^{+}22, TXG^{+}21, WGLZ20, ZZZ^{+}24a, ZZP23]. Accelerators [BPW^{+}23, NISJS21, PKRS23, RP20, WBS23, XMW^{+}22, YHD^{+}23]. Access [CCWW24, CZZM24, CLZ^{+}22b, HLVR21, HZX^{+}23, LHL23, Mei23, VMT^{+}20, XCH^{+}22, XZW^{+}24, YHT^{+}23, ZXL^{+}24]. Account [LWZ23b]. Accuracy [FC23]. Accurate [CHM^{+}20, HLO^{+}21, LYZC22, LWC^{+}22, LLR^{+}24, TW24, WXHZ20, ZZZ^{+}24a, ZHX^{+}22, ZSW^{+}22]. Achieving [AD25, OHWL21, WLH^{+}20a, YWH^{+}20, ZLC^{+}22]. Across [CLL22, KKS21, LYL^{+}20b, LLS^{+}24, WLL^{+}20]. Activation [FLW^{+}23]. Activations [XYW22]. Active [ASS^{+}24, SDHQ21]. Acute [sKW22]. Adaptability [GWLZ21]. Adaptation [CQZ^{+}21, GMI^{+}22]. AdaptChain [XXP^{+}23, ZHH^{+}24]. Adapting [LZWW22]. Adaptive [ABC^{+}24, CXÖ^{+}20, CWC^{+}22, CWC^{+}23, CLX^{+}23, CZZ^{+}22, EFME24, FZC^{+}22, HYL^{+}24, LAY21, LSW^{+}23, LPW^{+}20, LYGG20, LCX^{+}24, SGJ^{+}20, SFW^{+}24, SSKG21, SLY^{+}24, TWY^{+}20, TSLC23, WHM^{+}21, WXX^{+}24, XXP^{+}23, ZDK^{+}22, ZXGZ21, ZZQ^{+}21, ZW22a, ZW22b, ZHH^{+}24, ZTL24, ZZD^{+}24b, ZXW^{+}24]. Adaptively [CXW^{+}24]. Addictive [LHL^{+}22]. Address [ZWSF24].

Addressable [CKO^{+}21]. Addressing [CGH^{+}22, PM22]. ADMM [ZXG^{+}22]. Adoption [HMM22, LYDZ21]. ADRL [KMBR21]. Advanced [LCL^{+}24]. Advances [YLT^{+}21]. Advancing [HMM22, Par22]. Advantage [FWT^{+}24]. Adversarial [GMI^{+}22]. AESM2 [WMG^{+}23]. aeSpTV [CXÖ^{+}20]. Against [CSJB20, HO23, HYP^{+}22, MHZ^{+}22, WLZ^{+}23]. AGCM [CYZ^{+}23]. AGCM-3DLF [CYZ^{+}23]. Age [ZMY^{+}24]. Age-of-Event [ZMY^{+}24]. Agent [CQZ^{+}24, CLY^{+}24, GLP^{+}21, JHB24, LZZ^{+}23, LTH^{+}21, ZXGZ21, ZCL^{+}22]. Agent-Based [GLP^{+}21]. Agents [WIBD22]. Aggregate [ATF23]. Aggregates [MDC24]. Aggregation [BCVD23, CWC^{+}23, HHJ^{+}25, KLH^{+}20a, LLL^{+}23b, MHLM24]. Aggregation-Based [BCVD23]. Agile [SPZE20, WHM^{+}24]. Agnostic [IYAK23, WHLM21]. Agreement [KM23b]. AI [GXC^{+}23, OSF22, TCJ22b, YZC^{+}23]. Aided [LXW^{+}23]. AIDTN [YZC^{+}23]. AIMC [KBA^{+}24]. AIoT [HTL^{+}24]. Airborne [YLW^{+}22]. AIREBO [GDS^{+}23]. Alert [SGJ^{+}20]. Algebra [SZS^{+}23]. Algorithm [AA23, BM22, CA20a, DS23, DFLG21, FTYL20, GKK21, JSL^{+}25, KSB^{+}22, LWL^{+}23, MFYB22, QWHC21, SSH21, TSW^{+}21, TCT^{+}22, TSV21, WSX^{+}23, WCT21, YRBÇ22, ZJHS20, ZZZ^{+}24a, ZDL^{+}21, ZSH^{+}21, ZRXF23, ZSFX23, ZLCW23, HLVR21]. Algorithm-Based [ZDL^{+}21]. Algorithmic [FFQ^{+}22]. Algorithms [ABG20, ASH^{+}22, BK21, CFM^{+}21b, CLZ^{+}21, DA20, FC23, GLA20, HSY^{+}20, LCG^{+}21, LLDL23, MGG^{+}20, MKKS21, Pil23, RDBG25, RXL^{+}20, SLX^{+}21b, SCL^{+}21b, XJX24, YDL23, ZZH^{+}21, ZMP23]. Alignment [JZLZ24, WWJL24, ZZZ^{+}24a]. All-Flash [KZK^{+}19, KZK^{+}20]. Alleviating [ZLZ^{+}23]. Allocation [AA23, BBG22, CZH^{+}20b, CZJ^{+}22, HZW^{+}21, HCZ^{+}20, HND20, HLZ^{+}21, HYP^{+}22, LT20, LZZ^{+}23, LNX^{+}22, LLDL23, Nak21, RDBG25, SWGB23, SLZR21, SDBM23, WHL^{+}23, WZGM23, YLT^{+}21, YZWT20, ZWSF24, CZH^{+}20a]. Allocation-Based [Nak21]. Allocations [LK20, XCH^{+}22, YWH^{+}21]. Alternating [CFLY21, ZHQ^{+}23]. AMG [BCVD23]. among [SGJ^{+}20]. AMR [WPG^{+}24]. Analog [JLL^{+}22]. Analysis [CQZ^{+}21, CBL22, CGM21, GLA20, GZJ^{+}21, HLVR21, HWS^{+}24, JWZ^{+}24, KY22, LWL^{+}25, PK21, RT25, WDL^{+}20, WSM^{+}20, WZL^{+}23, WSLX22, ZTA^{+}21, ZYK^{+}22, ZMS^{+}22, ZFH22, ZF23]. Analytic [BM20, CLL22]. Analytics [CLL^{+}21, CC23, CCZW24, JQG^{+}22, LSL^{+}23, LCL^{+}20, OZCW22, PZZ^{+}22, WNL20, XXW^{+}24, ZYM^{+}20, ZLW^{+}21]. Analyzing [IATB20]. Anomaly [ASS^{+}24, BMMB22, FWT^{+}24, KMBR21, LLL^{+}21b, OSF22, PWX^{+}23, ZHX^{+}22]. Anomaly-Aware [KMBR21]. Anticipation [BMMB22]. APIs [ALAK20]. APMigration [TWY^{+}20]. App [XCH^{+}21b]. Applicable [SBM24]. Application [BBGY20, CA20b, JLW^{+}24, KMLE20, LHL23, Man22, Mei23, PK21, SMK^{+}23, SWOM20, WGBS23, ZXL^{+}24]. Application-Level [KMLE20]. Applications [AKG20, CFM^{+}21a, CBB^{+}20, CTBT21, DFP23, DDN^{+}22, DLMF22, DDX^{+}24, DMPR22, FTYL20, GGHP21, GXC^{+}23, HPB21, HW22, JBLJ23, JWZ^{+}24, KFS^{+}21, KEMC22, LLX^{+}22, LGH^{+}24, LK21, LMFK23, LZWW22, LLP^{+}23, LXL23, MGG^{+}20, NAL^{+}20, QWYG20, SdR^{+}21, SMCH20, SLHH23, TCT^{+}22, WHG^{+}22, WDJ21, XZL20, XYL^{+}21, XLL^{+}20a, XZL^{+}21, ZGM21, ZLJ^{+}23, ZMS^{+}22, ZLK^{+}22]. Approach [ACP^{+}22, AA24, BBGY20, CAAB20, CSZ^{+}23, DLLL22, DWZ^{+}24, GWLZ21, HCZ^{+}20, HXW^{+}20, KKP21, KMA^{+}20, LSW^{+}23, LTH^{+}21, LLL^{+}21c, LHX^{+}25, MZWX21, MLS21, MSSK21, OSF22, SdR^{+}21, SSKG21, WZHW22, WLM^{+}20, WGW^{+}23, XZL20, ZJS^{+}24, ZYX^{+}22]. Approaches [BPTV23, WZL^{+}23]. Approximate [ABC^{+}24, CL20b, DTN^{+}22, RP20]. Approximated [MLS21]. Approximation [BM22, HW22]. APQ [YHD^{+}23]. Arbitrary [XTH^{+}23]. Arbitration [NKP^{+}24]. Architectural [CLZP20]. Architecture [ASMA21, CGC^{+}22, GLW^{+}21, HLW^{+}20, JTX^{+}22, JLL^{+}22, LLY^{+}20, LYL^{+}20a, LHQ^{+}20, LLK24, LLL^{+}23b, MCT21, RCLJT22, SMSK21, WLF^{+}20, XMW^{+}22, YZL24, ZSL^{+}23, ZGNZ22, ZF23]. Architectures [GXW^{+}20, KHOI20, LMH^{+}20, LC20, LX23, MKJ^{+}22, QTR21, RP20, ZYL^{+}20, ZFY^{+}20, ZZY^{+}21, ZZG^{+}21a, ZHH^{+}24]. Archival [XHQC20]. Archive [VMT^{+}20]. Area [ACDK20, SKW23]. ARENA [TXG^{+}21]. Arithmetic [LLD22, NCB^{+}21, WRLS22, WLH^{+}23]. ARM [LMH^{+}20, MZC^{+}22b, WJZ^{+}24, LJZ^{+}20]. ARMv8 [YFD^{+}24]. Array [GHGW24, KZK^{+}19, KZK^{+}20, XMW^{+}22]. Arrays [QQD^{+}24]. Ary [ZLCW23]. Assessment [DFJ^{+}23, FXL^{+}23, RCW^{+}23]. Assignment [LLL^{+}21c]. Assignments [HYP^{+}22]. Assisted [CBB^{+}22, KOA22, LZDO24, YCZC22]. Astraea [YSG^{+}22]. Asymmetric [YZJ^{+}21]. Asynchronism [WGW^{+}23]. Asynchronous [HLB^{+}23, LLC^{+}22, LLDL23, TXG^{+}21, TKRB22, ZWK^{+}20, ZLD^{+}23, ZGG21, ZLRY22]. Athena [GGO21]. Atmospheric [CYZ^{+}23]. Atom [GDZ^{+}20, XSC^{+}23]. Atomic [CKO^{+}21]. Atomicity [OHWL21]. AtRec [WFY^{+}24]. Attack [CSJB20, LJZY20]. Attacks [MHZ^{+}22]. Attain [WJL^{+}24]. Attention [HYL^{+}23]. Attribute [LHPW20, WMG^{+}23]. Attribute-Based [LHPW20, WMG^{+}23]. Auction [LZW^{+}23]. Auction-Based [LZW^{+}23]. Auctions [KSP^{+}20]. Auditing [CMX^{+}20, LHC^{+}21, LCH23, ZZC^{+}23]. Augmented [CQZ^{+}21, QXL^{+}24, TCJ22b]. Austin [DPGG22]. Authorized [LWL^{+}22a]. Auto [DWZ^{+}24, GSH^{+}21, LPL23, SLY^{+}24, CGC^{+}22]. Auto-GNAS [CGC^{+}22]. Auto-Tiering [LPL23]. Auto-Tuning [DWZ^{+}24, GSH^{+}21, SLY^{+}24]. AutoDDL [CLG^{+}24]. Autogeneration [WGBS23]. Automated [JTX^{+}22, KHLZ20, LLT^{+}23, YHD^{+}23, ZZZ^{+}24b, ZMS^{+}22]. Automatic [CLG^{+}24, DMST20, LJZ^{+}20, MZC^{+}22b, WGLZ20, YZL^{+}20, ZLJ^{+}23]. Automotive [XZL20]. Autonomic [YWH^{+}21]. Autonomous [MLX23]. AutoRS [MLX23]. Autoscaling [CZL^{+}24, LZWW22]. Autotuning [DHM25]. Availability [BOGM21, LHPW20, ZXL^{+}24]. Availability-Aware [ZXL^{+}24]. Available [LyLG^{+}25]. Average [Mei23, YOM21]. Averaging [LBNN^{+}21, LCX^{+}24]. Avoiding [LBNN^{+}21]. Awakening [HCR^{+}22]. Aware [ABC^{+}24, BBGY20, CZJ^{+}22, CWL^{+}21, DQC^{+}21, DNLX25, FTYL20, GCL^{+}21, HPB21, HNKO20, HLZ^{+}20, HTB22, HND20, HYL^{+}24, HTL^{+}24, HZK^{+}24, HFW^{+}21, JHB24, JLL^{+}20, KMBR21, KS23, KOA22, KMM20, LLHJ20, LCCZ20a, LWC^{+}22, LyLG^{+}25, LXC^{+}22, LLL^{+}21c, LWL^{+}22b, MXS21, MTL^{+}20, MGBD24, NFP^{+}20, NLX^{+}22, OZCW22, PSS^{+}20, RDBG25, RXL^{+}20, SL20, SMCH20, SDHQ21, SZ20, TSV21, WZZ^{+}20, WZY^{+}22, WBS23, WJZ^{+}24, XJX24, XHQC20, XZL^{+}21, XXC^{+}23, YHS^{+}20, YLW^{+}22, YBY^{+}22, YZJ^{+}21, ZGM21, ZLL22a, ZYD^{+}23, ZLD^{+}23, ZMY^{+}24, ZXL^{+}24, ZSX^{+}20, ZGQ^{+}21, ZFH22, SFW^{+}24]. Azure [LMX^{+}22].

B [ZYL^{+}20]. Back [YWS^{+}23, YWZ^{+}20]. Back-Up [YWZ^{+}20]. Background [TKRB22]. Backtracking [WHC^{+}21]. Backtracking-Based [WHC^{+}21]. Backup [ZYF^{+}20]. Balance [LZF^{+}24]. Balanced [AD25, GZY21, LWZ23b, LWL^{+}25, LJH^{+}23, TWYL20, XZY^{+}24, YWH^{+}20]. Balancing [AAA21, Ans20, BSPM23, CAAB20, CC22, DFXY20, DLC^{+}21, KAA20, KBA^{+}24, KEMC22, LZWL22, LYZS24, LXC^{+}22, Nak21, SGJ^{+}20, SPPS20]. BALS [CFLY21]. Band [SPS^{+}24]. Bandit [LSW^{+}23]. Bandwidth [BPTV23, CFLL21, CLG^{+}24, NFP^{+}20, SLZR21, ZFH22]. Bandwidth-Aware [NFP^{+}20, ZFH22]. Bare [CLZ^{+}22a]. Bare-Metal [CLZ^{+}22a]. Barriers [LBNN^{+}21]. Based [AHSW23, ABG20, ALAK20, AKG20, Ans20, BPW^{+}23, BCVD23, CRZ^{+}23, CZL^{+}22, CLY^{+}24, CXL^{+}23, CLZ^{+}20, DZL^{+}21, FTYL20, GGA24, GLF^{+}21, GLP^{+}21, GHG^{+}20, GGL^{+}23, HCR^{+}22, HLW^{+}20, HLW^{+}21a, HZW^{+}21, HNKO20, HZW^{+}25, HYL^{+}24, HLZ^{+}21, HYP^{+}22, HLB^{+}23, JDD^{+}24, JLWS24, JLQ^{+}23, JLL^{+}22, JLY^{+}23b, JWZ^{+}23, KMBR21, KFEG21, KKA^{+}20, KM23a, LZS^{+}24, LJZY20, LYL^{+}20a, LWZ^{+}22, LLK22, LCZ^{+}23, LZZ^{+}23, LMZ^{+}20, LWX^{+}23, LCL^{+}24, LLL^{+}21c, LZM^{+}20, LCM^{+}20, LZW^{+}23, LLZ^{+}23, LHPW20, LWY^{+}22, MDC24, MWNK22, Nak21, NKP^{+}24, PZL^{+}22, PWX^{+}23, RT25, SSH21, SGH^{+}23, SLY^{+}23, SCA23, SWOM20, TXX^{+}21, TSW^{+}21, WLF^{+}20, WHM^{+}21, WJG^{+}21, WHC^{+}21, WZHW22, WMG^{+}23, WHM^{+}23, WHL^{+}23, WXZ24, WQZD25, WYW21, WHLM23, WZGM23, WUR^{+}24, XZJ^{+}20, XMW^{+}24, XRS^{+}23, YCA^{+}20, YLC^{+}23, YW20, YLW^{+}22, YRBÇ22, YSZL21, YBY^{+}22, YTL^{+}23, YHT^{+}23, ZJHS20, ZDK^{+}22, ZZH^{+}20a, ZQM^{+}22, ZHX^{+}22, ZPL^{+}22, ZZC^{+}23, ZGZ^{+}24, ZZN^{+}24, ZDL^{+}21, ZKP20, ZRXF23]. Based [ZJGD21, AKZ^{+}20, MRFP20, YHD^{+}23, ZZP23]. BASIR [AKG20]. Batch [CCZ^{+}21, LYZS24, LXL23, QXL^{+}24]. Batch-Processing [LXL23]. Batched [LRBV23, YS22]. Bayesian [CRZ^{+}23, GGL^{+}23, LCZ^{+}23, SGH^{+}23, ZZZ^{+}24b, FFQ^{+}22, KSZ24, KKP21, OSF22, SCA23, YJWM24]. Bayesian-Driven [ZZZ^{+}24b]. BeauForT [JBLJ23]. Behavior [CA20b]. Behavioral [LHL^{+}22]. Behaviors [SLG^{+}23]. Belady [WJL^{+}24]. Bench [DFP23]. Benchmark [DFP23, ZZG^{+}21a]. Benchmarking [JLW^{+}24, LGC^{+}22, ONP^{+}23]. Benchmarks [GM21]. Benefits [HAD^{+}22]. Benes [NKP^{+}24]. Best [KBA^{+}24, LYZC22, TGFPRA20]. Best-Effort [TGFPRA20]. Better [ÇSS21, LZF^{+}24]. between [LCX^{+}22]. Beyond [WJL^{+}24]. Bi [KFS^{+}21, LWZ^{+}22, LCX^{+}24]. Bi-Directional [LCX^{+}24]. Bi-Objective [KFS^{+}21, LWZ^{+}22]. Bidirection [GHGW24]. Bidirection-Sequence [GHGW24]. Bidirectional [CLZ^{+}22b]. Bifactor [BM22]. Big [DWZ^{+}24, NK21, WXZ24, XXW^{+}24]. Big-Data [WXZ24]. Billion [GDZ^{+}20]. Billion-Atom [GDZ^{+}20]. Bin [LT20]. Binarized [LS21]. Binary [CMLH20]. Bio [DSW^{+}23]. Bio-ESMD [DSW^{+}23]. Biological [DSW^{+}23]. Bipartite [FTYL20, WZL^{+}23]. BIRD [WYZ^{+}24]. Biscotti [SFYB21]. Bit [GLX^{+}22, LS21]. Bit-Tensor-Cores [LS21]. Bittide [LCIS24]. Black [DSZ^{+}24, NAL^{+}20]. Black-Box [DSZ^{+}24, NAL^{+}20]. Blanket [CRZ^{+}23, GGL^{+}23, LCZ^{+}23, SGH^{+}23, SCA23]. BLAS [ASH^{+}22, ZGZ^{+}23]. BLB [CWC^{+}22]. BLB-gcForest [CWC^{+}22]. Block [ABG20, YRBÇ22, ZZH^{+}20b]. Block-Based [ABG20, YRBÇ22].

Blockchain [JLWS24, LFZ^{+}21, LWZ23b, MZWX21, MMR^{+}21, SFYB21, WHL^{+}23, WXZ24, WPG^{+}22, XXP^{+}23, YHT^{+}23, ZZC^{+}23, ZRXF23]. Blockchain-Based [WHL^{+}23, WXZ24, ZZC^{+}23]. Blockchains [HYP^{+}22, XMW^{+}24]. Blocked [CFLY21]. Blocking [AHSW23, WLF^{+}22]. Blocks [GHM^{+}24]. Blockwise [BLYZ21]. BNN [GLW^{+}21]. Body [AMW^{+}21, DLY^{+}24]. Boltzmann [HLVR21, GVD^{+}22, LCL^{+}24]. Bond [GDS^{+}23]. Bond-Order [GDS^{+}23]. Boolean [YYZ^{+}20]. Booster [WLH^{+}23]. Boosting [AAK22, GSL^{+}20, GK21, HLW^{+}21b, ZF23, ZWSF24]. Boson [LGC^{+}22]. BOSSA [CYH^{+}21]. Both [PSS^{+}20, TRN^{+}21]. Bottleneck [QQD^{+}24]. Bound [CLZ^{+}21]. Boundaries [CLZ^{+}24]. Bounded [OHWL21, SPZE20, WPG^{+}24, ZLX^{+}20]. Bounds [DYFL21, WLL^{+}20]. Box [DSZ^{+}24, NAL^{+}20]. Brain [ZCL^{+}22]. Brain-Computer [ZCL^{+}22]. Breadth [PRL20]. Breadth-First [PRL20]. Breaking [LBNN^{+}21]. Bridging [LCX^{+}22]. Broadcasts [MK24]. Brokerage [LZJ^{+}20]. Bron [WCT21]. Budget [CZL^{+}24, CZZ^{+}22, FSF^{+}20, SMCH20]. Budget-Constrained [FSF^{+}20, SMCH20]. Buffer [DSCL21, HYL^{+}24, ZMP23]. Buffer-Aware [HYL^{+}24]. Buffered [Mei23]. Buffers [HCR^{+}22]. Bugs [AKG20, LGH^{+}24]. Build [LDL22]. Building [RCLJT22, ZZM^{+}23]. Burst [DZL^{+}21]. BurstBalancer [LZF^{+}24]. Bursty [NJG^{+}22, YYW^{+}20]. Business [YLL^{+}20]. Busy [LGH^{+}24]. Busy-Wait [LGH^{+}24]. Butterfly [WZL^{+}23]. Byte [CKO^{+}21, WJL^{+}24, GHM^{+}24]. Byte-Addressable [CKO^{+}21]. Byzantine [JBLJ23, KDREV21, MK24]. Byzantine-Resilient [KDREV21]. Byzantine-Tolerant [MK24].

C [SP20]. Cache [CGLC20, JYF^{+}24, LHC^{+}21, PSS^{+}20, SPPS20, SDHQ21, TXX^{+}21, WHM^{+}24, XTH^{+}23, ZZN^{+}24, ZYK^{+}22, ZZH^{+}20b]. Caching [ASMA21, CC23, GXC^{+}23, LYK20, LSL^{+}23, LWL^{+}25, WXH^{+}24, XCH^{+}21a, XCH^{+}22, XJX24, XTH^{+}23, YWH^{+}20, ZQM^{+}22, LYK20]. Calculations [HLNW22]. Calibration [WGW^{+}23]. California [LAG^{+}22]. Call [HWW^{+}23]. Camera [KKS21]. CAMIG [HTB22]. Canary [ZWL^{+}21]. Cap [KHLZ20]. Capability [ZHQ^{+}23]. Capacity [AMvBI22, LYGG20]. Capelin [AMvBI22]. Carbon [GZS^{+}25]. Carbon-Efficient [GZS^{+}25]. CARMA [LHPW20]. Cars [SLX20]. Cartesian [DA20]. Case [AMW^{+}21, CFLL21, OMD^{+}21]. Causal [MK24, SDZ21, ZJHS20]. Causality [PK21]. Causes [YYW^{+}20]. cCUDA [SNN^{+}20]. CD [WBS23]. CD-MSA [WBS23]. Cell [BTL^{+}22, GLF^{+}21]. Center [CGLC20, CLY^{+}24, LZF^{+}24, OAM^{+}25, TLQ^{+}20, ZDK^{+}22]. Centers [CWL^{+}21, HWW^{+}23, HFW^{+}21, JHB24, LyLG^{+}25, LWL^{+}22b, SL20, SLZR21, YTL^{+}24, YZWT20, ZZH^{+}20a]. Central [RCW^{+}23]. Centrality [MGG^{+}20]. Centric [DPGG22, DSW^{+}23, DHH^{+}22, FGH^{+}22, JBLJ23, JWZ^{+}24, KSW^{+}22, LAG^{+}22, LL22, PKJ^{+}22, TXG^{+}21, ZJH^{+}23, ZSL^{+}23, ZZD^{+}24b, ZSH^{+}21, ZCZ^{+}22]. CERT [XWDC23]. CERT-DF [XWDC23]. Certificates [LWL^{+}22a]. CGRA [MWL24, ZSW^{+}20]. CGRAs [KOA22, LMZ^{+}20, SDMC24]. Chain [LLL^{+}23a, LWZ23b]. Chaining [AA24]. Chains [LLJC21, YZS^{+}21]. Challenges [TSC^{+}24]. Changes [BGZR21]. Characterization [MTT^{+}22]. Characterizing [CQW^{+}20, CA20b, GTH22]. Charging [TWYL20]. Checkpointing [DMPR22]. Checkpoints [WHRL21]. CHEESE [CXL^{+}23]. Chip [GHG^{+}20, HCA24, MAOA22, WL20]. Chiplet [LLK22]. Chiplet-Based [LLK22]. Chips [CL20b, LKH23, ZHP^{+}23]. Cholesky [LLY^{+}20]. Chunking [JLY^{+}23b, XZJ^{+}20]. Chunks [ZFW^{+}20]. CIA [LCH23]. Circuit [LWY^{+}20]. Circulation [CYZ^{+}23]. Class [MTT^{+}22]. Classification [CA20b]. Clemson [DHH^{+}22]. CLIC [CCZW24]. Client [DLJ^{+}22, HLW^{+}21b, JBLJ23]. Client-Centric [JBLJ23]. Client-Level [DLJ^{+}22]. Clients [JLJ21, WHLM21]. Clique [WCT21]. Clock [PK21]. Cloud [AKZ^{+}20, AA23, AKG20, AMvBI22, CHM^{+}20, CMX^{+}20, CLL^{+}21, CZL^{+}22, CHY^{+}24, CZZ^{+}22, CLZ^{+}22b, DS22, DLLL22, FZC^{+}22, FWCB22, GXC^{+}23, HNKO20, HND20, JHB24, KSVR23, LJZY20, LL20, LLHJ20, LWZ^{+}22, LSW^{+}23, LCH23, LYZS24, LYDZ21, LPW^{+}20, LXL23, LPL23, LLL^{+}21c, LWL^{+}22b, LPH^{+}24, LWLX24, MLWX20, Nak21, OQCW20, RCW^{+}23, SWGB23, SMCH20, SLLL20, SSKG21, SLY^{+}23, TSV21, TCJ22b, WDL^{+}20, WLM^{+}20, WPG^{+}22, WXH^{+}24, XXC^{+}23, YS22, ZDK^{+}22, ZZD^{+}24a, ZYW^{+}23, ZLL^{+}22b, ZZH^{+}20b, ZLK^{+}22, ZRXF23, ZLR^{+}20, sKW22]. Cloud-Based [AKG20]. Cloud-Edge [CZL^{+}22, CHY^{+}24, CLZ^{+}22b, FZC^{+}22, KSVR23, WXH^{+}24, ZZD^{+}24a, ZYW^{+}23]. Cloud-Native [GXC^{+}23]. Cloud-Scale [LPH^{+}24]. Cloudlet [BM22]. Clouds [CJLW22, CZL^{+}24, FSF^{+}20, HTB22, IRB21, KMBR21, KPHA20, LHPW20, MRFP20, RXL^{+}20, TCT^{+}22, WDJ21, WLY^{+}20, WLH20b, XZL^{+}21, ZLT^{+}24]. CloudSentry [LPH^{+}24]. CloudSimPer [SZZY24]. Cluster [CZZY23, HKL^{+}20, KKA^{+}20, LXGY23, LZW^{+}23, Par21b, YWH^{+}20, ZGM21]. Cluster-Wide [HKL^{+}20]. Clustered [BLK^{+}20, DLJ^{+}22, FWT^{+}24, SZ20, ZLD^{+}23]. Clustering [CXL^{+}23, GLL^{+}20, LYZC22, MGBD24, WSX^{+}23, YZL^{+}20].

Clustering-Based [CXL^{+}23]. Clusters [CQW^{+}20, GCL^{+}22, KZK^{+}19, KZK^{+}20, PBC^{+}21, QLP^{+}23, TRN^{+}21, YHS^{+}20, YLL21, YSG^{+}22, ZW22b, ZFH22, ZF23]. CNN [DZS^{+}21, JLL^{+}22, JZLZ24, LLK22, YZSX23, ZZQ^{+}21, ZDL^{+}21]. CNNs [GWLZ21, KBA^{+}24, OMD^{+}21, XMW^{+}22]. Co [GWLZ21, HCA24, JYF^{+}24, LWZ^{+}22, SNN^{+}20, XYL^{+}21, SDHQ21]. Co-Active [SDHQ21]. Co-Design [HCA24, JYF^{+}24]. Co-Evolution [LWZ^{+}22]. Co-Reconfiguration [GWLZ21]. Co-Scheduling [SNN^{+}20]. Co-Verification [XYL^{+}21]. Coalition [NLX^{+}22]. Coarse [JMF22, SLX21a]. Coarse-Grained [SLX21a]. Code [GGO21, ZMS^{+}22, LLR^{+}24]. Coded [HFW^{+}21, LWC^{+}20, SL20, SPPS20, XLL^{+}20b, XHQC20, ZFH22, ZF23]. Codes [KZK^{+}19, KZK^{+}20, WGN^{+}23, WSLX22, dBMH21]. Codesign [CLG^{+}21]. Coding [HZW^{+}25, LXL^{+}24, Mei23, ZKL24, CSJB20]. CoFilter [CLZ^{+}22a]. Coflow [DGE^{+}24, LYL^{+}20b, SLZR21]. Coflows [DNLX25]. Cognizant [WVSL23]. Coherence [QJZF23]. Cold [HCR^{+}22]. Collaboration [HWS^{+}24, RCW^{+}23, WXH^{+}24]. Collaborative [CSZ^{+}23, FXL^{+}23, LCH23, LLL^{+}21c, SCYJ21, SDHQ21, XCH^{+}21a, YZJ^{+}21, ZZD^{+}24a, ZZQ^{+}21, ZLL^{+}22b]. Collecting [MLS21]. Collection [LLR^{+}24, ZSW^{+}22]. Collective [KLH^{+}20a, QZCZ21]. Color [ZSH^{+}21]. Color-Centric [ZSH^{+}21]. Coloring [ZSH^{+}21]. Combination [WCD^{+}24]. Combinatorial [ASH^{+}22, KSP^{+}20, LSW^{+}23]. Combined [RCW^{+}23]. Combining [Ans20, ZYF^{+}20]. ComboFunc [WCD^{+}24]. ComboTree [WYW^{+}22]. Comment [CSJB20, ST20]. Comments [DS23]. Commodity [CWL22]. Communication [CL20b, CXW^{+}24, GHM^{+}24, GTH22, HAD^{+}22, LLL^{+}24b, LZH^{+}24, MRFP20, MLS21, OS20, SCL21a, SYS^{+}22, TSLC23, WPZ^{+}21, WGQ^{+}22, WYW21, WYZ^{+}24, WUR^{+}24, XWDC23, XLZD24, YDL23, ZXGZ21, ZBB^{+}22]. Communication-Constrained [HAD^{+}22]. Communication-Efficient [CXW^{+}24, LZH^{+}24, WGQ^{+}22, XLZD24]. Communicational [LZQ^{+}23]. Communications [SZCL23]. Community [ZSFX23]. Compact [LLR^{+}24, XMW^{+}22]. Compaction [CBB^{+}22]. Compaction/Restoration [CBB^{+}22]. Comparative [RT25]. Comparing [LZS^{+}24]. Compensated [WGQ^{+}22]. Competition [Par21b]. Compilation [JMF22, LMZ^{+}20, ZSW^{+}20, ZCJ^{+}22]. Compiler [CBB^{+}22]. Compiler-Assisted [CBB^{+}22]. Completely [QHC20]. Completion [LXW^{+}23, SNK20, WXT^{+}24, ZLW20]. Complex [HLVR21, SRD^{+}20]. Complex-Objective [SRD^{+}20]. Complexity [KM23b]. Component [XSC^{+}23]. Components [ABG20]. Composite [SMCH20]. Compound [QHC20]. Compressed [HZX^{+}23]. Compression [BPW^{+}23, BLYZ21, JWZ^{+}23, SYS^{+}22, WLL^{+}20, WPG^{+}24, ZYF^{+}20, ZZS^{+}22, ZLX^{+}20]. Compressor [WYZ^{+}24]. Computation [BDWH24, BK21, CTL24, CFM^{+}21b, DFJ^{+}23, DLLL22, GZJ^{+}21, KM23b, LMH^{+}20, LCG^{+}21, MGG^{+}20, MHW^{+}21, MKKS21, MRFP20, QZCZ21, SCYJ21, SLX^{+}21b, SCL^{+}21b, TRN^{+}21, WPZ^{+}21, XXM^{+}20, ZZH^{+}21, ZLW^{+}23, lM20]. Computational [LZQ^{+}23, PH21]. Computations [ACP^{+}22, GNST21, GZY21, ZGG21]. Compute [AMvBI22, LX23, YZWT20]. Compute-Efficient [LX23]. Compute-Intensive [YZWT20]. Computer [ZSL^{+}23, ZCL^{+}22]. Computing [AB21, AMW^{+}21, BBGY20, BBG22, BOGM21, BM22, BMMB22, CXÖ^{+}20, CDvK^{+}22, CGM21, DS23, DLLL22, DFLG21, GGA24, GLP^{+}21, HKL^{+}20, HLW^{+}21a, HCZ^{+}20, HLL22, HLLL22, HLB^{+}23, JSL^{+}25, KHLZ20, KLJV24, LHC^{+}21, LLX^{+}22, LTZ^{+}23, LYZS24, LLL^{+}23a, LN24, LCL^{+}24, LZW^{+}23, LJH^{+}23, LWY^{+}22, MZC^{+}22a, Man22, MFYB22, MTL^{+}20, MHM22, MWNK22, MWL24, NDW^{+}21, PSK^{+}22, QWYG20, RDBG25, SCYJ21, SHZ^{+}23, SYT20, TXG^{+}21, TCJ22a, TCJ22b, VMT^{+}20, WDL^{+}20, WHM^{+}21, WHM^{+}23, WHM^{+}24, WHLM21, WDZ^{+}23, WXH^{+}24, XCH^{+}21b, XCH^{+}21a, XCH^{+}22, XZW^{+}24, XWDC23, XXW^{+}24, YWS^{+}23, YWL^{+}24, YHT^{+}23, ZZG^{+}21b, ZHP^{+}23, ZZZ^{+}24b, ZXL^{+}24, dBMH21]. Computing-Efficient [XWDC23]. Concepts [BFK^{+}23]. Concurrency [BADP22, HTB22, LZWW22]. Concurrency-Aware [HTB22]. Concurrent [CCWW24, GN22, HLNW22, SNN^{+}20, SBM24, SDMC24, TJD^{+}24, TGFPRA20, WSHJ23, YCZC22]. Conference [WR23]. Confidentiality [XYL^{+}21]. Configuration [CQZ^{+}21, CLL^{+}21, DWZ^{+}24, HLL22, LMZ^{+}20, LXL23, NFP^{+}20, WZZ^{+}20, WCN^{+}24, ZZD^{+}24b]. Configurations [GSH^{+}21]. Conflict [HCG^{+}23]. Conflicts [KPHA20]. Congestion [TWYL20]. Congestion-Balanced [TWYL20]. Conjugate [YYL^{+}24]. Connected [ABG20, LHXH22, SLX20]. Connectivity [LHXH22]. Conscious [ZZSC20]. Consensus [ACDK20, CYZ^{+}24, LFZ^{+}21, QWHC21, SSS20, WIBD22, ZXGZ21, ZKL24]. Considering [TRN^{+}21]. Consistency [CYZ^{+}24, FMP^{+}23, KMLE20, LCL^{+}20, MZWX21, SDZ21, YW20, sKW22]. Consistent [JLY^{+}23a, Nak21]. Consolidation [ZDK^{+}22]. Consortium [MZWX21]. Constrained [AP20, CZZ^{+}22, CIZ^{+}20, FSF^{+}20, HZW^{+}21, HAD^{+}22, KSP^{+}20, KMA^{+}20, MMR^{+}21, QWYG20, SMCH20, WYZ^{+}24, ZLZ^{+}23]. Constraint [CRZ^{+}23, GGL^{+}23, LCZ^{+}23, LLL^{+}21c, SGH^{+}23, SRD^{+}20, SCA23, TCT^{+}22]. Constraint-Based [CRZ^{+}23, GGL^{+}23, LCZ^{+}23, SGH^{+}23, SCA23]. Constraints [ZXGZ21]. Construction [JXX^{+}23, LMX^{+}24]. Consumption [Pil23, QWYG20, RT25, RCW^{+}23, SWOM20]. Container [LJZY20, LLP^{+}23, LWL^{+}22b, SPCT23, WCD^{+}24, ZTA^{+}21]. Container-Based [LJZY20]. Containerizing [ZZG^{+}21b]. Containers [CZR20]. Content [JLY^{+}23b, SCS^{+}23, TXX^{+}21, WJL^{+}24, XZJ^{+}20]. Content-Defined [JLY^{+}23b, XZJ^{+}20]. Content-Driven [TXX^{+}21]. Contention [BSPM23, YRQ23]. Context [XJX24]. Context-Aware [XJX24]. Contigs [GKK21]. Contiguous [KLH^{+}20a]. Continuum [FZC^{+}22]. Contract [AA24, FZD^{+}24, JLQ^{+}23, LLC^{+}22]. Contracts [LXGY23]. Control [CLZ^{+}20, CLZ^{+}22b, KHLZ20, LXC^{+}22, SLX20, ZXGZ21, ZTL24, ZYX^{+}22]. Controlled [DGE^{+}24]. Controller [HO23, RP20]. Convergence [MGG^{+}20]. Convex [LZH^{+}24]. ConvNet [LZZ21]. Convolution [LZW22b, MZC^{+}22b]. Convolutional [BPW^{+}23, FLW^{+}23, GLF^{+}21, LZWL22, LJW^{+}23, ZQM^{+}22, ZDL^{+}21]. Convolutional-Neural-Network [BPW^{+}23]. Convolutions [KK23, KS23]. Cooling [TRN^{+}21]. CoopEdge [YHT^{+}23]. Cooperative [CQZ^{+}24, CZZM24, KY22, SYT20, WBS23, YHT^{+}23, ZQM^{+}22]. Coordinate [HLNW22]. Coordinated [WHRL21, ZYM^{+}20]. Coordinating [LZWW22, SPCT23]. Coordination [DLMF22, HSH^{+}22]. COPA [YWZ^{+}20, HCR^{+}22]. Copy [ZZC^{+}23].

Core [BLK^{+}20, HSY^{+}20, HZJH23, JYF^{+}24, LX23, WC20, XWJ^{+}20, YFD^{+}24, YZL^{+}20, YZS^{+}21, ZFY^{+}20]. Cores [HFC^{+}23, HW22, LS21, NCB^{+}21, SJLN20, SLG^{+}23, WZHW22, YYL^{+}24]. Correlated [CYY^{+}22, CYF^{+}23]. Correlation [SWOM20]. Cost [CFLL21, CLG^{+}24, CZL^{+}24, CZZM24, DFXY20, DWZ^{+}24, FLPL22, HLL22, KKA^{+}20, LK21, LMFK23, LPL23, LN24, LLL^{+}21c, LWL^{+}22b, OZCW22, RZLT20, TCT^{+}22, WDL^{+}20, WLF^{+}20, WQZD25, WLH20b, XCH^{+}21b, XZW^{+}24, YWZ^{+}20, ZKL24, ZSX^{+}20, ZLK^{+}22]. Cost-Aware [LLL^{+}21c, LWL^{+}22b, OZCW22, ZSX^{+}20]. Cost-Based [KKA^{+}20]. Cost-Driven [LPL23]. Cost-Effective [CZZM24, WQZD25, XCH^{+}21b, XZW^{+}24, YWZ^{+}20]. Cost-Efficient [HLL22, TCT^{+}22]. Cost-Optimal [WLH20b]. Cost-Performance [WDL^{+}20]. Countdown [CBB^{+}20]. Counting [HWF^{+}22, PWZ^{+}21, YRBÇ22]. Coupled [KBA^{+}24, LWZ^{+}23a, YWS^{+}23]. Covers [KK24]. CPLNS [CQZ^{+}24]. CPU [KHLZ20, LyLG^{+}25, LZP24, QXL^{+}20, RT25, SdR^{+}21, TSW^{+}21, TJD^{+}24, ZZG^{+}21a, ZHP^{+}23, ZZN^{+}24, lM20]. CPU-FPGA [QXL^{+}20]. CPU-GPU [ZZG^{+}21a]. CPU/GPU [HLB^{+}23, TJD^{+}24]. CPUs [KK23, KPA^{+}20, LJZ^{+}20, LHZ^{+}23, MZC^{+}22b, WFY^{+}24, WJZ^{+}24, YFD^{+}24, ZGZ^{+}23, ZCHZ23]. Crash [KMLE20, LCL^{+}20]. Crashes [CEP22]. CREPE [SDMC24]. CREW [LWLX24]. Critical [LGH^{+}24]. Criticality [CBL22, HPB21]. Criticality- [HPB21]. Critique [BK21, CRZ^{+}23, CFM^{+}21b, DPGG22, DHH^{+}22, FGH^{+}22, GGL^{+}23, KSW^{+}22, LAG^{+}22, LL22, LCZ^{+}23, LCG^{+}21, MKKS21, PKJ^{+}22, SGH^{+}23, SCL^{+}21b, ZZH^{+}21, ZCZ^{+}22]. Crocus [HKL^{+}20]. Cross [CCZW24, HZB^{+}24, HHJ^{+}25, JLWS24, KLJV24, LJM^{+}23, LWLX24, SL20, XMW^{+}24, YXDL24, ZYX^{+}22]. Cross-Cloud [LWLX24]. Cross-Device [LJM^{+}23]. Cross-Layer [ZYX^{+}22]. Cross-Platform [CCZW24, HZB^{+}24, KLJV24]. Cross-Rack-Aware [SL20]. Cross-Shard [JLWS24, XMW^{+}24]. Cross-Silo [HHJ^{+}25, YXDL24]. Crowdsensing [LHL^{+}22]. Crowdsourced [ZYW^{+}23]. Crowdsourcing [XXWS24]. CSOCs [SGJ^{+}20]. CSP [GNST21]. CT [DPGG22, DHH^{+}22, FGH^{+}22, KSW^{+}22, LAG^{+}22, LL22, PKJ^{+}22, ZCZ^{+}22]. Cubes [ZLCW23]. Cuckoo [LGZ^{+}21]. CUDA [ZJHS20]. CUDA-Based [ZJHS20]. cuNH [GXW22]. cuPC [ZJHS20]. CURE [WL20]. Current [TSC^{+}24]. Customer [CZZ^{+}22, WZZ^{+}20]. Customizable [ACH^{+}20]. Customized [DLY^{+}24, LKH23]. Customizing [GZW^{+}22]. cuTensor [ZLWW20]. cuTensor-Tubal [ZLWW20]. Cuttlefish [CQZ^{+}21]. Cyber [MZW22, TMP23, ZMY^{+}24]. Cyber-Physical [MZW22, TMP23, ZMY^{+}24]. CycleCloud [LMX^{+}22].

D [CYZ^{+}23, DA20, GHG^{+}20, HJEV^{+}21, OMD^{+}21]. D2D [XJX24]. D3Q19 [HLVR21]. DAG [WZGM23]. DAOS [SHC^{+}22]. Dap [CWC^{+}23]. Dap-FL [CWC^{+}23]. Dark [HGA20]. Dark-Silicon [HGA20]. Data [AMvBI22, BM20, BGZR21, BPP21, CYY^{+}22, CYF^{+}23, CFM^{+}21a, CCWW24, CZH^{+}20a, CZH^{+}20b, CGLC20, CYH^{+}21, CLL22, CCZW24, CLY^{+}24, CLG^{+}21, CWL^{+}21, CGL^{+}22, CLZ^{+}22b, DWZ^{+}24, DLC^{+}21, DLJ^{+}22, DSW^{+}23, GLL^{+}20, GZJ^{+}21, GXC^{+}23, HZB^{+}24, HLZ^{+}20, HZW^{+}25, HLVR21, HLZ^{+}21, HZX^{+}23, HWW^{+}23, HFW^{+}21, JHB24, JQG^{+}22, JLY^{+}23b, KFS^{+}21, LC20, LHC^{+}21, LSW^{+}23, LCH23, LTZ^{+}23, LYDZ21, LyLG^{+}25, LJT^{+}25, LWC^{+}20, LPW^{+}20, LSY^{+}20, LLP^{+}23, LZF^{+}24, LWL^{+}22b, LLR^{+}24, LJH^{+}23, LCX^{+}24, LHX^{+}25, LHC^{+}25, MDC24, MCT21, NAL^{+}20, NK21, OZCW22, OAM^{+}25, PZZ^{+}22, PHY20, RDBG25, SKW23, SL20, SLZR21, SNK20, SHZ^{+}23, TXG^{+}21, TSV21, TLQ^{+}20, VMT^{+}20, WHG^{+}22, WVSL23, WXX^{+}24, WXZ24, WQZD25, WCW^{+}23, XZJ^{+}20, XCH^{+}21b, XCH^{+}21a, XCH^{+}22, XLL^{+}20b, XLL^{+}20a, XXW^{+}24, XHY^{+}24, YLC^{+}23, YW20, YTL^{+}24, YZWT20, ZDK^{+}22, ZZH^{+}20a, ZZC^{+}23, ZHH^{+}24, ZLP^{+}24, ZYM^{+}20, ZLRY22, ZLX^{+}20]. Data-Centric [TXG^{+}21]. Data-Driven [AMvBI22, BM20, CZH^{+}20a, CZH^{+}20b]. Data-Intensive [GXC^{+}23, NAL^{+}20, XLL^{+}20a]. Data-Parallel [GZJ^{+}21, HLZ^{+}21, KFS^{+}21, LC20]. Database [LHZ^{+}23]. Databases [NRB^{+}20, YYZ^{+}20]. Datacenter [CZR20, CFLL21, DFXY20, DNLX25, LTT^{+}20, LDZ^{+}24, SWGB23, SFW^{+}24]. Datacenters [AMvBI22, CLL22, FXL^{+}23, LYL^{+}20b, QJC^{+}24, SZZY24, YWZ^{+}20, ZSX^{+}20, ZLCL20]. Dataflow [DLY^{+}24, LIP^{+}21, PKRS23, ZGZ^{+}24]. Dataset [CLZ^{+}21]. Datasets [CTL24, WLY22, ZFW^{+}20]. Datastores [OHWL21, PGY^{+}22]. DC [WLP^{+}23]. DC-WAN [WLP^{+}23]. DCNN [ZGZ^{+}24]. DDoS [LJZY20]. Deadline [DNLX25, MTL^{+}20, TCT^{+}22, WBS23, ZLGZ23]. Deadline-Aware [DNLX25, MTL^{+}20, WBS23]. Deal [LN24]. Debugging [LZW^{+}22a]. Decentralized [CYH^{+}21, DNLX25, HHJ^{+}25, HAD^{+}22, HNKO20, JXX^{+}23, LNX^{+}22, LLDL23, LHC^{+}25, Man22, SQR^{+}21, TSLC23, WNL20, WGQ^{+}22, YHT^{+}23, ZZC^{+}23, ZWL^{+}21]. Decodable [Mei23]. Decomposition [AKA22, KAA21, LLL^{+}21a, SLKA23]. Decompression [PZZ^{+}22]. Decoupled [DZS^{+}21, RCLJT22]. Deduplication [CGL^{+}22, CLX^{+}23, HKL^{+}20, JLY^{+}23b, LYDZ21, LJH^{+}23, LHX^{+}25, LHC^{+}25, SHZ^{+}23, TXX^{+}21, XZJ^{+}20, XCL^{+}23, XXP^{+}23, ZFW^{+}20, ZYF^{+}20]. Deduplication-Based [TXX^{+}21]. Deep [BLYZ21, CQW^{+}20, CHM^{+}20, CWC^{+}22, CZZY23, CLG^{+}24, CHY^{+}24, CLY^{+}24, CLG^{+}21, CCZ^{+}21, CSZ^{+}23, GHM^{+}24, GLX^{+}22, GCL^{+}22, GZS^{+}25, HLO^{+}21, HLW^{+}21a, HYL^{+}23, JHB24, KMBR21, KK23, KSVR23, KOA22, LHRX20, LCX^{+}22, LLL^{+}24a, LTH^{+}21, LXC^{+}22, LPL23, LWX^{+}23, LYN^{+}20, MHW^{+}21, MHZ^{+}22, MZC^{+}22b, MHM22, OMD^{+}21, PZL^{+}22, PBC^{+}21, QZCZ21, SCL21a, TCJ22a, WLY22, WHM^{+}24, WWJL24, WXH^{+}24, XWDC23, YCA^{+}20, YS22, YSG^{+}22, YBY^{+}22, YZWT20, YDL23, YWL^{+}24, ZWL^{+}21, ZGQ^{+}21, ZLW^{+}23, ZRXF23, LWY^{+}22]. DeepBoot [CZZY23]. DeepCAT [DWZ^{+}24]. DeepSlicing [ZZQ^{+}21]. DeepTM [ZRC^{+}24]. Defeat [LJZY20]. Defect [GLF^{+}21]. Defense [ZYX^{+}22]. DeFi [LXGY23]. Defined [CWY^{+}23, HO23, JLY^{+}23b, LLL^{+}21b, WNA^{+}20, XZJ^{+}20, YWH^{+}21]. DegaFL [HHJ^{+}25]. Degree [CTBT21]. Delay [CZJ^{+}22, LLX^{+}22, RXL^{+}20]. Delay-Aware [CZJ^{+}22, RXL^{+}20]. Delay-Sensitive [LLX^{+}22]. Delaying [TGFPRA22]. Delays [OS20]. Delivery [LLP^{+}23, WJL^{+}24]. Delta [JWZ^{+}23, ZYF^{+}20]. Demand [CJLW22, NJG^{+}22, OAM^{+}25, SCS^{+}23, YZSX23, ZLT^{+}24, ZLCL20]. Demystifying [LN24]. Dense [LAC24]. Density [GLL^{+}20]. Dependable [WLY^{+}20]. Dependencies [AB21]. Dependent [CHY^{+}24, MLX23]. Deployed [LXGY23]. Deployment [CZZM24, FZC^{+}22, HWW^{+}23, JLW^{+}24, JTX^{+}22, LSY21, LLL^{+}23a, LLJC21, LWY^{+}22, SMCH20, SFW^{+}24, ZYD^{+}23, ZXL^{+}24]. DePo [MZW22]. Depthwise [LZW22b].

Derivation [BM20]. Descent [EFME24, LPG^{+}22, LCCZ20b, PHY20, ZKP20]. Design [CJLW22, HPB21, HZK^{+}24, HFW^{+}21, HC24, HCA24, JYF^{+}24, JWZ^{+}23, KK23, MTT^{+}22, ONP^{+}23, QXL^{+}20, SL20, TPV20, WL20, WVSL23, WHL^{+}23, WYZ^{+}24, XZJ^{+}20, ZFH22, ZF23, HBG^{+}22]. Designing [SJLN20, WLH20b]. Detection [BMMB22, CSZ^{+}23, DSCL21, FWT^{+}24, GLF^{+}21, GLL^{+}21, LLL^{+}21b, LPH^{+}24, NISJS21, OSF22, PWX^{+}23, PK21, ZHX^{+}22, ZSFX23]. Deterioration [WGW^{+}23]. Determinism [CTBT21]. Deterministic [GN22, XLL^{+}20b]. DeTraS [TGFPRA22]. Development [CQW^{+}20, PSK^{+}22]. Device [FC23, LWL^{+}24, LJM^{+}23, PLJK22, ZYD^{+}23]. Devices [CYZ^{+}24, HTL^{+}24, JZLZ24, LLS^{+}24, LDJX^{+}23, MAOA22, YZSX23]. DF [XWDC23]. DH [JJ22]. DH-SVRF [JJ22]. DHash [WLF^{+}22]. DHT [HNKO20]. DHT-Based [HNKO20]. Diabetic [MMGR23]. Diagnosability [LHXH22]. Diagnosis [ASS^{+}24, HLH22, SLHH23]. Diego [GGL^{+}23, LAG^{+}22]. DIESEL [WLY22]. Difference [ZLYL22]. Differential [HWS^{+}24]. Differentially [HLO^{+}21, ZLYL22, ZXG^{+}22]. Differentiated [FMP^{+}23]. Digital [JLL^{+}22, SYD^{+}24]. Digital/Analog [JLL^{+}22]. Dilithium [SYD^{+}24]. Dimensional [CYF^{+}23, CHM^{+}20, GLL^{+}20, LWL^{+}25, ZJS^{+}24]. Directed [LLDL23, WIBD22]. Directional [LCX^{+}24]. Directory [QJZF23]. Dirty [LZM^{+}20]. Disaggregated [LXL^{+}24, QJC^{+}24]. Discord [ZJGD21]. Discounting [LZJ^{+}20]. Discovery [CRZ^{+}23, GGL^{+}23, LCZ^{+}23, SGH^{+}23, SCA23]. Discrepancy [ZLD^{+}23]. Discrepancy-Aware [ZLD^{+}23]. Disjoint [KK24]. Disk [ZZH^{+}20a]. Dispatching [DZL^{+}21, MTL^{+}20]. Dissecting [RT25, SLG^{+}23]. Dissimilarity [ZCJY20]. Distance [MGG^{+}20, WXT^{+}24]. Distances [ZZH^{+}20b]. Distillation [BLYZ21, JBY^{+}23]. Distributed [AKA22, AKZ^{+}20, Ano20, ACC^{+}22, ASH^{+}22, CWC^{+}22, CLL22, CLG^{+}24, CYZ^{+}24, CWL^{+}21, CXL^{+}23, CLX^{+}23, CZL^{+}24, CDvK^{+}22, CWY^{+}23, CLZ^{+}20, CLZ^{+}24, DFJ^{+}23, DZL^{+}21, DTN^{+}22, DZS^{+}21, DSZ^{+}24, DLX^{+}24, FXL^{+}23, GMI^{+}22, GHM^{+}24, GLL^{+}20, GKK21, GBM20, GZJ^{+}21, GCL^{+}22, HND20, HSH^{+}22, HLS^{+}23, HC24, JLJ21, JHB24, JLW^{+}24, JWW^{+}22, JLY^{+}23a, JSL^{+}25, JQG^{+}22, KKS21, LLT^{+}23, LHRX20, LYL^{+}20b, LZJ^{+}20, LLC^{+}21, LWL^{+}22a, LPG^{+}22, LXGY23, LCLW21, LMX^{+}24, LWL^{+}25, LJT^{+}25, LPW^{+}20, LTH^{+}21, LHL23, LZH^{+}24, LLZ^{+}23, LLR^{+}24, LWL^{+}23, MGG^{+}20, MDM22, NDW^{+}21, OSF22, OQCW20, OZCW22, OHWL21, Par22, QZCZ21, SKW23, SGJ^{+}20, SCL21a, SKV^{+}20, SZZY24, SPS^{+}24, SDBM23, TW24, WHC^{+}21, WLY22, WMG^{+}23, WGN^{+}23, WLZ^{+}23, WLY^{+}20, WZL^{+}23, WNA^{+}20, WXH^{+}24, XXM^{+}20, XYW22, XZL20, XYL^{+}21, XWDC23, YLL21, YJWM24, YTL^{+}24, YSZL21, YDL23, YTL^{+}23, YYZ^{+}20, ZZD^{+}24a, ZXGZ21, ZXG^{+}22, ZW22a, ZPL^{+}22, ZJH^{+}23, ZYD^{+}23, ZYM^{+}20, ZJS^{+}24]. Distributed [ZLT^{+}24, ZSX^{+}20, ZCW^{+}20, ZWL^{+}21, ZGG21, ZGQ^{+}21]. Distributed-Memory [ASH^{+}22]. Distributing [JQG^{+}22]. Distribution [DLJ^{+}22, HZW^{+}25, KFS^{+}21, WYZ^{+}24, XCH^{+}21b, XLL^{+}20b]. Distributional [LYZS24, WHLM23]. Diverse [HZX^{+}23, ZMY^{+}24]. Diversity [BDS^{+}21]. Divide&Content [BSPM23]. Dividing [CZP^{+}23]. DL [WPZ^{+}21]. DL2 [PBC^{+}21]. DLRM [TJD^{+}24]. DLS [PLJK22]. DMA [LZDO24]. DMA-Assisted [LZDO24]. DNN [CZL^{+}22, DLX^{+}24, JWZ^{+}23, KY22, LLT^{+}23, LZQ^{+}23, LLK24, LLL^{+}24b, LMX^{+}24, PKRS23, WBS23, WUR^{+}24, WWJL24, XZL^{+}21, XXC^{+}23, YHD^{+}23, YZL24, ZHX^{+}22, ZW22b, ZYD^{+}23, ZLC^{+}22, ZCJ^{+}22, ZSL^{+}24, ZRC^{+}24]. DNN-Based [CZL^{+}22, ZHX^{+}22]. DNN-Driven [XZL^{+}21]. DNNs [LJT^{+}25, ZW22a]. Do [LZF^{+}24]. Docker [ZTA^{+}21]. Doctrine [XCL^{+}23]. Document [FWCB22]. Domain [CWY^{+}23, GMI^{+}22, LXQ^{+}24, LJW^{+}23]. Domain-Incremental [LXQ^{+}24]. Dominance [CTL24]. Dominant [CFLL21]. DONE [DTN^{+}22]. Downtime [CEP22]. DRFL [MMGR23]. Driven [AMvBI22, BM20, CZH^{+}20a, CZH^{+}20b, CDvK^{+}22, LHQ^{+}20, LPL23, MZC^{+}22a, OQCW20, PBC^{+}21, TXX^{+}21, UXL^{+}21, WLM^{+}20, XZL^{+}21, ZGM21, ZZZ^{+}24b]. Driving [MLX23]. DRL [ZDK^{+}22, ZMP23]. DRL-Based [ZDK^{+}22]. DRONE [ZJH^{+}23]. DS [ZXG^{+}22]. DS-ADMM [ZXG^{+}22]. DTN [YZC^{+}23]. Dual [KPHA20]. Duplicated [WDZ^{+}23]. Duplication [AA23, OS20]. Dynamic [BFK^{+}23, CAAB20, CZZY23, GWLZ21, HSY^{+}20, HZJH23, HCA24, HLNW22, KSB^{+}22, KEMC22, KPA^{+}20, LWL^{+}24, LK20, LNX^{+}22, LJT^{+}25, LMZ^{+}20, MWNK22, NFP^{+}20, NK21, NDW^{+}21, PK21, SMK^{+}23, SPCT23, WNL20, WLF^{+}22, WZL^{+}22, WVSL23, WZL^{+}23, ZZZ^{+}24a]. Dynamic/Static [GWLZ21]. Dynamically [MZW22]. Dynamics [DLY^{+}24, LCL^{+}24].

e-PoS [SQR^{+}21]. E2bird [CCZ^{+}21]. E3NE [GWG^{+}22]. EA [YSZL21]. EA-Based [YSZL21]. EcoFed [WUR^{+}24]. Economic [GLP^{+}21]. Economics [LHL^{+}22]. Economies [GLP^{+}21]. EDF [DYFL21]. Edge [AB21, BBGY20, BBG22, BM22, CZH^{+}20b, CZL^{+}22, CJLW22, CYZ^{+}24, CHY^{+}24, CGL^{+}22, CWL22, CXL^{+}23, CZL^{+}24, CZZM24, CLZ^{+}22b, DZL^{+}21, DLLL22, DTN^{+}22, FLPL22, FC23, FZC^{+}22, GWLX22, HLO^{+}21, HLW^{+}21a, HCG^{+}23, HCZ^{+}20, HLL22, HZW^{+}25, HXW^{+}20, JBY^{+}23, KSVR23, LLHJ20, LHC^{+}21, LLX^{+}22, LSL^{+}23, LZZ^{+}23, LNX^{+}22, LSY21, LXC^{+}22, LLL^{+}21c, LZW^{+}23, LJH^{+}23, LHX^{+}25, LHC^{+}25, LWY^{+}22, MLWX20, MZW22, MZC^{+}22a, MTL^{+}20, MHM22, MMR^{+}21, NDW^{+}21, OSF22, RDBG25, RXL^{+}20, RCW^{+}23, SCYJ21, SCS^{+}23, TCJ22b, WHM^{+}21, WGQ^{+}22, WHM^{+}23, WHM^{+}24, WQZD25, WHLM21, WXH^{+}24, XCH^{+}21b, XCH^{+}21a, XCH^{+}22, XCL^{+}23, XJX24, XZW^{+}24, XZL^{+}21, XXW^{+}24, YLC^{+}23, YHT^{+}23, ZZD^{+}24a, ZZG^{+}21b, ZQM^{+}22, ZHX^{+}22, ZYW^{+}23, ZXL^{+}24, ZLT^{+}24, ZLCW23, CZH^{+}20a]. Edge-Cloud [LLHJ20]. EdgeDR [CJLW22]. EdgeHydra [HZW^{+}25]. Edges [JQG^{+}22, YLL21]. Editor [Par20, Par21a]. Editorial [BZS21, Par21b, Par22, WR23, ZSP22]. EDP [SdR^{+}21]. EEPC [SLX20]. Effective [CGL^{+}22, CZZM24, DFJ^{+}23, LLJC21, SNN^{+}20, WYZY24, WQZD25, XCH^{+}21b, XZW^{+}24, YWZ^{+}20, ZXL^{+}24]. Effectiveness [KM23a]. Efficiency [HLW^{+}21b, KPA^{+}20, WYW^{+}22, ZWK^{+}20]. Efficiency-Boosting [HLW^{+}21b]. Efficient [ASMA21, ABBA23, BLK^{+}20, BPTV23, CXÖ^{+}20, CLL^{+}21, CZL^{+}22, CZP^{+}23, CCZW24, CXW^{+}24, CWL22, CLX^{+}23, CLZ^{+}22b, DS23, DSCL21, DDN^{+}22, DQC^{+}21, DFLG21, FLPL22, FZC^{+}22, GGZ^{+}20, GXW22, GHM^{+}24, GZJ^{+}21, GCL^{+}22, GZS^{+}25, GWLX22, HLW^{+}20, HLL22, HGA20, HTL^{+}24, HZX^{+}23, IRB21, JTX^{+}22, JYF^{+}24, JZLZ24, LLT^{+}23, LCZ^{+}20, LLL^{+}21b, LZX^{+}21, LYZC22, LZW^{+}22a, LLK22, LXL^{+}24, LGZ^{+}24, LLS^{+}24, LLK24, LLL^{+}24a, LGH^{+}24, LCLW21, LMX^{+}24, LWL^{+}25, LLL^{+}23b, LX23, LZH^{+}24, LWZ^{+}24, LZL^{+}24, LZM^{+}20, LWLX24, MWNK22, MWL24, NRB^{+}20, NLX^{+}22, QQD^{+}24, QLP^{+}23, RXL^{+}20, SLX20, SCL21a, SNK20, SJLN20, SZM20, TCT^{+}22, TWX22, WQKH20, WHC^{+}21, WGQ^{+}22, WBS23, WDCK23, WLH^{+}23, WXZ24, WPG^{+}22, WUR^{+}24, WWJL24, XWDC23, XLZD24, XLL^{+}20b, XWJ^{+}20, YLL21, YLC^{+}23, YZL24, YTL^{+}24, YZWT20, YT20, ZDK^{+}22, ZLWW20, ZCZ^{+}21, ZZG^{+}21b, ZJH^{+}23, ZZC^{+}23, ZLD^{+}23, ZZM^{+}23, ZSL^{+}23, ZXW^{+}24, ZSW^{+}22, ZLC^{+}22, ZCJ^{+}22, ZYX^{+}22, ZSX^{+}20, ZZH^{+}20b, ZLCL20, ZLRY22, ZRC^{+}24, ZLCW23, lM20]. Efficiently [GSH^{+}21]. Effort [TGFPRA20]. EiC [Par22]. Elastic [CCZ^{+}21, GLX^{+}22, GXC^{+}23, HYP^{+}22, QXL^{+}24, WDJ21, WPZ22, ZLR^{+}20]. ElasticBatch [QXL^{+}24]. Election [KM23b, SPZE20, SOI^{+}20]. Electric [KLH^{+}20b, TWYL20]. Electrical [ZGZ^{+}24]. Electroluminescence [GLF^{+}21]. Electrostatics [BDWH24]. Embedded [AMKS21, ASLPE20, ATF23, CBL22, GCL^{+}21, PZZ^{+}22, XZL20, YLW^{+}22]. Embedding [AAK22, FXL^{+}23, YLL21, ZLCW23]. Emerging [GWG^{+}22]. Empowering [YLS^{+}23]. Enable [TXG^{+}21]. Enabled [HTB22, MLWX20, RXL^{+}20]. Enabling [CCWW24, HKL^{+}20, HTL^{+}24, HZX^{+}23, LXL^{+}24, LGZ^{+}24, LZL^{+}24, LCL^{+}24, LJH^{+}23, LHC^{+}25, PGY^{+}22, TWX22, XXW^{+}24, YZL24, YYZ^{+}20, YHT^{+}23, ZZS^{+}22, ZCJ^{+}22, ZZSC20]. Encoded [PK21]. Encoding [CSJB20, GWG^{+}22, JSL^{+}25]. Encrypted [FWCB22, LYDZ21, SHZ^{+}23, WMG^{+}23, YYZ^{+}20]. Encryption [WLH^{+}23]. End [CLG^{+}21, DLLL22, GWG^{+}22, KSZ24, MLX23]. End-Edge [DLLL22]. End-to-End [CLG^{+}21, GWG^{+}22, KSZ24, MLX23]. Endpoint [LYL^{+}20b]. Endpoint-Flexible [LYL^{+}20b]. Energy [BBGY20, BLK^{+}20, CBB^{+}20, CZL^{+}22, IRB21, JHB24, KFS^{+}21, KPA^{+}20, KMM20, LLK22, LLK24, LWZ^{+}24, MWL24, Pil23, QWHC21, QWYG20, RT25, RCW^{+}23, SLX20, SWGB23, SNK20, SZZY24, SJLN20, SZ20, SWOM20, TRN^{+}21, WZL^{+}22, XZL^{+}21, YZL24, ZDK^{+}22, ZSL^{+}23, ZLT^{+}24]. Energy-Aware [BBGY20, JHB24, XZL^{+}21]. Energy-Efficient [BLK^{+}20, CZL^{+}22, LLK22, LLK24, LWZ^{+}24, MWL24, SLX20, SJLN20, YZL24, ZDK^{+}22]. Energy-Recycling [QWHC21]. Enforcing [LCL^{+}20]. Engine [DQC^{+}21, PZL^{+}22, RCLJT22]. Engineering [CWY^{+}23]. Engines [ALAK20, JJ22]. Enhanced [CCZ^{+}21, HLS^{+}23, WXT^{+}24, YS22]. Enhancement [XZL20]. Enhancing [LYK20, WXX^{+}24]. ENLARGE [QLP^{+}23]. EnosLib [CDvK^{+}22]. Ensemble [WHM^{+}23]. Ensembles [HLNW22]. Entry [QJZF23]. Enumeration [WHC^{+}21, WCT21]. Environment [CZJ^{+}22, HCZ^{+}20, JLQ^{+}23, JTX^{+}22, LJZY20, LHC^{+}21, MLX23, SMCH20, WGQ^{+}22, WQZD25]. Environment-Dependent [MLX23]. Environments [BOGM21, CZL^{+}22, CHY^{+}24, LLL^{+}21c, SPCT23, TLH22, TCJ22a, WVSL23, YRBÇ22]. Epistasis [NISJS21]. Era [AMN22, TLGA^{+}22, HLW^{+}20]. ERA-LSTM [HLW^{+}20]. Erasure [HZW^{+}25, HFW^{+}21, KZK^{+}19, KZK^{+}20, LXL^{+}24, LWC^{+}20, SL20, WCW^{+}23, XLL^{+}20b, XHQC20, ZKL24, ZFH22, ZF23]. Erasure-Coded [HFW^{+}21, LWC^{+}20, SL20, XLL^{+}20b, XHQC20, ZFH22, ZF23]. Errata [CZH^{+}20a, KZK^{+}20]. Error [CEP22, GLA20, WLL^{+}20, WGQ^{+}22, WPG^{+}24, YLL^{+}20, ZLX^{+}20]. Error-Bounded [WPG^{+}24]. Error-Compensated [WGQ^{+}22]. ESetStore [LWC^{+}20]. ESMD [DSW^{+}23]. ESMO [LSL^{+}23]. Essentiality [FXL^{+}23]. Estimation [CMLH20, GCL^{+}22, LLL^{+}23b, WLL^{+}20, WQKH20, WC20]. Estuary [JLWS24]. ETH [BK21, KSW^{+}22]. Ethernet [BDS^{+}21]. ETICA [ASMA21]. Evacuation [DZL^{+}21]. Evaluating [LSC^{+}20, MKJ^{+}22]. Evaluation [HFW^{+}21, HCA24, HLNW22, LMX^{+}22, SL20, WVSL23, WSLX22, XXW^{+}24, vV20]. Event [BPW^{+}23, KDREV21, LHQ^{+}20, MZW22, WLM^{+}20, ZWK^{+}20, ZMY^{+}24, ZTL24]. Event-Based [BPW^{+}23]. Event-Driven [LHQ^{+}20, WLM^{+}20]. Event-Triggered [KDREV21, ZTL24]. Events [ZLZ^{+}23]. Eventual [sKW22]. Evolution [DSZ^{+}24, LWZ^{+}22]. Evolving [LCL^{+}20]. EXA2PRO [PSK^{+}22]. Exact [KSZ24, MGG^{+}20]. Exascale [AMN22, KHOI20, SHC^{+}22, TLGA^{+}22, WHG^{+}22]. Excessive [IYAK23, SFW^{+}24]. Exchange [DMST20]. Exclusion [QQD^{+}24]. Execution [FZD^{+}24, IXS22, JLQ^{+}23, LMX^{+}24, LLC^{+}22, MWNK22, WDZ^{+}23]. Existence [QHC20]. Expanding [CLZ^{+}24]. Expansion [SYT20]. Expectation [AKZ^{+}20]. Expedited [ZSL^{+}24]. Expediting [ZYD^{+}23]. Expensive [MZW22]. Experiences [TSC^{+}24]. Experiment [CDvK^{+}22]. Experiment-Driven [CDvK^{+}22]. Explainable [KY22]. Exploitation [KFEG21, XHY^{+}24]. Exploiting [BADP22, FLW^{+}23, GSL^{+}20, LWZ^{+}24, PRL20, ZCJY20, ZKL24]. Exploration [CLL^{+}21, KBA^{+}24, SLY^{+}24, YCZC22]. Exploring [HC24, KZK^{+}19, KZK^{+}20, LDL22, LJZY20, LTT^{+}20, LHZ^{+}23, LHL23, PZZ^{+}22]. Expression [JLK^{+}20]. Extending [LTH^{+}21]. Extensible [CCZW24, PGY^{+}22]. Extension [SZS^{+}23]. Extensions [TLGA^{+}22]. Extra [LHXH22]. Extreme [LCX^{+}22, LCL^{+}24]. Extreme-Scale [LCX^{+}22, LCL^{+}24]. Extrinsic [WSM^{+}20].

FaaS [LMFK23]. Fabrics [NKP^{+}24].

Factorization [CFLY21, GLL22, HLB^{+}23, LLY^{+}20, ZXG^{+}22, ZKP20]. Factory [DFP23]. Failure [AB21, CKO^{+}21, HFW^{+}21, ZZH^{+}20a, ZSL^{+}24]. Failure-Atomic [CKO^{+}21]. Failures [HCR^{+}22, HO23]. Fair [ABBA23, BSPM23, DGE^{+}24, LYN^{+}20, SQR^{+}21, YSG^{+}22]. Fairness [CFLL21, CSZ^{+}23, HLW^{+}21b, XLZD24]. Falcon [ABBA23, LZS^{+}24]. Family [LDL22]. FarSpot [ZLK^{+}22]. Fast [GN22, HLH22, LMH^{+}20, LZZ21, LGZ^{+}24, LWC^{+}20, LZWW22, LLR^{+}24, NCB^{+}21, PLJK22, SBM24, TW24, WXHZ20, WHM^{+}21, XZJ^{+}20, YCZC22, ZKL24, WXX^{+}24]. Faster [DDX^{+}24, HSY^{+}20]. FastTuning [LGZ^{+}24]. Fault [GHG^{+}20, HZW^{+}25, JBLJ23, KZK^{+}19, KZK^{+}20, KM23b, LT20, LHXH22, SSS20, SLHH23, WHRL21, YOM21, ZGZ^{+}23, ZDL^{+}21, ZGG21, ZLCW23]. Fault-Tolerant [GHG^{+}20, HZW^{+}25, KZK^{+}19, KZK^{+}20, KM23b, LT20, SSS20, YOM21, ZGG21]. Faulty [ZHQ^{+}23]. FCDedup [SHZ^{+}23]. FCPP [ATF23]. FDIL [LXQ^{+}24]. Feasibility [CBL22, LYK20]. FeatherCNN [LMH^{+}20]. Feature [LWL^{+}23, ZZD^{+}24b]. Feature-Level [ZZD^{+}24b]. Fed [MGBD24]. Fed-RAC [MGBD24]. Federated [CLMW22, CWC^{+}23, CXW^{+}24, CXL^{+}23, CSZ^{+}23, DTN^{+}22, DLC^{+}21, DLJ^{+}22, FWT^{+}24, FLPL22, GWLX22, HHJ^{+}25, HAD^{+}22, HLW^{+}21b, HWS^{+}24, JWZ^{+}23, JBY^{+}23, LWL^{+}24, LWZ^{+}23a, LZW^{+}22a, LXQ^{+}24, LNX^{+}22, LCCZ20b, LJM^{+}23, LZW^{+}23, LCX^{+}24, LYN^{+}20, MHLM24, MHM22, MGBD24, MMGR23, NLX^{+}22, Pil23, QWHC21, SFYB21, SZCL23, TSLC23, UXL^{+}21, WHM^{+}23, WHL^{+}23, WXX^{+}24, WHLM21, WYW21, WGW^{+}23, WHLM23, WUR^{+}24, XLZD24, YFB^{+}23, YXDL24, ZLYL22, ZLD^{+}23, ZLP^{+}24, ZLRY22]. Federated-Learning [LZW^{+}22a].

FedGraph [CLMW22]. FedHAP [YXDL24]. FedMDS [ZLD^{+}23]. FedProf [WHLM23]. FedREM [LWL^{+}24]. FedSCR [WYW21]. FedVeca [LCX^{+}24]. Feluca [ZSH^{+}21]. FenceKV [TWX22]. FFT [LJZ^{+}20]. FHVAC [ZZD^{+}24b]. Field [GDZ^{+}20, GDS^{+}22]. FIFO [GN22]. File [ABBA23, CA20a, CGL^{+}22, FWCB22, HSH^{+}22, SKV^{+}20, ZZM^{+}23, ZCW^{+}20]. Filter [CLZ^{+}22a, LGZ^{+}21]. Filtering [WXHZ20]. Finding [AKG20, CQZ^{+}24]. Fine [HZW^{+}21, KHLZ20, LMFK23, LHZ^{+}23, PLJK22, WCW^{+}23, WSHJ23, YZSX23, ZZY^{+}21, ZLGZ23]. Fine-Grain [ZLGZ23]. Fine-Grained [HZW^{+}21, KHLZ20, LMFK23, LHZ^{+}23, WCW^{+}23, WSHJ23, YZSX23, ZZY^{+}21, PLJK22]. Finite [GGO21]. Fire [TGFPRA22]. Fireplug [NRB^{+}20]. First [PRL20]. Fitness [RCLJT22]. FL [CWC^{+}23]. Flash [KMLE20, KZK^{+}19, KZK^{+}20, ZCJY20]. Flexible [ACDK20, BCG23, DLJ^{+}22, FBD22, LDL22, LYL^{+}20b, LLS^{+}24, PLJK22, QXL^{+}20, QJZF23, WLH^{+}20a, ZKL24, ZCJ^{+}22]. Flexible-Scale [LDL22]. FlexRaft [ZKL24]. Flink [GSH^{+}21]. Flourishes [CWC^{+}23]. Flow [WNA^{+}20, YLS^{+}23, ZWK^{+}20, ZLL22a, ZSW^{+}22]. Fluctuated [ZDC^{+}23]. Fluid [LCL^{+}24]. FLUPS [BCG23]. Fog [BOGM21, DLMF22, Man22, SMK^{+}23, SHZ^{+}23, TCJ22a, WVSL23]. Fold3D [LZQ^{+}23]. Footprint [CBB^{+}20, LAC24]. Force [BDWH24, GDZ^{+}20, GDS^{+}22]. Forest [CWC^{+}22, LCM^{+}20, XWDC23]. Forest-Based [LCM^{+}20]. Fork [NAL^{+}20]. Format [CYZ^{+}23, SLKA23, ZZSC20]. Formation [NLX^{+}22]. Formats [QJZF23]. Forwarding [DQC^{+}21, JJ22, LLL^{+}21b, ZCL^{+}22]. Foundation [LLT^{+}23]. Fourier [BCG23]. FPGA [CCYC21, JMF22, LYL^{+}20a, LZP24, MWNK22, QXL^{+}20, QTR21, WGLZ20, ZZP23, ZKP20]. FPGA-Accelerated [LYL^{+}20a]. FPGA-Based [ZZP23, MWNK22]. FPGAs [BDWH24, GWG^{+}22, GWLZ21, HLS^{+}23, LZX^{+}21, SLX21a]. Fractional [EFME24]. Fragmented [ZFW^{+}20]. Frame [LSL^{+}23]. Framework [AD25, CRZ^{+}23, CXÖ^{+}20, CL20b, CGC^{+}22, DNLX25, FWCB22, GMI^{+}22, GHM^{+}24, GWG^{+}22, GCL^{+}21, GGL^{+}23, HZB^{+}24, HLB^{+}23, HCA24, JHB24, JWZ^{+}23, KFEG21, KLJV24, LLT^{+}23, LCZ^{+}23, LTZ^{+}23, LLL^{+}24a, LNX^{+}22, LZP24, MKKP22, PSK^{+}22, QLP^{+}23, SLX20, SGH^{+}23, SLY^{+}23, SCA23, SLY^{+}24, TMP23, TSLC23, WHC^{+}21, WJZ^{+}24, WLY^{+}20, XWDC23, YCA^{+}20, YLC^{+}23, YSZL21, YWH^{+}21, YRQ23, ZZD^{+}24a, ZCZ^{+}21, ZW22a, ZZS^{+}22, ZJH^{+}23, ZMP23, ZLD^{+}23, ZCJ^{+}22, ZGG21, ZJGD21]. Frameworks [DWZ^{+}24, vV20]. Free [LAC24, PHY20, XZY^{+}24, YWH^{+}20, ZSL^{+}21]. Freezing [CXW^{+}24]. Frequency [LJW^{+}23, SdR^{+}21, WC20]. Frequency-Domain [LJW^{+}23]. Freyr [YWL^{+}24]. Friend [LYZC22]. Friendly [LDJX^{+}23, TGFPRA22, ZLL^{+}22b]. Friendly-Fire [TGFPRA22]. FrodoKEM [GJCC21]. FRuDA [GMI^{+}22]. Frustrated [LCX^{+}22]. FT [ZGZ^{+}23, ZDL^{+}21]. FT-BLAS [ZGZ^{+}23]. FT-CNN [ZDL^{+}21]. Full [YCA^{+}20, YFD^{+}24, YTL^{+}23, QXL^{+}20]. Full-Spectrum [YFD^{+}24]. Full-Stack [YCA^{+}20]. Fully [AKZ^{+}20, GSL^{+}20, WLH^{+}23]. Function [DS22, HYL^{+}23, LLL^{+}23a, LLJC21, WCD^{+}24, WCN^{+}24, WDZ^{+}23, YLT^{+}21, YZS^{+}21]. Function-Reuse [DS22]. Functions [MSSK21, SP20]. Fundus [MMGR23]. Fusing [GK21]. Fusion [QJZF23, WSHJ23, ZSL^{+}21]. Future [TSC^{+}24]. Fuzziness [ZLR^{+}20]. Fuzzy [GLF^{+}21].

G [LZL^{+}24, PZL^{+}22, YCZC22]. G-Learned [LZL^{+}24]. G-SLIDE [PZL^{+}22]. GA [WLY^{+}20]. GA-Par [WLY^{+}20]. Galaxyfly [LDL22]. Game [DLLL22, HCZ^{+}20, HXW^{+}20, HWS^{+}24, LYDZ21, LLL^{+}21c, WSM^{+}20]. Game-Based [LLL^{+}21c]. Game-Theoretical [HCZ^{+}20]. Gamora [HYL^{+}24]. Gang [DYFL21]. Gap [LCX^{+}22]. Gateway [LPH^{+}24]. Gateways [SCS^{+}23]. Gathering [SNK20]. Gauss [AMKS21]. Gaussian [LGC^{+}22]. gcForest [CWC^{+}22]. GCN [HLZ^{+}21]. GCN-Based [HLZ^{+}21]. GEM [CZP^{+}23]. GEMM [AMW^{+}21, FBD22, KS23, WJZ^{+}24]. GEMM-Accelerated [KS23]. General [CYZ^{+}23, KAA21, SZS^{+}23, TPV20]. Generalized [HLS^{+}23, KKA^{+}20, WLF^{+}20]. Generating [GKK21]. Generation [BTL^{+}22, HFC^{+}23, LJZ^{+}20, LCX^{+}22, MZC^{+}22b]. Generic [BOGM21]. Genetic [JSL^{+}25]. Genome [CCYC21, WXZ24]. Genomic [GKK21, ZZZ^{+}24a]. Geo [CLL22, CZL^{+}24, HND20, JLW^{+}24, JQG^{+}22, LYL^{+}20b, LZJ^{+}20, LPW^{+}20, NRB^{+}20, OQCW20, OZCW22, SZZY24, SDZ21, WLY^{+}20, YTL^{+}24, ZYM^{+}20, ZSX^{+}20]. Geo-Distributed [CLL22, CZL^{+}24, HND20, JLW^{+}24, JQG^{+}22, LYL^{+}20b, LZJ^{+}20, LPW^{+}20, OQCW20, OZCW22, SZZY24, WLY^{+}20, YTL^{+}24, ZYM^{+}20, ZSX^{+}20]. Geo-Replicated [SDZ21]. Geo-Replication [NRB^{+}20]. GeoDeploy [JLW^{+}24]. Geodesic [lM20]. Geographically [YYZ^{+}20]. Geometries [HLVR21]. Georgia [PKJ^{+}22]. GeoScale [CZL^{+}24]. Geostatistical [ACP^{+}22]. Giant [LLT^{+}23]. gIM [SSH21]. Global [CGM21, DYFL21, DLC^{+}21, LBNN^{+}21, LWL^{+}22b, LCX^{+}24, SLY^{+}24, WLH^{+}20a, WNL20, YXDL24]. gMig [LZM^{+}20]. GML [GSH^{+}21]. GMRES [LLD22]. GNAS [CGC^{+}22]. GNN [LZZ^{+}23, LZP24, WGN^{+}23, WSHJ23, ZZP23]. GNN-Based [LZZ^{+}23]. GOSH [TCJ22a]. Gossip [HLW^{+}21a, HAD^{+}22]. Gossip-Based [HLW^{+}21a]. GossipFL [TSLC23]. Gossips [FMP^{+}23]. GPGPU [RP20, WC20]. GPGPUs [QZFZ20]. GPU [AAK22, BCVD23, CQW^{+}20, CMLH20, CZZY23, DMI^{+}23, DSCL21, GXW22, GM21, GCL^{+}22, HZX^{+}24, HJEV^{+}21, HLVR21, HW22, HZX^{+}23, HLB^{+}23, IXS22, JDD^{+}24, KSVR23, LC20, LSC^{+}20, LH22, LZL^{+}24, LYGG20, MDC24, NCB^{+}21, PZZ^{+}22, PZL^{+}22, QLP^{+}23, SSH21, SYD^{+}24, TSW^{+}21, TJD^{+}24, TPV20, WZL^{+}22, WSX^{+}23, WDCK23, WSHJ23, XXC^{+}23, YLL21, YCZC22, YLW^{+}22, YSG^{+}22, ZJHS20, ZLW20, ZZG^{+}21a, ZW22b, ZHP^{+}23, ZGNZ22, ZSH^{+}21, ZMS^{+}22, ZJGD21, ZLGZ23]. GPU-Accelerated [CMLH20, MDC24, ZMS^{+}22]. GPU-Based [PZL^{+}22]. GPUDirect [LSC^{+}20]. GPUs [ABG20, CLZP20, CFLY21, CTL24, FBD22, GLL22, GJCC21, HWF^{+}22, KS23, LZS^{+}24, LCCZ20a, LS21, LAY21, LTZ^{+}23, LZW22b, PWZ^{+}21, SNN^{+}20, SLY^{+}24, WQKH20, WLH^{+}23, WCT21, ZLWW20, ZCZ^{+}21]. gQoS [LYGG20]. Gradient [EFME24, HHJ^{+}25, LPG^{+}22, LCCZ20b, PHY20, WLZ^{+}23, YYL^{+}24, ZW22a, ZKP20, ZWL^{+}21]. Gradients [SCL21a, ZLRY22]. Grading [MMGR23]. Graft [WWJL24]. Grain [KAA21, ZLGZ23]. Grained [HZW^{+}21, JMF22, KHLZ20, LMFK23, LHZ^{+}23, SLX21a, WCW^{+}23, WSHJ23, YZSX23, ZZY^{+}21, PLJK22]. Graph [ASS20, AAK22, ACC^{+}22, CTBT21, CGC^{+}22, CLMW22, CC22, CC23, CWXC24, FTYL20, GZW^{+}22, GTH22, GZY21, HND20, HLLL22, HWF^{+}22, JWZ^{+}24, LZWL22, LTZ^{+}23, LXW^{+}23, LH22, LWL^{+}25, NRB^{+}20, SRD^{+}20, XXM^{+}20, XWJ^{+}20, YLL21, YCZC22, YTL^{+}24, ZZZ^{+}24a, ZYD^{+}23, ZSH^{+}21, ZSX^{+}20]. Graph-Centric [JWZ^{+}24]. Graph500 [GZW^{+}22]. GraphAGILE [ZZP23]. Graphs [BFK^{+}23, CMSV20, GBM20, HWW^{+}23, HSY^{+}20, LWZ^{+}23a, MAOA22, QHC20, WIBD22, WZL^{+}23, ZJH^{+}23, ZHQ^{+}23]. Greedy [YS22]. Green [YWZ^{+}20, ZLCL20]. GreenFlow [GZS^{+}25]. Greening [TRN^{+}21]. Grid [GGO21, YYL^{+}24]. Group [LBNN^{+}21, WQKH20, ZHQ^{+}23]. GRP [FSF^{+}20]. GRP-HEFT [FSF^{+}20]. gSoFa [GLL22]. Guarantee [HLW^{+}21b, KLH^{+}20b, YW20]. Guarantees [CQW^{+}20, WWJL24]. Guest [BZS21, Par21b, WR23, ZSP22]. Guided [GSH^{+}21, LWL^{+}24].

H5Intent [DHM25]. Hamiltonian [ZLCW23]. Handle [BGZR21]. Hard [ZLGZ23]. Hardware [AMW^{+}21, DMST20, DNKB20, FFQ^{+}22, JYF^{+}24, ONP^{+}23, TGFPRA20, TGFPRA22, XMW^{+}22, ZGNZ22]. Hardware-Software [JYF^{+}24]. Harnessing [DS22]. Harvesting [YWL^{+}24]. Hash [WLF^{+}22]. HashCache [WDZ^{+}23]. HashFlow [ZSW^{+}22]. Hashing [LC20, Nak21, YXDL24]. HDF5 [DHM25, SHC^{+}22]. HDFIT [ONP^{+}23]. HE-Booster [WLH^{+}23]. Heavy [LPH^{+}24]. Hedonic [NLX^{+}22]. HEFT [FSF^{+}20]. Heterogeneity [HPB21, HLZ^{+}20, MGBD24, YZJ^{+}21, ZGQ^{+}21]. Heterogeneity-Aware [HPB21, HLZ^{+}20, YZJ^{+}21, ZGQ^{+}21]. Heterogeneous [ASLPE20, AP20, CAAB20, CMSV20, DS23, DFLG21, HLW^{+}21a, HFC^{+}23, HNKO20, HXW^{+}20, HLLL22, JSL^{+}25, KFS^{+}21, KKA^{+}20, KLJV24, KHOI20, KMM20, LYL^{+}20a, LZP24, LJT^{+}25, LSY21, LCL^{+}24, MAOA22, PSK^{+}22, PLJK22, QWYG20, SZ20, TSW^{+}21, TCT^{+}22, WCD^{+}24, XLL^{+}20a, XMW^{+}22, YZSX23, YRBÇ22, YOM21, YZL^{+}20, ZZH^{+}20a, ZFY^{+}20, ZGM21, ZMP23, ZHH^{+}24, ZCL^{+}22, ZF23, ZRC^{+}24]. HeteroYARN [LYL^{+}20a]. HEVC [CIZ^{+}20]. HGNNs [XHY^{+}24]. HI [JDD^{+}24]. HI-Kyber [JDD^{+}24]. Hidden [WZY^{+}22]. Hidden-State-Aware [WZY^{+}22]. Hiding [AKA22]. Hierarchical [AP20, FLPL22, GGA24, HGA20, LNX^{+}22, LLL^{+}23b, NLX^{+}22, WXHZ20, XZY^{+}24, YFB^{+}23]. Hierarchical-Reduction [LLL^{+}23b]. Hierarchically [HZX^{+}23]. High [ACP^{+}22, ASS20, AMW^{+}21, BDS^{+}21, BMMB22, BPP21, CYF^{+}23, CLZ^{+}22a, CA20a, CXÖ^{+}20, CHM^{+}20, CCYC21, CWC^{+}22, DNKB20, GLL^{+}20, GLW^{+}21, GLP^{+}21, GXC^{+}23, JLK^{+}20, JDD^{+}24, JJ22, KLH^{+}20a, LZS^{+}24, LJZ^{+}20, LHQ^{+}20, LCZ^{+}20, LLK22, LLK24, LZP24, LAC24, MFYB22, MZC^{+}22b, MKKP22, PSK^{+}22, QZFZ20, RCLJT22, SYD^{+}24, SLLL20, WL20, WZY^{+}22, YZL24, ZZZ^{+}24a, ZGZ^{+}23, ZLW20, ZYL^{+}20, ZZS^{+}22, ZHQ^{+}23, ZCJ^{+}22, dBMH21]. High-Dimensional [CYF^{+}23, CHM^{+}20, GLL^{+}20]. High-Level [GXC^{+}23, dBMH21]. High-Performance [AMW^{+}21, BDS^{+}21, CLZ^{+}22a, CXÖ^{+}20, CWC^{+}22, DNKB20, GLW^{+}21, GLP^{+}21, JLK^{+}20, JDD^{+}24, JJ22, LJZ^{+}20, LHQ^{+}20, LLK22, LLK24, MZC^{+}22b, SLLL20, WL20, YZL24, ZZZ^{+}24a, ZZS^{+}22, dBMH21]. High-Precision [WZY^{+}22]. High-Productivity [ACP^{+}22]. High-Quality [ASS20]. High-Speed [CA20a]. High-Throughput [CCYC21, LZP24, MKKP22, RCLJT22, SYD^{+}24]. Higher [ZSW^{+}20]. Highly [CHM^{+}20, LWC^{+}22, MAOA22, YWZ^{+}20, ZCW^{+}20]. Highly-Variable [CHM^{+}20]. HiHGNN [XHY^{+}24]. History [WXT^{+}24]. HitGNN [LZP24]. Hitter [LPH^{+}24]. Holistic [HLZ^{+}20, JZLZ24, LCM^{+}20]. Homogeneous [YWS^{+}23]. Homomorphic [AKZ^{+}20, ÇSS21, WLH^{+}23]. Hone [LLC^{+}21]. Hop [SCYJ21]. Hopscotch [JYF^{+}24]. Horizontal [ZLP^{+}24]. Horus [YBY^{+}22]. Hotness [GZY21, HZK^{+}24]. Hotness-Aware [HZK^{+}24]. Hotspot [JLL^{+}20]. Hotspot-Aware [JLL^{+}20]. HPC [ASS^{+}24, BDS^{+}21, GGHP21, HZB^{+}24, HMM22, KFS^{+}21, KMA^{+}20, ONP^{+}23, TSC^{+}24, YS22, ZLJ^{+}23, ZMS^{+}22, ZLK^{+}22]. HPETC [WXT^{+}24]. HPPT [HGA20]. HPPT-NoC [HGA20]. HRHS [DNKB20]. HSA [WZY^{+}22]. HSA-Net [WZY^{+}22]. Hua [SCL^{+}21b]. Human [CCYC21]. HW [GWLZ21]. HW/SW [GWLZ21]. Hybrid [CXL^{+}23, GLF^{+}21, HLZ^{+}20, JLL^{+}20, JLL^{+}22, KMBR21, LWZ^{+}22, LLL^{+}24b, LLL^{+}24a, LYDZ21, LHL23, OMD^{+}21, TWY^{+}20, WWJL24, XWJ^{+}20, ZHP^{+}23, ZZD^{+}24b, ZGZ^{+}24, ZGQ^{+}21, ZSFX23]. HybridChain [TW24]. Hyper [LZZ21, LGZ^{+}24]. Hyper-Parameter [LGZ^{+}24]. Hyper-Parameters [LZZ21]. Hypercube [WLF^{+}20]. Hypergraph [XRS^{+}23]. Hypergraph-Based [XRS^{+}23]. Hypergraphs [HZJH23].

I/O [ASMA21, BPTV23, GCL^{+}21, HCR^{+}22, HLZ^{+}20, KLH^{+}20a, LCZ^{+}20, LZDO24, LXC24, SHC^{+}22, TKRB22, WHG^{+}22, WYZY24, XWJ^{+}20, YYW^{+}20]. I/O-Efficient [XWJ^{+}20]. IaaS [FSF^{+}20, LHPW20, WLH20b]. iBalancer [ZLL22a]. IBM [NFP^{+}20]. iCELIA [YCA^{+}20]. ICN [LYK20]. Identical [WZGM23]. Identification [HFW^{+}21]. Identifying [CTBT21]. Idle [YWL^{+}24]. IEEE [AMN22, Ano20]. iGniter [XXC^{+}23]. II [LSZ^{+}21]. IID [LCX^{+}24, ZLRY22]. III [ZRXF23]. IIoT [YLC^{+}23]. Image [CGH^{+}22, KY22, WLY22, XRS^{+}23]. Images [GLF^{+}21, MMGR23, ZTA^{+}21]. Imaging [WLM^{+}20, HBG^{+}22]. Imbalance [CGH^{+}22]. Imbalanced [DLC^{+}21, HYP^{+}22]. iMLBench [ZZG^{+}21a]. Immature [CXW^{+}24]. Impact [GGHP21, LK20, NKP^{+}24, OAM^{+}25, PM22, ZWK^{+}20, ZLZ^{+}23]. Impacts [YYW^{+}20].

Implementation [DSW^{+}23, HPB21, JDD^{+}24, KK23, SYD^{+}24, ZGZ^{+}23]. Implementations [GXW22, GLP^{+}21, WDCK23]. Implementing [IATB20]. Implications [ZTA^{+}21]. Improve [FSPE20, KPA^{+}20, LXC24, LHL23]. Improved [HFW^{+}21, Nak21, ZRXF23]. Improvement [LCM^{+}20]. Improving [CCZ^{+}21, DMI^{+}23, GWLZ21, KLH^{+}20a, KM23a, TWY^{+}20, TXX^{+}21, WHG^{+}22, WSHJ23, ZCJY20, ZFW^{+}20, ZYF^{+}20]. In-Band [SPS^{+}24]. In-Line [ZYF^{+}20]. In-Memory [JLL^{+}20, LHZ^{+}23, QXL^{+}20, XHQC20, GGA24]. In-Network [LYK20, LTT^{+}20, ZPL^{+}22]. In-Place [SSS20]. in-Server [ZLL22a]. In-Storage [JLK^{+}20]. Incentive [LHL^{+}22, WSM^{+}20, WHL^{+}23, ZLCL20]. Incomplete [HXW^{+}20]. Incremental [LXQ^{+}24, SMK^{+}23]. Independent [QHC20]. Index [Ano20, HZK^{+}24, LZL^{+}24, LZL^{+}24]. Indexing [SKV^{+}20, WYW^{+}22]. Induced [CEP22]. Industrial [ZYX^{+}22]. Inference [CZZY23, DZS^{+}21, FC23, GLW^{+}21, HZX^{+}24, JZLZ24, KM23a, LMH^{+}20, LLK22, LH22, LJW^{+}23, LLR^{+}24, MHZ^{+}22, QXL^{+}24, WWJL24, XZL^{+}21, XXC^{+}23, YZSX23, ZZQ^{+}21, ZZP23, ZCHZ23, ZLW^{+}23]. Influence [GK21, SSH21]. Influence-Maximization [GK21]. Information [DLMF22, DNLX25, HXW^{+}20, SLY^{+}23, UXL^{+}21, ZSFX23]. Information-Coordination [DLMF22]. Infrastructure [LJT^{+}25, Nak21, SMSK21, ZJS^{+}24]. Infrastructure-Less [ZJS^{+}24]. Infrastructures [VMT^{+}20]. Initiative [WR23]. Inline [HKL^{+}20]. Innovative [AMN22]. Input [LAY21, WJZ^{+}24]. Input-Aware [WJZ^{+}24]. Inspection [LIP^{+}21]. Inspired [HGA20, WQZD25]. InSS [HZX^{+}24]. Instability [SWGB23].

Instability-Resilient [SWGB23]. Instantly [Mei23]. Instructions [CBB^{+}22]. INT [SPS^{+}24, YLS^{+}23]. INT-Label [SPS^{+}24]. Integer [WRLS22]. Integer-Only [WRLS22]. Integrated [KAT^{+}20, SKV^{+}20, TMP23, ZZY^{+}21, ZZG^{+}21a]. Integrating [OS20]. Integration [AMW^{+}21]. Integrity [CA20a, DFJ^{+}23, LHC^{+}21, LCH23, ZZC^{+}23]. Intel [PRL20, SWOM20]. Intelligence [FC23, JBY^{+}23, LNX^{+}22]. Intelligent [CLMW22, GCL^{+}22, HZX^{+}24, LWZ^{+}22, LJM^{+}23, SCS^{+}23]. Intensive [GXC^{+}23, LK20, NAL^{+}20, XLL^{+}20a, YZWT20]. Intent [DHM25]. Inter [DNLX25, SFW^{+}24]. Inter-Datacenter [DNLX25, SFW^{+}24]. Interaction [MR24, ZCL^{+}22]. Interactions [GDS^{+}22]. Interconnect [LSC^{+}20]. Interconnection [HLH22, LDL22, WLF^{+}20]. Interconnects [LLK22]. Interface [LXC24, ZWL^{+}21]. Interfaces [KBA^{+}24]. Interference [LDZ^{+}24, MXS21, WYZY24, XXC^{+}23, YBY^{+}22, ZGM21]. Interference-Aware [MXS21, XXC^{+}23, YBY^{+}22, ZGM21]. Interleaved [MAOA22]. Intermittent [SLHH23]. Interprocedural [GZJ^{+}21]. Interval [RZLT20]. Intra [KLH^{+}20a, SFW^{+}24]. Intra- [SFW^{+}24]. Intra-Node [KLH^{+}20a]. Intrusion [CSZ^{+}23]. Intrusive [SdR^{+}21]. Invalidation [ZYK^{+}22]. Investigating [LYDZ21]. IO [BPTV23, HZB^{+}24]. IO-Sets [BPTV23]. IoT [CZL^{+}22, DLMF22, FWT^{+}24, LLX^{+}22, LXC^{+}22, LDJX^{+}23, LLR^{+}24, MMR^{+}21, WNA^{+}20, ZJS^{+}24]. IPPTS [DS23, DFLG21]. IrGEMM [WJZ^{+}24]. IRIS [KLJV24]. Irregular [LHL23, WJZ^{+}24]. Irrevocability [TGFPRA20]. ISA [SZS^{+}23, SZ20, YWS^{+}23]. Isolation [FSPE20]. Issue [AMN22]. Issuing [LWL^{+}22a]. Iterative [BCVD23, DMPR22, ZGG21].

Jacobi [AMKS21]. Jacobi-Embedded [AMKS21]. Job [AD25, BM20, KKA^{+}20, KMA^{+}20, LJM^{+}23, RZLT20, YZWT20]. Jobs [CLL22, GCL^{+}22, HND20, HLZ^{+}21, LZJ^{+}20, LYZS24, ZDC^{+}23, ZLR^{+}20]. Join [ABC^{+}24, NAL^{+}20]. Joint [CLY^{+}24, HWW^{+}23, JXX^{+}23, LSL^{+}23, LJT^{+}25, LSY21, RCW^{+}23, WHL^{+}23, WCD^{+}24, WCN^{+}24, XTH^{+}23, ZZG^{+}21b]. Jump [JLY^{+}23b]. Just [JMF22]. Just-In-Time [JMF22].

K-Asynchronous [ZLRY22]. K-Athena [GGO21]. KEM [GXW22]. Kerbosch [WCT21]. Kernel [CXÖ^{+}20, KOA22, SYT20, WSHJ23]. Kernels [CTBT21, FBD22, GK21, LJZ^{+}20, MZC^{+}22b, SNN^{+}20, WQKH20]. Key [ACH^{+}20, JLY^{+}23a, JLL^{+}20, LCLW21, LCM^{+}20, PM22, QXL^{+}20, SDZ21, TWX22, ZZM^{+}23]. Key-Value [LCLW21]. Khatri [AAA21]. KLNK [CLZ^{+}24]. Knowledge [BLYZ21, JBY^{+}23, MHLM24]. Known [CL20a, MD22]. Kokkos [TLGA^{+}22]. KV [QXL^{+}20]. KVM [IYAK23]. Kyber [GJCC21, JDD^{+}24].

Label [ABG20, SPS^{+}24]. Labeling [SPS^{+}24]. Labels [XYW22]. Lambda [SMSK21]. Landing [GLL^{+}21]. Landlord [SPCT23]. Large [CBB^{+}20, CLL^{+}21, CQZ^{+}24, CGH^{+}22, DSW^{+}23, DSZ^{+}24, GLP^{+}21, GZY21, GZJ^{+}21, HWF^{+}22, JWZ^{+}24, KSB^{+}22, KKA^{+}20, LYZC22, LZQ^{+}23, LH22, LHXH22, LMX^{+}24, LZF^{+}24, LHPW20, MCT21, OMD^{+}21, PH21, TSW^{+}21, WLP^{+}23, WZL^{+}23, YHS^{+}20, YZL^{+}20, ZZH^{+}20a, ZJH^{+}23, ZYW^{+}23, ZTA^{+}21, ZSL^{+}24, ZSX^{+}20, ZCW^{+}20, ZSFX23]. Large-Scale [CBB^{+}20, CLL^{+}21, CQZ^{+}24, CGH^{+}22, DSW^{+}23, DSZ^{+}24, GZY21, GZJ^{+}21, HWF^{+}22, JWZ^{+}24, KSB^{+}22, LYZC22, LHXH22, LZF^{+}24, PH21, TSW^{+}21, WLP^{+}23, YHS^{+}20, YZL^{+}20, ZJH^{+}23, ZYW^{+}23, ZTA^{+}21, ZSL^{+}24, ZCW^{+}20, ZSFX23]. Latency [AKA22, Ans20, CZR20, JLY^{+}23a, KAA20, LWZ23b, LGH^{+}24, LSY21, LLP^{+}23, NAL^{+}20, OHWL21, QXL^{+}20, SLG^{+}23, TLQ^{+}20, WQZD25, ZZQ^{+}21, ZLL22a, ZZP23]. Latency-Critical [LGH^{+}24]. Latency-Sensitive [LLP^{+}23]. Lattice [GVD^{+}22, HLVR21, LZS^{+}24, LCL^{+}24]. Lattice-Based [LZS^{+}24]. Launch [RZLT20]. Launching [DDX^{+}24]. law [ZJH^{+}23]. Layer [ALAK20, LFZ^{+}21, Mei23, ZBB^{+}22, ZYX^{+}22]. Layout [WHG^{+}22]. LB [LWZ23b]. LB-Chain [LWZ23b]. LB4OMP [KEMC22]. Leader [KM23b, SPZE20, SOI^{+}20]. Leadership [MTT^{+}22]. Leadership-Class [MTT^{+}22]. Leakage [WLZ^{+}23]. Leap [CYZ^{+}23]. Leap-Format [CYZ^{+}23]. Learned [LZL^{+}24, LZL^{+}24]. Learning [ASS^{+}24, AB21, CRZ^{+}23, CQW^{+}20, CZH^{+}20a, CZH^{+}20b, CHM^{+}20, CLMW22, CZZY23, CWC^{+}23, CLG^{+}24, CYZ^{+}24, CHY^{+}24, CXW^{+}24, CLY^{+}24, CLG^{+}21, CLZ^{+}21, CXL^{+}23, CIZ^{+}20, CCZ^{+}21, CSZ^{+}23, DTN^{+}22, DLC^{+}21, DLJ^{+}22, DSZ^{+}24, EFME24, FWT^{+}24, FLPL22, GHM^{+}24, GLL^{+}21, GCL^{+}22, GZS^{+}25, GSH^{+}21, GWLX22, GGL^{+}23, HLO^{+}21, HLW^{+}21a, HHJ^{+}25, HZW^{+}21, HAD^{+}22, HYL^{+}23, HYL^{+}24, HLW^{+}21b, HWS^{+}24, IRB21, JHB24, JXX^{+}23, JWZ^{+}23, KSZ24, KMBR21, KFEG21, KK23, KSVR23, KKP21, KOA22, LWL^{+}24, LHRX20, LWZ^{+}23a, LZJ^{+}20, LCCZ20a, LCX^{+}22, LZW^{+}22a, LCZ^{+}23, LZZ^{+}23, LYZS24, LXQ^{+}24, LNX^{+}22, LCCZ20b, LTH^{+}21, LXC^{+}22, LJM^{+}23, LPL23, LWX^{+}23, LZW^{+}23, LLZ^{+}23, LWY^{+}22, MHLM24, MZC^{+}22b, MHM22, MGBD24, MMGR23, NLX^{+}22, OMD^{+}21, PZL^{+}22, PBC^{+}21, Pil23, QXL^{+}24, QZCZ21, QWHC21, SGJ^{+}20, SFYB21, SCL21a, SGH^{+}23, SCA23, SDBM23, SZCL23, TW24, TSLC23, UXL^{+}21, WL20, WNL20, WGLZ20, WHM^{+}21, WLY22, WZL^{+}22, WZHW22, WHM^{+}23, WHL^{+}23, WHM^{+}24, WXX^{+}24, WLH20b]. Learning [WHLM21, WYW21, WHLM23, WYZ^{+}24, WUR^{+}24, WWJL24, WXH^{+}24, XJX24, XXM^{+}20, XYW22, XLZD24, YCA^{+}20, YS22, YFB^{+}23, YJWM24, YSG^{+}22, YBY^{+}22, YZWT20, YDL23, YWL^{+}24, ZJHS20, ZLYL22, ZZD^{+}24a, ZLWW20, ZZH^{+}20a, ZCZ^{+}21, ZZG^{+}21a, ZGM21, ZLD^{+}23, ZLP^{+}24, ZDC^{+}23, ZYX^{+}22, ZWL^{+}21, ZGQ^{+}21, ZLRY22, ZLW^{+}23]. Learning-Augmented [QXL^{+}24]. Learning-Based [HYL^{+}24, KMBR21, LWX^{+}23, LLZ^{+}23]. Learning-Driven [PBC^{+}21, ZGM21]. Least [CFLY21]. Ledger [LWL^{+}22a]. Legacy [CFM^{+}21a]. LegoSwap [LLS^{+}24]. Less [LZF^{+}24, ZJS^{+}24]. Level [ASMA21, AHSW23, BSPM23, DLJ^{+}22, DSZ^{+}24, GSL^{+}20, GXW^{+}20, GCL^{+}21, GXC^{+}23, HSH^{+}22, IATB20, KMLE20, SHZ^{+}23, ZSL^{+}21, ZZD^{+}24b, dBMH21]. Level-Based [AHSW23]. Leveraging [BPP21, HYL^{+}23, ZZH^{+}20b]. libEnsemble [HLNW22]. Liberator [LTZ^{+}23]. Library [BCG23, CBB^{+}20, CDvK^{+}22, HLNW22, KEMC22, SP20]. Light [IXS22]. Light-Weight [IXS22]. LightFed [GWLX22]. Lightweight [CZR20, CGL^{+}22, HLLL22, SZS^{+}23, SPS^{+}24, WYZ^{+}24, ZHX^{+}22]. Like [XRS^{+}23]. Limited [GBM20]. Line [PWX^{+}23, ZYF^{+}20]. Linear [BCVD23, CSJB20, HLS^{+}23, LRBV23, LAC24, PZL^{+}22, SZS^{+}23]. LINPACK [TSW^{+}21]. Liquid [GCL^{+}22]. List* [Ano21]. Live [CQZ^{+}21, CWXC24, HTB22, LZM^{+}20, ZYW^{+}23, ZZD^{+}24b]. Load [AAA21, Ans20, CAAB20, CGH^{+}22, DZL^{+}21, DFXY20, KBA^{+}24, KEMC22, LWZ23b, LYZS24, LXC^{+}22, MGG^{+}20, SPPS20, XZY^{+}24, YWH^{+}20, ZLL22a]. Load-Aware [ZLL22a]. Load-Balanced [LWZ23b, XZY^{+}24, YWH^{+}20]. Load-Balancing [LXC^{+}22]. Loads [SFW^{+}24]. Local [CGM21, DNLX25, GLL^{+}20, WXX^{+}24]. Local-Density [GLL^{+}20]. Locality [CWXC24, CWL^{+}21, FTYL20, HNKO20, KFEG21, KS23]. Locality-Aware [FTYL20, HNKO20, KS23]. Locality-Preserving [CWXC24]. Localization [LGH^{+}24]. Localized [DDN^{+}22]. Locally [WSLX22]. Location [SMCH20]. Location-Aware [SMCH20]. Lock [PHY20]. Lock-Free [PHY20]. Locking [GXW^{+}20]. LOCUS [CZJ^{+}22]. LOFS [CGL^{+}22]. Log [DDN^{+}22]. Logging [DDN^{+}22]. Logical [LCIS24]. Long [BDWH24, CTD^{+}23, CZZ^{+}22, HLW^{+}20]. Long-Range [BDWH24, CTD^{+}23]. Long-Term [CZZ^{+}22]. LoomIO [HSH^{+}22]. Loops [COE20, KKP21, MDM22]. Lossless [CTD^{+}23, TLQ^{+}20, WGBS23]. Lossy [SZCL23, SYS^{+}22, WLL^{+}20, WPG^{+}24, ZLX^{+}20]. Low [Ans20, CZR20, DFXY20, DWZ^{+}24, JLWS24, JLY^{+}23a, LJZY20, LWZ23b, OHWL21, QXL^{+}20, WL20, WLF^{+}20, WQZD25, XWDC23, YRQ23, ZZQ^{+}21, ZZP23, ZSFX23]. Low-Cost [DFXY20, DWZ^{+}24]. Low-Latency [LWZ23b, WQZD25, ZZP23]. Low-Memory [ZSFX23]. Low-Power [WL20]. Low-Rate [LJZY20]. LRU [ZYK^{+}22]. LSH [PZL^{+}22, ZPL^{+}22]. LSH-Based [ZPL^{+}22]. LSTM [HLW^{+}20].

Machine [BADP22, CLZ^{+}21, FZD^{+}24, GLL^{+}21, GSH^{+}21, HZW^{+}21, HWF^{+}22, IRB21, JXX^{+}23, KFEG21, KKP21, LZJ^{+}20, LZX^{+}21, LZZ^{+}23, LLZ^{+}23, MR24, RZLT20, WDL^{+}20, WNL20, WGLZ20, WZL^{+}22, XXM^{+}20, ZDK^{+}22, ZCZ^{+}21, ZZG^{+}21a, ZZD^{+}24b]. Machine-Centric [ZZD^{+}24b]. Machine-Learning-Based [KFEG21].

Machines [GKK21, HTB22, KPHA20]. Magnetohydrodynamics [GGO21]. Maintenance [HSY^{+}20, HZJH23]. Makespan [YOM21]. Making [KMLE20, SQR^{+}21]. Malicious [WIBD22]. Malleability [TSC^{+}24]. Malware [GLL^{+}21]. Manage [WHRL21]. Management [ASLPE20, BPTV23, CGLC20, CIZ^{+}20, HTB22, HTL^{+}24, HCA24, IRB21, JLL^{+}20, KSP^{+}20, LSY21, SGJ^{+}20, SDHQ21, TXX^{+}21, WZHW22, WYZY24, WNA^{+}20, YLC^{+}23, ZRC^{+}24]. Manager [BSPM23]. Many [KK24, KKS21, LX23, YZL^{+}20, YZS^{+}21, ZFY^{+}20]. Many-Camera [KKS21]. Many-Core [LX23, YZL^{+}20, YZS^{+}21, ZFY^{+}20]. Many-to-Many [KK24]. Manycore [LLY^{+}20]. Manycores [KSP^{+}20]. Mapped [PGY^{+}22]. Mapping [CCYC21, CZP^{+}23, KOA22, LZX^{+}21, ZMP23, ZGZ^{+}24]. Mapping-Aware [KOA22]. MapReduce [GGZ^{+}20, TRN^{+}21]. Marked [LGZ^{+}21]. Market [GLL^{+}21, ZLK^{+}22]. Market-Scale [GLL^{+}21]. Markets [AKG20]. Markov [CRZ^{+}23, GGL^{+}23, LCZ^{+}23, SGH^{+}23, SCA23]. Massive [DPGG22, DHH^{+}22, FGH^{+}22, KSW^{+}22, LAG^{+}22, LL22, PKJ^{+}22, ZCZ^{+}22]. Massively [BCG23, HYL^{+}20, KM23a]. Matching [JLK^{+}20, XLL^{+}20a]. Materials [WXZ24]. Mathematical [SP20]. Matrices [AHSW23]. Matricized [AAA21]. Matrix [AHSW23, CFLY21, HLB^{+}23, HC24, MKJ^{+}22, SNK20, YFD^{+}24, ZXG^{+}22, ZKP20]. Matrix-Multiple [HC24]. Matrix-Power-Vector [AHSW23]. Max [FLPL22]. Maximal [WCT21]. Maximization [AKZ^{+}20, CZZ^{+}22, GK21, MLWX20, SSH21, WZZ^{+}20]. Maximized [TWYL20]. Maximizing [FC23, LLX^{+}22, LLL^{+}23a]. MC [SNK20]. MCDS [TCJ22b]. MCFsyn [LGZ^{+}21]. MCMC [DMST20]. MD [BDWH24]. MEAN [XCL^{+}23]. Means [YZL^{+}20].

Measurement [LXW^{+}23, RT25, WXHZ20, WXT^{+}24, YLS^{+}23]. Measurements [WLP^{+}23]. MEC [CZJ^{+}22, LTH^{+}21]. Mechanism [CJLW22, GLF^{+}21, GHG^{+}20, HYL^{+}23, LCIS24, LHL^{+}22, LLZ^{+}23, WHL^{+}23, ZTL24, ZYX^{+}22, ZLCL20]. Mechanisms [BBG22, WSM^{+}20]. Medical [XRS^{+}23]. Medium [KAA21]. Medium-Grain [KAA21]. Meets [CLX^{+}23, XXM^{+}20]. Membership [MHZ^{+}22]. Memory [ASS20, ASH^{+}22, CZP^{+}23, CC23, CKO^{+}21, CLZ^{+}20, CLZ^{+}24, DPGG22, DHH^{+}22, FGH^{+}22, GGA24, GGZ^{+}20, GKK21, GN22, GBM20, HLW^{+}20, HZK^{+}24, JLL^{+}20, KSZ24, KSW^{+}22, LK20, LZZ21, LAG^{+}22, LL22, LTZ^{+}23, LZDO24, LXL^{+}24, LLS^{+}24, LLL^{+}24a, LHZ^{+}23, LHL23, LAC24, PGY^{+}22, PKJ^{+}22, QXL^{+}20, RP20, SBM24, SYT20, TWY^{+}20, TGFPRA20, TGFPRA22, WC20, XHQC20, ZZM^{+}23, ZSL^{+}23, ZXW^{+}24, ZCZ^{+}22, ZRC^{+}24, ZSFX23]. Memory-Centric [DPGG22, DHH^{+}22, FGH^{+}22, KSW^{+}22, LAG^{+}22, LL22, PKJ^{+}22, ZSL^{+}23, ZCZ^{+}22]. Memory-Efficient [GGZ^{+}20, LLL^{+}24a]. Memory-Mapped [PGY^{+}22]. MEMPHA [KHOI20]. Memristive [ZCL^{+}22]. MemXCT [DPGG22, DHH^{+}22, FGH^{+}22, KSW^{+}22, LAG^{+}22, LL22, LMX^{+}22, PKJ^{+}22, ZCZ^{+}22, HBG^{+}22]. Merak [LLT^{+}23]. Merging [LXC24, SCL21a]. Meshes [KK24]. Message [DDN^{+}22, KHOI20, KM23b]. Message-Passing [KHOI20]. Meta [WHM^{+}21]. Metadata [ZZM^{+}23, ZCW^{+}20]. Metal [CLZ^{+}22a]. Method [AMKS21, DTN^{+}22, GVD^{+}22, KOA22, LYZC22, LLL^{+}24b, LCL^{+}24, QQD^{+}24, RCW^{+}23, TWY^{+}20, WXZ24, YYL^{+}24, ZZSC20, lM20]. Methods [AMW^{+}21, LZX^{+}21, TXX^{+}21, ZGNZ22]. Metric [LDZ^{+}24]. MF [HLB^{+}23]. MG [SCL21a]. MG-WFBP [SCL21a].

Microarchitectures [IXS22].

Microbenchmarks [SLG^{+}23]. Microglia [ZL24]. Microservice [CZL^{+}24, DFP23, FZC^{+}22, HWW^{+}23, LWY^{+}22, SFW^{+}24, WDJ21, WLY^{+}20]. Middleware [WHRL21]. MIG [QXL^{+}24]. Migration [CWXC24, CLX^{+}23, HTB22, LWZ23b, LTH^{+}21, LZM^{+}20, TWY^{+}20]. Millimeter [GDZ^{+}20]. Millimeter-Scale [GDZ^{+}20]. Million [HFC^{+}23]. Millisecond [ZLL22a]. Min [FLPL22]. Min-Max [FLPL22]. Minimal [Pil23]. Minimization [LSY21]. Minimizing [KAA20, Mei23, XLL^{+}20a]. Minimum [CKS^{+}20, ZKL24]. Minimum-Cost [ZKL24]. Mining [MCT21, ZJGD21]. Minority [HXW^{+}20, ZZH^{+}20a]. Miosix [ATF23]. MIPD [ZW22a]. Miss [WJL^{+}24]. Mistakes [FWT^{+}24]. Mitaka [LZS^{+}24]. Mitigating [IYAK23, LLC^{+}21, NJG^{+}22]. Mitigation [TGFPRA22]. Mix [SZZY24]. Mixed [ACP^{+}22, CBL22, LLD22, QJC^{+}24]. Mixed-Criticality [CBL22]. Mixed-Precision [ACP^{+}22]. Mixing [XYW22, sKW22]. ML [AKA22, WPZ22]. MM* [HLH22]. MO [DQC^{+}21]. MO-Tree [DQC^{+}21]. Mobile [BBGY20, BBG22, DLC^{+}21, GLL^{+}21, HLL22, JBLJ23, JZLZ24, LLHJ20, LLS^{+}24, LZW^{+}23, LJH^{+}23, MLWX20, RXL^{+}20, TCJ22b, WIBD22, WHLM21, XZL^{+}21]. Mobility [WNA^{+}20]. Mode [BK21, CFM^{+}21b, LCG^{+}21, MKKS21, SLX^{+}21b, SCL^{+}21b, ZZH^{+}21]. Model [AA24, AD25, BM20, BOGM21, CYZ^{+}23, CZH^{+}20a, CZH^{+}20b, DZS^{+}21, GWLX22, HLH22, HLS^{+}23, JXX^{+}23, JWZ^{+}23, KHOI20, LZW^{+}22a, LWC^{+}22, LSL^{+}23, LJT^{+}25, LDJX^{+}23, LHPW20, MHLM24, MZWX21, MD22, NAL^{+}20, NK21, SZCL23, TLGA^{+}22, WHM^{+}23, WFY^{+}24, YLW^{+}22, ZLD^{+}23, ZLW^{+}23, ZLCW23]. Model-Based [WHM^{+}23]. Model-Parallel [HLS^{+}23]. Modeling [ACP^{+}22, COE20, LK21, LMFK23, TLH22, WDL^{+}20, WLL^{+}20, ZLJ^{+}23]. Modelling [FTYL20, TMP23]. Models [BFK^{+}23, DA20, GLP^{+}21, HLS^{+}23, JZLZ24, KAA21, LLT^{+}23, LHRX20, LCCZ20a, LZQ^{+}23, LMX^{+}24, LYN^{+}20, MHZ^{+}22, TCJ22a, ZXW^{+}24]. Modern [KK23, LSC^{+}20, LHZ^{+}23, MXS21, TSC^{+}24, ZCHZ23]. Modulo [SDMC24, ZSW^{+}20]. MoE [ZXW^{+}24]. Molecular [DLY^{+}24]. Moment [GVD^{+}22]. Momentum [LCCZ20b, YFB^{+}23]. Monetary [ZLK^{+}22]. Monitoring [MXS21, SMSK21]. Motif [ZJGD21]. MPAS [HFC^{+}23]. MPAS-A [HFC^{+}23]. MPI [CBB^{+}20, CTBT21, DDN^{+}22, DDX^{+}24, GGZ^{+}20, KLH^{+}20a, TPV20, WGBS23]. MPI-GPU [TPV20]. MPMoE [ZXW^{+}24]. MPSoC [SJLN20]. MPSoCs [YOM21]. MRAM [HCR^{+}22, YCA^{+}20]. MRCN [LKH23]. MSA [WBS23]. Multi [BCVD23, BLK^{+}20, CAAB20, CZH^{+}20a, CZH^{+}20b, CQZ^{+}24, CLY^{+}24, CWY^{+}23, COE20, CZZM24, DLY^{+}24, DSZ^{+}24, GM21, HCG^{+}23, HZX^{+}24, HZW^{+}21, HLB^{+}23, JHB24, JYF^{+}24, KSVR23, LK20, LFZ^{+}21, LWZ^{+}22, LSW^{+}23, LCH23, LZZ^{+}23, LLK24, LZP24, LTH^{+}21, LJM^{+}23, LLL^{+}23a, LWX^{+}23, LCM^{+}20, LGZ^{+}21, MZC^{+}22a, MAOA22, MXS21, MHM22, MWNK22, OQCW20, SCYJ21, SMCH20, SRD^{+}20, TLH22, TPV20, WZHW22, WMG^{+}23, WBS23, XCH^{+}22, XZW^{+}24, XLZD24, XSC^{+}23, YFB^{+}23, YFD^{+}24, YZL24, YSG^{+}22, YHT^{+}23, YYL^{+}24, ZXGZ21, ZZY^{+}21, ZW22b, ZLJ^{+}23, ZZC^{+}23, ZLZ^{+}23, ZDC^{+}23, ZXL^{+}24, ZJS^{+}24, ZWL^{+}21, ZCL^{+}22, ZF23, ZLR^{+}20]. Multi-Access [CZZM24, XCH^{+}22, XZW^{+}24, YHT^{+}23, ZXL^{+}24]. Multi-Agent [CQZ^{+}24, CLY^{+}24, JHB24, LZZ^{+}23, LTH^{+}21, ZXGZ21, ZCL^{+}22]. Multi-Body [DLY^{+}24]. Multi-Cloud [LSW^{+}23, LCH23, SMCH20].

Multi-Component [XSC^{+}23]. Multi-Constrained [ZLZ^{+}23]. Multi-Constraint [SRD^{+}20]. Multi-Copy [ZZC^{+}23]. Multi-Core [BLK^{+}20, JYF^{+}24, YFD^{+}24]. Multi-Cores [WZHW22, YYL^{+}24]. Multi-CPU [HLB^{+}23]. Multi-CPU/GPU [HLB^{+}23]. Multi-Data [CLY^{+}24]. Multi-Dimensional [ZJS^{+}24]. Multi-DNN [LLK24, YZL24]. Multi-Domain [CWY^{+}23]. Multi-FPGA [LZP24]. Multi-GPU [BCVD23, GM21, HZX^{+}24, ZW22b]. Multi-Grid [YYL^{+}24]. Multi-Hop [SCYJ21]. Multi-Interface [ZWL^{+}21]. Multi-Job [LJM^{+}23]. Multi-Layer [LFZ^{+}21]. Multi-Level [DSZ^{+}24]. Multi-Objective [HZW^{+}21, MAOA22]. Multi-OS [TLH22]. Multi-Owner [WMG^{+}23]. Multi-Parameter [ZLJ^{+}23]. Multi-Party [LGZ^{+}21]. Multi-Query [ZZY^{+}21]. Multi-Replica [LCH23]. Multi-Resolution [TPV20]. Multi-Server [ZDC^{+}23]. Multi-Socket [COE20]. Multi-SP [HCG^{+}23]. Multi-Stage [LCM^{+}20, ZLR^{+}20]. Multi-Stripe [ZF23]. Multi-Swarm [LWZ^{+}22]. Multi-Task [CZH^{+}20a, CZH^{+}20b, MHM22, MWNK22, SCYJ21, XLZD24]. Multi-Tenancy [WBS23]. Multi-Tenant [MXS21, YSG^{+}22]. Multi-Threaded [LK20]. Multi-Tier [KSVR23, LLL^{+}23a, YFB^{+}23]. Multi-Tiered [OQCW20]. Multi-Type [LWX^{+}23]. Multi-User [MZC^{+}22a, WMG^{+}23]. Multi-Workflow [LWZ^{+}22]. Multi-Zone [GM21]. Multicast [CWY^{+}23, JJ22, LKH23]. Multicasting [MLWX20, RXL^{+}20]. Multicasts [MK24]. Multicore [ASLPE20, CC22, FSPE20, KPA^{+}20, YZJ^{+}21]. Multicores [BSPM23]. Multidimensional [CTL24, LLL^{+}24b]. Multilayer [SMK^{+}23]. Multimedia [DS22]. Multipath [DFXY20]. Multipathing [BDS^{+}21]. Multiple [CCWW24, HAD^{+}22, HTB22, HO23, HC24, ZZZ^{+}24b]. Multiplication [AHSW23, HC24, MFYB22, MKJ^{+}22]. Multiplications [YFD^{+}24]. Multiprocessor [KMM20, WZGM23, HCA24]. Multiserver [AD25, WZZ^{+}20]. Multiserver-Job [AD25]. Multitasking [ZGNZ22]. Multithreaded [KEMC22]. Mutual [UXL^{+}21]. MZI [NKP^{+}24]. MZI-Based [NKP^{+}24].

n [DDN^{+}22]. Nanyang [LL22]. NAS [GM21, YSZL21]. National [SCL^{+}21b]. Native [GXC^{+}23, MFYB22]. Near [CEP22, CZP^{+}23, CLG^{+}24, LLJC21, WGBS23, ZZS^{+}22]. Near-Lossless [WGBS23]. Near-Memory [CZP^{+}23]. Near-Optimal [CLG^{+}24, LLJC21]. Near-Zero [CEP22]. Necessary [CBL22]. Neighborhood [CQZ^{+}24, GTH22]. NeoFlow [ZCJ^{+}22]. Net [WZY^{+}22]. NetSHa [ZPL^{+}22]. Network [ACDK20, AA23, BPW^{+}23, BLYZ21, CRZ^{+}23, CSJB20, CL20b, CMLH20, CLL22, CWL^{+}21, CGL^{+}22, FXL^{+}23, GLF^{+}21, GCL^{+}22, GHG^{+}20, GGL^{+}23, HCG^{+}23, HYL^{+}23, KAT^{+}20, KKS21, LYK20, LKH23, LLHJ20, LZWL22, LCZ^{+}23, LXW^{+}23, LLL^{+}24a, LCL^{+}20, LH22, LTT^{+}20, LXC^{+}22, LLP^{+}23, LLL^{+}23a, MHZ^{+}22, Mei23, MZWX21, OZCW22, PLJK22, PWX^{+}23, QZFZ20, SGH^{+}23, SPS^{+}24, SCA23, TSV21, WL20, WXT^{+}24, YLT^{+}21, YJWM24, YWH^{+}21, YTL^{+}23, ZPL^{+}22, ZTL24]. Network-Aware [CWL^{+}21]. Network-Based [YTL^{+}23]. Network-Efficient [GCL^{+}22]. Network-Integrated [KAT^{+}20]. Network-on-Chip [GHG^{+}20, WL20]. Network-on-Chips [CL20b, LKH23]. Network-Wide [SPS^{+}24]. Networked [SYT20]. Networks [BDS^{+}21, CGLC20, CFLL21, CXL^{+}23, CWY^{+}23, CZZM24, DZL^{+}21, FFQ^{+}22, FLW^{+}23, FLPL22, GWG^{+}22, GK21, GLX^{+}22, HO23, HND20, HLH22, KSZ24, KSB^{+}22, LDL22, LLHJ20, LS21, LLL^{+}21b, LHXH22, LTT^{+}20, LJW^{+}23, LLDL23, MLWX20, MHW^{+}21, MHM22, MMR^{+}21, MDM22, NDW^{+}21, OSF22, SPPS20, SLHH23, SYS^{+}22, TLQ^{+}20, WZY^{+}22, WRLS22, WJL^{+}24, XJX24, XTH^{+}23, YFB^{+}23, YLL^{+}20, ZQM^{+}22, ZDL^{+}21, ZJS^{+}24, ZWL^{+}21, ZCL^{+}22, ZRXF23, ZSFX23]. Neural [BPW^{+}23, BLYZ21, CMLH20, CQZ^{+}21, CGC^{+}22, FFQ^{+}22, FLW^{+}23, GLF^{+}21, GWG^{+}22, GLX^{+}22, HW22, LS21, LZX^{+}21, LZWL22, LWC^{+}22, LLL^{+}24a, LH22, LJW^{+}23, MHW^{+}21, MHZ^{+}22, MHM22, PLJK22, QZFZ20, RCLJT22, SYS^{+}22, WRLS22, XRS^{+}23, ZTL24, ZDL^{+}21, ZL24]. Neural-Like [XRS^{+}23]. Neuromorphic [YWS^{+}23, ZHP^{+}23]. Neuron [ZCL^{+}22]. NeuroSpector [PKRS23]. Neutralization [SBM24]. NewHope [GXW22, GJCC21]. Newton [DTN^{+}22]. Newton-type [DTN^{+}22]. Next [BTL^{+}22]. NFV [MLWX20, RXL^{+}20, ZHH^{+}24]. NFV-Enabled [RXL^{+}20, MLWX20]. NIOT [ZCHZ23]. NITI [WRLS22]. NoC [HGA20, RP20, ZGZ^{+}24]. NoCs [WLH^{+}20a]. Node [FXL^{+}23, GHG^{+}20, KLH^{+}20a]. Noisiness [MRFP20]. Non [CTBT21, KLH^{+}20a, KMM20, LSW^{+}23, SdR^{+}21, WLF^{+}22, ZLRY22, LCX^{+}24]. Non-Blocking [WLF^{+}22]. Non-Contiguous [KLH^{+}20a]. Non-Determinism [CTBT21]. Non-IID [ZLRY22, LCX^{+}24]. Non-Intrusive [SdR^{+}21]. Non-Preemptive [KMM20]. Non-Stationary [LSW^{+}23]. Normal [BK21, CFM^{+}21b, LCG^{+}21, MKKS21, SLX^{+}21b, SCL^{+}21b, ZZH^{+}21]. Note [Par20, Par21a]. Novel [JDD^{+}24, LLL^{+}21a, LX23, LCM^{+}20, QWHC21, TMP23, WLF^{+}20, ZXG^{+}22, ZCHZ23, ZGZ^{+}24]. NSGA [LSZ^{+}21, ZRXF23]. NSGA-II [LSZ^{+}21]. NSGA-III [ZRXF23]. NUMA [BSPM23]. Number [CKS^{+}20]. Numeric [SLG^{+}23]. Numerical [TPV20, XRS^{+}23]. NV [LSC^{+}20]. NV-SLI [LSC^{+}20]. NVGraph [LCL^{+}20]. NVLink [LSC^{+}20]. NVMe [SDHQ21]. NVMeoF [YZC^{+}23]. NVMM [LCL^{+}20]. NVRAM [CLZP20]. NVSwitch [LSC^{+}20].

O [ASMA21, BPTV23, GCL^{+}21, HCR^{+}22, HLZ^{+}20, KLH^{+}20a, LCZ^{+}20, LZDO24, LXC24, SHC^{+}22, TKRB22, WHG^{+}22, WYZY24, YYW^{+}20]. O-Efficient [XWJ^{+}20]. O3BNN [GLW^{+}21]. O3BNN-R [GLW^{+}21]. Object [HSH^{+}22, KKS21]. Object-Level [HSH^{+}22]. Objective [CAAB20, HZW^{+}21, KFS^{+}21, LWZ^{+}22, LCX^{+}24, MAOA22, SRD^{+}20]. Oblivious [CCWW24]. Octans [YZS^{+}21]. Off [AA24, DLMF22, IATB20, WSLX22]. Off-Chaining [AA24]. Offline [WHM^{+}24, ZZD^{+}24a]. Offline-Online [WHM^{+}24]. Offline-Transfer-Online [ZZD^{+}24a]. Offload [MZW22]. Offloading [CZL^{+}22, CHY^{+}24, DLLL22, FC23, HXW^{+}20, MZC^{+}22a, MHW^{+}21, NJG^{+}22, QZCZ21, SCYJ21, WHM^{+}21, XZL^{+}21, ZLW^{+}23]. OfpCNN [YZSX23]. On-Demand [NJG^{+}22, SCS^{+}23, YZSX23]. On-Edge [CZH^{+}20b, CZH^{+}20a]. On-Line [PWX^{+}23]. Online [ABBA23, BGZR21, CYY^{+}22, CMSV20, CJLW22, CGL^{+}22, DLMF22, DWZ^{+}24, LT20, LZJ^{+}20, LMZ^{+}20, LLL^{+}23a, LWX^{+}23, LZH^{+}24, LWL^{+}23, MZC^{+}22a, MTL^{+}20, WHG^{+}22, WZHW22, WHM^{+}24, WLH20b, XCH^{+}21a, XJX24, YS22, ZZD^{+}24a]. Only [CXW^{+}24, WRLS22]. Onto [ZMP23, GLL^{+}21]. Opca [CCWW24]. Open [DFP23, VMT^{+}20]. Open-Access [VMT^{+}20]. Open-Source [DFP23]. OpenCL [WQKH20]. OpenKMC [XSC^{+}23]. OpenMP [SdR^{+}21, WJG^{+}21, YRQ23]. OpenStack [JTX^{+}22]. Operational [WYW^{+}22]. Operations [KAA20, LZW22b, WLF^{+}22, ZLWW20]. Operators [CWL^{+}21]. Opportunities [LJZY20, TSC^{+}24]. Optical [GHG^{+}20, ZGZ^{+}24]. Optical-Electrical [ZGZ^{+}24]. Optimal [CLG^{+}24, DMPR22, FTYL20, LZH^{+}24, LLJC21, OS20, SGJ^{+}20, SOI^{+}20, WVSL23, WLH20b, WSLX22, YZS^{+}21]. Optimised [SMSK21]. Optimistic [CCWW24, SDZ21, XMW^{+}24]. Optimization [ABBA23, BBGY20, CLY^{+}24, DLLL22, DZS^{+}21, DSZ^{+}24, FLPL22, GDS^{+}22, KFS^{+}21, KKP21, LBNN^{+}21, LLL^{+}21a, LWZ^{+}22, LK21, LMFK23, LTH^{+}21, LSZ^{+}21, LXL23, LZH^{+}24, LHX^{+}25, MKKP22, NK21, PKRS23, RCW^{+}23, SLLL20, SLY^{+}24, WDL^{+}20, WZL^{+}22, WCN^{+}24, WGW^{+}23, YS22, YDL23, ZCHZ23, ZMY^{+}24, ZJS^{+}24, ZRXF23, ZLX^{+}20, HBG^{+}22]. Optimizations [FFQ^{+}22, SWOM20]. Optimize [SdR^{+}21]. Optimized [ABG20, KMM20, LLZ^{+}23, YOM21, YZC^{+}23]. Optimizing [CLL22, HLZ^{+}21, LZW22b, TSW^{+}21, WYZY24, WPG^{+}24, YFD^{+}24, YYL^{+}24, ZFY^{+}20, ZYM^{+}20, ZLK^{+}22, ZWSF24]. Orchestrating [TJD^{+}24]. Orchestration [HKL^{+}20, MSSK21, PLJK22, RCW^{+}23, WLY^{+}20]. Orchestrator [HZX^{+}24]. Order [GDS^{+}23, GLW^{+}21, IXS22, MDC24]. Ordered [HZK^{+}24]. Ordering [MK24]. Oriented [KMA^{+}20, LKH23, LTT^{+}20, LZWW22, LYGG20]. Orthogonal [ZZS^{+}22]. OS-Level [BSPM23]. Out-of-Core [XWJ^{+}20]. Out-of-Memory [LTZ^{+}23]. Out-Of-Order [IXS22, GLW^{+}21, MDC24]. Output [FLW^{+}23]. Outsourced [WCW^{+}23]. Overall [ZCJY20]. Overcome [HCR^{+}22]. Overflow [DSCL21]. Overhead [XWDC23, ZZSC20]. Overlapped [LZM^{+}20]. Overlapping [WPZ^{+}21].

Overlay [JMF22, LLP^{+}23, ZZP23]. Overload [LPH^{+}24]. Oversubscribed [KMA^{+}20]. Oversubscription [YHS^{+}20]. OWebSync [JLJ21]. Owner [WMG^{+}23].

P [TLQ^{+}20, XRS^{+}23, ZL24]. P-PFC [TLQ^{+}20]. P2P [HNKO20, WSM^{+}20]. P4SGD [HLS^{+}23]. Pache [CGLC20]. Packed [ZFW^{+}20]. Packet [CLZ^{+}22a, CGLC20, JJ22, LCZ^{+}20, Mei23, NJG^{+}22]. Packing [LT20]. Page [CLZ^{+}24, HCR^{+}22, LZM^{+}20, TWY^{+}20]. PageRank [WQZD25]. PageRank-Inspired [WQZD25]. Pages [ZCJY20]. Paired [KK24]. PaKman [GKK21]. PALE [SPZE20]. Parabolic [ZTL24]. Paradigm [LZZ^{+}23, ZSH^{+}21]. Parallel [AMKS21, Ano20, BM20, BCG23, BLK^{+}20, BLYZ21, BK21, BADP22, CRZ^{+}23, CFM^{+}21a, CL20a, CFLY21, CGC^{+}22, CQZ^{+}24, CTL24, CHY^{+}24, CFM^{+}21b, CLZ^{+}21, COE20, CKS^{+}20, CGM21, DMST20, DA20, DLY^{+}24, EFME24, FZD^{+}24, GK21, GM21, GZJ^{+}21, GGL^{+}23, HPB21, HCG^{+}23, HLZ^{+}20, HLZ^{+}21, HSY^{+}20, HLLL22, HLS^{+}23, HC24, JWW^{+}22, JLG^{+}23, JWZ^{+}24, KAA21, KFS^{+}21, KSB^{+}22, KKP21, KM23a, LK20, LC20, LBNN^{+}21, LIP^{+}21, LLL^{+}21a, LCZ^{+}23, LLL^{+}24b, LLL^{+}24a, LCG^{+}21, LLC^{+}22, MHW^{+}21, MKKS21, Par22, QWYG20, SLX20, SLX21a, SLX^{+}21b, SGH^{+}23, SCA23, SCL^{+}21b, TXG^{+}21, TKRB22, XZL20, XYL^{+}21, YJWM24, ZJHS20, ZZH^{+}21, ZCZ^{+}21, ZHP^{+}23, ZLC^{+}22, ZYX^{+}22, ZLGZ23, lM20]. Parallelism [DZS^{+}21, GSL^{+}20, LLT^{+}23, LGZ^{+}24, LH22, LJT^{+}25, LWZ^{+}24, MHZ^{+}22, OMD^{+}21, PRL20, SZM20, WCN^{+}24, XZY^{+}24, XHY^{+}24, ZFY^{+}20, ZXW^{+}24, ZLW^{+}23]. Parallelisms [ZMY^{+}24]. Parallelization [CYZ^{+}23, DPGG22, DHH^{+}22, FGH^{+}22, GDS^{+}22, GM21, KSW^{+}22, LAG^{+}22, LL22, LSZ^{+}21, LLJC21, PHY20, PKJ^{+}22, WGLZ20, ZGM21, ZCZ^{+}22]. Parallelizing [LZQ^{+}23]. Parameter [LGZ^{+}24, WPZ^{+}21, WPZ22, ZW22b, ZLJ^{+}23]. Parameters [CXW^{+}24, LZZ21]. PaRSEC [ACP^{+}22]. Partial [LWLX24, SCYJ21, YW20, ZLW^{+}23]. Participants [MGBD24]. Particle [BTL^{+}22]. Particle-in-Cell [BTL^{+}22]. Partition [GZY21, YWH^{+}20, ZWL^{+}21]. Partitioned [ZLCW23]. Partitioning [ASS20, DA20, GBM20, KAA21, KOA22, LGZ^{+}24, LWL^{+}25, PSS^{+}20, SMK^{+}23, SLX21a, SRD^{+}20, SZ20, WJG^{+}21, WUR^{+}24, YZSX23, ZSX^{+}20]. Partitioning-Based [WJG^{+}21, WUR^{+}24]. Party [LGZ^{+}21, vDIDB23]. Passing [KHOI20]. Path [BDS^{+}21, CQZ^{+}24, KK24, ZLCW23]. Paths [KSB^{+}22]. Pattern [GVD^{+}22, JLK^{+}20, LMZ^{+}20, ZLW20, ZJGD21]. Pattern-Based [LMZ^{+}20]. Patterns [GTH22, HLVR21]. PaVM [FZD^{+}24]. PBFT [LFZ^{+}21]. PC [ZJHS20]. PCIe [LSC^{+}20, TSW^{+}21]. PCIe-Based [TSW^{+}21]. PDEs [ZTL24]. Peak [ASLPE20, MR24]. Peak-Power [ASLPE20]. Pebbles [BPP21]. Peking [CFM^{+}21b, SGH^{+}23, FGH^{+}22]. Per-Flow [YLS^{+}23]. Per-Switch [YLS^{+}23]. Perceived [CZJ^{+}22, WZZ^{+}20]. Perception [HJEV^{+}21]. Performance [ASS^{+}24, AMW^{+}21, BDS^{+}21, BMMB22, BK21, CLZ^{+}22a, CXÖ^{+}20, CWC^{+}22, CFM^{+}21b, COE20, DNKB20, DNLX25, DLX^{+}24, GSL^{+}20, GLA20, GLW^{+}21, GTH22, GLP^{+}21, GGO21, HGA20, IATB20, JLK^{+}20, JDD^{+}24, JJ22, JWZ^{+}24, KFS^{+}21, KLJV24, LJZ^{+}20, LHQ^{+}20, LCZ^{+}20, LLK22, LLK24, LXC24, LRBV23, LK21, LMFK23, LCG^{+}21, LMX^{+}22, LHZ^{+}23, LHL23, LAC24, LCM^{+}20, LLZ^{+}23, MFYB22, MKKS21, MZC^{+}22b, MTT^{+}22, MMR^{+}21, OAM^{+}25, PM22, PSS^{+}20, QZFZ20, SGJ^{+}20, SLX^{+}21b, SLLL20, SCL^{+}21b, SWOM20, TWY^{+}20, TXX^{+}21, WDL^{+}20, WHG^{+}22, WL20, WQKH20, WC20, WYZY24, YHS^{+}20, YZL24, YLW^{+}22, ZZZ^{+}24a, ZGZ^{+}23, ZLW20, ZCJY20, ZFW^{+}20, ZYF^{+}20, ZZH^{+}21, ZZS^{+}22, ZLJ^{+}23, ZSW^{+}20, ZTA^{+}21, ZCJ^{+}22, ZLX^{+}20, dBMH21]. Performance-Aware [YHS^{+}20]. Performance-Portable [KLJV24]. Performance-Time [LLZ^{+}23]. Performant [BCG23, FBD22, MAOA22]. Period [ZMY^{+}24]. Periodic [KMM20, TSV21]. Periodical [ZLR^{+}20]. Permissioned [HYP^{+}22]. Persistence [CLZP20]. Persistent [CKO^{+}21, LZDO24, WYW^{+}22, ZFW^{+}20, ZZM^{+}23, ZZN^{+}24]. Personalised [MHM22]. Personalization [XLZD24]. Personalized [JBY^{+}23]. Perspective [LHRX20, LHL^{+}22, QZCZ21, XLL^{+}20a]. Pervasive [NDW^{+}21]. Pessimistic [LHXH22]. PetaKV [ZZM^{+}23]. Petrel [ZGQ^{+}21]. PetscSF [ZBB^{+}22]. PFC [CTD^{+}23, TLQ^{+}20]. PFC-Relay [CTD^{+}23]. Phase [CA20b, NK21, PSS^{+}20]. Phase-Aware [PSS^{+}20]. Phi [PRL20]. Photon [LGC^{+}22]. Photonic [LLK22, LLK24]. Photovoltaic [GLF^{+}21]. Physical [LL20, MZW22, TMP23, ZMY^{+}24]. Piggybacking [JQG^{+}22]. Pilot [MTT^{+}22]. PIM [JLL^{+}22]. Pipeline [XZY^{+}24, ZXW^{+}24, ZLC^{+}22]. PISTIS [KDREV21, LWL^{+}22a]. Place [SSS20]. Placement [BBGY20, BM22, CZJ^{+}22, HLZ^{+}20, HLL22, HYL^{+}23, LZJ^{+}20, LSW^{+}23, LZZ^{+}23, LPW^{+}20, Man22, NDW^{+}21, SMK^{+}23, SCS^{+}23, SDMC24, TSV21, WVSL23, WQZD25, WCD^{+}24, YZS^{+}21, ZLZ^{+}23]. Plane [GSL^{+}20]. Planes [QQD^{+}24]. Planetary [BK21, CFM^{+}21b, LCG^{+}21, MKKS21, SLX^{+}21b, SCL^{+}21b, ZZH^{+}21]. Platform [CXÖ^{+}20, CQW^{+}20, CCZW24, HZB^{+}24, KKS21, KHLZ20, KLJV24, LZP24, TJD^{+}24, ZJS^{+}24]. Platformn [LMX^{+}22]. Platforms [ASMA21, CMSV20, COE20, GBM20, GXC^{+}23, HLW^{+}21a, HZX^{+}23, KFS^{+}21, MTT^{+}22, WZGM23, WYZ^{+}24, ZMP23]. PM [ZZN^{+}24]. PM-Based [ZZN^{+}24]. POCLib [ZZS^{+}22]. Point [SLY^{+}23]. Policies [WLH20b]. Policy [MKKP22, OQCW20, ZWSF24]. Policy-Driven [OQCW20]. Pollution [CSJB20]. Polynomial [WLH^{+}23]. Popularity [XXWS24]. Population [SOI^{+}20]. Portable [GN22, GGO21, KLJV24, LRBV23, SP20, YT20]. Portfolios [AMvBI22]. PoS [SQR^{+}21]. Post [GXW22, SYD^{+}24, SZM20, WDCK23]. Post-Quantum [GXW22, SYD^{+}24, SZM20, WDCK23]. PostMan [NJG^{+}22]. Potential [DS22, DLLL22, GDS^{+}23]. Power [AHSW23, ASLPE20, CIZ^{+}20, DDN^{+}22, HZW^{+}21, HGA20, HND20, HTL^{+}24, KLH^{+}20b, KMA^{+}20, LCZ^{+}20, WL20, WZHW22, XCH^{+}22, YWZ^{+}20, ZJH^{+}23, ZLT^{+}24, DDN^{+}22]. Power-Aware [HND20]. Power-Constrained [CIZ^{+}20, HZW^{+}21, KMA^{+}20]. Power-Efficient [DDN^{+}22]. Power-law [ZJH^{+}23]. Power-Performance [HGA20]. POWER8 [NFP^{+}20]. Powercap [HZW^{+}21]. Powered [HTL^{+}24, SZZY24]. PPOAccel [MKKP22]. PQC [GJCC21]. Practical [CLZ^{+}22b, FWCB22, KMLE20, QZCZ21, SPZE20, ZYW^{+}23]. Practice [BFK^{+}23, CZH^{+}20a, CZH^{+}20b, PH21]. Pre [ZXW^{+}24]. Pre-Trained [ZXW^{+}24]. Precision [ACP^{+}22, LLD22, WZY^{+}22]. Preconditioned [YYL^{+}24]. Preconditioner [BCVD23]. PredCom [MLS21]. Predictable [XXC^{+}23]. Predicted [MD22]. Predicting [LPG^{+}22]. Prediction [ACP^{+}22, CHM^{+}20, HLZ^{+}21, IRB21, LWC^{+}22, NAL^{+}20, RCLJT22, SSKG21, WZY^{+}22, WLP^{+}23, YLC^{+}23, YLL^{+}20, YBY^{+}22, ZZH^{+}20a]. Prediction-Based [YLC^{+}23, YBY^{+}22]. Predictive [CZP^{+}23, MXS21, MLS21, TLQ^{+}20]. Preemptive [CZR20, DYFL21, KMM20, YLW^{+}22]. Preface [HMM22, PH21, WR23]. Prefetch [NFP^{+}20]. Prefix [CGH^{+}22]. Preloading [HYL^{+}24]. Preprocessing [CLG^{+}21]. Presence [LWL^{+}24]. Preserving [AKZ^{+}20, CWXC24, LZW^{+}22a, LCLW21, LYN^{+}20, MHW^{+}21, MCT21, ST20, XXWS24, ZLP^{+}24, vDIDB23]. Prevent [LyLG^{+}25]. Preventive [HO23]. Pricing [BBG22, CYY^{+}22, CFLL21]. Primitives [DMI^{+}23, ZLWW20]. Prioritization [LTT^{+}20]. Prioritized [GZY21, ZMP23]. Priority [CLZ^{+}20, HO23, WXT^{+}24]. Priority-Based [CLZ^{+}20]. Privacy [AKZ^{+}20, HWS^{+}24, LZW^{+}22a, LCLW21, LYN^{+}20, MHW^{+}21, MCT21, ST20, XXWS24, ZLP^{+}24, vDIDB23]. Privacy-Preserving [AKZ^{+}20, LZW^{+}22a, LCLW21, LYN^{+}20, MHW^{+}21, MCT21, ST20, ZLP^{+}24]. Private [CYY^{+}22, CYF^{+}23, HLO^{+}21, SFYB21, VMT^{+}20, ZLYL22, ZXG^{+}22]. Proactive [SCS^{+}23, WXH^{+}24]. Probabilistic [KKP21, OHWL21, YW20]. Processing [BFK^{+}23, BPP21, CL20a, HYL^{+}20, HZK^{+}24, JLK^{+}20, LLC^{+}21, LXL23, MZW22, NJG^{+}22, TW24, XWJ^{+}20, XMW^{+}24, YCZC22, YTL^{+}24, ZWK^{+}20, ZZY^{+}21, ZZS^{+}22, ZJH^{+}23, ZDC^{+}23, ZJS^{+}24, ZSX^{+}20, vV20]. Processing-in-Memory [HZK^{+}24]. Processor [ABC^{+}24, YWS^{+}23]. Processors [CL20a, CKS^{+}20, FSPE20, JLG^{+}23, SZ20, SWOM20, YZJ^{+}21]. Procurement [AMvBI22]. Product [AAA21, CXÖ^{+}20, ST20, vDIDB23]. Production [ASS^{+}24]. Productive [KFEG21]. Productivity [ACP^{+}22, PSK^{+}22]. Profiles [GGHP21]. Profiling [WHLM23]. Profit [CZZ^{+}22, WZZ^{+}20]. Program [GZJ^{+}21]. Programmable [HLS^{+}23]. Programming [ATF23, TLGA^{+}22, ZZZ^{+}24a]. Programs [KHOI20, LK20]. Proof [QWHC21, SQR^{+}21]. Proof-of-Stake [SQR^{+}21]. Proofs [CYH^{+}21, LL20, MHLM24]. Propagation [GVD^{+}22]. Properties [CLZ^{+}21, YTL^{+}23]. Proposed [ST20]. Protection [AA24, LPH^{+}24]. Proteus [DLX^{+}24]. Protocol [JLWS24, KDREV21, LGZ^{+}21, vDIDB23]. Protocols [DDN^{+}22, ST20, SOI^{+}20]. Prototypes [YXDL24]. PROV [HZB^{+}24]. PROV-IO [HZB^{+}24]. Provenance [HZB^{+}24]. Providers [LCH23]. Providing [CTD^{+}23]. Provisioning [CZZ^{+}22, FSF^{+}20, LLHJ20, XZW^{+}24, XXC^{+}23]. Proximal [MKKP22]. Proxy [WGBS23]. Pruned [JZLZ24]. Pruning [JXX^{+}23, WQKH20, YHD^{+}23]. Pseudo [YLW^{+}22]. Pseudo-Preemptive [YLW^{+}22]. Pub [DQC^{+}21]. Pub/Sub [DQC^{+}21]. Public [VMT^{+}20]. Publishing [MHZ^{+}22]. Push [XXWS24]. Pyxis [QJC^{+}24].

Q [LWY^{+}22]. QAT [LHQ^{+}20]. QoS [ABC^{+}24, CQW^{+}20, LWC^{+}22, LLL^{+}21c, LYGG20, RCW^{+}23, SFW^{+}24, WZY^{+}22, ZZZ^{+}24b]. QoS-Aware [SFW^{+}24, ABC^{+}24]. QoS-Oriented [LYGG20]. Quality [ASS20]. Quantization [GLX^{+}22, JWZ^{+}23, YHD^{+}23]. Quantization-Based [JWZ^{+}23]. Quantized [ZXG^{+}22]. Quantum [GXW22, LWY^{+}20, LCX^{+}22, SYD^{+}24, SSKG21, SZM20, WSM^{+}20, WDCK23]. Queries [CYY^{+}22, XXW^{+}24, YYZ^{+}20]. Query [TWX22, WPG^{+}22, YCZC22, ZZY^{+}21]. Queued [LXC24]. Queueing [COE20, MZWX21]. Queues [GN22, SDBM23]. Queuing [AA24].

Quorum [YW20]. Quorum-Based [YW20]. QWEB [LHQ^{+}20].

R [CKO^{+}21, GLW^{+}21]. R&D [AMN22]. R-tree [CKO^{+}21]. RAC [MGBD24]. Race [PK21]. Rack [DFXY20, SL20, ZF23]. RADAR [HZK^{+}24]. RADICAL [MTT^{+}22]. RADICAL-Pilot [MTT^{+}22]. RAID [LXC24]. RAM [GGA24]. Random [CLZ^{+}20, HZX^{+}23]. Range [BDWH24, CTD^{+}23, TWX22]. Rank [ZLWW20]. Rao [AAA21]. Rapid [JMF22]. Rapidly [NJG^{+}22]. RAS [HFW^{+}21]. Rate [LJZY20]. Ratio [WDL^{+}20, WLL^{+}20, WJL^{+}24]. Ray [DPGG22, DHH^{+}22, FGH^{+}22, KSW^{+}22, LAG^{+}22, LL22, PKJ^{+}22, ZCZ^{+}22, HBG^{+}22]. RDF [YCZC22]. RDMA [CTD^{+}23, YCZC22]. RDMA-Assisted [YCZC22]. Re [JQG^{+}22, WWJL24]. Re-Alignment [WWJL24]. Re-Distributing [JQG^{+}22]. REACT [JLK^{+}20]. Reactive [GDZ^{+}20, GDS^{+}22]. Read [CCYC21, CZP^{+}23, OHWL21, PM22]. Read-Performance [PM22]. Real [BLK^{+}20, CMLH20, CHY^{+}24, CKS^{+}20, CBL22, DNKB20, JLG^{+}23, JSL^{+}25, KLH^{+}20b, KDREV21, KMM20, MLX23, WIBD22, WZGM23, XZL20, YLW^{+}22, YZC^{+}23, ZLGZ23]. Real-Time [BLK^{+}20, CMLH20, CHY^{+}24, CKS^{+}20, CBL22, DNKB20, JLG^{+}23, KLH^{+}20b, KDREV21, KMM20, MLX23, WZGM23, XZL20, YLW^{+}22, YZC^{+}23, ZLGZ23]. Real-Valued [WIBD22]. Reality [CQZ^{+}21]. Realizations [ACC^{+}22]. Realizing [GLP^{+}21, XCL^{+}23]. Reclamation [GN22, SBM24]. Recognition [SLY^{+}23]. Recommendation [WFY^{+}24]. Recommender [EFME24]. Reconciliation [LGZ^{+}21]. Reconciling [KPHA20]. Reconfigurable [CZP^{+}23, JJ22, TXG^{+}21, WLH^{+}20a]. Reconfiguration [DLMF22, GWLZ21, WLH^{+}20a]. Reconfigurations [PM22]. Reconfiguring [QQD^{+}24]. Reconstruction [DPGG22, DHH^{+}22, FGH^{+}22, KSW^{+}22, LAG^{+}22, LL22, PKJ^{+}22, ZCZ^{+}22]. Record [ZSW^{+}22]. Recovery [CEP22, LWC^{+}20, LAC24, XLL^{+}20b, ZKL24, ZSL^{+}24]. Recurrent [SYS^{+}22]. Recurring [LXL23]. Recursive [GHGW24, SLX21a]. Recycling [QWHC21]. Redesign [GDS^{+}23]. Redesigning [XSC^{+}23]. Reduce [CBB^{+}20, KAA20, LWLX24, SPCT23]. Reduced [LAC24, PHY20]. Reducing [LK20, TLQ^{+}20, ZFW^{+}20]. Reduction [LLL^{+}23b, WYW21, YLC^{+}23]. Reductions [NCB^{+}21]. Redundancy [XZY^{+}24, YWH^{+}20]. Redundancy-Free [XZY^{+}24, YWH^{+}20]. Refactor [GDS^{+}22]. Regime [KM23a]. Regional [WQZD25]. Registers [SZS^{+}23]. Registration [CGH^{+}22]. Regression [LYZC22]. Regret [LZH^{+}24]. Regret-Optimal [LZH^{+}24]. Regular [JLK^{+}20, WLF^{+}22, YTL^{+}23]. Regularized [GLW^{+}21]. Rehabilitating [WGW^{+}23]. ReHy [JLL^{+}22]. Reinforcement [CHY^{+}24, CLY^{+}24, CIZ^{+}20, HYL^{+}23, JHB24, KMBR21, LZZ^{+}23, LYZS24, LTH^{+}21, LXC^{+}22, LPL23, LWX^{+}23, LLZ^{+}23, QZCZ21, SGJ^{+}20, SDBM23, WL20, WHM^{+}21, WZHW22, WHM^{+}23, WHM^{+}24, WXH^{+}24, YS22, YZWT20, YWL^{+}24, ZZD^{+}24a, ZYX^{+}22]. Relative [JSL^{+}25, ZLX^{+}20]. Relative-Error-Bounded [ZLX^{+}20]. Relay [CTD^{+}23]. Reliability [ASLPE20, KMM20, LL20, LLHJ20, ONP^{+}23, WHLM21, XYL^{+}21]. Reliability-Agnostic [WHLM21]. Reliability-Aware [LLHJ20]. Reliable [WL20, ZCW^{+}20]. Relieving [HCG^{+}23]. Reloaded [PWZ^{+}21]. RENDA [TSV21]. Renewable [JHB24, SWGB23, SZZY24]. Reordering [CC22]. Reorganization [WHG^{+}22]. Repair [WSLX22, ZFH22, ZF23]. Repair-Scaling [WSLX22]. Repairable [WSLX22]. Replacement [WHM^{+}24, ZYK^{+}22]. Replay [LXQ^{+}24, ZMP23]. Replica [DMST20, LCH23, LPW^{+}20]. Replicated [PM22, SKW23, SDZ21]. Replication [AP20, BM20, BADP22, CYH^{+}21, HNKO20, NRB^{+}20]. Representation [GVD^{+}22, WHLM23]. Reproducibility [BK21, CFM^{+}21b, HMM22, LCG^{+}21, LMX^{+}22, MKKS21, Par22, PH21, SLX^{+}21b, SCL^{+}21b, WR23, ZZH^{+}21, HBG^{+}22]. Repurposing [IXS22]. Reputation [NLX^{+}22, RCW^{+}23]. Reputation-Aware [NLX^{+}22]. Request [HWW^{+}23]. Requests [KLH^{+}20a]. Requirements [LWX^{+}23]. ReRAM [HLW^{+}20, JLL^{+}22, YHD^{+}23]. ReRAM-Based [YHD^{+}23, HLW^{+}20, JLL^{+}22]. ReRAMs [LJW^{+}23]. Research [CQW^{+}20, CDvK^{+}22, Par22]. Reservation [GGHP21]. Resettable [PK21]. Reshaping [ZLT^{+}24]. Resilient [AB21, KDREV21, SWGB23, WIBD22]. Resistant [HZK^{+}24]. Resizing [JYF^{+}24]. Resolution [TPV20]. Resource [AP20, BBG22, BOGM21, BSPM23, CLY^{+}24, CZZ^{+}22, CIZ^{+}20, FSF^{+}20, FZC^{+}22, GCL^{+}22, HKL^{+}20, HLZ^{+}21, HYP^{+}22, KMBR21, KSP^{+}20, LT20, LZZ^{+}23, LNX^{+}22, LSY21, LLDL23, LLZ^{+}23, MGBD24, MMR^{+}21, RXL^{+}20, SMK^{+}23, TSV21, WPZ22, WHL^{+}23, WCD^{+}24, WCN^{+}24, WYZ^{+}24, XXC^{+}23, YHS^{+}20, YLT^{+}21]. Resource-Aware [MGBD24]. Resource-Constrained [AP20, MMR^{+}21, WYZ^{+}24]. Resources [LyLG^{+}25, YWL^{+}24, ZLR^{+}20]. Response [CJLW22, ZLCL20]. Responsiveness [CCZ^{+}21]. Restart [CL20a]. Restarted [LLD22]. Restoration [CBB^{+}22]. Restore [ZFW^{+}20, ZYF^{+}20]. Retargeting [NISJS21]. Retention [HCR^{+}22]. Rethinking [FWT^{+}24, LZQ^{+}23].

Retinopathy [MMGR23]. Retraining [KY22]. Retrievability [CYH^{+}21]. Retrieval [FWCB22, WXZ24, YXDL24]. Retrospecting [LyLG^{+}25]. Reusability [XHY^{+}24]. Reuse [DS22, GHG^{+}20, LTZ^{+}23, ZZH^{+}20b]. Revenue [LLL^{+}23a, MZC^{+}22a, ZXL^{+}24]. Revenue-Driven [MZC^{+}22a]. Revenue-Effective [ZXL^{+}24]. Reverse [SDMC24]. Reverse-Modulo-Scheduling [SDMC24]. Reversible [GNST21]. Reviewers [Ano21]. Revisiting [HZJH23, IYAK23, ZZN^{+}24]. RHDOFS [LWL^{+}23]. Ring [TXG^{+}21]. Rings [MCT21, WLH^{+}20a]. Ripple [LHC^{+}25]. RIS [SSH21]. RIS-Based [SSH21]. Risk [HFW^{+}21, WZZ^{+}20]. Risk-Aware [HFW^{+}21, WZZ^{+}20]. RIVA [CA20a]. RLPTO [LLZ^{+}23]. RLQ [SDBM23]. RLTiering [LPL23]. RMWPaxos [SSS20]. RNA [MR24]. Robin [Ans20]. Robotic [AA23]. Robust [CA20a, DNLX25, JBLJ23, LHX^{+}25, NRB^{+}20, XZW^{+}24, XWDC23, ZSW^{+}20]. Roll [DDN^{+}22]. Rollback [DDN^{+}22, LAC24]. Rollback-Free [LAC24]. Round [Ans20]. Round-Robin [Ans20]. Route [LLL^{+}23b]. Routing [BDS^{+}21, GHG^{+}20, HYL^{+}23, HWW^{+}23, LKH23, LWX^{+}23, MGG^{+}20, SLX21a, SLZR21, XTH^{+}23, ZYM^{+}20, ZCL^{+}22]. RTGPU [ZLGZ23]. Run [CBB^{+}20]. Run-Time [CBB^{+}20]. run}Data [JQG^{+}22]. Runtime [ASS^{+}24, AA24, HPB21, LIP^{+}21, MXS21, SdR^{+}21, YTL^{+}23, ZZSC20]. Rusty [MXS21].

SaberLDA [LCCZ20a]. Safety [XZL20]. Sampling [CLMW22, LGC^{+}22, WXX^{+}24, ZLW20, ZMY^{+}24]. San [GGL^{+}23, LAG^{+}22]. SaPus [ZW22b]. Satellite [KY22, XXW^{+}24]. Satisfaction [LLX^{+}22]. SC-CGRA [MWL24]. SC19 [Par21b].

Scalability [ÇSS21, CLZ^{+}21, DMI^{+}23, WYW^{+}22]. Scalable [AKA22, Ans20, DDX^{+}24, GLL22, GKK21, HKL^{+}20, JLK^{+}20, KKS21, KZK^{+}19, KZK^{+}20, LFZ^{+}21, LLK22, LZWL22, LPW^{+}20, LWX^{+}23, LWL^{+}23, MSSK21, PGY^{+}22, QJZF23, SRD^{+}20, SWOM20, TPV20, WPZ^{+}21, WPZ22, ZBB^{+}22, ZLC^{+}22]. Scalar [ST20, vDIDB23]. Scale [ACH^{+}20, CBB^{+}20, CLL^{+}21, CQZ^{+}24, CGH^{+}22, DSW^{+}23, DSZ^{+}24, GDZ^{+}20, GLP^{+}21, GZY21, GLL^{+}21, GZJ^{+}21, HWF^{+}22, JWZ^{+}24, KSB^{+}22, LDL22, LYZC22, LCX^{+}22, LHXH22, LZF^{+}24, LCL^{+}24, LPH^{+}24, LHPW20, MCT21, PH21, TSW^{+}21, WLP^{+}23, YHS^{+}20, YZL^{+}20, ZJH^{+}23, ZYW^{+}23, ZTA^{+}21, ZSL^{+}24, ZCW^{+}20, ZSFX23]. Scale-Out [ACH^{+}20]. Scaling [ATF23, ASH^{+}22, CLY^{+}24, HFC^{+}23, HYL^{+}20, KMBR21, LZJ^{+}20, OMD^{+}21, WC20, WCD^{+}24, WSLX22, XXP^{+}23, ZZZ^{+}24b, HBG^{+}22]. Scan [CGH^{+}22]. Scattering [CZP^{+}23]. SCC [CRZ^{+}23, GGL^{+}23, LCZ^{+}23, SGH^{+}23, BK21, CFM^{+}21b, DPGG22, DHH^{+}22, FGH^{+}22, KSW^{+}22, LAG^{+}22, LL22, LCG^{+}21, MKKS21, PKJ^{+}22, SCL^{+}21b, ZZH^{+}21, ZCZ^{+}22]. Scenarios [AMvBI22]. Schedule [LMX^{+}24, ZDC^{+}23]. Scheduler [DNKB20, GZS^{+}25, IYAK23, JLY^{+}23a, PBC^{+}21, YSG^{+}22, YZJ^{+}21]. Scheduling [BGZR21, BLK^{+}20, CMSV20, CQW^{+}20, CZR20, CZZY23, CLY^{+}24, CLG^{+}21, CWL^{+}21, CKS^{+}20, DGE^{+}24, DZL^{+}21, DS23, DFLG21, DYFL21, DNLX25, FSF^{+}20, FSPE20, FTYL20, GHM^{+}24, GCL^{+}22, HZX^{+}24, JHB24, JLG^{+}23, JSL^{+}25, KKA^{+}20, KLH^{+}20b, KKP21, KY22, KS23, KMM20, KMA^{+}20, LYL^{+}20b, LLC^{+}21, LWZ^{+}22, LSL^{+}23, LYZS24, LLL^{+}24b, LyLG^{+}25, LJM^{+}23, LWL^{+}22b, LLZ^{+}23, MLX23, MAOA22, MTL^{+}20, OS20, PKRS23, Pil23, QXL^{+}24, QJC^{+}24, QWYG20, RZLT20, SNN^{+}20, SLZR21, SFW^{+}24, SDMC24, TCT^{+}22, TCJ22a, TCJ22b, WDL^{+}20, WDJ21, WJG^{+}21, WPZ22, WBS23, WYZY24, WNA^{+}20, YS22, YLW^{+}22, YBY^{+}22, ZZG^{+}21b, ZLL22a, ZYW^{+}23, ZYM^{+}20, ZSW^{+}20, ZFH22, ZLR^{+}20, ZLGZ23]. Scheme [CSJB20, CGLC20, CZZM24, CLZ^{+}22b, FSF^{+}20, HLZ^{+}20, HFW^{+}21, HLW^{+}21b, JDD^{+}24, LCH23, SZM20, SDHQ21]. Schemes [KY22, XHQC20]. Science [PH21]. Scientific [DS23, DFLG21, HZB^{+}24, SMSK21, ZLX^{+}20]. Scores [CTL24]. SDN [HTB22, WNA^{+}20]. SDN-Enabled [HTB22]. SDVN [CSZ^{+}23]. Seamless [JLJ21]. Search [CGC^{+}22, CQZ^{+}24, GZW^{+}22, LGZ^{+}24, LCLW21, PRL20, RCLJT22, SKV^{+}20, WMG^{+}23, ZPL^{+}22]. Section [HMM22, Par21b, PH21, WR23]. Secure [CWC^{+}23, FWCB22, GWLX22, MHZ^{+}22, SFYB21, TW24, WXZ24, YHT^{+}23, ZCZ^{+}21]. Securing [LXGY23, WLZ^{+}23, WCW^{+}23]. Security [MSSK21, WVSL23]. Security-and-Time [WVSL23]. Seemingly [WJL^{+}24]. Segmentation [XRS^{+}23]. Seidel [AMKS21]. Seismic [HYL^{+}20, WLM^{+}20]. SEIZE [LIP^{+}21]. Selecting [LZZ21]. Selection [DMST20, HLW^{+}21b, LWL^{+}23, ZLP^{+}24, ZZSC20]. Selective [WHLM23, YWH^{+}20]. Self [DLC^{+}21, HTL^{+}24, JBY^{+}23, ZW22b]. Self-Adaptive [ZW22b]. Self-Balancing [DLC^{+}21]. Self-Knowledge [JBY^{+}23]. Self-Powered [HTL^{+}24]. Semantic [SZS^{+}23]. Semi [CYZ^{+}24, WHRL21, ZLD^{+}23]. Semi-Asynchronous [ZLD^{+}23]. Semi-Coordinated [WHRL21]. Semi-Supervised [CYZ^{+}24]. Send [KAA20]. Sensitive [LLX^{+}22, LLP^{+}23]. Sentinels [DFJ^{+}23]. Separable [LZW22b]. Separation [TWX22]. Sequence [GHGW24, ZZZ^{+}24a]. Sequence-to-Graph [ZZZ^{+}24a]. Sequences [SSS20]. Serial [LXC24]. Series [WSX^{+}23, WR23]. Server [CZZM24, HLL22, WPZ^{+}21, WPZ22, ZWK^{+}20, ZLL22a, ZDC^{+}23]. Serverless [KHLZ20, LK21, LN24, NLX^{+}22, WCD^{+}24, WCN^{+}24, WLM^{+}20, WDZ^{+}23, YWL^{+}24]. Servers [BM20, CLZ^{+}22a]. Service [CZJ^{+}22, LLHJ20, LLX^{+}22, LLL^{+}23a, LWX^{+}23, LLJC21, MD22, NDW^{+}21, OAM^{+}25, SMCH20, SKV^{+}20, XZW^{+}24, YZS^{+}21, ZCW^{+}20]. Services [AB21, CCZ^{+}21, WSM^{+}20, WPG^{+}22]. Serving [WWJL24]. Set [LGZ^{+}21]. Sets [BPTV23]. Setting [HO23]. SF-Sketch [LSY^{+}20]. SFC [LSY21, RCW^{+}23]. SGCNAX [LZWL22]. SGD [HLB^{+}23, WLZ^{+}23]. SGD-Based [HLB^{+}23]. SGD Tucker [LLL^{+}21a]. ShanghaiTech [LCZ^{+}23]. Shape [WSX^{+}23]. Shaped [GBM20]. Shard [JLWS24, XMW^{+}24, XMW^{+}24]. Sharded [BADP22]. Sharding [HYP^{+}22, JLWS24, LWZ23b, XMW^{+}24]. Sharding-Based [HYP^{+}22, XMW^{+}24]. Shared [ASS20, CC23, CLZ^{+}20, CLZ^{+}24, KSZ24]. Shared-Memory [ASS20, CC23]. Sharer [QJZF23]. Sharing [CFLL21, CLZ^{+}22b, HZX^{+}24, LYGG20, RDBG25, RXL^{+}20, WCW^{+}23, ZHH^{+}24, ZLW^{+}21]. Sharing-Aware [RDBG25]. Shift [DLJ^{+}22]. Short [CCYC21, HLW^{+}20, HYL^{+}24]. Short-Read [CCYC21]. Short-Term [HLW^{+}20]. Shortest [KSB^{+}22]. Shutdown [AKG20]. Signature [SYD^{+}24, SZM20, WDCK23]. Signatures [LZS^{+}24]. Significant [GLX^{+}22]. Silhouette [CLL^{+}21]. Silicon [HGA20]. Silo [HHJ^{+}25, YXDL24]. SIMD [CBB^{+}22, ZYL^{+}20]. Simeuro [ZHP^{+}23]. Similarity [LCLW21, LHL23]. Simple [BPTV23, SBM24]. Simplified [MRFP20]. SIMT [GXW^{+}20]. Simulating [DLX^{+}24, SZZY24]. Simulation [DSW^{+}23, GDZ^{+}20, GDS^{+}22, KM23a, LWY^{+}20, MRFP20, QZFZ20, QLP^{+}23, TMP23]. Simulation-Based [KM23a]. Simulations [BTL^{+}22, GLP^{+}21, LCX^{+}22, LCL^{+}24, WPG^{+}24, XSC^{+}23]. Simulator [ZHP^{+}23]. Simultaneous [ASLPE20]. Single [AAK22, CL20a, HWF^{+}22, KSB^{+}22, SZ20]. Single-ISA [SZ20]. Single-Source [KSB^{+}22]. Size [Ans20, WQKH20, YT20]. Size-Based [Ans20]. Sized [GHM^{+}24]. Sketch [LSY^{+}20, PWX^{+}23, ZWL^{+}21, LSY^{+}20]. Sketches [BPP21]. SketchINT [YLS^{+}23]. Skew [GGZ^{+}20, HZK^{+}24]. Skew-Resistant [HZK^{+}24]. Skew-Tolerant [GGZ^{+}20]. Skipping [WDZ^{+}23]. SLA [LyLG^{+}25]. Slack [CBB^{+}20]. Slanted [HJEV^{+}21]. Slark [DNLX25]. SLEEF [SP20]. SLI [LSC^{+}20]. Slice [KPHA20]. Slicing [HCG^{+}23]. SLIDE [PZL^{+}22]. SLO [GCL^{+}21, LZWW22, WWJL24]. SLO-Aware [GCL^{+}21]. SLO-Oriented [LZWW22]. Slowdown [DGE^{+}24]. Small [LZZ21]. Smart [AA24, CZL^{+}22, FZD^{+}24, JLQ^{+}23, LXGY23, LLC^{+}22]. SmartTuning [LZZ21]. SMT [FSPE20, LyLG^{+}25]. SMT-Aware [LyLG^{+}25]. Snapshots [JWZ^{+}23]. SNN [LWZ^{+}24, QLP^{+}23]. Socket [COE20]. SoCs [JYF^{+}24]. Software [CWY^{+}23, HO23, JYF^{+}24, LLL^{+}21b, LZM^{+}20, RT25, SPCT23, WNA^{+}20, YWH^{+}21]. Software-Based [LZM^{+}20, RT25]. Software-Defined [CWY^{+}23, LLL^{+}21b, YWH^{+}21]. Solutions [YYW^{+}20]. Solve [ZSL^{+}21]. Solver [LX23, LAC24, TPV20]. Solvers [BCVD23, LRBV23]. Some [QHC20]. Sophisticated [TJD^{+}24]. Sorter [GHGW24]. Sorting [ÇSS21]. Source [DFP23, KSB^{+}22]. Sources [CTBT21]. Sova [YWH^{+}21]. SP [HCG^{+}23]. Space [HC24, LGZ^{+}24]. Spanning [QHC20]. Spark [FTYL20]. Sparse [AKA22, AHSW23, CXÖ^{+}20, CFLY21, GLL22, HC24, KAA21, LLY^{+}20, LLL^{+}21a, LXW^{+}23, LRBV23, LH22, SZS^{+}23, SLKA23, ZSL^{+}21, ZSFX23]. Sparsification [PZL^{+}22, WGQ^{+}22, ZW22a]. Sparsified [TSLC23]. Sparsity [FLW^{+}23, JZLZ24, LCCZ20a, LAY21]. Sparsity-Aware [LCCZ20a]. Spatial [ABC^{+}24, MKJ^{+}22, SLY^{+}23, XXWS24]. Spatially [YYW^{+}20]. SpatialSSJP [ABC^{+}24]. Spatio [HZX^{+}24]. Spatio-Temporal [HZX^{+}24]. Spatiotemporal [AB21, DQC^{+}21]. Spatiotemporal-Aware [DQC^{+}21]. Special [AMN22, HMM22, Par21b, PH21, WR23]. Spectrum [YFD^{+}24]. Speculative [JLY^{+}23b, YHS^{+}20]. Speed [CA20a, ZXG^{+}22]. Speeding [CFM^{+}21a]. Speeds [ZDC^{+}23]. Speedup [TLH22]. SpGEMM [DA20]. SPHINCS [SZM20]. Spiking [GWG^{+}22, QZFZ20, ZL24]. Spin [LCX^{+}22]. Spinning [IYAK23]. Spite [WIBD22]. Split [CWXC24, CXL^{+}23, SLHH23, XYW22]. Split-Star [SLHH23]. Splitting [AD25, CWC^{+}22, GWLX22, JLWS24]. SpMSpV [LAY21]. SpMV [LAY21, ZZSC20]. SpMV/SpMSpV [LAY21]. SPOC [ST20]. Sporadic [DYFL21]. Spot [AKG20, ZLK^{+}22]. Sprawl [SPCT23]. Springald [MDC24]. SPRINT [LLK22]. Squares [CFLY21]. SR [LXQ^{+}24]. SR-FDIL [LXQ^{+}24]. SSA [GHGW24]. SSD [LXC24, ZCJY20, ZZH^{+}20b, ZWSF24]. SSDs [GSL^{+}20, SLLL20, SDHQ21]. SSRAID [LXC24]. Stable [XZL20, ZLRY22]. Stack [YCA^{+}20]. Stackelberg [HWS^{+}24]. Stage [LSY^{+}20, LCM^{+}20, LPH^{+}24, ZMY^{+}24, ZSH^{+}21, ZLR^{+}20]. Stake [SQR^{+}21]. Stale [ZLRY22]. Staleness [OHWL21]. Stamps [CL20a]. Standard [SP20]. Standardizing [ALAK20]. Star [SLHH23]. State [BADP22, JLWS24, WZY^{+}22, ZLW^{+}21].

Stateful [CLZ^{+}22a, MSSK21]. Static [ABC^{+}24, AA23, GWLZ21, GZJ^{+}21, ZWSF24]. Stationary [LSW^{+}23]. Stealing [CGH^{+}22]. Stencil [SLY^{+}24, SWOM20]. Stencil-Based [SWOM20]. Step [WGW^{+}23]. Steps [HAD^{+}22, WCN^{+}24]. Stereo [CMLH20]. Stixels [HJEV^{+}21]. Stochastic [BBGY20, BOGM21, EFME24, LBNN^{+}21, LLL^{+}21a, LPG^{+}22, MWL24, PHY20, ZKP20]. Stopping [XZL20]. Storage [CCWW24, CMX^{+}20, CGL^{+}22, CLX^{+}23, HKL^{+}20, HNKO20, JLK^{+}20, KMLE20, KAT^{+}20, LL20, LSW^{+}23, LCH23, LWC^{+}20, LPL23, MDM22, OQCW20, SLLL20, TXX^{+}21, WXZ24, WCW^{+}23, XZJ^{+}20, XCL^{+}23, XLL^{+}20b, ZZC^{+}23, ZTA^{+}21, ZLT^{+}24, ZZH^{+}20b, ZRXF23]. Storages [LPW^{+}20]. Store [LCM^{+}20, QXL^{+}20, YW20, ZZM^{+}23]. Stores [ACH^{+}20, JLY^{+}23a, JLL^{+}20, LCLW21, PM22, SDZ21, TGFPRA22, XHQC20]. Stragglers [LLC^{+}21]. Strategies [AP20, DMPR22, DSZ^{+}24, GGHP21, HCA24, SZCL23, TWYL20]. Strategy [CC23, CGL^{+}22, CLX^{+}23, DLLL22, LLL^{+}21a, LXC24, LDZ^{+}24, LWL^{+}22b, SCS^{+}23, XWJ^{+}20, ZW22b, ZGZ^{+}24]. Stream [ABC^{+}24, BGZR21, LLC^{+}21, SZS^{+}23, ZZY^{+}21, ZZZ^{+}24b, ZLW^{+}21, vV20, HWF^{+}22]. Stream-Static [ABC^{+}24]. Streaming [BFK^{+}23, HYL^{+}24, LWL^{+}23, PHY20, XXW^{+}24, ZFY^{+}20, ZGM21, ZYW^{+}23, ZZD^{+}24b]. Streams [BPP21, LSY^{+}20, MDC24]. Stripe [LXC24, ZF23]. Stripe-Queued [LXC24]. Stripe-Threaded [LXC24]. Strong [OMD^{+}21, sKW22]. Structural [ZSFX23]. Structure [DZS^{+}21, WYW^{+}22, WYW21, YJWM24, ZJHS20, ZSFX23]. Structure-Based [WYW21]. Structured [GGO21, JWW^{+}22, Nak21]. Structures [SKW23]. STT [GGA24, HCR^{+}22, YCA^{+}20].

STT-MRAM [HCR^{+}22]. STT-MRAM-Based [YCA^{+}20]. STT-RAM-Based [GGA24]. Student [Par21b]. Study [AMW^{+}21]. Sub [CWC^{+}22, PZL^{+}22, ZLL22a, DQC^{+}21]. Sub-Forest [CWC^{+}22]. Sub-Linear [PZL^{+}22]. Sub-Millisecond [ZLL22a]. Subgraph [WHC^{+}21, ZJH^{+}23]. Subgraph-Centric [ZJH^{+}23]. Subspace [GLL^{+}20]. Subutai [CFM^{+}21a]. Suite [KDREV21, ZZG^{+}21a]. Sunway [DSW^{+}23, GDZ^{+}20, GDS^{+}23, HFC^{+}23, HYL^{+}20, LLY^{+}20, LWY^{+}20, LCX^{+}22, LGC^{+}22, LSZ^{+}21, XSC^{+}23, ZLJ^{+}23]. SunwayLB [LCL^{+}24]. Supercomputer [DSW^{+}23, GZW^{+}22, GDS^{+}23, HFC^{+}23, HYL^{+}20, LCX^{+}22, XSC^{+}23, YZL^{+}20, ZLJ^{+}23]. Supercomputers [LCL^{+}24, YYW^{+}20]. Supercomputing [GGZ^{+}20, JTX^{+}22]. Supermarket [MD22]. Supervised [CYZ^{+}24]. Support [CLZP20, SYT20, YLL^{+}20]. Supported [ZGNZ22]. Suppressing [LDZ^{+}24]. Supremacy [LWY^{+}20]. Surrogate [TCJ22a]. Surveillance [CWL22]. Survey [CA20b, QTR21, YDL23, ZSL^{+}23, ZGNZ22]. SVM [JWW^{+}22]. SVRF [JJ22]. SW [GWLZ21]. Swapping [LLS^{+}24]. Swarm [LWZ^{+}22]. Swift [ZSL^{+}24]. Swing [CTD^{+}23]. Switch [CLZ^{+}22a, HLS^{+}23, YLS^{+}23]. Switch-Accelerated [CLZ^{+}22a]. Switches [DFXY20]. Switching [NKP^{+}24]. swMPAS [HFC^{+}23]. swMPAS-A [HFC^{+}23]. Symbiotic [FSPE20]. Symbolic [GLL22]. Symmetric [SLKA23]. Synapse [BPW^{+}23]. Synchronization [CFM^{+}21a, CGM21, DMI^{+}23, JLJ21, LGH^{+}24, ZSL^{+}21, ZHH^{+}24, ZGQ^{+}21]. Synchronization-Free [ZSL^{+}21]. Synchronize [CXW^{+}24]. Synchrony [LCIS24]. Synergistic [LXQ^{+}24]. Synthesis [dBMH21]. System [CYH^{+}21, CZZY23, CCZW24, CLY^{+}24, CLZ^{+}24, CSZ^{+}23, DSW^{+}23, FC23, GWLX22, HPB21, HLLL22, HCA24, KAT^{+}20, KMM20, LZZ21, LCX^{+}22, LWC^{+}20, LMZ^{+}20, LSZ^{+}21, LPL23, MAOA22, OQCW20, OZCW22, PLJK22, PSS^{+}20, QXL^{+}24, QXL^{+}20, QTR21, SFYB21, SWGB23, SHZ^{+}23, YLL21, YLW^{+}22, YLL^{+}20, YZC^{+}23, ZYF^{+}20, ZXGZ21, ZZM^{+}23, ZMY^{+}24, ZLC^{+}22, ZCL^{+}22]. System-on-Chip [HCA24]. Systematic [PKRS23]. Systematically [GLL^{+}21]. Systems [ASS^{+}24, AA23, ASLPE20, ATF23, ASH^{+}22, BBG22, BFK^{+}23, BMMB22, CAAB20, CZL^{+}22, CC22, CWL22, CLX^{+}23, COE20, CBL22, DS22, DS23, DQC^{+}21, DFLG21, DLC^{+}21, EFME24, FWCB22, GGZ^{+}20, HKL^{+}20, HZB^{+}24, HZW^{+}21, HNKO20, HLZ^{+}20, HSH^{+}22, HZK^{+}24, JSL^{+}25, KSVR23, KKA^{+}20, KLH^{+}20b, KMA^{+}20, LL20, LIP^{+}21, LXL^{+}24, LCL^{+}20, LXC^{+}22, LHL23, LZW^{+}23, MZW22, MXS21, MWNK22, PZZ^{+}22, PSK^{+}22, Par22, QWYG20, RDBG25, SKW23, SKV^{+}20, SYT20, TMP23, TSW^{+}21, TSC^{+}24, TCJ22b, WJG^{+}21, WMG^{+}23, WCW^{+}23, WHLM21, WPG^{+}22, XZJ^{+}20, XZL20, XYL^{+}21, XLL^{+}20b, XLL^{+}20a, XRS^{+}23, YS22, YLL^{+}20, YBY^{+}22, YZS^{+}21, YDL23, ZFW^{+}20, ZZH^{+}20a, ZHH^{+}24, ZTA^{+}21, ZL24, ZYX^{+}22, ZCW^{+}20, Ano20]. Systolic [GHGW24, XMW^{+}22].

T [AKG20, LYK20]. T-BASIR [AKG20]. T-Caching [LYK20]. Tables [WLF^{+}22]. TAC [WPG^{+}24]. Tackling [MGBD24]. Tag [CSJB20, QJZF23]. Tag-Sharer-Fusion [QJZF23]. TaihuLight [DSW^{+}23, HYL^{+}20, LWY^{+}20, LGC^{+}22, LSZ^{+}21, GDZ^{+}20]. Tail [TLQ^{+}20, ZLL22a]. Taking [FWT^{+}24, MR24]. Tardiness [DYFL21, XLL^{+}20a]. Target [PSS^{+}20]. Targets [ZZZ^{+}24b]. Task [ALAK20, CMSV20, CL20a, CZH^{+}20a, CZH^{+}20b, CLY^{+}24, FTYL20, GBM20, HPB21, HLLL22, LZZ^{+}23, LH22, LTH^{+}21, LLL^{+}21c, LLZ^{+}23, MZC^{+}22a, MAOA22, MHZ^{+}22, MTL^{+}20, MHM22, MWNK22, OS20, QWYG20, RDBG25, SCYJ21, SZ20, TCJ22a, WHM^{+}21, WJG^{+}21, WZGM23, XLZD24, XXWS24, ZZG^{+}21b, ZYM^{+}20, ZJS^{+}24]. Task-Aware [SZ20]. Task-Based [ALAK20]. Task-Parallel [HPB21]. Taskflow [HLLL22]. Taskgraph [YRQ23]. Tasking [YRQ23]. Tasks [AP20, CZZY23, CHY^{+}24, CKS^{+}20, DYFL21, JLG^{+}23, KMM20, LZQ^{+}23, QJC^{+}24, WJG^{+}21, WLY22, WSHJ23, YS22, ZLGZ23]. Taxonomic [LHRX20]. TC [HWF^{+}22]. TC-Stream [HWF^{+}22]. TCSA [LGH^{+}24]. TDM [HGA20]. TDTA [WZGM23]. Team [BK21, CRZ^{+}23, CFM^{+}21b, GGL^{+}23, LCZ^{+}23, LCG^{+}21, MKKS21, SGH^{+}23, SCL^{+}21b, ZZH^{+}21, DPGG22, DHH^{+}22, FGH^{+}22, KSW^{+}22, LAG^{+}22, LL22, PKJ^{+}22, ZCZ^{+}22]. Tech [PKJ^{+}22]. Technique [BGZR21]. Techniques [CA20b, LC20, QTR21, ZFH22]. Technological [LL22]. TEE [LWL^{+}22a, LXGY23]. Telecommunications [YLL^{+}20]. Telemetry [SPS^{+}24]. Temperature [DMST20, KSP^{+}20]. Temperature-Constrained [KSP^{+}20]. Template [KSB^{+}22]. Tempo [SLY^{+}23]. Tempo-Spatial [SLY^{+}23]. Temporal [HZX^{+}24, LWZ^{+}24, YTL^{+}23, ZLYL22, ZQM^{+}22]. Temporal-Unrolled [LWZ^{+}24]. Tenancy [WBS23]. Tenant [MXS21, YSG^{+}22]. Tensor [AAA21, AKA22, CXÖ^{+}20, GHM^{+}24, HW22, KAA21, LS21, LXW^{+}23, MFYB22, NCB^{+}21, NISJS21, PWX^{+}23, SLG^{+}23, WXT^{+}24, ZLWW20, ZLW20, ZRC^{+}24]. TensorGEMM [LMH^{+}20]. Tensorox [HW22]. Term [CZZ^{+}22, HLW^{+}20]. Texas [DPGG22]. Text [HZX^{+}23]. TFormer [LDJX^{+}23]. TGNN [XZY^{+}24]. Their [GGHP21]. Theorems [LDZ^{+}24]. Theoretic [DLLL22]. Theoretical [HCZ^{+}20]. Theory [LYDZ21, XLL^{+}20a]. Thermal [HCA24, IRB21]. Thrashing [CLZ^{+}20]. Thread [FSPE20, GXW^{+}20, GCL^{+}21, ZSL^{+}21]. Thread-Level [GXW^{+}20, ZSL^{+}21]. Threaded [LK20, LXC24]. Threads [IATB20, SdR^{+}21, TKRB22]. Threats [WLZ^{+}23]. Three [ZMY^{+}24]. Three-Stage [ZMY^{+}24]. Throughput [CCYC21, CCZ^{+}21, LKH23, LZS^{+}24, LPG^{+}22, LZP24, MLWX20, MKKP22, NK21, RCLJT22, SYD^{+}24, SLG^{+}23, ZYL^{+}20]. Throughput-Oriented [LKH23]. TianHe [JTX^{+}22, GZW^{+}22]. TianheGraph [GZW^{+}22]. Tied [WJG^{+}21]. Tier [KSVR23, LPL23, LLL^{+}23a, YFB^{+}23]. Tiered [OQCW20]. Tiering [LPL23]. Tightly [YWS^{+}23]. Tightly-Coupled [YWS^{+}23]. Tile [KBA^{+}24]. Tiled [MKJ^{+}22]. Time [BLK^{+}20, CBB^{+}20, CL20a, CMLH20, CHY^{+}24, CKS^{+}20, CBL22, DNKB20, JMF22, JLG^{+}23, KLH^{+}20b, KPHA20, KDREV21, KMM20, LLL^{+}23b, LLZ^{+}23, MLX23, Mei23, PSS^{+}20, SPZE20, SOI^{+}20, WSX^{+}23, WVSL23, WZGM23, XZL20, YLW^{+}22, YZC^{+}23, ZLGZ23]. Time-Optimal [SOI^{+}20]. Timed [MDM22]. Timely [XXW^{+}24]. Times [AAA21, MD22]. Timescale [CLY^{+}24]. Timestamped [ZLW^{+}21]. TLC [ZCJY20]. Token [LTT^{+}20]. Token-Oriented [LTT^{+}20]. Tolerance [JBLJ23, WHRL21, ZDL^{+}21]. Tolerant [GGZ^{+}20, GHG^{+}20, HZW^{+}25, KZK^{+}19, KZK^{+}20, KM23b, LT20, MK24, SSS20, YOM21, ZGZ^{+}23, ZHQ^{+}23, ZGG21]. Tomography [HBG^{+}22]. Tool [ZMS^{+}22]. Top [DFXY20]. Top-of-Rack [DFXY20]. Topic [LCCZ20a]. Topology [JXX^{+}23, LWC^{+}22, WZGM23, XTH^{+}23, ZYD^{+}23]. Topology-Aware [LWC^{+}22, ZYD^{+}23]. Topology-Based [WZGM23].

TowerSketch [YLS^{+}23]. Trace [VMT^{+}20]. Traces [MLS21, SYS^{+}22]. Tracing [WGBS23]. Tracking [CWL22, KKS21, ZXGZ21]. Trade [DLMF22, IATB20, WSLX22]. Trade-Off [DLMF22, IATB20, WSLX22]. Tradeoffs [AMW^{+}21]. Trading [CYY^{+}22, CYF^{+}23, HGA20]. Traffic [CWY^{+}23, LZF^{+}24, NJG^{+}22, PWX^{+}23, WXHZ20, WLP^{+}23, XHQC20]. Traffic-Aware [XHQC20]. Trained [ZXW^{+}24]. Training [CZZY23, CLZ^{+}21, DLX^{+}24, HLS^{+}23, JWW^{+}22, JLL^{+}22, LLT^{+}23, LHRX20, LZZ21, LZQ^{+}23, LLL^{+}24b, LLL^{+}24a, LZP24, LJT^{+}25, MHW^{+}21, MHZ^{+}22, OMD^{+}21, PLJK22, TJD^{+}24, WPZ^{+}21, WZL^{+}22, WGQ^{+}22, WPZ22, WRLS22, WGN^{+}23, WFY^{+}24, WXX^{+}24, WSHJ23, XZY^{+}24, XYW22, ZW22a, ZW22b, ZYD^{+}23, ZLC^{+}22, ZCJ^{+}22, ZSL^{+}24, ZRC^{+}24]. Transaction [HYP^{+}22, TW24, XXP^{+}23, XMW^{+}24]. Transactional [TGFPRA20, TGFPRA22]. Transactions [Ano20]. Transcoding [CIZ^{+}20]. Transfer [ABBA23, CZH^{+}20a, CZH^{+}20b, ZZD^{+}24a, ZZH^{+}20a]. Transferrable [DWZ^{+}24]. Transfers [CA20a, CLL22, NK21]. Transform [BCG23, SNK20]. Transformation [LMZ^{+}20, lM20]. Transformations [dBMH21]. Transformers [ZCHZ23]. Transient [CEP22]. Transient-Error-Induced [CEP22]. Translator [LZX^{+}21]. Transmission [LDJX^{+}23, SZCL23]. Transmission-Friendly [LDJX^{+}23]. Transparency [PH21]. Transparent [TKRB22]. Transport [DFXY20]. Travel [LLL^{+}23b]. Traversal [CWXC24]. Tree [GBM20, ZZN^{+}24, CKO^{+}21, DQC^{+}21, ZYL^{+}20]. Tree-Shaped [GBM20]. Trees [QHC20]. Trial [DMST20, GLA20]. Triangle [HWF^{+}22, PWZ^{+}21, YRBÇ22].

Triangular [ZSL^{+}21]. Tridiagonal [LX23]. Triggered [KDREV21, ZTL24]. Trillion [XSC^{+}23]. Trillion-Atom [XSC^{+}23]. Tripartite [LXW^{+}23]. True [AAA21]. Trust [PWZ^{+}21]. Trusted [LWL^{+}22a, MHLM24]. TrustZone [JLQ^{+}23]. TrustZone-Based [JLQ^{+}23]. Truthful [ZLCL20]. TSC [JLQ^{+}23]. TSC-VEE [JLQ^{+}23]. Tsing [SCL^{+}21b]. Tsinghua [CRZ^{+}23, ZZH^{+}21, ZCZ^{+}22]. Tubal [ZLWW20, ZLW20, ZLWW20]. Tubal-Rank [ZLWW20]. Tubal-Sampling [ZLW20]. Tucker [LLL^{+}21a, SLKA23]. Tuning [CWC^{+}23, DWZ^{+}24, GSH^{+}21, LGZ^{+}24, SdR^{+}21, SLY^{+}24, WJZ^{+}24, YT20, ZMS^{+}22]. Tuple [LLC^{+}21]. Turbo [WNL20]. TurboMGNN [WSHJ23]. Turing [LS21]. Turnaround [PSS^{+}20]. Twins [DFJ^{+}23]. Two [ASMA21, CLY^{+}24, LWL^{+}25, LSY^{+}20, LPL23, LPH^{+}24, NK21, SHZ^{+}23, ZSH^{+}21]. Two-Dimensional [LWL^{+}25]. Two-Level [ASMA21, SHZ^{+}23]. Two-Phase [NK21]. Two-Stage [LSY^{+}20, LPH^{+}24, ZSH^{+}21]. Two-Tier [LPL23]. Two-Timescale [CLY^{+}24]. Type [LWX^{+}23, DTN^{+}22]. Types [sKW22].

Ubiquitous [WGLZ20]. UC [GGL^{+}23].

UFC2 [ZLL^{+}22b]. Ultra

[CZP^{+}23, QXL^{+}20]. Ultra-Efficient

[CZP^{+}23]. Ultra-Low-Latency [QXL^{+}20]. UMA-MF [HLB^{+}23]. Unaligned [SLLL20]. Unattainable [WJL^{+}24]. Unbounded [ZLRY22]. Uncertainties [LHX^{+}25]. Understanding [NKP^{+}24]. Undervolting [KPA^{+}20]. Undirected [GK21]. Unequal [CC23, GHM^{+}24]. Unequal-Sized [GHM^{+}24]. Unicast [JJ22]. Unicast/Multicast [JJ22]. Unicasts [MK24]. Unified [HLB^{+}23]. Uniformly [GHGW24]. Universal [WYW^{+}22]. University [LCG^{+}21, MKKS21, CRZ^{+}23, CFM^{+}21b, DPGG22, DHH^{+}22, FGH^{+}22, LAG^{+}22, LL22, LCZ^{+}23, SGH^{+}23, SCL^{+}21b, ZZH^{+}21, ZCZ^{+}22]. Unknown [CL20a, XXWS24]. Unpredictability [LWL^{+}24]. Unreliable [XCL^{+}23]. Unrolled [LWZ^{+}24]. Unsupervised [AKA22]. Untrustworthy [SJLN20]. Unused [HW22]. Upcoming [GGHP21]. Update [XWJ^{+}20, ZW22b]. Updates [LCLW21, SL20]. Updating [KSB^{+}22]. Urban [WNA^{+}20]. US-Byte [GHM^{+}24]. Usable [CMX^{+}20]. User [CZJ^{+}22, DHM25, HCZ^{+}20, IATB20, LLX^{+}22, MZC^{+}22a, WMG^{+}23, XCH^{+}22, ZLL^{+}22b]. User-Friendly [ZLL^{+}22b]. User-Level [IATB20]. User-Perceived [CZJ^{+}22]. Users [CCWW24]. Userspace [PGY^{+}22]. Using [ASS^{+}24, AMvBI22, CHY^{+}24, COE20, CIZ^{+}20, DDN^{+}22, GJCC21, HWF^{+}22, IRB21, JLW^{+}24, KLH^{+}20b, LYZS24, LH22, LJW^{+}23, LPL23, MMGR23, MWL24, SHC^{+}22, SYS^{+}22, TKRB22, TCJ22a, WLH^{+}20a, WL20, WRLS22, WSX^{+}23, WCT21, WHRL21, YCZC22, ZZZ^{+}24a, ZGZ^{+}24, ZLT^{+}24, SGJ^{+}20]. Utility [HNKO20]. Utility- [HNKO20]. Utilization [ZLGZ23]. Utilize [WLH20b].

Validation [FZD^{+}24]. Value [ACH^{+}20, JLY^{+}23a, JLL^{+}20, KMA^{+}20, LCLW21, LCM^{+}20, PM22, QXL^{+}20, SDZ21, TWX22, WZZ^{+}20, WQZD25, ZZM^{+}23]. Value- [WZZ^{+}20]. Value-Oriented [KMA^{+}20]. Valued [WIBD22]. Variability [OAM^{+}25]. Variable [CHM^{+}20, CA20b]. Variance [PHY20]. Variance-Reduced [PHY20]. Varied [LAY21]. vCPU [IYAK23, WYZY24]. Vector [AHSW23, CXÖ^{+}20, HC24, MGG^{+}20, MFYB22, PK21]. Vectorisation [PRL20]. Vectorization [GDS^{+}22, GK21]. Vectorized [LCX^{+}24, SP20]. Vectors [LAY21]. VEE [JLQ^{+}23]. Vehicles [TWYL20]. Velocity [BGZR21, BPP21]. Verifiable [WPG^{+}22]. Verification [CA20a, LZM^{+}20, XYL^{+}21, YTL^{+}23]. Versa [YZL24]. Versa-DNN [YZL24]. Versatile [YZL24]. Versioned [SKW23]. Vertical [ZLP^{+}24]. Very [KKA^{+}20]. vGPU [LZM^{+}20]. Via [CTBT21, DZS^{+}21, GSH^{+}21, HLZ^{+}21, MHZ^{+}22, WHC^{+}21, ZWL^{+}21, ZGQ^{+}21, CRZ^{+}23, CYZ^{+}23, CZR20, CC22, CTD^{+}23, DGE^{+}24, FFQ^{+}22, GSL^{+}20, GGL^{+}23, HW22, JQG^{+}22, JBY^{+}23, LZJ^{+}20, LS21, LCZ^{+}23, LWZ23b, LHXH22, LCCZ20b, NJG^{+}22, PZL^{+}22, RCLJT22, SFW^{+}24, SGH^{+}23, SBM24, SPS^{+}24, SYT20, SCA23, SLG^{+}23, TWY^{+}20, TXX^{+}21, WNL20, WHM^{+}24, WWJL24, XXW^{+}24, YLC^{+}23, YZWT20, YWL^{+}24, ZFW^{+}20, ZZH^{+}20b, ZLW^{+}23]. Video [CQZ^{+}21, CWL22, HYL^{+}24, LSL^{+}23, SLY^{+}23, XJX24, ZZD^{+}24b]. Violations [LyLG^{+}25]. Virtual [FXL^{+}23, FZD^{+}24, HTB22, HYL^{+}23, JLQ^{+}23, JLG^{+}23, KPHA20, MSSK21, WDL^{+}20, YWH^{+}21, ZDK^{+}22]. Virtualization [GCL^{+}21, KSVR23, LYGG20, QTR21, YLT^{+}21]. Virtualized [ASMA21, ZLC^{+}22]. ViT [LDJX^{+}23]. ViTrack [CWL22]. VLSI [QQD^{+}24]. VM [IYAK23, ZLZ^{+}23]. VM-Agnostic [IYAK23]. Vol [Ano20]. Volume [GGO21, KLH^{+}20a, KAA20, LZJ^{+}20]. Volume-Discounting [LZJ^{+}20]. Voluminous [BPP21]. VPIC [BTL^{+}22]. vPipe [ZLC^{+}22]. VQL [WPG^{+}22].

Wait [AD25, LBNN^{+}21, LGH^{+}24]. Wait-Avoiding [LBNN^{+}21]. WAMP [YLW^{+}22]. WAN [WLP^{+}23]. Warp [ZSL^{+}21]. Warp-Level [ZSL^{+}21]. Warsaw [MKKS21]. Washington [LCG^{+}21]. WASP [HTL^{+}24]. Wavelet [SNK20]. Web [JBLJ23, LHQ^{+}20, LZWW22, JLJ21]. WEED [SNK20]. WEED-MC [SNK20]. Weight [IXS22]. Weighted [YOM21]. Welfare [TWYL20]. Welfare-Maximized [TWYL20]. Well [OHWL21].

Well-Bounded [OHWL21]. WFBP [SCL21a]. Which [KBA^{+}24]. While [KAA20]. Whole [CCYC21]. Wide [ACDK20, HKL^{+}20, SKW23, SPS^{+}24]. Wide-Area [SKW23]. Widely [SBM24]. Wiera [OQCW20]. Win [LN24]. Win-Win [LN24]. Window [MDC24]. Window-Based [MDC24]. Wireless [FLPL22, MDM22, ZMY^{+}24]. Wisely [SCL21a]. Within [LDZ^{+}24]. Without [PZZ^{+}22]. Work [CGH^{+}22, WQKH20]. Work-Group [WQKH20]. Work-Stealing [CGH^{+}22]. Workflow [CLG^{+}21, DS23, DFLG21, FSF^{+}20, JHB24, JSL^{+}25, LWZ^{+}22, MRFP20, TCT^{+}22, TCJ22b, VMT^{+}20]. Workflows [BGZR21, RCLJT22]. Workgroup [YT20]. Working [LZZ21]. Workload [CC22, HTL^{+}24, KFS^{+}21, LZWL22, SDBM23, SDHQ21, YLW^{+}22, ZGM21]. Workload-Aware [HTL^{+}24, SDHQ21, YLW^{+}22]. Workloads [CHM^{+}20, GZS^{+}25, SSKG21, TSV21, WZL^{+}22]. Worldwide [FMP^{+}23]. WPaxos [ACDK20]. Write [LXC24]. Writes [SLLL20]. Wukong [YCZC22].

X [DPGG22, DHH^{+}22, FGH^{+}22, KSW^{+}22, LAG^{+}22, LL22, PKJ^{+}22, XMW^{+}24, ZCZ^{+}22, HBG^{+}22]. X-Ray [DPGG22, DHH^{+}22, FGH^{+}22, KSW^{+}22, LAG^{+}22, LL22, PKJ^{+}22, ZCZ^{+}22, HBG^{+}22]. X-Shard [XMW^{+}24]. x86 [KPA^{+}20, LJZ^{+}20, ZGZ^{+}23, WJZ^{+}24]. Xeon [PRL20, SWOM20]. XMSS [WDCK23].

YARN [LYL^{+}20a]. YuenyeungSpTRSV [ZSL^{+}21].

Zero [AD25, CEP22, MHLM24]. Zero-Knowledge [MHLM24]. Zero-Wait [AD25]. Zone [GM21]. Zurich [BK21, KSW^{+}22].

References

Alirezazadeh:2023:SAA

[AA23] Saeid Alirezazadeh and Luís A. Alexandre. Static algorithm allocation with duplication in robotic network cloud systems. IEEE Transactions on Parallel and Distributed Systems, 34 (6):1897–1908, June 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ali:2024:CSC

[AA24] Isra M. Ali and Mohamed M. Abdallah. On off-chaining smart contract runtime protection: a queuing model approach. IEEE Transactions on Parallel and Distributed Systems, 35(8):1345–1359, August 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Abubaker:2021:TLB

[AAA21] N. Abubaker, S. Acer, and C. Aykanat. True load balancing for matricized tensor times Khatri–Rao product. IEEE Transactions on Parallel and Distributed Systems, 32(8): 1974–1986, August 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Aljundi:2022:BGE

[AAK22] Amro Alabsi Aljundi, Taha Atahan Akyildiz, and Kamer Kaya. Boosting graph embedding on a single GPU. IEEE Transactions on Parallel and Distributed Systems, 33(11):3092– 3105, November 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Aral:2021:LSF

[AB21] A. Aral and I. Brandi. Learning spatiotemporal failure dependencies for resilient edge computing services. IEEE Transactions on Parallel and Distributed Systems, 32(7):1578– 1590, July 2021. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Arifuzzaman:2023:FFE

[ABBA23] Md Arifuzzaman, Brian Bockelman, James Basney, and Engin Arslan. Falcon: Fair and efficient online file transfer optimization. IEEE Transactions on Parallel and Distributed Systems, 34(8):2265– 2278, August 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

AlJawarneh:2024:SQA

[ABC^{+}24] Isam Mashhour Al Jawarneh, Paolo Bellavista, Antonio Corradi, Luca Foschini, and Rebecca Montanari. SpatialSSJP: QoS-aware adaptive approximate stream-static spatial join processor. IEEE Transactions on Parallel and Distributed Systems, 35(1):73–88, January 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Allegretti:2020:OBB

[ABG20] S. Allegretti, F. Bolelli, and C. Grana. Optimized blockbased algorithms to label connected components on GPUs. IEEE Transactions on Parallel and Distributed Systems, 31(2): 423–438, February 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Augustine:2022:DGR

[ACC^{+}22] John Augustine, Keerti Choudhary, Avi Cohen, David Peleg, Sumathi Sivasubramaniam, and Suman Sourav. Distributed graph realizations. IEEE Transactions on Parallel and Distributed Systems, 33 (6):1321–1337, June 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ailijiang:2020:WWA

[ACDK20] A. Ailijiang, A. Charapko, M. Demirbas, and T. Kosar. WPaxos: Wide area network flexible consensus. IEEE Transactions on Parallel and Distributed Systems, 31(1):211– 223, January 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Anwar:2020:CSK

[ACH^{+}20] A. Anwar, Y. Cheng, H. Huang, J. Han, H. Sim, D. Lee, F. Douglis, and A. R. Butt. Customizable scale-out key– value stores. IEEE Transactions on Parallel and Distributed Systems, 31(9):2081– 2096, September 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Abdulah:2022:AGM

[ACP^{+}22] Sameh Abdulah, Qinglei Cao, Yu Pei, George Bosilca, Jack Dongarra, Marc G. Genton, David E. Keyes, Hatem Ltaief, and Ying Sun. Accelerating geostatistical modeling and prediction with mixedprecision computations: a highproductivity approach with PaRSEC. IEEE Transactions on Parallel and Distributed Systems, 33(4):964–976, April 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Anselmi:2025:BSF

[AD25] Jonatha Anselmi and Josu Doncel. Balanced splitting: a framework for achieving zerowait in the multiserver-job model. IEEE Transactions on Parallel and Distributed Systems, 36(1):43–54, January 2025. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Alappat:2023:LBB

[AHSW23] Christie Alappat, Georg Hager, Olaf Schenk, and Gerhard Wellein. Level-based blocking for sparse matrices: Sparse matrix-power-vector multiplication. IEEE Transactions on Parallel and Distributed Systems, 34(2):581–597, February 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Abubaker:2022:SUM

[AKA22] Nabil Abubaker, M. Ozan Karsavuran, and Cevdet Aykanat. Scalable unsupervised ML: Latency hiding in distributed sparse tensor decomposition. IEEE Transactions on Parallel and Distributed Systems, 33(11):3028–3040, November 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Alourani:2020:BFS

[AKG20] A. Alourani, A. D. Kshemkalyani, and M. Grechanik. T-BASIR: Finding shutdown bugs for cloud-based applications in cloud spot markets. IEEE Transactions on Parallel and Distributed Systems, 31(8): 1912–1924, August 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Alabdulatif:2020:FHB

[AKZ^{+}20] A. Alabdulatif, I. Khalil, A. Y. Zomaya, Z. Tari, and X. Yi. Fully homomorphic based privacy-preserving distributed expectation maximization on cloud. IEEE Transactions on Parallel and Distributed Systems, 31(11):2668– 2681, November 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Alomairy:2020:ALS

[ALAK20] R. Alomairy, H. Ltaief, M. Abduljabbar, and D. Keyes. Abstraction layer for standardizing APIs of task-based engines. IEEE Transactions on Parallel and Distributed Systems, 31(11):2482–2495, November 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ahmadi:2021:PJE

[AMKS21] A. Ahmadi, F. Manganiello, A. Khademi, and M. C. Smith. A parallel Jacobi-embedded Gauss–Seidel method. IEEE Transactions on Parallel and Distributed Systems, 32(6): 1452–1464, June 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Alam:2022:ISI

[AMN22] Sadaf R. Alam, Lois Curfman McInnes, and Kengo Nakajima. IEEE special issue on innovative R&D toward the exascale era. IEEE Transactions on Parallel and Distributed Systems, 33(4):736–738, April 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Andreadis:2022:CDD

[AMvBI22] Georgios Andreadis, Fabian Mastenbroek, Vincent van Beek, and Alexandru Iosup. Capelin: Data-driven compute capacity procurement for cloud datacenters using portfolios of scenarios. IEEE Transactions on Parallel and Distributed Systems, 33(1):26–39, January 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Asri:2021:HAI

[AMW^{+}21] M. Asri, D. Malhotra, J. Wang, G. Biros, L. K. John, and A. Gerstlauer. Hardware accelerator integration tradeoffs for high-performance computing: a case study of GEMM acceleration in N-body methods. IEEE Transactions on Parallel and Distributed Systems, 32(8): 2035–2048, August 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Anonymous:2020:IIT

[Ano20] Anonymous. 2019 index IEEE Transactions on Parallel and Distributed Systems vol. 30. IEEE Transactions on Parallel and Distributed Systems, 31 (2):1–33, February 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Anonymous:2021:RL

[Ano21] Anonymous. 2020 reviewers list*. IEEE Transactions on Parallel and Distributed Systems, 32(5):1270–1276, May 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Anselmi:2020:CSB

[Ans20] J. Anselmi. Combining sizebased load balancing with round-robin for scalable low latency. IEEE Transactions on Parallel and Distributed Systems, 31(4):886–896, April 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ao:2020:RCR

[AP20] W. C. Ao and K. Psounis.

Resource-constrained replication strategies for hierarchical and heterogeneous tasks. IEEE Transactions on Parallel and Distributed Systems, 31(4):793– 804, April 2020. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Azad:2022:CBS

[ASH^{+}22] Ariful Azad, Oguz Selvitopi, Md Taufique Hussain, John R. Gilbert, and Aydın Buluç. Combinatorial BLAS 2.0: Scaling combinatorial algorithms on distributed-memory systems. IEEE Transactions on Parallel and Distributed Systems, 33(4):989–1001, April 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ansari:2020:SMP

[ASLPE20] M. Ansari, J. Saber-Latibari, M. Pasandideh, and A. Ejlali. Simultaneous management of peak-power and reliability in heterogeneous multicore embedded systems. IEEE Transactions on Parallel and Distributed Systems, 31(3):623– 633, March 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ahmadian:2021:EET

[ASMA21] S. Ahmadian, R. Salkhordeh, O. Mutlu, and H. Asadi. ET-ICA: Efficient two-level I/O caching architecture for virtualized platforms. IEEE Transactions on Parallel and Distributed Systems, 32(10):2415– 2433, October 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Akhremtsev:2020:HQS

[ASS20] Y. Akhremtsev, P. Sanders, and C. Schulz. High-quality shared-memory graph partitioning. IEEE Transactions on Parallel and Distributed Systems, 31(11):2710– 2722, November 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Aksar:2024:RPA

[ASS^{+}24] Burak Aksar, Efe Sencan, Benjamin Schwaller, Omar Aaziz, Vitus J. Leung, Jim Brandt, Brian Kulis, Manuel Egele, and Ayse K. Coskun. Runtime performance anomaly diagnosis in production HPC systems using active learning. IEEE Transactions on Parallel and Distributed Systems, 35(4):693– 706, April 2024. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Audrito:2023:PFS

[ATF23] Giorgio Audrito, Federico Terraneo, and William Fornaciari. FCPP+Miosix: Scaling aggregate programming to embedded systems. IEEE Transactions on Parallel and Distributed Systems, 34(3):869–880, March 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Burgos:2022:ECS

[BADP22] Aldenio Burgos, Eduardo Alchieri, Fernando Dotti, and Fernando Pedone. Exploiting concurrency in sharded parallel state machine replication. IEEE Transactions on Parallel and Distributed Systems, 33(9): 2133–2147, September 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bahreini:2022:MRA

[BBG22] Tayebeh Bahreini, Hossein Badri, and Daniel Grosu. Mechanisms for resource allocation and pricing in mobile edge computing systems. IEEE Transactions on Parallel and Distributed Systems, 33 (3):667–682, March 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Badri:2020:EAA

[BBGY20] H. Badri, T. Bahreini, D. Grosu, and K. Yang. Energy-aware application placement in mobile edge computing: a stochastic optimization approach. IEEE Transactions on Parallel and Distributed Systems, 31(4):909– 922, April 2020. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Balty:2023:FFP

[BCG23] Pierre Balty, Philippe Chatelain, and Thomas Gillis. FLUPS — a flexible and performant massively parallel Fourier transform library. IEEE Transactions on Parallel and Distributed Systems, 34(7):2011– 2024, July 2023. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bernaschi:2023:MGA

[BCVD23] Massimo Bernaschi, Alessandro Celestini, Flavio Vella, and Pasqua D’Ambra. A multi-GPU aggregation-based AMG preconditioner for iterative linear solvers. IEEE Transactions on Parallel and Distributed Systems, 34(8):2365– 2376, August 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Besta:2021:HPR

[BDS^{+}21] M. Besta, J. Domke, M. Schneider, M. Konieczny, S. D. Girolamo, T. Schneider, A. Singla, and T. Hoefler. Highperformance routing with multipathing and path diversity in Ethernet and HPC networks. IEEE Transactions on Parallel and Distributed Systems, 32 (4):943–959, April 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bandara:2024:LRM

[BDWH24] Sahan Bandara, Anthony Ducimo, Chunshu Wu, and Martin Herbordt. Long-range MD electrostatics force computation on FPGAs. IEEE Transactions on Parallel and Distributed Systems, 35(10): 1690–1707, October 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Besta:2023:PSP

[BFK^{+}23] Maciej Besta, Marc Fischer, Vasiliki Kalavri, Michael Kapralov, and Torsten Hoefler. Practice of streaming processing of dynamic graphs: Concepts, models, and systems. IEEE Transactions on Parallel and Distributed Systems, 34 (6):1860–1876, June 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Barika:2021:OST

[BGZR21] M. Barika, S. Garg, A. Y.

Zomaya, and R. Ranjan. Online scheduling technique to handle data velocity changes in stream workflows. IEEE Transactions on Parallel and Distributed Systems, 32(8):2115– 2130, August 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Burger:2021:CPN

[BK21] Manuel Burger and Jan Kleine.

Critique of planetary normal mode computation: Parallel algorithms, performance, and reproducibility by SCC Team From ETH Zurich. IEEE Transactions on Parallel and Distributed Systems, 32(11): 2627–2630, November 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). See [SLX^{+}21b].

Bhuiyan:2020:EEP

[BLK^{+}20] A. Bhuiyan, D. Liu, A. Khan, A. Saifullah, N. Guan, and Z. Guo. Energy-efficient parallel real-time scheduling on clustered multi-core. IEEE Transactions on Parallel and Distributed Systems, 31(9):2097– 2111, September 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Blakeney:2021:PBK

[BLYZ21] C. Blakeney, X. Li, Y. Yan, and Z. Zong. Parallel blockwise knowledge distillation for deep neural network compression. IEEE Transactions on Parallel and Distributed Systems, 32 (7):1765–1776, July 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bajunaid:2020:DDD

[BM20] N. Bajunaid and D. A. Menascé. Data-driven derivation of an analytic model for parallel servers with job replication. IEEE Transactions on Parallel and Distributed Systems, 31(10):2435–2452, October 2020. CODEN ITDSEO.

ISSN 1045-9219 (print), 1558-2183 (electronic).

Bhatta:2022:BAA

[BM22] Dixit Bhatta and Lena Mashayekhy.

A bifactor approximation algorithm for cloudlet placement in edge computing. IEEE Transactions on Parallel and Distributed Systems, 33(8):1787– 1798, August 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Borghesi:2022:ADA

[BMMB22] Andrea Borghesi, Martin Molan, Michela Milano, and Andrea Bartolini. Anomaly detection and anticipation in high performance computing systems. IEEE Transactions on Parallel and Distributed Systems, 33 (4):739–750, April 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Battula:2021:GSM

[BOGM21] S. K. Battula, M. M. O’Reilly, S. Garg, and J. Montgomery. A generic stochastic model for resource availability in fog computing environments. IEEE Transactions on Parallel and Distributed Systems, 32(4):960– 974, April 2021. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Buddhika:2021:PLS

[BPP21] T. Buddhika, S. L. Pallickara, and S. Pallickara. Pebbles: Leveraging sketches for processing voluminous, high velocity data streams. IEEE Transactions on Parallel and Distributed Systems, 32(8):2005– 2020, August 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Boito:2023:ISS

[BPTV23] Francieli Boito, Guillaume Pallez, Luan Teylo, and Nicolas Vidal. IO-Sets: Simple and efficient approaches for I/O bandwidth management. IEEE Transactions on Parallel and Distributed Systems, 34(10): 2783–2796, October 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bamberg:2023:SCE

[BPW^{+}23] Lennart Bamberg, Arash Pourtaherian, Luc Waeijen, Anupam Chahar, and Orlando Moreira. Synapse compression for event-based convolutionalneural-network accelerators. IEEE Transactions on Parallel and Distributed Systems, 34 (4):1227–1240, April 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bilbao:2023:DFL

[BSPM23] Carlos Bilbao, Juan Carlos Saez, and Manuel Prieto-Matias. Divide&Content: a fair OS-level resource manager for contention balancing on NUMA multicores. IEEE Transactions on Parallel and Distributed Systems, 34(11):2928– 2945, November 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bird:2022:VNG

[BTL^{+}22] Robert Bird, Nigel Tan, Scott V. Luedtke, Stephen Lien Harrell, Michela Taufer, and Brian Albright. VPIC 2.0: Next generation particle-in-cell simulations. IEEE Transactions on Parallel and Distributed Systems, 33(4):952– 963, April 2022. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Balaji:2021:GE

[BZS21] P. Balaji, J. Zhai, and M. Si.

Guest editorial. IEEE Transactions on Parallel and Distributed Systems, 32(7):1511– 1512, July 2021. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Charyyev:2020:RRI

[CA20a] B. Charyyev and E. Arslan.

RIVA: Robust integrity verification algorithm for high-speed file transfers. IEEE Transactions on Parallel and Distributed Systems, 31(6):1387– 1399, June 2020. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Criswell:2020:SPC

[CA20b] K. Criswell and T. Adegbija.

A survey of phase classification techniques for characterizing variable application behavior. IEEE Transactions on Parallel and Distributed Systems, 31(1):224–236, January 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cabrera:2020:DMO

[CAAB20] A. Cabrera, A. Acosta, F. Almeida, and V. Blanco. A dynamic multi objective approach for dynamic load balancing in heterogeneous systems. IEEE Transactions on Parallel and Distributed Systems, 31(10): 2421–2434, October 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cesarini:2020:CSR

[CBB^{+}20] D. Cesarini, A. Bartolini, A. Borghesi, C. Cavazzoni, M. Luisier, and L. Benini. Countdown slack: a run-time library to reduce energy footprint in large-scale MPI applications. IEEE Transactions on Parallel and Distributed Systems, 31(11):2696– 2709, November 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cebrian:2022:CAC

[CBB^{+}22] Juan M. Cebrian, Thibaud Balem, Adrián Barredo, Marc Casas, Miquel Moretó, Alberto Ros, and Alexandra Jimborean. Compiler-assisted compaction/ restoration of SIMD instructions. IEEE Transactions on Parallel and Distributed Systems, 33(4):779–791, April 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chwa:2022:NFA

[CBL22] Hoon Sung Chwa, Hyeongboo Baek, and Jinkyu Lee. Necessary feasibility analysis for mixed-criticality real-time embedded systems. IEEE Transactions on Parallel and Distributed Systems, 33(7):1520– 1537, July 2022. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2022:WBG

[CC22] YuAng Chen and Yeh-Ching Chung. Workload balancing via graph reordering on multicore systems. IEEE Transactions on Parallel and Distributed Systems, 33(5):1231– 1245, May 2022. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2023:UCS

[CC23] YuAng Chen and Yeh-Ching Chung. An unequal caching strategy for shared-memory graph analytics. IEEE Transactions on Parallel and Distributed Systems, 34(3):955– 967, March 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Che:2024:OEO

[CCWW24] Yuezhi Che, Dazhao Cheng, Xiao Wang, and Rujia Wang. Opca: Enabling optimistic concurrent access for multiple users in oblivious data storage. IEEE Transactions on Parallel and Distributed Systems, 35(11): 1891–1903, November 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2021:HTF

[CCYC21] Y.-L. Chen, B.-Y. Chang, C.- H. Yang, and T.-D. Chiueh. A high-throughput FPGA accelerator for short-read mapping of the whole human genome. IEEE Transactions on Parallel and Distributed Systems, 32 (6):1465–1478, June 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cui:2021:EEE

[CCZ^{+}21] W. Cui, Q. Chen, H. Zhao, M. Wei, X. Tang, and M. Guo. E2bird: Enhanced elastic batch for improving responsiveness and throughput of deep learning services. IEEE Transactions on Parallel and Distributed Systems, 32(6):1307– 1321, June 2021. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2024:CEE

[CCZW24] Qixiang Chen, Zhijun Chen, Kai Zhang, and X. Sean Wang. CLIC: an extensible and efficient cross-platform data analytics system. IEEE Transactions on Parallel and Distributed Systems, 35(1):34–45, January 2024. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cherrueau:2022:ELE

[CDvK^{+}22] Ronan-Alexandre Cherrueau, Marie Delavergne, Alexandre van Kempen, Adrien Lebre, Dimitri Pertin, Javier Rojas Balderrama, Anthony Simonet, and Matthieu Simonin. EnosLib: a library for experiment-driven research in distributed computing. IEEE Transactions on Parallel and Distributed Systems, 33(6): 1464–1477, June 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2022:NZD

[CEP22] Chao Chen, Greg Eisenhauer, and Santosh Pande. Near-zero downtime recovery from transient-error-induced crashes. IEEE Transactions on Parallel and Distributed Systems, 33(4):765–778, April 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2021:CPB

[CFLL21] L. Chen, Y. Feng, B. Li, and B. Li. A case for pricing bandwidth: Sharing datacenter networks with cost dominant fairness. IEEE Transactions on Parallel and Distributed Systems, 32(5):1256– 1269, May 2021. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2021:BBA

[CFLY21] J. Chen, J. Fang, W. Liu, and C. Yang. BALS: Blocked alternating least squares for parallel sparse matrix factorization on GPUs. IEEE Transactions on Parallel and Distributed Systems, 32(9):2291– 2302, September 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cataldo:2021:SSL

[CFM^{+}21a] R. Cataldo, R. Fernandes, K. J. M. Martin, J. Silveira, G. Sanchez, J. Sepúlveda, C. Marcon, and J.-P. Diguet. Subutai: Speeding up legacy parallel applications through data synchronization. IEEE Transactions on Parallel and Distributed Systems, 32(5): 1102–1116, May 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cheng:2021:CPN

[CFM^{+}21b] Yihua Cheng, Zejia Fan, Jing Mai, Yifan Wu, Pengcheng Xu, Yuxuan Yan, Zhenxin Fu, and Yun Liang. Critique of planetary normal mode computation: Parallel algorithms, performance, and reproducibility by SCC Team From Peking University. IEEE Transactions on Parallel and Distributed Systems, 32(11):2643– 2645, November 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). See [SLX^{+}21b].

Chen:2022:AGP

[CGC^{+}22] Jiamin Chen, Jianliang Gao, Yibo Chen, Babatounde Moctard Oloulade, Tengfei Lyu, and Zhao Li. Auto-GNAS: a parallel graph neural architecture search framework. IEEE Transactions on Parallel and Distributed Systems, 33(11): 3117–3128, November 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Copik:2022:WSP

[CGH^{+}22] Marcin Copik, Tobias Grosser, Torsten Hoefler, Paolo Bientinesi, and Benjamin Berkels. Work-stealing prefix scan: Addressing load imbalance in large-scale image registration. IEEE Transactions on Parallel and Distributed Systems, 33 (3):523–535, March 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cheng:2022:LLO

[CGL^{+}22] Geyao Cheng, Deke Guo, Lailong Luo, Junxu Xia, and Siyuan Gu. LOFS: a lightweight online file storage strategy for effective data deduplication at network edge. IEEE Transactions on Parallel and Distributed Systems, 33(10):2263– 2276, October 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2020:PPM

[CGLC20] T. Chen, X. Gao, T. Liao, and G. Chen. Pache: a packet management scheme of cache in data center networks. IEEE Transactions on Parallel and Distributed Systems, 31(2): 253–265, February 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cicirelli:2021:AGL

[CGM21] F. Cicirelli, A. Giordano, and C. Mastroianni. Analysis of global and local synchronization in parallel computing. IEEE Transactions on Parallel and Distributed Systems, 32 (5):988–1000, May 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2020:TAP

[CHM^{+}20] Z. Chen, J. Hu, G. Min, A. Y. Zomaya, and T. El-Ghazawi. Towards accurate prediction for high-dimensional and highly-variable cloud workloads with deep learning. IEEE Transactions on Parallel and Distributed Systems, 31(4):923– 934, April 2020. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2024:RTO

[CHY^{+}24] Xing Chen, Shengxi Hu, Chujia Yu, Zheyi Chen, and Geyong Min. Real-time offloading for dependent and parallel tasks in cloud-edge environments using deep reinforcement learning. IEEE Transactions on Parallel and Distributed Systems, 35 (3):391–404, March 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Costero:2020:RMP

[CIZ^{+}20] L. Costero, A. Iranfar, M. Zapater, F. D. Igual, K. Olcoz, and D. Atienza. Resource management for powerconstrained HEVC transcoding using reinforcement learning. IEEE Transactions on Parallel and Distributed Systems, 31(12):2834–2850, December 2020. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2022:EOM

[CJLW22] Shutong Chen, Lei Jiao, Fangming Liu, and Lin Wang. EdgeDR: an online mechanism design for demand response in edge clouds. IEEE Transactions on Parallel and Distributed Systems, 33(2):343– 358, February 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cho:2021:FAB

[CKO^{+}21] S. Cho, W. Kim, S. Oh, C. Kim, K. Koh, and B. Nam. Failureatomic byte-addressable R-tree for persistent memory. IEEE Transactions on Parallel and Distributed Systems, 32(3):601– 614, March 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cho:2020:SPR

[CKS^{+}20] H. Cho, C. Kim, J. Sun, A. Easwaran, J. Park, and B. Choi. Scheduling parallel real-time tasks on the minimum number of processors. IEEE Transactions on Parallel and Distributed Systems, 31(1): 171–186, January 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Champati:2020:SRT

[CL20a] J. P. Champati and B. Liang. Single restart with time stamps for parallel task processing with known and unknown processors. IEEE Transactions on Parallel and Distributed Systems, 31(1):187–200, January 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2020:ACF

[CL20b] Y. Chen and A. Louri. An approximate communication framework for network-on-chips. IEEE Transactions on Parallel and Distributed Systems, 31(6):1434–1446, June 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cheng:2021:AEE

[CLG^{+}21] Y. Cheng, D. Li, Z. Guo, B. Jiang, J. Geng, W. Bai, J. Wu, and Y. Xiong. Accelerating end-to-end deep learning workflow with codesign of data preprocessing and scheduling. IEEE Transactions on Parallel and Distributed Systems, 32 (7):1802–1814, July 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2024:AAD

[CLG^{+}24] Jinfan Chen, Shigang Li, Ran Guo, Jinhui Yuan, and Torsten Hoefler. AutoDDL: Automatic distributed deep learning with near-optimal bandwidth cost. IEEE Transactions on Parallel and Distributed Systems, 35(8): 1331–1344, August 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2021:SEC

[CLL^{+}21] Y. Chen, L. Lin, B. Li, Q. Wang, and Q. Zhang. Silhouette: Efficient cloud configuration exploration for largescale analytics. IEEE Transactions on Parallel and Distributed Systems, 32(8):2049– 2061, August 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2022:ONT

[CLL22] Li Chen, Shuhao Liu, and Baochun Li. Optimizing network transfers for data analytic jobs across geo-distributed datacenters. IEEE Transactions on Parallel and Distributed Systems, 33(2):403–414, February 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2022:FFG

[CLMW22] Fahao Chen, Peng Li, Toshiaki Miyazaki, and Celimuge Wu. FedGraph: Federated graph learning with intelligent sampling. IEEE Transactions on Parallel and Distributed Systems, 33(8):1775–1786, August 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cheng:2023:WDM

[CLX^{+}23] Geyao Cheng, Lailong Luo, Junxu Xia, Deke Guo, and Yuchen Sun. When deduplication meets migration: an efficient and adaptive strategy in distributed storage systems. IEEE Transactions on Parallel and Distributed Systems, 34 (10):2749–2766, October 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2024:TTJ

[CLY^{+}24] Shuangwu Chen, Jiangming Li, Qifeng Yuan, Huasen He, Sen Li, and Jian Yang. Twotimescale joint optimization of task scheduling and resource scaling in multi-data center system based on multi-agent deep reinforcement learning. IEEE Transactions on Parallel and Distributed Systems, 35(12): 2331–2346, December 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ci:2020:RPB

[CLZ^{+}20] Y. Ci, M. R. Lyu, Z. Zhang, D. Zuo, and X. Yang. Random priority-based thrashing control for distributed shared memory. IEEE Transactions on Parallel and Distributed Systems, 31(3):663–674, March 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cheng:2021:WDP

[CLZ^{+}21] D. Cheng, S. Li, H. Zhang, F. Xia, and Y. Zhang. Why dataset properties bound the scalability of parallel machine learning training algorithms. IEEE Transactions on Parallel and Distributed Systems, 32 (7):1702–1712, July 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cao:2022:CHP

[CLZ^{+}22a] Jiamin Cao, Ying Liu, Yu Zhou, Lin He, Chen Sun, Yangyang Wang, and Mingwei Xu. CoFilter: High-performance switch-accelerated stateful packet filter for bare-metal servers. IEEE Transactions on Parallel and Distributed Systems, 33(9): 2249–2262, September 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cui:2022:PEB

[CLZ^{+}22b] Jie Cui, Bei Li, Hong Zhong, Geyong Min, Yan Xu, and Lu Liu. A practical and efficient bidirectional access control scheme for cloud-edge data sharing. IEEE Transactions on Parallel and Distributed Systems, 33(2):476–488, February 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ci:2024:KEP

[CLZ^{+}24] Yi-Wei Ci, Michael R. Lyu, Zhan Zhang, De-Cheng Zuo, and Xiao-Zong Yang. KLNK: Expanding page boundaries in a distributed shared memory system. IEEE Transactions on Parallel and Distributed Systems, 35(9):1524– 1535, September 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2020:ASN

[CLZP20] S. Chen, L. Liu, W. Zhang, and L. Peng. Architectural support for NVRAM persistence in GPUs. IEEE Transactions on Parallel and Distributed Systems, 31(5):1107– 1120, May 2020. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2020:GAR

[CMLH20] G. Chen, H. Meng, Y. Liang, and K. Huang. GPUaccelerated real-time stereo estimation with binary neural network. IEEE Transactions on Parallel and Distributed Systems, 31(12):2896–2907, December 2020. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Canon:2020:OST

[CMSV20] L. Canon, L. Marchal, B. Simon, and F. Vivien. Online scheduling of task graphs on heterogeneous platforms. IEEE Transactions on Parallel and Distributed Systems, 31(3):721– 732, March 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2020:TUC

[CMX^{+}20] F. Chen, F. Meng, T. Xiang, H. Dai, J. Li, and J. Qin. Towards usable cloud storage auditing. IEEE Transactions on Parallel and Distributed Systems, 31(11):2605– 2617, November 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cho:2020:PMP

[COE20] Y. Cho, S. Oh, and B. Egger.

Performance modeling of parallel loops on multi-socket platforms using queueing systems. IEEE Transactions on Parallel and Distributed Systems, 31(2): 318–331, February 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2020:DLR

[CQW^{+}20] Z. Chen, W. Quan, M. Wen, J. Fang, J. Yu, C. Zhang, and L. Luo. Deep learning research and development platform: Characterizing and scheduling with QoS guarantees on GPU clusters. IEEE Transactions on Parallel and Distributed Systems, 31(1):34– 50, January 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2021:CNC

[CQZ^{+}21] N. Chen, S. Quan, S. Zhang, Z. Qian, Y. Jin, J. Wu, W. Li, and S. Lu. Cuttlefish: Neural configuration adaptation for video analysis in live augmented reality. IEEE Transactions on Parallel and Distributed Systems, 32(4):830– 841, April 2021. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2024:CCP

[CQZ^{+}24] Kai Chen, Qingjun Qu, Feng Zhu, Zhengming Yi, and Wenjie Tang. CPLNS: Cooperative parallel large neighborhood search for large-scale multiagent path finding. IEEE Transactions on Parallel and Distributed Systems, 35(11): 2069–2086, November 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cao:2023:CPF

[CRZ^{+}23] Juncheng Cao, Kaiyuan Rong, Mingshu Zhai, Zeyu Song, Yanyu Ren, Yuxi Zhu, Wentao Han, and Jidong Zhai. Critique of A Parallel Framework for Constraint-Based Bayesian Network Learning via Markov Blanket Discovery by SCC Team From Tsinghua University. IEEE Transactions on Parallel and Distributed Systems, 34(6):1723–1726, June 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). See [SCA23].

Chang:2020:CTE

[CSJB20] J. Chang, B. Shao, Y. Ji, and G. Bian. Comment on A Tag Encoding Scheme Against Pollution Attack to Linear Network Coding. IEEE Transactions on Parallel and Distributed Systems, 31(11):2618– 2619, November 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cetin:2021:HSB

[ÇSS21] G. S. Çetin, E. Sava, and B. Sunar. Homomorphic sorting with better scalability. IEEE Transactions on Parallel and Distributed Systems, 32 (4):760–771, April 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cui:2023:CID

[CSZ^{+}23] Jie Cui, Hu Sun, Hong Zhong, Jing Zhang, Lu Wei, Irina Bolodurina, and Debiao He. Collaborative intrusion detection system for SDVN: a fairness federated deep learning approach. IEEE Transactions on Parallel and Distributed Systems, 34(9): 2512–2528, September 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chapp:2021:IDS

[CTBT21] Dylan Chapp, Nigel Tan, Sanjukta Bhowmick, and Michela Taufer. Identifying degree and sources of non-determinism in MPI applications via graph kernels. IEEE Transactions on Parallel and Distributed Systems, 32(12):2936–2952, December 2021. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2023:SPL

[CTD^{+}23] Yanqing Chen, Chen Tian, Jiaqing Dong, Song Feng, Xu Zhang, Chang Liu, Peiwen Yu, Nai Xia, Wanchun Dou, and Guihai Chen. Swing: Providing long-range lossless RDMA via PFC-Relay. IEEE Transactions on Parallel and Distributed Systems, 34(1):63– 75, January 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2024:PCD

[CTL24] Wei-Mei Chen, Hsin-Hung Tsai, and Joon Fong Ling. Parallel computation of dominance scores for multidimensional datasets on GPUs. IEEE Transactions on Parallel and Distributed Systems, 35(6):764– 776, June 2024. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2022:BGH

[CWC^{+}22] Zexi Chen, Ting Wang, Haibin Cai, Subrota Kumar Mondal, and Jyoti Prakash Sahoo. BLBgcForest: a high-performance distributed deep forest with adaptive sub-forest splitting. IEEE Transactions on Parallel and Distributed Systems, 33(11):3141–3152, November 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2023:DFF

[CWC^{+}23] Qian Chen, Zilong Wang, Jiawei Chen, Haonan Yan, and Xiaodong Lin. Dap-FL: Federated learning flourishes by adaptive tuning and secure aggregation. IEEE Transactions on Parallel and Distributed Systems, 34(6):1923– 1941, June 2023. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cheng:2021:NAL

[CWL^{+}21] L. Cheng, Y. Wang, Q. Liu, D. H. J. Epema, C. Liu, Y. Mao, and J. Murphy. Network-aware locality scheduling for distributed data operators in data centers. IEEE Transactions on Parallel and Distributed Systems, 32(6): 1494–1510, June 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cheng:2022:VET

[CWL22] Linsong Cheng, Jiliang Wang, and Yinghui Li. ViTrack: Efficient tracking on the edge for commodity video surveillance systems. IEEE Transactions on Parallel and Distributed Systems, 33(3):723–735, March 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2024:LPG

[CWXC24] Rong Chen, Xingda Wei, Xiating Xie, and Haibo Chen. Locality-preserving graph traversal with split live migration. IEEE Transactions on Parallel and Distributed Systems, 35 (10):1810–1825, October 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chiang:2023:DMT

[CWY^{+}23] Sheng-Hao Chiang, Chih-Hang Wang, De-Nian Yang, Wanjiun Liao, and Wen-Tsuen Chen. Distributed multicast traffic engineering for multidomain software-defined networks. IEEE Transactions on Parallel and Distributed Systems, 34(2):446–462, February 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cheng:2023:CDC

[CXL^{+}23] Zhipeng Cheng, Xiaoyu Xia, Minghui Liwang, Xuwei Fan, Yanglong Sun, Xianbin Wang, and Lianfen Huang. CHEESE: Distributed clustering-based hybrid federated split learning over edge networks. IEEE Transactions on Parallel and Distributed Systems, 34(12): 3174–3191, December 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2020:AAE

[CXÖ^{+}20] Y. Chen, G. Xiao, M. T. Özsu, C. Liu, A. Y. Zomaya, and T. Li. aeSpTV: an adaptive and efficient framework for sparse tensor–vector product kernel on a high-performance computing platform. IEEE Transactions on Parallel and Distributed Systems, 31(10):2329– 2345, October 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2024:SOI

[CXW^{+}24] Chen Chen, Hong Xu, Wei Wang, Baochun Li, Bo Li, Li Chen, and Gong Zhang. Synchronize only the immature parameters: Communicationefficient federated learning by freezing parameters adaptively. IEEE Transactions on Parallel and Distributed Systems, 35 (7):1155–1173, July 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cai:2023:TCD

[CYF^{+}23] Hui Cai, Yuanyuan Yang, Weibei Fan, Fu Xiao, and Yanmin Zhu. Towards correlated data trading for highdimensional private data. IEEE Transactions on Parallel and Distributed Systems, 34(3): 1047–1059, March 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2021:BDS

[CYH^{+}21] D. Chen, H. Yuan, S. Hu, Q. Wang, and C. Wang. BOSSA: a decentralized system for proofs of data retrievability and replication. IEEE Transactions on Parallel and Distributed Systems, 32(4):786– 798, April 2021. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cai:2022:OPT

[CYY^{+}22] Hui Cai, Fan Ye, Yuanyuan Yang, Yanmin Zhu, Jie Li, and Fu Xiao. Online pricing and trading of private data in correlated queries. IEEE Transactions on Parallel and Distributed Systems, 33(3):569– 585, March 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cao:2023:PAA

[CYZ^{+}23] Hang Cao, Liang Yuan, He Zhang, Yunquan Zhang, Baodong Wu, Kun Li, Shigang Li, Minghua Zhang, Pengqi Lu, and Junmin Xiao. AGCM-3DLF: Accelerating atmospheric general circulation model via 3-D parallelization and leap-format. IEEE Transactions on Parallel and Distributed Systems, 34 (3):766–780, March 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2024:DSS

[CYZ^{+}24] Hao-Rui Chen, Lei Yang, Xinglin Zhang, Jiaxing Shen, and Jiannong Cao. Distributed semi-supervised learning with consensus consistency on edge devices. IEEE Transactions on Parallel and Distributed Systems, 35(2):310–323, February 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2020:EEM

[CZH^{+}20a] Q. Chen, Z. Zheng, C. Hu, D. Wang, and F. Liu. Errata to On-Edge Multi-Task Transfer Learning: Model and Practice With Data-Driven Task Allocation. IEEE Transactions on Parallel and Distributed Systems, 31(11):2569, November 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2020:EMT

[CZH^{+}20b] Q. Chen, Z. Zheng, C. Hu, D. Wang, and F. Liu. Onedge multi-task transfer learning: Model and practice with data-driven task allocation. IEEE Transactions on Parallel and Distributed Systems, 31 (6):1357–1371, June 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2022:LUP

[CZJ^{+}22] Yu Chen, Sheng Zhang, Yibo Jin, Zhuzhong Qian, Mingjun Xiao, Jidong Ge, and Sanglu Lu. LOCUS: User-perceived delay-aware service placement and user allocation in MEC environment. IEEE Transactions on Parallel and Distributed Systems, 33(7):1581– 1592, July 2022. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2022:EEO

[CZL^{+}22] Xing Chen, Jianshan Zhang, Bing Lin, Zheyi Chen, Katinka Wolter, and Geyong Min. Energy-efficient offloading for DNN-based smart IoT systems in cloud-edge environments. IEEE Transactions on Parallel and Distributed Systems, 33 (3):683–697, March 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cheng:2024:GMA

[CZL^{+}24] Ke Cheng, Sheng Zhang, Meizhao Liu, Yingcheng Gu, Liu Wei, Huanyu Cheng, Kai Liu, Yu Song, Xiaohang Shi, Andong Zhu, and Lei Tang. GeoScale: Microservice autoscaling with cost budget in geo-distributed edge clouds. IEEE Transactions on Parallel and Distributed Systems, 35 (4):646–662, April 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2023:GUE

[CZP^{+}23] Longlong Chen, Jianfeng Zhu, Guiqiang Peng, Mingxu Liu, Shaojun Wei, and Leibo Liu. GEM: Ultra-efficient nearmemory reconfigurable acceleration for read mapping by dividing and predictive scattering. IEEE Transactions on Parallel and Distributed Systems, 34(12):3059–3072, December 2023. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2020:PLL

[CZR20] W. Chen, X. Zhou, and J. Rao. Preemptive and low latency datacenter scheduling via lightweight containers. IEEE Transactions on Parallel and Distributed Systems, 31(12): 2749–2762, December 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cong:2022:CAR

[CZZ^{+}22] Peijin Cong, Zhixing Zhang, Junlong Zhou, Xin Liu, Yao Liu, and Tongquan Wei. Customer adaptive resource provisioning for long-term cloud profit maximization under constrained budget. IEEE Transactions on Parallel and Distributed Systems, 33(6):1373– 1392, June 2022. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cong:2024:CES

[CZZM24] Rong Cong, Zhiwei Zhao, Linyuanqi Zhang, and Geyong Min. Cost-effective server deployment for multi-access edge networks: a cooperative scheme. IEEE Transactions on Parallel and Distributed Systems, 35(9):1583– 1597, September 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2023:DDS

[CZZY23] Zhenqian Chen, Xinkui Zhao, Chen Zhi, and Jianwei Yin. DeepBoot: Dynamic scheduling system for training and inference deep learning tasks in GPU cluster. IEEE Transactions on Parallel and Distributed Systems, 34(9):2553– 2567, September 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Demirci:2020:CPM

[DA20] G. V. Demirci and C. Aykanat. Cartesian partitioning models for 2D and 3D parallel SpGEMM algorithms. IEEE Transactions on Parallel and Distributed Systems, 31(12): 2763–2775, December 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

deFineLicht:2021:THL

[dBMH21] J. de Fine Licht, M. Besta, S. Meierhans, and T. Hoefler. Transformations of highlevel synthesis codes for highperformance computing. IEEE Transactions on Parallel and Distributed Systems, 32(5): 1014–1029, May 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Dichev:2022:PLR

[DDN^{+}22] Kiril Dichev, Daniele De Sensi, Dimitrios S. Nikolopoulos, Kirk W. Cameron, and Ivor Spence. Power Log n Roll:

Power-efficient localized rollback for MPI applications using message logging protocols. IEEE Transactions on Parallel and Distributed Systems, 33 (6):1276–1288, June 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Dong:2024:FSM

[DDX^{+}24] Yong Dong, Yiqin Dai, Min Xie, Kai Lu, Ruibo Wang, Juan Chen, Mingtian Shao, and Zheng Wang. Faster and scalable MPI applications launching. IEEE Transactions on Parallel and Distributed Systems, 35(2):264–279, February 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

DeCapitanidiVimercati:2023:STE

[DFJ^{+}23] Sabrina De Capitani di Vimercati, Sara Foresti, Sushil Jajodia, Stefano Paraboschi, Pierangela Samarati, and Roberto Sassi. Sentinels and twins: Effective integrity assessment for distributed computation. IEEE Transactions on Parallel and Distributed Systems, 34(1):108–122, January 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Djigal:2021:IEA

[DFLG21] H. Djigal, J. Feng, J. Lu, and J. Ge. IPPTS: an efficient algorithm for scientific workflow scheduling in heterogeneous computing systems.

IEEE Transactions on Parallel and Distributed Systems, 32 (5):1057–1071, May 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Detti:2023:PBO

[DFP23] Andrea Detti, Ludovico Funari, and Luca Petrucci. Bench: an open-source factory of benchmark microservice applications. IEEE Transactions on Parallel and Distributed Systems, 34 (3):968–980, March 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Dong:2020:LCD

[DFXY20] E. Dong, X. Fu, M. Xu, and Y. Yang. Low-cost datacenter load balancing with multipath transport and top-of-rack switches. IEEE Transactions on Parallel and Distributed Systems, 31(10):2232– 2247, October 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

DePellegrini:2024:FCS

[DGE^{+}24] Francesco De Pellegrini, Vaibhav Kumar Gupta, Rachid El Azouzi, Serigne Gueye, Cedric Richier, and Jeremie Leguay. Fair coflow scheduling via controlled slowdown. IEEE Transactions on Parallel and Distributed Systems, 35(12):2347– 2360, December 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Dube:2022:CMM

[DHH^{+}22] Griffin Dube, Cavender Holt, John Hollowell, Sarah Placke, Sansriti Ranjan, Nikolas Heitzig, and Jon Calhoun. Critique of MemXCT: Memory-centric X-ray CT reconstruction with massive parallelization by SCC Team from Clemson University. IEEE Transactions on Parallel and Distributed Systems, 33(9): 2054–2057, September 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Devarajan:2025:HAH

[DHM25] Hariharan Devarajan, Gerd Heber, and Kathryn Mohror. H5Intent: Autotuning HDF5 with user intent. IEEE Transactions on Parallel and Distributed Systems, 36(2):108– 119, February 2025. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Duan:2021:SBF

[DLC^{+}21] M. Duan, D. Liu, X. Chen, R. Liu, Y. Tan, and L. Liang. Self-balancing federated learning with global imbalanced data in mobile systems. IEEE Transactions on Parallel and Distributed Systems, 32(1):59– 71, January 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Duan:2022:FCF

[DLJ^{+}22] Moming Duan, Duo Liu, Xinyuan Ji, Yu Wu, Liang Liang, Xianzhang Chen, Yujuan Tan, and Ao Ren. Flexible clustered federated learning for client-level data distribution shift. IEEE Transactions on Parallel and Distributed Systems, 33(11):2661– 2674, November 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ding:2022:PGT

[DLLL22] Yan Ding, Kenli Li, Chubo Liu, and Keqin Li. A potential game theoretic approach to computation offloading strategy optimization in end-edge– cloud computing. IEEE Transactions on Parallel and Distributed Systems, 33(6):1503– 1519, June 2022. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Donassolo:2022:ORI

[DLMF22] Bruno Donassolo, Arnaud Legrand, Panayotis Mertikopoulos, and Ilhem Fajjari. Online reconfiguration of IoT applications in the fog: The information-coordination tradeoff. IEEE Transactions on Parallel and Distributed Systems, 33(5):1156–1172, May 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Duan:2024:PSP

[DLX^{+}24] Jiangfei Duan, Xiuhong Li, Ping Xu, Xingcheng Zhang, Shengen Yan, Yun Liang, and Dahua Lin. Proteus: Simulating the performance of distributed DNN training. IEEE Transactions on Parallel and Distributed Systems, 35(10): 1867–1878, October 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Deng:2024:AMB

[DLY^{+}24] Quan Deng, Qiang Liu, Ming Yuan, Xiaohui Duan, Lin Gan, Jinzhe Yang, Wenlai Zhao, Zhenxiang Zhang, Guiming Wu, Wayne Luk, Haohuan Fu, and Guangwen Yang. Acceleration of multi-body molecular dynamics with customized parallel dataflow. IEEE Transactions on Parallel and Distributed Systems, 35(12):2297– 2314, December 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Dalmia:2023:ISG

[DMI^{+}23] Preyesh Dalmia, Rohan Mahapatra, Jeremy Intan, Dan Negrut, and Matthew D. Sinclair. Improving the scalability of GPU synchronization primitives. IEEE Transactions on Parallel and Distributed Systems, 34(1):275–290, January 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Du:2022:OCS

[DMPR22] Yishu Du, Loris Marchal, Guillaume Pallez, and Yves Robert. Optimal checkpointing strategies for iterative applications. IEEE Transactions on Parallel and Distributed Systems, 33 (3):507–522, March 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Dabiri:2020:REM

[DMST20] K. Dabiri, M. Malekmohammadi, A. Sheikholeslami, and H. Tamura. Replica exchange MCMC hardware with automatic temperature selection and parallel trial. IEEE Transactions on Parallel and Distributed Systems, 31(7):1681– 1692, July 2020. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Derafshi:2020:HHP

[DNKB20] D. Derafshi, A. Norollah, M. Khosroanjam, and H. Beitollahi. HRHS: a high-performance real-time hardware scheduler. IEEE Transactions on Parallel and Distributed Systems, 31 (4):897–908, April 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Dong:2025:SPR

[DNLX25] Xiaodong Dong, Lihai Nie, Zheli Liu, and Yang Xiang. Slark: a performance robust decentralized inter-datacenter deadline-aware coflows scheduling framework with local information. IEEE Transactions on Parallel and Distributed Systems, 36(2):197–211, February 2025. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Davis:2022:CMM

[DPGG22] Brock Davis, Juan Paez, Jack Gaither, and Joe A. Garcia. Critique of MemXCT: Memory-centric X-ray CT reconstruction with massive parallelization by SCC Team from the University of Texas at Austin. IEEE Transactions on Parallel and Distributed Systems, 33(9):2062– 2065, September 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ding:2021:MTE

[DQC^{+}21] T. Ding, S. Qian, J. Cao, G. Xue, Y. Zhu, J. Yu, and M. Li. MO-Tree: an efficient forwarding engine for spatiotemporal-aware Pub/Sub systems. IEEE Transactions on Parallel and Distributed Systems, 32(4):855–866, April 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Denninnart:2022:HPF

[DS22] Chavit Denninnart and Mohsen Amini Salehi. Harnessing the potential of function-reuse in multimedia cloud systems. IEEE Transactions on Parallel and Distributed Systems, 33(3):617– 629, March 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Devaraj:2023:CIE

[DS23] Rajesh Devaraj and Arnab Sarkar. Comments on IPPTS: an efficient algorithm for scientific workflow scheduling in heterogeneous computing systems. IEEE Transactions on Parallel and Distributed Systems, 34 (3):810–811, March 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Di:2021:EBO

[DSCL21] B. Di, J. Sun, H. Chen, and D. Li. Efficient buffer overflow detection on GPU. IEEE Transactions on Parallel and Distributed Systems, 32 (5):1161–1177, May 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Duan:2023:PBE

[DSW^{+}23] Xiaohui Duan, Qi Shao, Junben Weng, Bertil Schmidt, Lin Gan, Guohui Li, Haohuan Fu, Wei Xue, Weiguo Liu, and Guangwen Yang. Bio-ESMD: a data centric implementation for large-scale biological system simulation on Sunway Taihu-Light supercomputer. IEEE Transactions on Parallel and Distributed Systems, 34(3):881– 893, March 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Duan:2024:DES

[DSZ^{+}24] Qiqi Duan, Chang Shao, Guochen Zhou, Minghan Zhang, Qi Zhao, and Yuhui Shi. Distributed evolution strategies with multi-level learning for large-scale black-box optimization. IEEE Transactions on Parallel and Distributed Systems, 35(11):2087– 2101, November 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Dinh:2022:DDA

[DTN^{+}22] Canh T. Dinh, Nguyen H. Tran, Tuan Dung Nguyen, Wei Bao, Amir Rezaei Balef, Bing B. Zhou, and Albert Y. Zomaya. DONE: Distributed approximate Newton-type method for federated edge learning. IEEE Transactions on Parallel and Distributed Systems, 33(11): 2648–2660, November 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Dou:2024:DLC

[DWZ^{+}24] Hui Dou, Yilun Wang, Yiwen Zhang, Pengfei Chen, and Zibin Zheng. DeepCAT+: a lowcost and transferrable online configuration auto-tuning approach for big data frameworks. IEEE Transactions on Parallel and Distributed Systems, 35(11):2114–2131, November 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Dong:2021:TBS

[DYFL21] Zheng Dong, Kecheng Yang, Nathan Fisher, and Cong Liu. Tardiness bounds for sporadic gang tasks under preemptive global EDF scheduling. IEEE Transactions on Parallel and Distributed Systems, 32(12):

2867–2879, December 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Deng:2021:BLE

[DZL^{+}21] S. Deng, C. Zhang, C. Li, J. Yin, S. Dustdar, and A. Y. Zomaya. Burst load evacuation based on dispatching and scheduling in distributed edge networks. IEEE Transactions on Parallel and Distributed Systems, 32(8):1918– 1932, August 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Du:2021:MPO

[DZS^{+}21] J. Du, X. Zhu, M. Shen, Y. Du, Y. Lu, N. Xiao, and X. Liao. Model parallelism optimization for distributed inference via decoupled CNN structure. IEEE Transactions on Parallel and Distributed Systems, 32(7):1665–1676, July 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Elahi:2024:PFS

[EFME24] Fatemeh Elahi, Mahmood Fazlali, Hadi Tabatabaee Malazi, and Mehdi Elahi. Parallel fractional stochastic gradient descent with adaptive learning for recommender systems. IEEE Transactions on Parallel and Distributed Systems, 35 (3):470–483, March 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Faingnaert:2022:FPG

[FBD22] Thomas Faingnaert, Tim Besard, and Bjorn De Sutter. Flexible performant GEMM kernels on GPUs. IEEE Transactions on Parallel and Distributed Systems, 33(9):2230– 2248, September 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Fresa:2023:OAM

[FC23] Andrea Fresa and Jaya Prakash Champati. Offloading algorithms for maximizing inference accuracy on edge device in an edge intelligence system. IEEE Transactions on Parallel and Distributed Systems, 34 (7):2025–2039, July 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Fan:2022:ABN

[FFQ^{+}22] Hongxiang Fan, Martin Ferianc, Zhiqiang Que, Xinyu Niu, Miguel Rodrigues, and Wayne Luk. Accelerating Bayesian neural networks via algorithmic and hardware optimizations. IEEE Transactions on Parallel and Distributed Systems, 33(12):3387–3399, December 2022. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Fan:2022:CMM

[FGH^{+}22] Zejia Fan, Yuchen Gu, Zhewen Hao, Yueyang Pan, Pengcheng Xu, Yuxuan Yan, Fangyuan Yang, Zhenxin Fu, and Yun Liang. Critique of MemXCT: Memory-centric X-ray CT reconstruction with massive parallelization by SCC Team from Peking University. IEEE Transactions on Parallel and Distributed Systems, 33(9): 2032–2034, September 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Feng:2022:MMC

[FLPL22] Jie Feng, Lei Liu, Qingqi Pei, and Keqin Li. Min-max cost optimization for efficient hierarchical federated learning in wireless edge networks. IEEE Transactions on Parallel and Distributed Systems, 33(11): 2687–2700, November 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Fan:2023:ACN

[FLW^{+}23] Zhihua Fan, Wenming Li, Zhen Wang, Tianyu Liu, Haibin Wu, Yanhuan Liu, Meng Wu, Xinxin Wu, Xiaochun Ye, Dongrui Fan, Ninghui Sun, and Xuejun An. Accelerating convolutional neural networks by exploiting the sparsity of output activation. IEEE Transactions on Parallel and Distributed Systems, 34(12):3253– 3265, December 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Frey:2023:DCW

[FMP^{+}23] Davide Frey, Achour Mostefaoui, Matthieu Perrin, Pierre-Louis Roman, and François Taïani. Differentiated consistency for worldwide gossips. IEEE Transactions on Parallel and Distributed Systems, 34 (1):1–15, January 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Faragardi:2020:GHB

[FSF^{+}20] H. R. Faragardi, M. R. Saleh Sedghpour, S. Fazliahmadi, T. Fahringer, and N. Rasouli. GRP-HEFT: a budgetconstrained resource provisioning scheme for workflow scheduling in IaaS clouds. IEEE Transactions on Parallel and Distributed Systems, 31 (6):1239–1254, June 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Feliu:2020:TII

[FSPE20] J. Feliu, J. Sahuquillo, S. Petit, and L. Eeckhout. Thread isolation to improve symbiotic scheduling on SMT multicore processors. IEEE Transactions on Parallel and Distributed Systems, 31(2):359–373, February 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Fu:2020:OLA

[FTYL20] Z. Fu, Z. Tang, L. Yang, and C. Liu. An optimal localityaware task scheduling algorithm based on bipartite graph modelling for Spark applications. IEEE Transactions on Parallel and Distributed Systems, 31(10):2406–2420, October 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Fu:2022:PFS

[FWCB22] Junsong Fu, Na Wang, Baojiang Cui, and Bharat K. Bhargava. A practical framework for secure document retrieval in encrypted cloud file systems. IEEE Transactions on Parallel and Distributed Systems, 33 (5):1246–1261, May 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Fan:2024:TAM

[FWT^{+}24] Jiamin Fan, Kui Wu, Guoming Tang, Yang Zhou, and Shengqiang Huang. Taking advantage of the mistakes: Rethinking clustered federated learning for IoT anomaly detection. IEEE Transactions on Parallel and Distributed Systems, 35(6):707–721, June 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Fan:2023:NEA

[FXL^{+}23] Weibei Fan, Fu Xiao, Mengjie Lv, Lei Han, Junchang Wang, and Xin He. Node essentiality assessment and distributed collaborative virtual network embedding in datacenters. IEEE Transactions on Parallel and Distributed Systems, 34(4): 1265–1280, April 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Fu:2022:ARE

[FZC^{+}22] Kaihua Fu, Wei Zhang, Quan Chen, Deze Zeng, and Minyi Guo. Adaptive resource efficient microservice deployment in cloud-edge continuum. IEEE Transactions on Parallel and Distributed Systems, 33(8): 1825–1840, August 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Fang:2024:PPV

[FZD^{+}24] Yaozheng Fang, Zhiyuan Zhou, Surong Dai, Jinni Yang, Hui Zhang, and Ye Lu. PaVM: a parallel virtual machine for smart contract execution and validation. IEEE Transactions on Parallel and Distributed Systems, 35(1):186–202, January 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gou:2020:PTS

[GBM20] C. Gou, A. Benoit, and L. Marchal. Partitioning tree-shaped task graphs for distributed platforms with limited memory. IEEE Transactions on Parallel and Distributed Systems, 31 (7):1533–1544, July 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gong:2021:TLS

[GCL^{+}21] X. Gong, D. Cao, Y. Li, X. Liu, Y. Li, J. Zhang, and T. Li. A thread level SLOaware I/O framework for embedded virtualization. IEEE Transactions on Parallel and Distributed Systems, 32(3):500– 513, March 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gu:2022:LIR

[GCL^{+}22] Rong Gu, Yuquan Chen, Shuai Liu, Haipeng Dai, Guihai Chen, Kai Zhang, Yang Che, and Yihua Huang. Liquid: Intelligent resource estimation and network-efficient scheduling for deep learning jobs on distributed GPU clusters. IEEE Transactions on Parallel and Distributed Systems, 33(11): 2808–2820, November 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gao:2022:ORF

[GDS^{+}22] Ping Gao, Xiaohui Duan, Bertil Schmidt, Wusheng Zhang, Lin Gan, Haohuan Fu, Wei Xue, Weiguo Liu, and Guangwen Yang. Optimization of reactive force field simulation: Refactor, parallelization, and vectorization for interactions. IEEE Transactions on Parallel and Distributed Systems, 33(2):359– 373, February 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gao:2023:RAA

[GDS^{+}23] Ping Gao, Xiaohui Duan, Bertil Schmidt, Wubing Wan, Jiaxu Guo, Wusheng Zhang, Lin Gan, Haohuan Fu, Wei Xue, Weiguo Liu, and Guangwen Yang. Redesign and accelerate the AIREBO bond-order potential on the new Sunway supercomputer. IEEE Transactions on Parallel and Distributed Systems, 34(12):3117– 3132, December 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gao:2020:MSB

[GDZ^{+}20] P. Gao, X. Duan, T. Zhang, M. Zhang, B. Schmidt, X. Zhang, H. Sun, W. Zhang, L. Gan, W. Xue, H. Fu, W. Liu, and G. Yang. Millimeter-scale and billion-atom reactive force field simulation on Sunway Taihulight. IEEE Transactions on Parallel and Distributed Systems, 31(12):2954–2967, December 2020. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gajaria:2024:SRB

[GGA24] Dhruv Gajaria, Kevin Antony Gomez, and Tosiron Adegbija. STT-RAM-based hierarchical in-memory computing. IEEE Transactions on Parallel and Distributed Systems, 35(9): 1615–1629, September 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gainaru:2021:PUH

[GGHP21] A. Gainaru, B. Goglin, V. Honoré, and G. Pallez. Profiles of upcoming HPC applications and their impact on reservation strategies. IEEE Transactions on Parallel and Distributed Systems, 32(5):1178– 1190, May 2021. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gupta:2023:CPF

[GGL^{+}23] Arunav Gupta, John Ge, John Li, Zihao Kong, Kaiwen He, Matthew Mikhailov, Bryan Chin, Xiaochen Li, Max Apodaca, Paul Rodriguez, Mahidar Tatineni, Mary Thomas, and Santosh Bhatt. Critique of A Parallel Framework for Constraint-Based Bayesian Network Learning via Markov Blanket Discovery by SCC Team From UC San Diego. IEEE Transactions on Parallel and Distributed Systems, 34 (6):1727–1730, June 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). See [SCA23].

Grete:2021:KAP

[GGO21] P. Grete, F. W. Glines, and B. W. O’Shea. K-Athena: a performance portable structured grid finite volume magnetohydrodynamics code. IEEE Transactions on Parallel and Distributed Systems, 32(1):85– 97, January 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gao:2020:MES

[GGZ^{+}20] T. Gao, Y. Guo, B. Zhang, P. Cicotti, Y. Lu, P. Balaji, and M. Taufer. Memory-efficient and skew-tolerant MapReduce over MPI for supercomputing systems. IEEE Transactions on Parallel and Distributed Systems, 31(12):2734–2748, December 2020. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Guo:2020:FTR

[GHG^{+}20] P. Guo, W. Hou, L. Guo, W. Sun, C. Liu, H. Bao, L. H. K. Duong, and W. Liu. Fault-tolerant routing mechanism in 3D optical networkon-chip based on node reuse. IEEE Transactions on Parallel and Distributed Systems, 31 (3):547–564, March 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gao:2024:SUR

[GHGW24] Teng Gao, Lan Huang, Shang Gao, and Kangping Wang. SSA: a uniformly recursive bidirection-sequence systolic sorter array. IEEE Transactions on Parallel and Distributed Systems, 35(10):1721– 1734, October 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gao:2024:UBE

[GHM^{+}24] Yunqi Gao, Bing Hu, Mahdi Boloursaz Mashhadi, A-Long Jin, Pei Xiao, and Chunming Wu. US-Byte: an efficient communication framework for scheduling unequal-sized tensor blocks in distributed deep learning. IEEE Transactions on Parallel and Distributed Systems, 35(1): 123–139, January 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gupta:2021:PAU

[GJCC21] N. Gupta, A. Jati, A. K. Chauhan, and A. Chattopadhyay. PQC acceleration using GPUs: FrodoKEM, NewHope, and Kyber. IEEE Transactions on Parallel and Distributed Systems, 32(3):575–586, March 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gokturk:2021:BPI

[GK21] G. Göktürk and K. Kaya. Boosting parallel influencemaximization kernels for undirected networks with fusing and vectorization. IEEE Transactions on Parallel and Distributed Systems, 32(5):1001– 1013, May 2021. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ghosh:2021:PSA

[GKK21] P. Ghosh, S. Krishnamoorthy, and A. Kalyanaraman. PaKman: a scalable algorithm for generating genomic contigs on distributed memory machines. IEEE Transactions on Parallel and Distributed Systems, 32 (5):1191–1209, May 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gaveau:2020:PAT

[GLA20] J. Gaveau, C. J. Le Martret, and M. Assaad. Performance analysis of trial and error algorithms. IEEE Transactions on Parallel and Distributed Systems, 31(6):1343–1356, June 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ge:2021:HFC

[GLF^{+}21] C. Ge, Z. Liu, L. Fang, H. Ling, A. Zhang, and C. Yin. A hybrid fuzzy convolutional neural network based mechanism for photovoltaic cell defect detection with electroluminescence images. IEEE Transactions on Parallel and Distributed Systems, 32(7):1653–1664, July 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Geng:2020:LDS

[GLL^{+}20] Y. Geng, Q. Li, M. Liang, C. Chi, J. Tan, and H. Huang. Local-density subspace distributed clustering for highdimensional data. IEEE Transactions on Parallel and Distributed Systems, 31(8):1799– 1814, August 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gong:2021:SLM

[GLL^{+}21] L. Gong, H. Lin, Z. Li, F. Qian, Y. Li, X. Ma, and Y. Liu. Systematically landing machine learning onto marketscale mobile malware detection. IEEE Transactions on Parallel and Distributed Systems, 32 (7):1615–1628, July 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gaihre:2022:GSS

[GLL22] Anil Gaihre, Xiaoye Sherry Li, and Hang Liu. gSoFa: Scalable sparse symbolic LU factorization on GPUs. IEEE Transactions on Parallel and Distributed Systems, 33(4):1015– 1026, April 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gill:2021:HPC

[GLP^{+}21] A. Gill, M. Lalith, S. Poledna, M. Hori, K. Fujita, and T. Ichimura. High-performance computing implementations of agent-based economic models for realizing 1:1 scale simulations of large economies. IEEE Transactions on Parallel and Distributed Systems, 32(8): 2101–2114, August 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Geng:2021:ORO

[GLW^{+}21] T. Geng, A. Li, T. Wang, C. Wu, Y. Li, R. Shi, W. Wu, and M. Herbordt. O3BNN-R: an out-of-order architecture for high-performance and regularized BNN inference. IEEE Transactions on Parallel and Distributed Systems, 32(1):199– 213, January 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gong:2022:ESB

[GLX^{+}22] Cheng Gong, Ye Lu, Kunpeng Xie, Zongming Jin, Tao Li, and Yanzhi Wang. Elastic significant bit quantization and acceleration for deep neural networks. IEEE Transactions on Parallel and Distributed Systems, 33(11):3178– 3193, November 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gonzlez:2021:MGP

[GM21] M. González and E. Morancho. Multi-GPU parallelization of the NAS multi-zone parallel benchmarks. IEEE Transactions on Parallel and Distributed Systems, 32(1):229– 241, January 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gan:2022:FFD

[GMI^{+}22] Shaoduo Gan, Akhil Mathur, Anton Isopoussu, Fahim Kawsar, Nadia Berthouze, and Nicholas D. Lane. FRuDA: Framework for distributed adversarial domain adaptation. IEEE Transactions on Parallel and Distributed Systems, 33(11):3153– 3164, November 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Giersch:2022:FPC

[GN22] Oliver Giersch and Jörg Nolte.

Fast and portable concurrent FIFO queues with deterministic memory reclamation. IEEE Transactions on Parallel and Distributed Systems, 33(3):604– 616, March 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Galindo:2021:RCC

[GNST21] C. Galindo, N. Nishida, J. Silva, and S. Tamarit. Reversible CSP computations. IEEE Transactions on Parallel and Distributed Systems, 32 (6):1425–1436, June 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Guo:2021:GEA

[GSH^{+}21] Yijin Guo, Huasong Shan, Shixin Huang, Kai Hwang, Jianping Fan, and Zhibin Yu. GML: Efficiently autotuning Flink’s configurations via guided machine learning. IEEE Transactions on Parallel and Distributed Systems, 32(12):2921–2935, December 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gao:2020:BPS

[GSL^{+}20] C. Gao, L. Shi, K. Liu, C. J. Xue, J. Yang, and Y. Zhang. Boosting the performance of SSDs via fully exploiting the plane level parallelism. IEEE Transactions on Parallel and Distributed Systems, 31(9): 2185–2200, September 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ghosh:2022:CPG

[GTH22] Sayan Ghosh, Nathan R. Tallent, and Mahantesh Halappanavar. Characterizing performance of graph neighborhood communication patterns. IEEE Transactions on Parallel and Distributed Systems, 33(4):915– 928, April 2022. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gounley:2022:PPM

[GVD^{+}22] John Gounley, Madhurima Vardhan, Erik W. Draeger, Pedro Valero-Lara, Shirley V. Moore, and Amanda Randles. Propagation pattern for moment representation of the lattice Boltzmann method. IEEE Transactions on Parallel and Distributed Systems, 33(3):642– 653, March 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gerlinghoff:2022:EEE

[GWG^{+}22] Daniel Gerlinghoff, Zhehui Wang, Xiaozhe Gu, Rick Siow Mong Goh, and Tao Luo. E3NE: an end-to-end framework for accelerating spiking neural networks with emerging neural encoding on FPGAs. IEEE Transactions on Parallel and Distributed Systems, 33(11):3207–3219, November 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Guo:2022:LES

[GWLX22] Jialin Guo, Jie Wu, Anfeng Liu, and Neal N. Xiong. LightFed: an efficient and secure federated edge learning system on model splitting. IEEE Transactions on Parallel and Distributed Systems, 33(11):2701– 2713, November 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gong:2021:IHS

[GWLZ21] L. Gong, C. Wang, X. Li, and X. Zhou. Improving HW/SW adaptability for accelerating CNNs on FPGAs through a dynamic/static co-reconfiguration approach. IEEE Transactions on Parallel and Distributed Systems, 32(7):1854–1865, July 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gu:2023:HLD

[GXC^{+}23] Rong Gu, Zhihao Xu, Yang Che, Xu Wang, Haipeng Dai, Kai Zhang, Bin Fan, Haojun Hou, Li Yi, Yu Ding, Yihua Huang, and Guihai Chen. High-level data abstraction and elastic data caching for dataintensive AI applications on cloud-native platforms. IEEE Transactions on Parallel and Distributed Systems, 34(11): 2946–2964, November 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gao:2020:TLL

[GXW^{+}20] L. Gao, Y. Xu, R. Wang, Z. Luan, Z. Yu, and D. Qian. Thread-level locking for SIMT architectures. IEEE Transactions on Parallel and Distributed Systems, 31(5):1121– 1136, May 2020. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gao:2022:CEG

[GXW22] Yiwen Gao, Jia Xu, and Hongbing Wang. cuNH: Efficient GPU implementations of post-quantum KEM NewHope. IEEE Transactions on Parallel and Distributed Systems, 33 (3):551–568, March 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gu:2021:TEL

[GZJ^{+}21] R. Gu, Z. Zuo, X. Jiang, H. Yin, Z. Wang, L. Wang, X. Li, and Y. Huang. Towards efficient large-scale interprocedural program static analysis on distributed dataparallel computation. IEEE Transactions on Parallel and Distributed Systems, 32(4):867– 883, April 2021. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gu:2025:GCE

[GZS^{+}25] Diandian Gu, Yihao Zhao, Peng Sun, Xin Jin, and Xuanzhe Liu. GreenFlow: a carbon-efficient scheduler for deep learning workloads. IEEE Transactions on Parallel and Distributed Systems, 36(2):168– 184, February 2025. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gan:2022:TCG

[GZW^{+}22] Xinbiao Gan, Yiming Zhang, Ruibo Wang, Tiejun Li, Tiaojie Xiao, Ruigeng Zeng, Jie Liu, and Kai Lu. TianheGraph: Customizing graph search for Graph500 on Tianhe supercomputer. IEEE Transactions on Parallel and Distributed Systems, 33(4):941– 951, April 2022. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gong:2021:ALS

[GZY21] S. Gong, Y. Zhang, and G. Yu. Accelerating large-scale prioritized graph computations by hotness balanced partition. IEEE Transactions on Parallel and Distributed Systems, 32 (4):746–759, April 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hashemi:2022:BMG

[HAD^{+}22] Abolfazl Hashemi, Anish Acharya, Rudrajit Das, Haris Vikalo, Sujay Sanghavi, and Inderjit Dhillon. On the benefits of multiple gossip steps in communication-constrained decentralized federated learning. IEEE Transactions on Parallel and Distributed Systems, 33(11):2727–2739, November 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

lu:2022:MDO

[HBG^{+}22] Mert Hidayetğlu, Tekin Biçer, Simon Garcia de Gonzalo, Bin Ren, Doğa Gürsoy, Rajkumar Kettimuthu, Ian T. Foster, and Wen-Mei W. Hwu. MemXCT: Design, optimization, scaling, and reproducibility of X-ray tomography imaging. IEEE Transactions on Parallel and Distributed Systems, 33(9): 2014–2031, September 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Huang:2024:EDS

[HC24] Hua Huang and Edmond Chow.

Exploring the design space of distributed parallel sparse matrix-multiple vector multiplication. IEEE Transactions on Parallel and Distributed Systems, 35(11):1977– 1988, November 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Huang:2024:EFD

[HCA24] Darong Huang, Luis Costero, and David Atienza. An evaluation framework for dynamic thermal management strategies in 3D MultiProcessor systemon-chip co-design. IEEE Transactions on Parallel and Distributed Systems, 35(11):2161– 2176, November 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Han:2023:MSN

[HCG^{+}23] Rongxin Han, Dezhi Chen, Song Guo, Jingyu Wang, Qi Qi, Lu Lu, and Jianxin Liao. Multi-SP network slicing parallel relieving edge network conflict. IEEE Transactions on Parallel and Distributed Systems, 34(11):2860–2875, November 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hadizadeh:2022:CCP

[HCR^{+}22] Mostafa Hadizadeh, Elham Cheshmikhani, Maysam Rahmanpour, Onur Mutlu, and Hossein Asadi. CoPA: Cold page awakening to overcome retention failures in STT-MRAM based I/O buffers. IEEE Transactions on Parallel and Distributed Systems, 33(10):2304– 2317, October 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

He:2020:GTA

[HCZ^{+}20] Q. He, G. Cui, X. Zhang, F. Chen, S. Deng, H. Jin, Y. Li, and Y. Yang. A game-theoretical approach for user allocation in edge computing environment. IEEE Transactions on Parallel and Distributed Systems, 31(3):515– 529, March 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hao:2023:SSM

[HFC^{+}23] Xiaoyu Hao, Tao Fang, Junshi Chen, Jun Gu, Jiawang Feng, Hong An, and Chun Zhao. swMPAS-A: Scaling MPAS-A to 39 million heterogeneous cores on the new generation Sunway supercomputer. IEEE Transactions on Parallel and Distributed Systems, 34(1): 141–153, January 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Huang:2021:DER

[HFW^{+}21] W. Huang, J. Fang, S. Wan, C. Xie, and X. He. Design and evaluation of a risk-aware failure identification scheme for improved RAS in erasure-coded data centers. IEEE Transactions on Parallel and Distributed Systems, 32(1):16–30, January 2021. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hesham:2020:HND

[HGA20] S. Hesham, D. Goehringer, and M. A. Abd El Ghany. HPPT-NoC: a dark-silicon inspired hierarchical TDM NoC with efficient power-performance trading. IEEE Transactions on Parallel and Distributed Systems, 31(3):675–694, March 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Han:2025:DDG

[HHJ^{+}25] Jialiang Han, Yudong Han, Xiang Jing, Gang Huang, and Yun Ma. DegaFL: Decentralized gradient aggregation for cross-silo federated learning. IEEE Transactions on Parallel and Distributed Systems, 36(2): 212–225, February 2025. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hernandez-Juarez:2021:PSS

[HJEV^{+}21] D. Hernandez-Juarez, A. Espinosa, D. Vazquez, A. M. Lopez, and J. C. Moure. 3D perception with slanted stixels on GPU. IEEE Transactions on Parallel and Distributed Systems, 32(10):2434– 2447, October 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hamandawana:2020:CEC

[HKL^{+}20] P. Hamandawana, A. Khan, C. Lee, S. Park, and Y. Kim. Crocus: Enabling computing resource orchestration for inline cluster-wide deduplication on scalable storage systems. IEEE Transactions on Parallel and Distributed Systems, 31(8): 1740–1753, August 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Huang:2023:MUM

[HLB^{+}23] Yizhi Huang, Yan Liu, Yang Bai, Si Chen, and Renfa Li. UMA-MF: a unified multi-CPU/GPU asynchronous computing framework for SGDbased matrix factorization. IEEE Transactions on Parallel and Distributed Systems, 34(11):2978–2993, November 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Huang:2022:FDI

[HLH22] Yanze Huang, Limei Lin, and Sun-Yuan Hsieh. A fast f(r, k + 1)/k-diagnosis for interconnection networks under MM* model. IEEE Transactions on Parallel and Distributed Systems, 33(7):1593– 1604, July 2022. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

He:2022:CES

[HLL22] Zhenli He, Kenli Li, and Keqin Li. Cost-efficient server configuration and placement for mobile edge computing. IEEE Transactions on Parallel and Distributed Systems, 33(9):2198– 2212, September 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Huang:2022:TLP

[HLLL22] Tsung-Wei Huang, Dian-Lun Lin, Chun-Xun Lin, and Yibo Lin. Taskflow: a lightweight parallel and heterogeneous task graph computing system. IEEE Transactions on Parallel and Distributed Systems, 33(6): 1303–1320, June 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hudson:2022:LLC

[HLNW22] Stephen Hudson, Jeffrey Larson, John-Luke Navarro, and Stefan M. Wild. libEnsemble: a library to coordinate the concurrent evaluation of dynamic ensembles of calculations. IEEE Transactions on Parallel and Distributed Systems, 33(4):977–988, April 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Han:2021:ADP

[HLO^{+}21] R. Han, D. Li, J. Ouyang, C. H. Liu, G. Wang, D. Wu, and L. Y. Chen. Accurate differentially private deep learning on the edge. IEEE Transactions on Parallel and Distributed Systems, 32(9):2231– 2247, September 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Huang:2023:PPS

[HLS^{+}23] Hongjing Huang, Yingtao Li, Jie Sun, Xueying Zhu, Jie Zhang, Liang Luo, Jialin Li, and Zeke Wang. P4SGD: Programmable switch enhanced model-parallel training on generalized linear models on distributed FPGAs. IEEE Transactions on Parallel and Distributed Systems, 34(8):2311– 2324, August 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Herschlag:2021:AGD

[HLVR21] G. Herschlag, S. Lee, J. S. Vetter, and A. Randles. Analysis of GPU data access patterns on complex geometries for the D3Q19 Lattice Boltzmann Algorithm. IEEE Transactions on Parallel and Distributed Systems, 32(10):2400–2414, October 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Han:2020:ELE

[HLW^{+}20] J. Han, H. Liu, M. Wang, Z. Li, and Y. Zhang. ERA-LSTM: an efficient ReRAM-based architecture for long short-term memory. IEEE Transactions on Parallel and Distributed Systems, 31(6):1328–1342, June 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Han:2021:AGB

[HLW^{+}21a] R. Han, S. Li, X. Wang, C. H. Liu, G. Xin, and L. Y. Chen. Accelerating gossipbased deep learning in heterogeneous edge computing platforms. IEEE Transactions on Parallel and Distributed Systems, 32(7):1591–1602, July 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Huang:2021:EBC

[HLW^{+}21b] T. Huang, W. Lin, W. Wu, L. He, K. Li, and A. Y. Zomaya. An efficiency-boosting client selection scheme for federated learning with fairness guarantee. IEEE Transactions on Parallel and Distributed Systems, 32(7):1552–1564, July 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

He:2020:HHA

[HLZ^{+}20] S. He, Z. Li, J. Zhou, Y. Yin, X. Xu, Y. Chen, and X. Sun. A holistic heterogeneity-aware data placement scheme for hybrid parallel I/O systems. IEEE Transactions on Parallel and Distributed Systems, 31 (4):830–842, April 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hu:2021:ORA

[HLZ^{+}21] Z. Hu, D. Li, D. Zhang, Y. Zhang, and B. Peng. Optimizing resource allocation for data-parallel jobs via GCNbased prediction. IEEE Transactions on Parallel and Distributed Systems, 32(9):2188– 2201, September 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Harrell:2022:AAR

[HMM22] Stephen Lien Harrell, Scott Michael, and Carlos Maltzahn. Advancing adoption of reproducibility in HPC: a preface to the special section. IEEE Transactions on Parallel and Distributed Systems, 33(9): 2011–2013, September 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hosseinalipour:2020:PAA

[HND20] S. Hosseinalipour, A. Nayak, and H. Dai. Power-aware allocation of graph jobs in geo-distributed cloud networks.

IEEE Transactions on Parallel and Distributed Systems, 31 (4):749–765, April 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hassanzadeh-Nazarabadi:2020:DUL

[HNKO20] Y. Hassanzadeh-Nazarabadi, A. Küpçü, and O. Ozkasap. Decentralized utility- and localityaware replication for heterogeneous DHT-based P2P cloud storage systems. IEEE Transactions on Parallel and Distributed Systems, 31(5):1183– 1193, May 2020. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

He:2023:PPS

[HO23] Fujun He and Eiji Oki. Preventive priority setting against multiple controller failures in software defined networks. IEEE Transactions on Parallel and Distributed Systems, 34(8): 2352–2364, August 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Han:2021:DIC

[HPB21] M. Han, J. Park, and W. Baek.

Design and implementation of a criticality- and heterogeneityaware runtime system for taskparallel applications. IEEE Transactions on Parallel and Distributed Systems, 32(5): 1117–1132, May 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hua:2022:LOL

[HSH^{+}22] Yusheng Hua, Xuanhua Shi, Kang He, Hai Jin, Wei Xie, Ligang He, and Yong Chen. LoomIO: Object-level coordination in distributed file systems. IEEE Transactions on Parallel and Distributed Systems, 33(8):1799–1810, August 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hua:2020:FPC

[HSY^{+}20] Q. Hua, Y. Shi, D. Yu, H. Jin, J. Yu, Z. Cai, X. Cheng, and H. Chen. Faster parallel core maintenance algorithms in dynamic graphs. IEEE Transactions on Parallel and Distributed Systems, 31(6):1287– 1300, June 2020. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

He:2022:CCA

[HTB22] TianZhang He, Adel N. Toosi, and Rajkumar Buyya. CAMIG: Concurrency-aware live migration management of multiple virtual machines in SDNenabled clouds. IEEE Transactions on Parallel and Distributed Systems, 33(10):2318– 2331, October 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hou:2024:WEP

[HTL^{+}24] Xiaofeng Hou, Xuehan Tang, Jiacheng Liu, Chao Li, Luhong Liang, and Kwang-Ting Cheng.

WASP: Efficient power management enabling workloadaware, self-powered AIoT devices. IEEE Transactions on Parallel and Distributed Systems, 35(8):1400–1414, August 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ho:2022:TAG

[HW22] Nhut-Minh Ho and Weng-Fai Wong. Tensorox: Accelerating GPU applications via neural approximation on unused tensor cores. IEEE Transactions on Parallel and Distributed Systems, 33(2):429–443, February 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Huang:2022:TSL

[HWF^{+}22] Jianqiang Huang, Haojie Wang, Xiang Fei, Xiaoying Wang, and Wenguang Chen. TC-Stream: Large-scale graph triangle counting on a single machine using GPUs. IEEE Transactions on Parallel and Distributed Systems, 33(11): 3067–3078, November 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Huang:2024:CFL

[HWS^{+}24] Guangjing Huang, Qiong Wu, Peng Sun, Qian Ma, and Xu Chen. Collaboration in federated learning with differential privacy: a Stackelberg game analysis. IEEE Transactions on Parallel and Distributed Systems, 35(3):455–469, March 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hu:2023:JDR

[HWW^{+}23] Yi Hu, Hao Wang, Liangyuan Wang, Menglan Hu, Kai Peng, and Bharadwaj Veeravalli. Joint deployment and request routing for microservice call graphs in data centers. IEEE Transactions on Parallel and Distributed Systems, 34(11): 2994–3011, November 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hu:2020:HEO

[HXW^{+}20] M. Hu, Z. Xie, D. Wu, Y. Zhou, X. Chen, and L. Xiao. Heterogeneous edge offloading with incomplete information: a minority game approach. IEEE Transactions on Parallel and Distributed Systems, 31(9): 2139–2154, September 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hu:2020:MSS

[HYL^{+}20] Y. Hu, H. Yang, Z. Luan, L. Gan, G. Yang, and D. Qian. Massively scaling seismic processing on Sunway TaihuLight supercomputer. IEEE Transactions on Parallel and Distributed Systems, 31(5):1194– 1208, May 2020. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

He:2023:LDR

[HYL^{+}23] Nan He, Song Yang, Fan Li, Stojan Trajanovski, Liehuang Zhu, Yu Wang, and Xiaoming Fu. Leveraging deep reinforcement learning with attention mechanism for virtual network function placement and routing. IEEE Transactions on Parallel and Distributed Systems, 34(4):1186–1201, April 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hou:2024:GLB

[HYL^{+}24] Biao Hou, Song Yang, Fan Li, Liehuang Zhu, Lei Jiao, Xu Chen, and Xiaoming Fu. Gamora: Learning-based buffer-aware preloading for adaptive short video streaming. IEEE Transactions on Parallel and Distributed Systems, 35(11):2132–2146, November 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Huang:2022:ERA

[HYP^{+}22] Huawei Huang, Zhengyu Yue, Xiaowen Peng, Liuding He, Wuhui Chen, Hong-Ning Dai, Zibin Zheng, and Song Guo. Elastic resource allocation against imbalanced transaction assignments in sharding-based permissioned blockchains. IEEE Transactions on Parallel and Distributed Systems, 33(10):

2372–2385, October 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Han:2024:PIC

[HZB^{+}24] Runzhou Han, Mai Zheng, Suren Byna, Houjun Tang, Bin Dong, Dong Dai, Yong Chen, Dongkyun Kim, Joseph Hassoun, and David Thorsley. PROV-IO^{+}: a cross-platform provenance framework for scientific data on HPC systems. IEEE Transactions on Parallel and Distributed Systems, 35 (5):844–861, May 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hua:2023:RCM

[HZJH23] Qiang-Sheng Hua, Xiaohui Zhang, Hai Jin, and Hong Huang. Revisiting core maintenance for dynamic hypergraphs. IEEE Transactions on Parallel and Distributed Systems, 34(3):981–994, March 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hua:2024:RSR

[HZK^{+}24] Yifan Hua, Shengan Zheng, Weihan Kong, Cong Zhou, Kaixin Huang, Ruoyan Ma, and Linpeng Huang. RADAR: a skew-resistant and hotnessaware ordered index design for processing-in-memory systems. IEEE Transactions on Parallel and Distributed Systems, 35(9): 1598–1614, September 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hao:2021:FGP

[HZW^{+}21] M. Hao, W. Zhang, Y. Wang, G. Lu, F. Wang, and A. V. Vasilakos. Fine-grained powercap allocation for powerconstrained systems based on multi-objective machine learning. IEEE Transactions on Parallel and Distributed Systems, 32(7):1789–1801, July 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

He:2025:EFT

[HZW^{+}25] Qiang He, Guobiao Zhang, Jiawei Wang, Ruikun Luo, Xiaohai Dai, Yuchong Hu, Feifei Chen, Hai Jin, and Yun Yang. EdgeHydra: Faulttolerant edge data distribution based on erasure coding. IEEE Transactions on Parallel and Distributed Systems, 36 (1):29–42, January 2025. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hu:2023:EER

[HZX^{+}23] Yihua Hu, Feng Zhang, Yifei Xia, Zhiming Yao, Letian Zeng, Haipeng Ding, Zhewei Wei, Xiao Zhang, Jidong Zhai, Xiaoyong Du, and Siqi Ma. Enabling efficient random access to hierarchically compressed text data on diverse GPU platforms. IEEE Transactions on Parallel and Distributed Systems, 34(10):2699–2717, October 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Han:2024:IIS

[HZX^{+}24] Ziyi Han, Ruiting Zhou, Chengzhong Xu, Yifan Zeng, and Renli Zhang. InSS: an intelligent scheduling orchestrator for multi-GPU inference with spatio-temporal sharing. IEEE Transactions on Parallel and Distributed Systems, 35 (10):1735–1748, October 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Iwasaki:2020:APT

[IATB20] S. Iwasaki, A. Amer, K. Taura, and P. Balaji. Analyzing the performance trade-off in implementing user-level threads. IEEE Transactions on Parallel and Distributed Systems, 31(8): 1859–1877, August 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ilager:2021:TPE

[IRB21] S. Ilager, K. Ramamohanarao, and R. Buyya. Thermal prediction for efficient energy management of clouds using machine learning. IEEE Transactions on Parallel and Distributed Systems, 32(5):1044– 1056, May 2021. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Iliakis:2022:RGM

[IXS22] Konstantinos Iliakis, Sotirios Xydis, and Dimitrios Soudris. Repurposing GPU microarchitectures with light-weight outof-order execution. IEEE Transactions on Parallel and Distributed Systems, 33(2):388– 402, February 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ishiguro:2023:RVA

[IYAK23] Kenta Ishiguro, Naoki Yasuno, Pierre-Louis Aublin, and Kenji Kono. Revisiting VM-agnostic KVM vCPU scheduler for mitigating excessive vCPU spinning. IEEE Transactions on Parallel and Distributed Systems, 34(10):2615–2628, October 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jannes:2023:PBR

[JBLJ23] Kristof Jannes, Emad Heydari Beni, Bert Lagaisse, and Wouter Joosen. BeauForT: Robust Byzantine fault tolerance for client-centric mobile web applications. IEEE Transactions on Parallel and Distributed Systems, 34(4):1241– 1252, April 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jin:2023:PEI

[JBY^{+}23] Hai Jin, Dongshan Bai, Dezhong Yao, Yutong Dai, Lin Gu, Chen Yu, and Lichao Sun. Personalized edge intelligence via federated self-knowledge distillation. IEEE Transactions on Parallel and Distributed Systems, 34(2): 567–580, February 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ji:2024:HKN

[JDD^{+}24] Xinyi Ji, Jiankuo Dong, Tonggui Deng, Pinchang Zhang, Jiafeng Hua, and Fu Xiao. HI-Kyber: a novel highperformance implementation scheme of Kyber based on GPU. IEEE Transactions on Parallel and Distributed Systems, 35(6):722–736, June 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jayanetti:2024:MAD

[JHB24] Amanda Jayanetti, Saman Halgamuge, and Rajkumar Buyya. Multi-agent deep reinforcement learning framework for renewable energy-aware workflow scheduling on distributed cloud data centers. IEEE Transactions on Parallel and Distributed Systems, 35(4):604– 615, April 2024. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jin:2022:DSR

[JJ22] Zhu Jin and Wen-Kang Jia. DH-SVRF: a reconfigurable unicast/multicast forwarding for high-performance packet forwarding engines. IEEE Transactions on Parallel and Distributed Systems, 33(5): 1262–1275, May 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jiang:2023:SPR

[JLG^{+}23] Xu Jiang, Haochun Liang, Nan Guan, Yue Tang, Lei Qiao, and Yi Wang. Scheduling parallel real-time tasks on virtual processors. IEEE Transactions on Parallel and Distributed Systems, 34(1):33–47, January 2023. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jannes:2021:OSS

[JLJ21] K. Jannes, B. Lagaisse, and W. Joosen. OWebSync: Seamless synchronization of distributed Web clients. IEEE Transactions on Parallel and Distributed Systems, 32(9): 2338–2351, September 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jeong:2020:RSH

[JLK^{+}20] W. S. Jeong, C. Lee, K. Kim, M. K. Yoon, W. Jeon, M. Jung, and W. W. Ro. REACT: Scalable and high-performance regular expression pattern matching accelerator for in-storage processing. IEEE Transactions on Parallel and Distributed Systems, 31(5):1137– 1151, May 2020. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jin:2020:HAH

[JLL^{+}20] H. Jin, Z. Li, H. Liu, X. Liao, and Y. Zhang. Hotspot-aware hybrid memory management for in-memory key–value stores. IEEE Transactions on Parallel and Distributed Systems, 31 (4):779–792, April 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jin:2022:RRB

[JLL^{+}22] Hai Jin, Cong Liu, Haikun Liu, Ruikun Luo, Jiahong Xu, Fubing Mao, and Xiaofei Liao. ReHy: a ReRAMbased digital/analog hybrid PIM architecture for accelerating CNN training. IEEE Transactions on Parallel and Distributed Systems, 33(11):2872– 2884, November 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jian:2023:TVT

[JLQ^{+}23] Zhaolong Jian, Ye Lu, Youyang Qiao, Yaozheng Fang, Xueshuo Xie, Dayi Yang, Zhiyuan Zhou, and Tao Li. TSC-VEE: a TrustZone-based smart contract virtual execution environment. IEEE Transactions on Parallel and Distributed Systems, 34(6):1773–1788, June 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jha:2024:GGD

[JLW^{+}24] Devki Nandan Jha, Yinhao Li, Zhenyu Wen, Graham Morgan, Prem Prakash Jayaraman, Maciej Koutny, Omer F. Rana, and Rajiv Ranjan. GeoDeploy: Geo-distributed application deployment using benchmarking. IEEE Transactions on Parallel and Distributed Systems, 35(12):2361–2374, December 2024. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jia:2024:ELC

[JLWS24] Linpeng Jia, Yanxiu Liu, Keyuan Wang, and Yi Sun. Estuary: a low cross-shard blockchain sharding protocol based on state splitting. IEEE Transactions on Parallel and Distributed Systems, 35(3):405– 420, March 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jiang:2023:CLL

[JLY^{+}23a] Wanchun Jiang, Haoyang Li, Yulong Yan, Fa Ji, Jiawei Huang, Jianxin Wang, and Tong Zhang. Consistent low latency scheduler for distributed key–value stores. IEEE Transactions on Parallel and Distributed Systems, 34(12):3012– 3027, December 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jin:2023:ACD

[JLY^{+}23b] Xiaozhong Jin, Haikun Liu, Chencheng Ye, Xiaofei Liao, Hai Jin, and Yu Zhang. Accelerating content-defined chunking for data deduplication based on speculative jump. IEEE Transactions on Parallel and Distributed Systems, 34(9): 2568–2579, September 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jain:2022:CGF

[JMF22] Abhishek Kumar Jain, Douglas L. Maskell, and Suhaib A. Fahmy. Coarse grained FPGA overlay for rapid just-in-time accelerator compilation. IEEE Transactions on Parallel and Distributed Systems, 33(6): 1478–1490, June 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jin:2022:RRD

[JQG^{+}22] Yibo Jin, Zhuzhong Qian, Song Guo, Sheng Zhang, Lei Jiao, and Sanglu Lu. runData: Re-distributing data via piggybacking for geo-distributed data analytics over edges. IEEE Transactions on Parallel and Distributed Systems, 33(1):40– 55, January 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jiang:2025:RRE

[JSL^{+}25] Junqiang Jiang, Zhifang Sun, Ruiqi Lu, Li Pan, and Zebo Peng. Real relative encoding genetic algorithm for workflow scheduling in heterogeneous distributed computing systems. IEEE Transactions on Parallel and Distributed Systems, 36(1):1–14, January 2025. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jiang:2022:EAD

[JTX^{+}22] Bingting Jiang, Zhuo Tang, Xiong Xiao, Jing Yao, Ronghui Cao, and Kenli Li. Efficient and automated deployment architecture for Open-Stack in TianHe SuperComputing environment. IEEE Transactions on Parallel and Distributed Systems, 33(8):1811– 1824, August 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jiang:2022:PDS

[JWW^{+}22] Jiantong Jiang, Zeyi Wen, Zeke Wang, Bingsheng He, and Jian Chen. Parallel and distributed structured SVM training. IEEE Transactions on Parallel and Distributed Systems, 33(5):1084–1096, May 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jin:2023:DQB

[JWZ^{+}23] Haoyu Jin, Donglei Wu, Shuyu Zhang, Xiangyu Zou, Sian Jin, Dingwen Tao, Qing Liao, and Wen Xia. Design of a quantization-based DNN delta compression framework for model snapshots and federated learning. IEEE Transactions on Parallel and Distributed Systems, 34(3):923–937, March 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jin:2024:GCP

[JWZ^{+}24] Yuyang Jin, Haojie Wang, Runxin Zhong, Chen Zhang, Xia Liao, Feng Zhang, and Jidong Zhai. Graph-centric performance analysis for largescale parallel applications. IEEE Transactions on Parallel and Distributed Systems, 35 (7):1221–1238, July 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jiang:2023:JMP

[JXX^{+}23] Zhida Jiang, Yang Xu, Hongli Xu, Lun Wang, Chunming Qiao, and Liusheng Huang. Joint model pruning and topology construction for accelerating decentralized machine learning. IEEE Transactions on Parallel and Distributed Systems, 34(10):2827– 2842, October 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jiang:2024:HHS

[JYF^{+}24] Zhe Jiang, Kecheng Yang, Nathan Fisher, Nan Guan, Neil C. Audsley, and Zheng Dong. Hopscotch: a hardwaresoftware co-design for efficient cache resizing on multi-core SoCs. IEEE Transactions on Parallel and Distributed Systems, 35(1):89–104, January 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jin:2024:EIP

[JZLZ24] Yuyang Jin, Runxin Zhong, Saiqin Long, and Jidong Zhai. Efficient inference for pruned CNN models on mobile devices with holistic sparsity alignment. IEEE Transactions on Parallel and Distributed Systems, 35(11):2208– 2223, November 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Karsavuran:2020:ROS

[KAA20] M. O. Karsavuran, S. Acer, and C. Aykanat. Reduce operations: Send volume balancing while minimizing latency. IEEE Transactions on Parallel and Distributed Systems, 31 (6):1461–1473, June 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Karsavuran:2021:PMG

[KAA21] M. O. Karsavuran, S. Acer, and C. Aykanat. Partitioning models for general medium-grain parallel sparse tensor decomposition. IEEE Transactions on Parallel and Distributed Systems, 32(1):147–159, January 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kettaneh:2020:NIS

[KAT^{+}20] I. Kettaneh, A. Alquraan, H. Takruri, S. Yang, A. C. Arpaci-Dusseau, R. H. Arpaci-Dusseau, and S. Al-Kiswany. The network-integrated storage system. IEEE Transactions on Parallel and Distributed Systems, 31(2):486–500, February 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Klein:2024:WCB

[KBA^{+}24] Joshua Klein, Irem Boybat, Giovanni Ansaloni, Marina Zapater, and David Atienza. Which coupled is best coupled? an exploration of AIMC tile interfaces and load balancing for CNNs. IEEE Transactions on Parallel and Distributed Systems, 35(10):1780–1795, October 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kozhaya:2021:PET

[KDREV21] D. Kozhaya, J. Decouchant, V. Rahli, and P. Esteves-Verissimo. PISTIS: an eventtriggered real-time Byzantineresilient protocol suite. IEEE Transactions on Parallel and Distributed Systems, 32(9): 2277–2290, September 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Korndorfer:2022:LDL

[KEMC22] Jonas H. Müller Korndörfer, Ahmed Eleliemy, Ali Mohammed, and Florina M. Ciorba. LB4OMP: a dynamic load balancing library for multithreaded applications. IEEE Transactions on Parallel and Distributed Systems, 33(4):830– 841, April 2022. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kayraklioglu:2021:MLB

[KFEG21] E. Kayraklioglu, E. Favry, and T. El-Ghazawi. A machinelearning-based framework for productive locality exploitation. IEEE Transactions on Parallel and Distributed Systems, 32(6):1409–1424, June 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Khaleghzadeh:2021:BOO

[KFS^{+}21] H. Khaleghzadeh, M. Fahad, A. Shahid, R. R. Manumachu, and A. Lastovetsky. Biobjective optimization of dataparallel applications on heterogeneous HPC platforms for performance and energy through workload distribution. IEEE Transactions on Parallel and Distributed Systems, 32(3):543– 560, March 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kim:2020:AFG

[KHLZ20] Y. K. Kim, M. R. Hoseiny-Farahabady, Y. C. Lee, and A. Y. Zomaya. Automated fine-grained CPU cap control in serverless computing platform. IEEE Transactions on Parallel and Distributed Systems, 31 (10):2289–2301, October 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Koohi:2020:MME

[KHOI20] S. Z. Koohi, N. A. W. A. Hamid, M. Othman, and G. Ibragimov. MEMPHA: Model of exascale messagepassing programs on heterogeneous architectures. IEEE Transactions on Parallel and Distributed Systems, 31(11): 2570–2581, November 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kelefouras:2023:DID

[KK23] Vasilios Kelefouras and Georgios Keramidas. Design and implementation of deep learning 2D convolutions on modern CPUs. IEEE Transactions on Parallel and Distributed Systems, 34(12):3104–3116, December 2023. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Keshavarz-Kohjerdi:2024:PMM

[KK24] Fatemeh Keshavarz-Kohjerdi. Paired many-to-many 2-disjoint path covers in meshes. IEEE Transactions on Parallel and Distributed Systems, 35(10): 1854–1866, October 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

KhudaBukhsh:2020:GCB

[KKA^{+}20] W. R. KhudaBukhsh, S. Kar, B. Alt, A. Rizk, and H. Koeppl. Generalized cost-based job scheduling in very large heterogeneous cluster systems. IEEE Transactions on Parallel and Distributed Systems, 31(11): 2594–2604, November 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kim:2021:PML

[KKP21] K.-R. Kim, Y. Kim, and S. Park. A probabilistic machine learning approach to scheduling parallel loops with Bayesian optimization. IEEE Transactions on Parallel and Distributed Systems, 32(7): 1815–1827, July 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Khochare:2021:SPD

[KKS21] A. Khochare, A. Krishnan, and Y. Simmhan. A scalable platform for distributed object tracking across a manycamera network. IEEE Transactions on Parallel and Distributed Systems, 32(6):1479– 1493, June 2021. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kang:2020:IMC

[KLH^{+}20a] Q. Kang, S. Lee, K. Hou, R. Ross, A. Agrawal, A. Choudhary, and W. Liao. Improving MPI collective I/O for high volume non-contiguous requests with intra-node aggregation. IEEE Transactions on Parallel and Distributed Systems, 31(11):2682– 2695, November 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kim:2020:PGE

[KLH^{+}20b] E. Kim, Y. Lee, L. He, K. G. Shin, and J. Lee. Power guarantee for electric systems using real-time scheduling. IEEE Transactions on Parallel and Distributed Systems, 31(8): 1783–1798, August 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kim:2024:IPP

[KLJV24] Jungwon Kim, Seyong Lee, Beau Johnston, and Jeffrey S. Vetter. IRIS: a performanceportable framework for crossplatform heterogeneous computing. IEEE Transactions on Parallel and Distributed Systems, 35(10):1796–1809, October 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kulkarni:2023:IES

[KM23a] Sourabh Kulkarni and Csaba Andras Moritz. Improving effectiveness of simulation-based inference in the massively parallel regime. IEEE Transactions on Parallel and Distributed Systems, 34(4):1100–1114, April 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kumar:2023:MCF

[KM23b] Manish Kumar and Anisur Rahaman Molla. On the message complexity of fault-tolerant computation: Leader election and agreement. IEEE Transactions on Parallel and Distributed Systems, 34(4):1115– 1127, April 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kumbhare:2020:VOJ

[KMA^{+}20] N. Kumbhare, A. Marathe, A. Akoglu, H. J. Siegel, G. Abdulla, and S. Hariri. A value-oriented job scheduling approach for power-constrained and oversubscribed HPC systems. IEEE Transactions on Parallel and Distributed Systems, 31(6):1419–1433, June 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kardani-Moghaddam:2021:AHA

[KMBR21] S. Kardani-Moghaddam, R. Buyya, and K. Ramamohanarao. ADRL: a hybrid anomaly-aware deep reinforcement learning-based resource scaling in clouds. IEEE Transactions on Parallel and Distributed Systems, 32 (3):514–526, March 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kang:2020:MAL

[KMLE20] D. H. Kang, C. Min, S. Lee, and Y. I. Eom. Making applicationlevel crash consistency practical on flash storage. IEEE Transactions on Parallel and Distributed Systems, 31(5):1009– 1020, May 2020. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kumar:2020:RAE

[KMM20] N. Kumar, J. Mayank, and A. Mondal. Reliability aware energy optimized scheduling of non-preemptive periodic realtime tasks on heterogeneous multiprocessor system. IEEE Transactions on Parallel and Distributed Systems, 31(4):871– 885, April 2020. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kojima:2022:MAK

[KOA22] Takuya Kojima, Ayaka Ohwada, and Hideharu Amano. Mappingaware kernel partitioning method for CGRAs assisted by deep learning. IEEE Transactions on Parallel and Distributed Systems, 33(5):1213–1230, May 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Koutsovasilis:2020:DUI

[KPA^{+}20] P. Koutsovasilis, K. Parasyris, C. D. Antonopoulos, N. Bellas, and S. Lalis. Dynamic undervolting to improve energy efficiency on multicore x86 CPUs. IEEE Transactions on Parallel and Distributed Systems, 31(12):2851–2864, December 2020. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kim:2020:RTS

[KPHA20] T. Kim, C. H. Park, J. Huh, and J. Ahn. Reconciling time slice conflicts of virtual machines with dual time slice for clouds. IEEE Transactions on Parallel and Distributed Systems, 31(10):2453–2465, October 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kim:2023:LAS

[KS23] Hyeonjin Kim and William J.

Song. LAS: Locality-aware scheduling for GEMM-accelerated convolutions in GPUs. IEEE Transactions on Parallel and Distributed Systems, 34(5): 1479–1494, May 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Khanda:2022:PAT

[KSB^{+}22] Arindam Khanda, Sriram Srinivasan, Sanjukta Bhowmick, Boyana Norris, and Sajal K. Das. A parallel algorithm template for updating single-source shortest paths in large-scale dynamic networks. IEEE Transactions on Parallel and Distributed Systems, 33(4):929– 940, April 2022. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Khdr:2020:CAT

[KSP^{+}20] H. Khdr, M. Shafique, S. Pagani, A. Herkersdorf, and J. Henkel. Combinatorial auctions for temperatureconstrained resource management in manycores. IEEE Transactions on Parallel and Distributed Systems, 31(7): 1605–1620, July 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kennedy:2023:MTG

[KSVR23] Jason Kennedy, Vishal Sharma, Blesson Varghese, and Carlos Reaño. Multi-tier GPU virtualization for deep learning in cloud-edge systems. IEEE Transactions on Parallel and Distributed Systems, 34 (7):2107–2123, July 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kleine:2022:CMM

[KSW^{+}22] Jan Kleine, Rahul Steiger, Simon Wachter, Emir man, Simon Jacob, and Dario Romaniello. Critique of MemXCT: Memory-centric Xray CT reconstruction with massive parallelization by SCC Team from ETH Zürich. IEEE Transactions on Parallel and Distributed Systems, 33(9): 2039–2042, September 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Karan:2024:EEB

[KSZ24] Subhadeep Karan, Zainul Abideen Sayed, and Jaroslaw Zola. Endto-end Bayesian networks exact learning in shared memory. IEEE Transactions on Parallel and Distributed Systems, 35 (4):634–645, April 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kim:2022:CSS

[KY22] Woo-Joong Kim and Chan-Hyun Youn. Cooperative scheduling schemes for explainable DNN acceleration in satellite image analysis and retraining. IEEE Transactions on Parallel and Distributed Systems, 33(7):1605–1618, July 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Koh:2019:EFT

[KZK^{+}19] S. Koh, J. Zhang, M. Kwon, J. Yoon, D. Donofrio, N. S. Kim, and M. Jung. Exploring fault-tolerant erasure codes for scalable all-flash array clusters. IEEE Transactions on Parallel and Distributed Systems, 30 (6):1312–1330, June 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). See errata [KZK^{+}20].

Koh:2020:EEF

[KZK^{+}20] S. Koh, J. Zhang, M. Kwon, J. Yoon, D. Donofrio, N. S. Kim, and M. Jung. Errata to exploring fault-tolerant erasure codes for scalable all-flash array clusters. IEEE Transactions on Parallel and Distributed Systems, 31(6):1460, June 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). See [KZK^{+}19].

Loreti:2024:RFR

[LAC24] Daniela Loreti, Marcello Artioli, and Anna Ciampolini. Rollback-free recovery for a high performance dense linear solver with reduced memory footprint. IEEE Transactions on Parallel and Distributed Systems, 35(7):1307– 1319, July 2024. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2022:CMMa

[LAG^{+}22] Xiaochen Li, Maximilian Apodaca, Arunav Gupta, Zihao Kong, Hongyi Pan, Hongyu Zhou, Mary Thomas, Martin Kandes, Zhaoyi Li, Mahidhar Tatineni, and Lewis Carroll. Critique of MemXCT: Memory-centric X-ray CT reconstruction with massive parallelization by SCC Team from University of California San Diego. IEEE Transactions on Parallel and Distributed Systems, 33(9):2043– 2046, September 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2021:ASS

[LAY21] M. Li, Y. Ao, and C. Yang. Adaptive SpMV/SpMSpV on GPUs for input vectors of varied sparsity. IEEE Transactions on Parallel and Distributed Systems, 32(7):1842– 1853, July 2021. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2021:BGB

[LBNN^{+}21] S. Li, T. Ben-Nun, G. Nadiradze, S. D. Girolamo, N. Dryden, D. Alistarh, and T. Hoefler. Breaking (global) barriers in parallel stochastic optimization with wait-avoiding group averaging. IEEE Transactions on Parallel and Distributed Systems, 32(7):1725– 1739, July 2021. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lessley:2020:DPH

[LC20] B. Lessley and H. Childs. Dataparallel hashing techniques for GPU architectures. IEEE Transactions on Parallel and Distributed Systems, 31(1):237– 250, January 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2020:SSA

[LCCZ20a] K. Li, J. Chen, W. Chen, and J. Zhu. SaberLDA: Sparsityaware learning of topic models on GPUs. IEEE Transactions on Parallel and Distributed Systems, 31(9):2112– 2124, September 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2020:AFL

[LCCZ20b] W. Liu, L. Chen, Y. Chen, and W. Zhang. Accelerating federated learning via momentum gradient descent. IEEE Transactions on Parallel and Distributed Systems, 31(8):1754– 1766, August 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2021:CPN

[LCG^{+}21] David Liu, Matthew Cinnamon, Thorne Garvin, Andrei Karavanov, Sungchan Park, Darius Strobeck, and Andrew Lumsdaine. Critique of planetary normal mode computation: Parallel algorithms, performance, and reproducibility by SCC Team From University of Washington. IEEE Transactions on Parallel and Distributed Systems, 32(11):2639– 2642, November 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). See [SLX^{+}21b].

Li:2023:CCI

[LCH23] Tengfei Li, Jianfeng Chu, and Liang Hu. CIA: a collaborative integrity auditing scheme for cloud data with multireplica on multi-cloud storage providers. IEEE Transactions on Parallel and Distributed Systems, 34(1):154–162, January 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lall:2024:LSB

[LCIS24] Sanjay Lall, Călin Caşcaval, Martin Izzard, and Tammo Spalink. Logical synchrony and the bittide mechanism. IEEE Transactions on Parallel and Distributed Systems, 35(11): 1936–1948, November 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lim:2020:NEC

[LCL^{+}20] S. Lim, T. Coy, Z. Lu, B. Ren, and X. Zhang. NVGraph: Enforcing crash consistency of evolving network analytics in NVMM systems. IEEE Transactions on Parallel and Distributed Systems, 31(6):1255– 1269, June 2020. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2024:SEE

[LCL^{+}24] Zhao Liu, Xuesen Chu, Xiaojing Lv, Hongsong Meng, Hanyue Liu, Guanghui Zhu, Haohuan Fu, and Guangwen Yang. SunwayLB: Enabling extreme-scale lattice Boltzmann method based computing fluid dynamics simulations on advanced heterogeneous supercomputers. IEEE Transactions on Parallel and Distributed Systems, 35(2):324–337, February 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lin:2021:PPS

[LCLW21] W. Lin, H. Cui, B. Li, and C. Wang. Privacy-preserving similarity search with efficient updates in distributed keyvalue stores. IEEE Transactions on Parallel and Distributed Systems, 32(5):1072– 1084, May 2021. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lu:2020:NMS

[LCM^{+}20] Z. Lu, Q. Cao, F. Mei, H. Jiang, and J. Li. A novel multi-stage forest-based key–value store for holistic performance improvement. IEEE Transactions on Parallel and Distributed Systems, 31(4):856–870, April 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2022:BGB

[LCX^{+}22] Mingfan Li, Junshi Chen, Qian Xiao, Fei Wang, Qingcai Jiang, Xuncheng Zhao, Rongfen Lin, Hong An, Xiao Liang, and Lixin He. Bridging the gap between deep learning and frustrated quantum spin system for extreme-scale simulations on new generation of Sunway supercomputer. IEEE Transactions on Parallel and Distributed Systems, 33(11):2846– 2859, November 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Luo:2024:FFV

[LCX^{+}24] Ping Luo, Jieren Cheng, N. Xiong, Zhenhao Liu, and Jie Wu. FedVeca: Federated vectorized averaging on Non-IID data with adaptive bi-directional global objective. IEEE Transactions on Parallel and Distributed Systems, 35(11):2102–2113, November 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2020:TPE

[LCZ^{+}20] X. Li, W. Cheng, T. Zhang, F. Ren, and B. Yang. Towards power efficient high performance packet I/O. IEEE Transactions on Parallel and Distributed Systems, 31(4):981– 996, April 2020. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2023:CPF

[LCZ^{+}23] Guancheng Li, Songhui Cao, Chuyi Zhao, Siyuan Zhang, Yuchen Ji, Haotian Jing, Zecheng Li, Jiajun Cheng, Yiwei Yang, and Shu Yin. Critique of A Parallel Framework for Constraint-Based Bayesian Network Learning via Markov Blanket Discovery by SCC Team From ShanghaiTech University. IEEE Transactions on Parallel and Distributed Systems, 34(6):1716–1719, June 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). See [SCA23].

Lu:2023:TTF

[LDJX^{+}23] Zhichao Lu, Chuntao Ding, Felix Juefei-Xu, Vishnu Naresh Boddeti, Shangguang Wang, and Yun Yang. TFormer: a transmission-friendly ViT model for IoT devices. IEEE Transactions on Parallel and Distributed Systems, 34(2):598– 610, February 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lei:2022:EGF

[LDL22] Fei Lei, Dezun Dong, and Xiangke Liao. Exploring the galaxyfly family to build flexible-scale interconnection networks. IEEE Transactions on Parallel and Distributed Systems, 33(5):1054–1068, May 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2024:SIW

[LDZ^{+}24] Yuhang Liu, Xin Deng, Jiapeng Zhou, Mingyu Chen, and Yungang Bao. Suppressing the interference within a datacenter: Theorems, metric and strategy. IEEE Transactions on Parallel and Distributed Systems, 35 (5):732–750, May 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2021:SML

[LFZ^{+}21] W. Li, C. Feng, L. Zhang, H. Xu, B. Cao, and M. A. Imran. A scalable multilayer PBFT consensus for blockchain. IEEE Transactions on Parallel and Distributed Systems, 32(5):1146–1160, May 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2022:BPG

[LGC^{+}22] Yuxuan Li, Lin Gan, Mingcheng Chen, Yaojian Chen, Haitian Lu, Chaoyang Lu, Jianwei Pan, Haohuan Fu, and Guangwen Yang. Benchmarking 50photon Gaussian boson sampling on the Sunway Taihu-Light. IEEE Transactions on Parallel and Distributed Systems, 33(6):1357–1372, June 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2024:TEL

[LGH^{+}24] Ning Li, Jianmei Guo, Bo Huang, Yuyang Li, Yilei Zhang, Chengdong Li, and Wenxin Huang. TCSA: Efficient localization of busy-wait synchronization bugs for latency-critical applications. IEEE Transactions on Parallel and Distributed Systems, 35(2):297–309, February 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Luo:2021:MMP

[LGZ^{+}21] Lailong Luo, Deke Guo, Yawei Zhao, Ori Rottenstreich, Richard T. B. Ma, and Xueshan Luo. MCFsyn: a multiparty set reconciliation protocol with the marked cuckoo filter. IEEE Transactions on Parallel and Distributed Systems, 32(11):2705–2718, November 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2024:FEF

[LGZ^{+}24] Xiaqing Li, Qi Guo, Guangyan Zhang, Siwei Ye, Guanhua He, Yiheng Yao, Rui Zhang, Yifan Hao, Zidong Du, and Weimin Zheng. FastTuning: Enabling fast and efficient hyperparameter tuning with partitioning and parallelism of search space. IEEE Transactions on Parallel and Distributed Systems, 35(7):1174– 1188, July 2024. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lin:2022:ALS

[LH22] Dian-Lun Lin and Tsung-Wei Huang. Accelerating large sparse neural network inference using GPU task graph parallelism. IEEE Transactions on Parallel and Distributed Systems, 33(11):3041– 3052, November 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2021:ACD

[LHC^{+}21] B. Li, Q. He, F. Chen, H. Jin, Y. Xiang, and Y. Yang. Auditing cache data integrity in the edge computing environment. IEEE Transactions on Parallel and Distributed Systems, 32 (5):1210–1223, May 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Luo:2025:RED

[LHC^{+}25] Ruikun Luo, Qiang He, Feifei Chen, Song Wu, Hai Jin, and Yun Yang. Ripple: Enabling decentralized data deduplication at the edge. IEEE Transactions on Parallel and Distributed Systems, 36(1):55–66, January 2025. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2022:AIM

[LHL^{+}22] Jiaqi Liu, Shiyue Huang, Deng Li, Sheng Wen, and Hui Liu. Addictive incentive mechanism in crowdsensing from the perspective of behavioral economics. IEEE Transactions on Parallel and Distributed Systems, 33(5):1109–1127, May 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2023:EMA

[LHL23] Wenjie Liu, Xubin He, and Qing Liu. Exploring memory access similarity to improve irregular application performance for distributed hybrid memory systems. IEEE Transactions on Parallel and Distributed Systems, 34(3):797– 809, March 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lv:2020:ABA

[LHPW20] H. Lv, J. Hillston, P. Piho, and H. Wang. An attributebased availability model for large scale IaaS clouds with CARMA. IEEE Transactions on Parallel and Distributed Systems, 31(3):733–748, March 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2020:QHP

[LHQ^{+}20] J. Li, X. Hu, D. Qian, C. Wei, G. McFadden, B. Will, P. Yu, W. Li, and H. Guan. QWEB: High-performance event-driven web architecture with QAT acceleration. IEEE Transactions on Parallel and Distributed Systems, 31(11):2633– 2649, November 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Langer:2020:DTD

[LHRX20] M. Langer, Z. He, W. Rahayu, and Y. Xue. Distributed training of deep learning models: a taxonomic perspective. IEEE Transactions on Parallel and Distributed Systems, 31(12):2802–2818, December 2020. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Luo:2025:EDD

[LHX^{+}25] Ruikun Luo, Qiang He, Mengxi Xu, Feifei Chen, Song Wu, Jing Yang, Yuan Gao, and Hai Jin. Edge data deduplication under uncertainties: a robust optimization approach. IEEE Transactions on Parallel and Distributed Systems, 36 (1):84–95, January 2025. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lin:2022:PFD

[LHXH22] Limei Lin, Yanze Huang, Li Xu, and Sun-Yuan Hsieh. A pessimistic fault diagnosability of large-scale connected networks via extra connectivity. IEEE Transactions on Parallel and Distributed Systems, 33(2):415– 428, February 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2023:EFG

[LHZ^{+}23] Zhuan Liu, Ruichen Han, Yansong Zhang, Yu Zhang, Xi Tang, Gang Deng, Tao Zhong, Roman Dementiev, Yunfei Lu, and Mingjian Que. Exploring fine-grained inmemory database performance for modern CPUs. IEEE Transactions on Parallel and Distributed Systems, 34(6):1757– 1772, June 2023. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2021:SRI

[LIP^{+}21] Y. Li, M. Interlandi, F. Psallidas, W. Wang, and C. Zaniolo. SEIZE: Runtime inspection for parallel dataflow systems. IEEE Transactions on Parallel and Distributed Systems, 32(4):842–854, April 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Luo:2023:EBD

[LJH^{+}23] Ruikun Luo, Hai Jin, Qiang He, Song Wu, and Xiaoyu Xia. Enabling balanced data deduplication in mobile edge computing. IEEE Transactions on Parallel and Distributed Systems, 34 (5):1420–1431, May 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2023:MJI

[LJM^{+}23] Ji Liu, Juncheng Jia, Beichen Ma, Chendi Zhou, Jingbo Zhou, Yang Zhou, Huaiyu Dai, and Dejing Dou. Multijob intelligent scheduling with cross-device federated learning. IEEE Transactions on Parallel and Distributed Systems, 34(2): 535–551, February 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ling:2025:JDD

[LJT^{+}25] Zhi Ling, Xiaofeng Jiang, Xiaobin Tan, Huasen He, Shiyin Zhu, and Jian Yang. Joint dynamic data and model parallelism for distributed training of DNNs over heterogeneous infrastructure. IEEE Transactions on Parallel and Distributed Systems, 36(2):150– 167, February 2025. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2023:FDI

[LJW^{+}23] Bosheng Liu, Zhuoshen Jiang, Yalan Wu, Jigang Wu, Xiaoming Chen, Peng Liu, Qingguo Zhou, and Yinhe Han. Frequency-domain inference acceleration for convolutional neural networks using ReRAMs. IEEE Transactions on Parallel and Distributed Systems, 34(12):3133– 3146, December 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2020:AGH

[LJZ^{+}20] Z. Li, H. Jia, Y. Zhang, T. Chen, L. Yuan, and R. Vuduc. Automatic generation of high-performance FFT kernels on Arm and x86 CPUs. IEEE Transactions on Parallel and Distributed Systems, 31(8): 1925–1941, August 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2020:ENO

[LJZY20] Z. Li, H. Jin, D. Zou, and B. Yuan. Exploring new opportunities to defeat lowrate DDoS attack in containerbased cloud environment. IEEE Transactions on Parallel and Distributed Systems, 31(3):695– 706, March 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Langr:2020:RII

[LK20] Daniel Langr and Marin Kočička. Reducing the impact of intensive dynamic memory allocations in parallel multithreaded programs. IEEE Transactions on Parallel and Distributed Systems, 31(5): 1152–1164, May 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lin:2021:MOP

[LK21] C. Lin and H. Khazaei. Modeling and optimization of performance and cost of serverless applications. IEEE Transactions on Parallel and Distributed Systems, 32(3):615– 632, March 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lee:2023:MTO

[LKH23] Young Sik Lee, Yong Wook Kim, and Tae Hee Han. MRCN: Throughput-oriented multicast routing for customized network-on-chips. IEEE Transactions on Parallel and Distributed Systems, 34(1):163– 179, January 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2020:PPR

[LL20] L. Li and L. Lazos. Proofs of physical reliability for cloud storage systems. IEEE Transactions on Parallel and Distributed Systems, 31(5):1048– 1065, May 2020. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2022:CMMb

[LL22] Shenggui Li and Bu-Sung Lee. Critique of MemXCT: Memory-centric X-ray CT reconstruction with massive parallelization by SCC Team from Nanyang Technological University. IEEE Transactions on Parallel and Distributed Systems, 33(9):2058– 2061, September 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2021:HMS

[LLC^{+}21] W. Li, D. Liu, K. Chen, K. Li, and H. Qi. Hone: Mitigating stragglers in distributed stream processing with tuple scheduling. IEEE Transactions on Parallel and Distributed Systems, 32(8):2021–2034, August 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2022:PAS

[LLC^{+}22] Jian Liu, Peilun Li, Raymond Cheng, N. Asokan, and Dawn Song. Parallel and asynchronous smart contract execution. IEEE Transactions on Parallel and Distributed Systems, 33(5):1097–1108, May 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lindquist:2022:ARG

[LLD22] Neil Lindquist, Piotr Luszczek, and Jack Dongarra. Accelerating restarted GMRES with mixed precision arithmetic. IEEE Transactions on Parallel and Distributed Systems, 33(4):1027–1037, April 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lu:2023:AAD

[LLDL23] Qingguo Lü, Xiaofeng Liao, Shaojiang Deng, and Huaqing Li. Asynchronous algorithms for decentralized resource allocation over directed networks.

IEEE Transactions on Parallel and Distributed Systems, 34 (1):16–32, January 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2020:RAN

[LLHJ20] J. Li, W. Liang, M. Huang, and X. Jia. Reliability-aware network service provisioning in mobile edge-cloud networks. IEEE Transactions on Parallel and Distributed Systems, 31 (7):1545–1558, July 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Luo:2021:EPN

[LLJC21] J. Luo, J. Li, L. Jiao, and J. Cai. On the effective parallelization and near-optimal deployment of service function chains. IEEE Transactions on Parallel and Distributed Systems, 32(5):1238–1255, May 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2022:PSH

[LLK22] Yuan Li, Ahmed Louri, and Avinash Karanth. SPRINT: a high-performance, energyefficient, and scalable chipletbased accelerator with photonic interconnects for CNN inference. IEEE Transactions on Parallel and Distributed Systems, 33(10):2332–2345, October 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2024:HPE

[LLK24] Yuan Li, Ahmed Louri, and Avinash Karanth. A highperformance and energy-efficient photonic architecture for multi-DNN acceleration. IEEE Transactions on Parallel and Distributed Systems, 35(1):46– 58, January 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2021:STN

[LLL^{+}21a] H. Li, Z. Li, K. Li, J. S. Rellermeyer, L. Chen, and K. Li. SGD Tucker: a novel stochastic optimization strategy for parallel sparse Tucker decomposition. IEEE Transactions on Parallel and Distributed Systems, 32(7):1828–1841, July 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2021:EFA

[LLL^{+}21b] Qi Li, Yunpeng Liu, Zhuotao Liu, Peng Zhang, and Chunhui Pang. Efficient forwarding anomaly detection in softwaredefined networks. IEEE Transactions on Parallel and Distributed Systems, 32(11):2676– 2690, November 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Long:2021:GBA

[LLL^{+}21c] S. Long, W. Long, Z. Li, K. Li, Y. Xia, and Z. Tang. A game-based approach for costaware task assignment with QoS constraint in collaborative edge and cloud environments. IEEE Transactions on Parallel and Distributed Systems, 32 (7):1629–1640, July 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2023:RMO

[LLL^{+}23a] Haolin Liu, Saiqin Long, Zhetao Li, Yu Fu, Yong Zuo, and Xinglin Zhang. Revenue maximizing online service function chain deployment in multitier computing network. IEEE Transactions on Parallel and Distributed Systems, 34(3):781– 796, March 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2023:EHR

[LLL^{+}23b] Zhao Liu, MengQuan Li, Min-Can Li, Lei Liao, and Kenli Li. An efficient hierarchicalreduction architecture for aggregation in route travel time estimation. IEEE Transactions on Parallel and Distributed Systems, 34(9):2541– 2552, September 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2024:MEH

[LLL^{+}24a] Dongsheng Li, Shengwei Li, Zhiquan Lai, Yongquan Fu, Xiangyu Ye, Lei Cai, and Linbo Qiao. A memory-efficient hybrid parallel framework for deep neural network training. IEEE Transactions on Parallel and Distributed Systems, 35 (4):577–591, April 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2024:MCS

[LLL^{+}24b] Shengwei Li, Kai Lu, Zhiquan Lai, Weijie Liu, Keshi Ge, and Dongsheng Li. A multidimensional communication scheduling method for hybrid parallel DNN training. IEEE Transactions on Parallel and Distributed Systems, 35(8):1415– 1428, August 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2023:ADD

[LLP^{+}23] Hao Liu, Wenxin Li, Yiren Pang, Renjie Pei, Yitao Hu, Yuan Liu, Lide Suo, and Keqiu Li. Accelerating data delivery of latency-sensitive applications in container overlay network. IEEE Transactions on Parallel and Distributed Systems, 34(12):3046–3058, December 2023. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lu:2024:CFA

[LLR^{+}24] Huali Lu, Feng Lyu, Ju Ren, Huaqing Wu, Conghao Zhou, Zhongyuan Liu, Yaoxue Zhang, and Xuemin Shen. CODE^{+}+: Fast and accurate inference for compact distributed IoT data collection. IEEE Transactions on Parallel and Distributed Systems, 35(11):2006– 2022, November 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2024:FEM

[LLS^{+}24] Changlong Li, Yu Liang, Liang Shi, Chao Wang, Chun Jason Xue, and Xuehai Zhou. Flexible and efficient memory swapping across mobile devices with LegoSwap. IEEE Transactions on Parallel and Distributed Systems, 35(1):140–153, January 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lai:2023:MED

[LLT^{+}23] Zhiquan Lai, Shengwei Li, Xudong Tang, Keshi Ge, Weijie Liu, Yabo Duan, Linbo Qiao, and Dongsheng Li. Merak: an efficient distributed DNN training framework with automated 3D parallelism for giant foundation models. IEEE Transactions on Parallel and Distributed Systems, 34(5):1466– 1478, May 2023. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2022:MUS

[LLX^{+}22] Jing Li, Weifa Liang, Wenzheng Xu, Zichuan Xu, Xiaohua Jia, Wanlei Zhou, and Jin Zhao. Maximizing user service satisfaction for delay-sensitive IoT applications in edge computing. IEEE Transactions on Parallel and Distributed Systems, 33 (5):1199–1212, May 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2020:ASC

[LLY^{+}20] M. Li, Y. Liu, H. Yang, Z. Luan, L. Gan, G. Yang, and D. Qian. Accelerating sparse Cholesky factorization on Sunway manycore architecture. IEEE Transactions on Parallel and Distributed Systems, 31(7):1636–1650, July 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lu:2023:RRL

[LLZ^{+}23] Xiaofeng Lu, Chao Liu, Senhao Zhu, Yilu Mao, Pietro Lio, and Pan Hui. RLPTO: a reinforcement learning-based performance-time optimized task and resource scheduling mechanism for distributed machine learning. IEEE Transactions on Parallel and Distributed Systems, 34(12):3266– 3279, December 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

laus:2020:EMP

[lM20] D. laus and D. Mongus. Efficient method for parallel computation of geodesic transformation on CPU. IEEE Transactions on Parallel and Distributed Systems, 31(4):935– 947, April 2020. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lin:2023:FGP

[LMFK23] Changyuan Lin, Nima Mahmoudi, Caixiang Fan, and Hamzeh Khazaei. Fine-grained performance and cost modeling and optimization for FaaS applications. IEEE Transactions on Parallel and Distributed Systems, 34(1):180–194, January 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lan:2020:FFI

[LMH^{+}20] H. Lan, J. Meng, C. Hundt, B. Schmidt, M. Deng, X. Wang, W. Liu, Y. Qiao, and S. Feng. FeatherCNN: Fast inference computation with TensorGEMM on ARM architectures. IEEE Transactions on Parallel and Distributed Systems, 31(3):580– 594, March 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2022:RPE

[LMX^{+}22] Yuchen Liu, Yixuan Meng, Kaiyuan Xu, Zijun Xu, Tianyuan Wu, Yiwei Yang, and Shu Yin. Reproducibility: Performance evaluation of MemXCT on Azure CycleCloud Platformn. IEEE Transactions on Parallel and Distributed Systems, 33(9): 2047–2049, September 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lin:2024:ESC

[LMX^{+}24] Zhiqi Lin, Youshan Miao, Guanbin Xu, Cheng Li, Olli Saarikivi, Saeed Maleki, and Fan Yang. Efficient schedule construction for distributed execution of large DNN models. IEEE Transactions on Parallel and Distributed Systems, 35(12):2375–2391, December 2024. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2020:PBD

[LMZ^{+}20] L. Liu, X. Man, J. Zhu, S. Yin, and S. Wei. Pattern-based dynamic compilation system for CGRAs with online configuration transformation. IEEE Transactions on Parallel and Distributed Systems, 31(12): 2981–2994, December 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2024:DCS

[LN24] Fangming Liu and Yipei Niu.

Demystifying the cost of serverless computing: Towards a win-win deal. IEEE Transactions on Parallel and Distributed Systems, 35(1):59–72, January 2024. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lim:2022:DEI

[LNX^{+}22] Wei Yang Bryan Lim, Jer Shyuan Ng, Zehui Xiong, Jiangming Jin, Yang Zhang, Dusit Niyato, Cyril Leung, and Chunyan Miao. Decentralized edge intelligence: a dynamic resource allocation framework for hierarchical federated learning. IEEE Transactions on Parallel and Distributed Systems, 33 (3):536–550, March 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2022:PTD

[LPG^{+}22] Zhuojin Li, Marco Paolieri, Leana Golubchik, Sung-Han Lin, and Wumo Yan. Predicting throughput of distributed stochastic gradient descent. IEEE Transactions on Parallel and Distributed Systems, 33(11):2900–2912, November 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lu:2024:CTS

[LPH^{+}24] Jianyuan Lu, Tian Pan, Shan He, Mao Miao, Guangzhe Zhou, Yining Qi, Shize Zhang, Enge Song, Xiaoqing Sun, Huaiyi Zhao, Biao Lyu, and Shunmin Zhu. CloudSentry: Two-stage heavy hitter detection for cloud-scale gateway overload protection. IEEE Transactions on Parallel and Distributed Systems, 35(4):616– 633, April 2024. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2023:RCD

[LPL23] Mingyu Liu, Li Pan, and Shijun Liu. RLTiering: a cost-driven auto-tiering system for two-tier cloud storage using deep reinforcement learning. IEEE Transactions on Parallel and Distributed Systems, 34(2):73– 90, February 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2020:SAD

[LPW^{+}20] K. Liu, J. Peng, J. Wang, W. Liu, Z. Huang, and J. Pan. Scalable and adaptive data replica placement for geo-distributed cloud storages. IEEE Transactions on Parallel and Distributed Systems, 31 (7):1575–1587, July 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liegeois:2023:PPB

[LRBV23] Kim Liegeois, Sivasankaran Rajamanickam, and Luc Berger-Vergiat. Performance portable batched sparse linear solvers. IEEE Transactions on Parallel and Distributed Systems, 34 (5):1524–1535, May 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2021:ABN

[LS21] A. Li and S. Su. Accelerating binarized neural networks via bit-tensor-cores in Turing GPUs. IEEE Transactions on Parallel and Distributed Systems, 32(7):1878– 1891, July 2021. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2020:EMG

[LSC^{+}20] A. Li, S. L. Song, J. Chen, J. Li, X. Liu, N. R. Tallent, and K. J. Barker. Evaluating modern GPU interconnect: PCIe, NVLink, NV-SLI, NVSwitch and GPUDirect. IEEE Transactions on Parallel and Distributed Systems, 31(1):94–110, January 2020. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2023:EJF

[LSL^{+}23] Ting Li, Jiyan Sun, Yinlong Liu, Xu Zhang, Dali Zhu, Zhaorui Guo, and Liru Geng. ESMO: Joint frame scheduling and model caching for edge video analytics. IEEE Transactions on Parallel and Distributed Systems, 34(8):2295– 2310, August 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2023:ADP

[LSW^{+}23] Li Li, Jiajie Shen, Bochun Wu, Yangfan Zhou, Xin Wang, and Keqin Li. Adaptive data placement in multi-cloud storage: a non-stationary combinatorial bandit approach. IEEE Transactions on Parallel and Distributed Systems, 34(11):2843– 2859, November 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2020:SST

[LSY^{+}20] L. Liu, Y. Shen, Y. Yan, T. Yang, M. Shahzad, B. Cui, and G. Xie. SF-Sketch: a twostage sketch for data streams. IEEE Transactions on Parallel and Distributed Systems, 31 (10):2263–2276, October 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2021:JSD

[LSY21] Y. Liu, X. Shang, and Y. Yang.

Joint SFC deployment and resource management in heterogeneous edge for latency minimization. IEEE Transactions on Parallel and Distributed Systems, 32(8):2131–2143, August 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2021:PON

[LSZ^{+}21] X. Liu, J. Sun, L. Zheng, S. Wang, Y. Liu, and T. Wei. Parallelization and optimization of NSGA-II on Sunway TaihuLight system. IEEE Transactions on Parallel and Distributed Systems, 32(4):975– 987, April 2021. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2020:FTP

[LT20] C. Li and X. Tang. On faulttolerant bin packing for online resource allocation. IEEE Transactions on Parallel and Distributed Systems, 31(4):817– 829, April 2020. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2021:DTM

[LTH^{+}21] C. Liu, F. Tang, Y. Hu, K. Li, Z. Tang, and K. Li. Distributed task migration optimization in MEC by extending multi-agent deep reinforcement learning approach. IEEE Transactions on Parallel and Distributed Systems, 32(7):1603–1614, July 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2020:ETO

[LTT^{+}20] K. Liu, B. Tian, C. Tian, B. Li, Q. Wang, J. Zheng, J. Sun, Y. Gao, W. Wang, G. Chen, W. Dou, Y. Jiang, H. Zhou, J. Jiang, F. Zhang, and G. Zhang. Exploring tokenoriented in-network prioritization in datacenter networks. IEEE Transactions on Parallel and Distributed Systems, 31 (5):1223–1238, May 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2023:LDR

[LTZ^{+}23] Shiyang Li, Ruiqi Tang, Jingyu Zhu, Ziyi Zhao, Xiaoli Gong, Wenwen Wang, Jin Zhang, and Pen-Chung Yew. Liberator: a data reuse framework for out-of-memory graph computing on GPUs. IEEE Transactions on Parallel and Distributed Systems, 34(6):1954– 1967, June 2023. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2020:EEC

[LWC^{+}20] C. Liu, Q. Wang, X. Chu, Y. Leung, and H. Liu. ESet-Store: an erasure-coded storage system with fast data recovery. IEEE Transactions on Parallel and Distributed Systems, 31(9): 2001–2016, September 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2022:TAN

[LWC^{+}22] Jiahui Li, Hao Wu, Jiapei Chen, Qiang He, and Ching-Hsien Hsu. Topology-aware neural model for highly accurate QoS prediction. IEEE Transactions on Parallel and Distributed Systems, 33(7): 1538–1552, July 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2022:PIT

[LWL^{+}22a] Zecheng Li, Haotian Wu, Lap Hou Lao, Songtao Guo, Yuanyuan Yang, and Bin Xiao. Pistis: Issuing trusted and authorized certificates with distributed ledger and TEE. IEEE Transactions on Parallel and Distributed Systems, 33(7): 1636–1649, July 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Long:2022:GCA

[LWL^{+}22b] Saiqin Long, Wen Wen, Zhetao Li, Kenli Li, Rong Yu, and Jiang Zhu. A global cost-aware container scheduling strategy in cloud data centers. IEEE Transactions on Parallel and Distributed Systems, 33(11): 2752–2766, November 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Luo:2023:RDO

[LWL^{+}23] Chuan Luo, Sizhao Wang, Tianrui Li, Hongmei Chen, Jiancheng Lv, and Zhang Yi. RHDOFS: a distributed online algorithm towards scalable streaming feature selection. IEEE Transactions on Parallel and Distributed Systems, 34(6):1830–1847, June 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lan:2024:FGF

[LWL^{+}24] Linsi Lan, Junbo Wang, Zhi Li, Krishna Kant, and Wanquan Liu. FedREM: Guided federated learning in the presence of dynamic device unpredictability. IEEE Transactions on Parallel and Distributed Systems, 35(7):1189– 1206, July 2024. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lin:2025:TDB

[LWL^{+}25] Shuai Lin, Rui Wang, Yongkun Li, Yinlong Xu, and John C. S. Lui. Two-dimensional balanced partitioning and efficient caching for distributed graph analysis. IEEE Transactions on Parallel and Distributed Systems, 36(2):133–149, February 2025. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Luo:2024:ECC

[LWLX24] Shouxi Luo, Renyi Wang, Ke Li, and Huanlai Xing. Efficient cross-cloud partial reduce with CREW. IEEE Transactions on Parallel and Distributed Systems, 35(11):2224– 2238, November 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2023:SDR

[LWX^{+}23] Chenyi Liu, Pingfei Wu, Mingwei Xu, Yuan Yang, and Nan Geng. Scalable deep reinforcement learning-based online routing for multi-type service requirements. IEEE Transactions on Parallel and Distributed Systems, 34(8):2337– 2351, August 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2020:QSC

[LWY^{+}20] R. Li, B. Wu, M. Ying, X. Sun, and G. Yang. Quantum supremacy circuit simulation on Sunway TaihuLight. IEEE Transactions on Parallel and Distributed Systems, 31 (4):805–816, April 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lv:2022:MDE

[LWY^{+}22] Wenkai Lv, Quan Wang, Pengfei Yang, Yunqing Ding, Bijie Yi, Zhenyi Wang, and Chengmin Lin. Microservice deployment in edge computing based on Deep Q learning. IEEE Transactions on Parallel and Distributed Systems, 33(11):2968–2978, November 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2022:MSC

[LWZ^{+}22] Huifang Li, Danjing Wang, MengChu Zhou, Yushun Fan, and Yuanqing Xia. Multiswarm co-evolution based hybrid intelligent optimization for bi-objective multi-workflow scheduling in the cloud. IEEE Transactions on Parallel and Distributed Systems, 33(9): 2183–2197, September 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lei:2023:FLC

[LWZ^{+}23a] Runze Lei, Pinghui Wang, Junzhou Zhao, Lin Lan, Jing Tao, Chao Deng, Junlan Feng, Xidian Wang, and Xiaohong Guan. Federated learning over coupled graphs. IEEE Transactions on Parallel and Distributed Systems, 34(4):1159–1172, April 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2023:LCL

[LWZ23b] Mingzhe Li, Wei Wang, and Jin Zhang. LB-Chain: Loadbalanced and low-latency blockchain sharding via account migration. IEEE Transactions on Parallel and Distributed Systems, 34 (10):2797–2810, October 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2024:ETU

[LWZ^{+}24] Fangxin Liu, Zongwu Wang, Wenbo Zhao, Ning Yang, Yongbiao Chen, Shiyuan Huang, Haomin Li, Tao Yang, Songwen Pei, Xiaoyao Liang, and Li Jiang. Exploiting temporalunrolled parallelism for energyefficient SNN acceleration. IEEE Transactions on Parallel and Distributed Systems, 35 (10):1749–1764, October 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2023:NCE

[LX23] Kan Liu and Wei Xue. A novel compute-efficient tridiagonal solver for many-core architectures. IEEE Transactions on Parallel and Distributed Systems, 34(1):195–206, January 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2022:DRL

[LXC^{+}22] Qingzhi Liu, Tiancong Xia, Long Cheng, Merijn van Eijk, Tanir Ozcelebi, and Ying Mao. Deep reinforcement learning for load-balancing aware network control in IoT edge systems. IEEE Transactions on Parallel and Distributed Systems, 33 (6):1491–1502, June 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2024:SSQ

[LXC24] Peixuan Li, Ping Xie, and Qiang Cao. SSRAID: a stripequeued and stripe-threaded merging I/O strategy to improve write performance of serial interface SSD RAID. IEEE Transactions on Parallel and Distributed Systems, 35(10): 1841–1853, October 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2023:SDS

[LXGY23] Zecheng Li, Bin Xiao, Songtao Guo, and Yuanyuan Yang. Securing deployed smart contracts and DeFi with distributed TEE cluster. IEEE Transactions on Parallel and Distributed Systems, 34(3):828– 842, March 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2023:CCO

[LXL23] Yang Liu, Huanle Xu, and Wing Cheong Lau. Cloud configuration optimization for recurring batch-processing applications. IEEE Transactions on Parallel and Distributed Systems, 34(5):1495–1507, May 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2024:EEE

[LXL^{+}24] Qiliang Li, Liangliang Xu, Yongkun Li, Min Lyu, Wei Wang, Pengfei Zuo, and Yinlong Xu. Enabling efficient erasure coding in disaggregated memory systems. IEEE Transactions on Parallel and Distributed Systems, 35(1):154– 168, January 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2024:SFS

[LXQ^{+}24] Yichen Li, Wenchao Xu, Yining Qi, Haozhao Wang, Ruixuan Li, and Song Guo. SR-FDIL: Synergistic replay for federated domain-incremental learning. IEEE Transactions on Parallel and Distributed Systems, 35(11):1879–1890, November 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2023:TGA

[LXW^{+}23] Xiaocan Li, Kun Xie, Xin Wang, Gaogang Xie, Kenli Li, Jiannong Cao, Dafang Zhang, and Jigang Wen. Tripartite graph aided tensor completion for sparse network measurement. IEEE Transactions on Parallel and Distributed Systems, 34(1):48–62, January 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liang:2021:IAH

[LYDZ21] X. Liang, Z. Yan, R. H. Deng, and Q. Zheng. Investigating the adoption of hybrid encrypted cloud data deduplication with game theory. IEEE Transactions on Parallel and Distributed Systems, 32(3):587– 600, March 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lu:2020:GQO

[LYGG20] Q. Lu, J. Yao, H. Guan, and P. Gao. gQoS: a QoS-oriented GPU virtualization with adaptive capacity sharing. IEEE Transactions on Parallel and Distributed Systems, 31(4):843– 855, April 2020. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lee:2020:CEF

[LYK20] S. Lee, I. Yeom, and D. Kim. T-Caching: Enhancing feasibility of in-network caching in ICN. IEEE Transactions on Parallel and Distributed Systems, 31 (7):1486–1498, July 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2020:HHF

[LYL^{+}20a] R. Li, Q. Yang, Y. Li, X. Gu, W. Xiao, and K. Li. HeteroYARN: a heterogeneous FPGA-accelerated architecture based on YARN. IEEE Transactions on Parallel and Distributed Systems, 31(12):2968– 2980, December 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2020:EFC

[LYL^{+}20b] W. Li, X. Yuan, K. Li, H. Qi, X. Zhou, and R. Xu. Endpointflexible coflow scheduling across geo-distributed datacenters. IEEE Transactions on Parallel and Distributed Systems, 31 (10):2466–2481, October 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liao:2025:RAC

[LyLG^{+}25] Haoyu Liao, Tong yu Liu, Jianmei Guo, Bo Huang, Dingyu Yang, and Jonathan Ding. Retrospecting available CPU resources: SMT-aware scheduling to prevent SLA violations in data centers. IEEE Transactions on Parallel and Distributed Systems, 36(1):67–83, January 2025. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lyu:2020:TFP

[LYN^{+}20] L. Lyu, J. Yu, K. Nandakumar, Y. Li, X. Ma, J. Jin, H. Yu, and K. S. Ng. Towards fair and privacy-preserving federated deep models. IEEE Transactions on Parallel and Distributed Systems, 31(11):2524– 2541, November 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2022:AEL

[LYZC22] Kun Li, Liang Yuan, Yunquan Zhang, and Gongwei Chen. An accurate and efficient large-scale regression method through best friend clustering. IEEE Transactions on Parallel and Distributed Systems, 33(11):3129–3140, November 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2024:BJL

[LYZS24] Tiangang Li, Shi Ying, Yishi Zhao, and Jianga Shang. Batch jobs load balancing scheduling in cloud computing using distributional reinforcement learning. IEEE Transactions on Parallel and Distributed Systems, 35(1):169–185, January 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2024:DAP

[LZDO24] Dingding Li, Weijie Zhang, Mianxiong Dong, and Kaoru Ota. DMA-assisted I/O for persistent memory. IEEE Transactions on Parallel and Distributed Systems, 35(5):829– 843, May 2024. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2024:BDL

[LZF^{+}24] Zirui Liu, Yikai Zhao, Zhuochen Fan, Tong Yang, Xiaodong Li, Ruwen Zhang, Kaicheng Yang, Zihan Jiang, Zheng Zhong, Yi Huang, Cong Liu, Jing Hu, Gaogang Xie, and Bin Cui. BurstBalancer: Do less, better balance for large-scale data center traffic. IEEE Transactions on Parallel and Distributed Systems, 35(6):777– 794, June 2024. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2024:CER

[LZH^{+}24] Jiandong Liu, Lan Zhang, Fengxiang He, Chi Zhang, Shanyang Jiang, and Xiang-Yang Li. Communicationefficient regret-optimal distributed online convex optimization. IEEE Transactions on Parallel and Distributed Systems, 35(11):2270– 2283, November 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2020:OPS

[LZJ^{+}20] X. Li, R. Zhou, L. Jiao, C. Wu, Y. Deng, and Z. Li. Online placement and scaling of geo-distributed machine learning jobs via volume-discounting brokerage. IEEE Transactions on Parallel and Distributed Systems, 31(4):948–966, April 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2024:GLI

[LZL^{+}24] Jiesong Liu, Feng Zhang, Lv Lu, Chang Qi, Xiaoguang Guo, Dong Deng, Guoliang Li, Huanchen Zhang, Jidong Zhai, Hechen Zhang, Yuxing Chen, Anqun Pan, and Xiaoyong Du. G-Learned Index: Enabling efficient learned index on GPU. IEEE Transactions on Parallel and Distributed Systems, 35 (6):795–812, June 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lu:2020:GEV

[LZM^{+}20] Q. Lu, X. Zheng, J. Ma, Y. Dong, Z. Qi, J. Yao, B. He, and H. Guan. gMig:

Efficient vGPU live migration with overlapped softwarebased dirty page verification. IEEE Transactions on Parallel and Distributed Systems, 31 (5):1209–1222, May 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lin:2024:HHT

[LZP24] Yi-Chien Lin, Bingyi Zhang, and Viktor K. Prasanna. Hit-GNN: High-throughput GNN training framework on CPU + multi-FPGA heterogeneous platform. IEEE Transactions on Parallel and Distributed Systems, 35(5):707–719, May 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2023:FRP

[LZQ^{+}23] Fanxin Li, Shixiong Zhao, Yuhao Qing, Xusheng Chen, Xiuxian Guan, Sen Wang, Gong Zhang, and Heming Cui. Fold3D: Rethinking and parallelizing computational and communicational tasks in the training of large DNN models. IEEE Transactions on Parallel and Distributed Systems, 34 (5):1432–1449, May 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lee:2024:HTL

[LZS^{+}24] Wai-Kong Lee, Raymond K. Zhao, Ron Steinfeld, Amin Sakzad, and Seong Oun Hwang. High throughput lattice-based signatures on GPUs: Comparing Falcon and Mitaka. IEEE Transactions on Parallel and Distributed Systems, 35(4):675– 692, April 2024. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2022:PPE

[LZW^{+}22a] Anran Li, Lan Zhang, Junhao Wang, Feng Han, and Xiang-Yang Li. Privacy-preserving efficient federated-learning model debugging. IEEE Transactions on Parallel and Distributed Systems, 33(10):2291– 2303, October 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lu:2022:ODS

[LZW22b] Gangzhao Lu, Weizhe Zhang, and Zheng Wang. Optimizing depthwise separable convolution operations on GPUs. IEEE Transactions on Parallel and Distributed Systems, 33 (1):70–87, January 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lu:2023:ABC

[LZW^{+}23] Renhao Lu, Weizhe Zhang, Yan Wang, Qiong Li, Xiaoxiong Zhong, Hongwei Yang, and Desheng Wang. Auctionbased cluster federated learning in mobile edge computing systems. IEEE Transactions on Parallel and Distributed Systems, 34(4):1145–1158, April 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2022:SSG

[LZWL22] Jiajun Li, Hao Zheng, Ke Wang, and Ahmed Louri. SGCNAX: a scalable graph convolutional neural network accelerator with workload balancing. IEEE Transactions on Parallel and Distributed Systems, 33(11): 2834–2845, November 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2022:CFC

[LZWW22] Jianshu Liu, Shungeng Zhang, Qingyang Wang, and Jinpeng Wei. Coordinating fast concurrency adapting with autoscaling for SLO-oriented web applications. IEEE Transactions on Parallel and Distributed Systems, 33(12):3349–3362, December 2022. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2021:EMM

[LZX^{+}21] Q. Li, X. Zhang, J. Xiong, W.- M. Hwu, and D. Chen. Efficient methods for mapping neural machine translator on FPGAs. IEEE Transactions on Parallel and Distributed Systems, 32 (7):1866–1877, July 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2021:SSH

[LZZ21] X. Li, G. Zhang, and W. Zheng. SmartTuning: Selecting hyperparameters of a ConvNet system for fast training and small working memory. IEEE Transactions on Parallel and Distributed Systems, 32(7):1690– 1701, July 2021. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2023:TPR

[LZZ^{+}23] Yihong Li, Xiaoxi Zhang, Tianyu Zeng, Jingpu Duan, Chuan Wu, Di Wu, and Xu Chen. Task placement and resource allocation for edge machine learning: a GNNbased multi-agent reinforcement learning paradigm. IEEE Transactions on Parallel and Distributed Systems, 34(12): 3073–3089, December 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Mann:2022:DAP

[Man22] Zoltán Ádám Mann. Decentralized application placement in fog computing. IEEE Transactions on Parallel and Distributed Systems, 33(12):3262– 3273, December 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Mack:2022:PMO

[MAOA22] Joshua Mack, Samet E. Arda, Umit Y. Ogras, and Ali Akoglu. Performant, multi-objective scheduling of highly interleaved task graphs on heterogeneous system on chip devices. IEEE Transactions on Parallel and Distributed Systems, 33(9): 2148–2162, September 2022.

CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Merani:2021:RPA

[MCT21] M. L. Merani, D. Croce, and I. Tinnirello. Rings for privacy: an architecture for large scale privacy-preserving data mining. IEEE Transactions on Parallel and Distributed Systems, 32(6):1340–1352, June 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Mitzenmacher:2022:SMK

[MD22] Michael Mitzenmacher and Matteo Dell’Amico. The supermarket model with known and predicted service times. IEEE Transactions on Parallel and Distributed Systems, 33(11):2740–2751, November 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Mencagli:2024:SGA

[MDC24] Gabriele Mencagli, Patrizio Dazzi, and Massimo Coppola. Springald: GPU-accelerated window-based aggregates over out-of-order data streams. IEEE Transactions on Parallel and Distributed Systems, 35(9): 1657–1671, September 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Mukherjee:2022:TLD

[MDM22] Anandarup Mukherjee, Pallav Kumar Deb, and Sudip Misra. Timed loops for distributed storage in wireless networks. IEEE Transactions on Parallel and Distributed Systems, 33 (3):698–709, March 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Mei:2023:MAP

[Mei23] Zhonghui Mei. Minimizing the average packet access time of the application layer for buffered instantly decodable network coding. IEEE Transactions on Parallel and Distributed Systems, 34(3):1035– 1046, March 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Martinez-Ferrer:2022:NTV

[MFYB22] Pedro J. Martinez-Ferrer, A. N.

Yzelman, and Vicenç Beltran. A native tensor–vector multiplication algorithm for high performance computing. IEEE Transactions on Parallel and Distributed Systems, 33(12): 3363–3374, December 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Mishra:2024:FRR

[MGBD24] Rahul Mishra, Hari Prabhat Gupta, Garvit Banga, and Sajal K. Das. Fed-RAC: Resource-aware clustering for tackling heterogeneity of participants in federated learning.

IEEE Transactions on Parallel and Distributed Systems, 35 (7):1207–1220, July 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Maccari:2020:EDL

[MGG^{+}20] L. Maccari, L. Ghiro, A. Guerrieri, A. Montresor, and R. L. Cigno. Exact distributed load centrality computation: Algorithms, convergence, and applications to distance vector routing. IEEE Transactions on Parallel and Distributed Systems, 31(7):1693–1706, July 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ma:2024:TMA

[MHLM24] Renwen Ma, Kai Hwang, Mo Li, and Yiming Miao. Trusted model aggregation with zero-knowledge proofs in federated learning. IEEE Transactions on Parallel and Distributed Systems, 35(11): 2284–2296, November 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Mills:2022:MTF

[MHM22] Jed Mills, Jia Hu, and Geyong Min. Multi-task federated learning for personalised deep neural networks in edge computing. IEEE Transactions on Parallel and Distributed Systems, 33(3):630–641, March 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Mao:2021:PPC

[MHW^{+}21] Y. Mao, W. Hong, H. Wang, Q. Li, and S. Zhong. Privacypreserving computation offloading for parallel deep neural networks training. IEEE Transactions on Parallel and Distributed Systems, 32(7):1777– 1788, July 2021. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Mao:2022:SDN

[MHZ^{+}22] Yunlong Mao, Wenbo Hong, Boyu Zhu, Zhifei Zhu, Yuan Zhang, and Sheng Zhong. Secure deep neural network models publishing against membership inference attacks via training task parallelism. IEEE Transactions on Parallel and Distributed Systems, 33(11): 3079–3091, November 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Misra:2024:BTC

[MK24] Anshuman Misra and Ajay D.

Kshemkalyani. Byzantinetolerant causal ordering for unicasts, multicasts, and broadcasts. IEEE Transactions on Parallel and Distributed Systems, 35(5):814–828, May 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Moon:2022:ESA

[MKJ^{+}22] Gordon Euhyun Moon, Hyoukjun Kwon, Geonhwa Jeong, Prasanth Chatarasi, Sivasankaran Rajamanickam, and Tushar Krishna. Evaluating spatial accelerator architectures with tiled matrix–matrix multiplication. IEEE Transactions on Parallel and Distributed Systems, 33 (4):1002–1014, April 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Meng:2022:PHT

[MKKP22] Yuan Meng, Sanmukh Kuppannagari, Rajgopal Kannan, and Viktor Prasanna. PPOAccel: a high-throughput acceleration framework for proximal policy optimization. IEEE Transactions on Parallel and Distributed Systems, 33(9):2066– 2078, September 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Masiak:2021:CPN

[MKKS21] Marek Masiak, Iwona Kotlarska, Ukasz Kondraciuk, and Maciej Szpindler. Critique of planetary normal mode computation: Parallel algorithms, performance, and reproducibility by SCC Team From University of Warsaw. IEEE Transactions on Parallel and Distributed Systems, 32(11):2635– 2638, November 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). See [SLX^{+}21b].

Miwa:2021:PPA

[MLS21] S. Miwa, I. Laguna, and M. Schulz. PredCom: a predictive approach to collecting approximated communication traces. IEEE Transactions on Parallel and Distributed Systems, 32(1):45–58, January 2021. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ma:2020:TMN

[MLWX20] Y. Ma, W. Liang, J. Wu, and Z. Xu. Throughput maximization of NFV-enabled multicasting in mobile edge cloud networks. IEEE Transactions on Parallel and Distributed Systems, 31(2):393–407, February 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ma:2023:AED

[MLX23] Jialiang Ma, Li Li, and Chengzhong Xu. AutoRS: Environment-dependent realtime scheduling for end-to-end autonomous driving. IEEE Transactions on Parallel and Distributed Systems, 34(12): 3238–3252, December 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Mohan:2023:DFL

[MMGR23] N. Jagan Mohan, R. Murugan, Tripti Goel, and Parthapratim Roy. DRFL: Federated learning in diabetic retinopathy grading using fundus images. IEEE Transactions on Parallel and Distributed Systems, 34 (6):1789–1801, June 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Misra:2021:BEP

[MMR^{+}21] S. Misra, A. Mukherjee, A. Roy, N. Saurabh, Y. Rahulamathavan, and M. Rajarajan. Blockchain at the edge: Performance of resource-constrained IoT networks. IEEE Transactions on Parallel and Distributed Systems, 32(1):174– 183, January 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Mondal:2024:TRR

[MR24] Chiranjeb Mondal and Sanjay Rajopadhye. Taking RNA– RNA interaction to machine peak. IEEE Transactions on Parallel and Distributed Systems, 35(6):737–749, June 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Matha:2020:SWS

[MRFP20] R. Mathá, S. Ristov, T. Fahringer, and R. Prodan. Simplified workflow simulation on clouds based on computation and communication noisiness. IEEE Transactions on Parallel and Distributed Systems, 31(7): 1559–1574, July 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Moradi:2021:SSA

[MSSK21] N. Moradi, A. Shameli-Sendi, and A. Khajouei. A scalable stateful approach for virtual security functions orchestration. IEEE Transactions on Parallel and Distributed Systems, 32 (6):1383–1394, June 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Meng:2020:ODA

[MTL^{+}20] J. Meng, H. Tan, X. Li, Z. Han, and B. Li. Online deadlineaware task dispatching and scheduling in edge computing. IEEE Transactions on Parallel and Distributed Systems, 31 (6):1270–1286, June 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Merzky:2022:DPC

[MTT^{+}22] Andre Merzky, Matteo Turilli, Mikhail Titov, Aymen Al-Saadi, and Shantenu Jha. Design and performance characterization of RADICAL-Pilot on leadership-class platforms. IEEE Transactions on Parallel and Distributed Systems, 33 (4):818–829, April 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Mou:2024:SCE

[MWL24] Di Mou, Bo Wang, and Dajiang Liu. SC-CGRA: an energyefficient CGRA using stochastic computing. IEEE Transactions on Parallel and Distributed Systems, 35(11):2023– 2038, November 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Minhas:2022:EDM

[MWNK22] Umar Ibrahim Minhas, Roger Woods, Dimitrios S. Nikolopoulos, and Georgios Karakonstantis. Efficient, dynamic multitask execution on FPGA-based computing systems. IEEE Transactions on Parallel and Distributed Systems, 33(3):710– 722, March 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Masouros:2021:RRI

[MXS21] D. Masouros, S. Xydis, and D. Soudris. Rusty: Runtime interference-aware predictive monitoring for modern multi-tenant systems. IEEE Transactions on Parallel and Distributed Systems, 32(1):184– 198, January 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ma:2022:TRD

[MZC^{+}22a] Zhi Ma, Sheng Zhang, Zhiqi Chen, Tao Han, Zhuzhong Qian, Mingjun Xiao, Ning Chen, Jie Wu, and Sanglu Lu. Towards revenue-driven multiuser online task offloading in edge computing. IEEE Transactions on Parallel and Distributed Systems, 33(5):1185– 1198, May 2022. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Meng:2022:AGH

[MZC^{+}22b] Jintao Meng, Chen Zhuang, Peng Chen, Mohamed Wahib, Bertil Schmidt, Xiao Wang, Haidong Lan, Dou Wu, Minwen Deng, Yanjie Wei, and Shengzhong Feng. Automatic generation of high-performance convolution kernels on ARM CPUs for deep learning. IEEE Transactions on Parallel and Distributed Systems, 33(11): 2885–2899, November 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ma:2022:DDO

[MZW22] Meng Ma, Jingbin Zhang, and Ping Wang. DePo: Dynamically offload expensive event processing to the edge of cyberphysical systems. IEEE Transactions on Parallel and Distributed Systems, 33(9):2120– 2132, September 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Meng:2021:CBC

[MZWX21] T. Meng, Y. Zhao, K. Wolter, and C.-Z. Xu. On consortium blockchain consistency: a queueing network model approach. IEEE Transactions on Parallel and Distributed Systems, 32(6):1369–1382, June 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Nakatani:2021:SAB

[Nak21]  Y. Nakatani.  Structured

allocation-based consistent hashing with improved balancing for cloud infrastructure. IEEE Transactions on Parallel and Distributed Systems, 32(9): 2248–2261, September 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Nguyen:2020:BBF

[NAL^{+}20] M. Nguyen, S. Alesawi, N. Li, H. Che, and H. Jiang. A blackbox fork–join latency prediction model for data-intensive applications. IEEE Transactions on Parallel and Distributed Systems, 31(9):1983– 2000, September 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Navarro:2021:GTC

[NCB^{+}21] C. A. Navarro, R. Carrasco, R. J. Barrientos, J. A. Riquelme, and R. Vega. GPU tensor cores for fast arithmetic reductions. IEEE Transactions on Parallel and Distributed Systems, 32(1):72–84, January 2021. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ning:2021:DDS

[NDW^{+}21] Z. Ning, P. Dong, X. Wang, S. Wang, X. Hu, S. Guo, T. Qiu, B. Hu, and R. Y. K. Kwok. Distributed and dynamic service placement in pervasive edge computing networks. IEEE Transactions on Parallel and Distributed Systems, 32(6):1277–1292, June 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Navarro:2020:BAD

[NFP^{+}20] C. Navarro, J. Feliu, S. Petit, M. E. Gómez, and J. Sahuquillo. Bandwidthaware dynamic prefetch configuration for IBM POWER8. IEEE Transactions on Parallel and Distributed Systems, 31(8): 1970–1982, August 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Nobre:2021:RTA

[NISJS21] R. Nobre, A. Ilic, S. Santander-Jiménez, and L. Sousa. Retargeting tensor accelerators for epistasis detection. IEEE Transactions on Parallel and Distributed Systems, 32(9): 2160–2174, September 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Niu:2022:PRM

[NJG^{+}22] Yipei Niu, Panpan Jin, Jian Guo, Yikai Xiao, Rong Shi, Fangming Liu, Chen Qian, and Yang Wang. PostMan: Rapidly mitigating bursty traffic via on-demand offloading of packet processing. IEEE Transactions on Parallel and Distributed Systems, 33(2):374– 387, February 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Nine:2021:TPD

[NK21] M. S. Q. Z. Nine and T. Kosar. A two-phase dynamic throughput optimization model for big data transfers. IEEE Transactions on Parallel and Distributed Systems, 32(2):269– 280, February 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Navaridas:2024:UIA

[NKP^{+}24] Javier Navaridas, Markos Kynigos, Jose A. Pascual, Mikel Luján, Jose Miguel-Alonso, and John Goodacre. Understanding the impact of arbitration in MZI-based Beneš switching fabrics. IEEE Transactions on Parallel and Distributed Systems, 35(2):338–348, February 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ng:2022:RAH

[NLX^{+}22] Jer Shyuan Ng, Wei Yang Bryan Lim, Zehui Xiong, Xianbin Cao, Jiangming Jin, Dusit Niyato, Cyril Leung, and Chunyan Miao. Reputation-aware hedonic coalition formation for efficient serverless hierarchical federated learning. IEEE Transactions on Parallel and Distributed Systems, 33(11):2675– 2686, November 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Neiheiser:2020:FER

[NRB^{+}20] R. Neiheiser, L. Rech, M. Bravo, L. Rodrigues, and M. Correia. Fireplug: Efficient and robust geo-replication of graph databases. IEEE Transactions on Parallel and Distributed Systems, 31(8):1942–1953, August 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Olliaro:2025:ISD

[OAM^{+}25] Diletta Olliaro, Adityo Anggraito, Marco Ajmone Marsan, Simonetta Balsamo, and Andrea Marin. The impact of service demand variability on data center performance. IEEE Transactions on Parallel and Distributed Systems, 36(2):120– 132, February 2025. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ouyang:2021:APA

[OHWL21] L. Ouyang, Y. Huang, H. Wei, and J. Lu. Achieving probabilistic atomicity with wellbounded staleness and low read latency in distributed datastores. IEEE Transactions on Parallel and Distributed Systems, 32(4):815–829, April 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Oyama:2021:CSS

[OMD^{+}21] Y. Oyama, N. Maruyama, N. Dryden, E. McCarthy, P. Harrington, J. Balewski, S. Matsuoka, P. Nugent, and B. Van Essen. The case for strong scaling in deep learning: Training large 3D CNNs with hybrid parallelism. IEEE Transactions on Parallel and Distributed Systems, 32(7): 1641–1652, July 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Omland:2023:HHD

[ONP^{+}23] Patrik Omland, Alessio Netti, Yang Peng, Andrea Baldovin, Michael Paulitsch, Gustavo Espinosa, Jorge Parra, Gereon Hinz, and Alois Knoll. HPC hardware design reliability benchmarking with HDFIT. IEEE Transactions on Parallel and Distributed Systems, 34 (3):995–1006, March 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Oh:2020:WPD

[OQCW20] K. Oh, N. Qin, A. Chandra, and J. Weissman. Wiera: Policy-driven multi-tiered geodistributed cloud storage system. IEEE Transactions on Parallel and Distributed Systems, 31(2):294–305, February 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Orr:2020:ITD

[OS20] M. Orr and O. Sinnen. Integrating task duplication in optimal task scheduling with communication delays. IEEE Transactions on Parallel and Distributed Systems, 31(10): 2277–2288, October 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Odiathevar:2022:BAD

[OSF22] Murugaraj Odiathevar, Winston K. G. Seah, and Marcus Frean. A Bayesian approach to distributed anomaly detection in edge AI networks. IEEE Transactions on Parallel and Distributed Systems, 33(12):3306–3320, December 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Oh:2022:NCA

[OZCW22] Kwangsung Oh, Minmin Zhang, Abhishek Chandra, and Jon Weissman. Network cost-aware geo-distributed data analytics system. IEEE Transactions on Parallel and Distributed Systems, 33(6):1407–1420, June 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Parashar:2020:EN

[Par20] M. Parashar. Editor’s note.

IEEE Transactions on Parallel and Distributed Systems, 31(2): 251–252, February 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Parashar:2021:EN

[Par21a] M. Parashar. Editor’s note.

IEEE Transactions on Parallel and Distributed Systems, 32 (4):743–745, April 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Parashar:2021:GES

[Par21b] Manish Parashar. Guest editorial: Special section on SC19 Student Cluster Competition. IEEE Transactions on Parallel and Distributed Systems, 32 (11):2606, November 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Parashar:2022:EEA

[Par22] Manish Parashar. EiC editorial: Advancing reproducibility in parallel and distributed systems research. IEEE Transactions on Parallel and Distributed Systems, 33(9):2010, September 2022. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Peng:2021:DDL

[PBC^{+}21] Y. Peng, Y. Bao, Y. Chen, C. Wu, C. Meng, and W. Lin. DL2: a deep learning-driven scheduler for deep learning clusters. IEEE Transactions on Parallel and Distributed Systems, 32(8):1947–1960, August 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Peng:2022:ESE

[PGY^{+}22] Ivy B. Peng, Maya B. Gokhale, Karim Youssef, Keita Iwabuchi, and Roger Pearce. Enabling scalable and extensible memory-mapped datastores in userspace. IEEE Transactions on Parallel and Distributed Systems, 33(4):866– 877, April 2022. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Plale:2021:TRP

[PH21] Beth Plale and Stephen Lien Harrell. Transparency and reproducibility practice in largescale computational science: a preface to the special section. IEEE Transactions on Parallel and Distributed Systems, 32(11):2607–2608, November 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Peng:2020:LFP

[PHY20] Y. Peng, Z. Hao, and X. Yun.

Lock-free parallelization for variance-reduced stochastic gradient descent on streaming data. IEEE Transactions on Parallel and Distributed Systems, 31(9):2220– 2231, September 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Pilla:2023:SAF

[Pil23] Laércio Lima Pilla. Scheduling algorithms for federated learning with minimal energy consumption. IEEE Transactions on Parallel and Distributed Systems, 34(4):1215– 1226, April 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Pozzetti:2021:REV

[PK21] T. Pozzetti and A. D. Kshemkalyani. Resettable encoded vector clock for causality analysis with an application to dynamic race detection. IEEE Transactions on Parallel and Distributed Systems, 32(4):772–785, April 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Prindle:2022:CMM

[PKJ^{+}22] Nicole Prindle, Ali Kazmi, Aman Jain, Albert Chen, Marissa Sorkin, Sudhanshu Agarwal, Richard Vuduc, and Vijay Thakkar. Critique of MemXCT: Memory-centric Xray CT reconstruction with massive parallelization by SCC Team from Georgia Tech. IEEE Transactions on Parallel and Distributed Systems, 33(9): 2035–2038, September 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Park:2023:NSO

[PKRS23] Chanho Park, Bogil Kim, Sungmin Ryu, and William J. Song. NeuroSpector: Systematic optimization of dataflow scheduling in DNN accelerators. IEEE Transactions on Parallel and Distributed Systems, 34(8):2279–2294, August 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Park:2022:DFF

[PLJK22] Pyeongsu Park, Jaewoon Lee, Heetaek Jeong, and Jangwoo Kim. DLS: a fast and flexible neural network training system with fine-grained heterogeneous device orchestration. IEEE Transactions on Parallel and Distributed Systems, 33(11):3194–3206, November 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Papaioannou:2022:ARP

[PM22] Antonis Papaioannou and Kostas Magoutis. Addressing the read-performance impact of reconfigurations in replicated key–value stores. IEEE Transactions on Parallel and Distributed Systems, 33(9):2106– 2119, September 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Paredes:2020:EPV

[PRL20] M. Paredes, G. Riley, and M. Luján. Exploiting parallelism and vectorisation in breadth-first search for the Intel Xeon Phi. IEEE Transactions on Parallel and Distributed Systems, 31(1):111– 128, January 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Papadopoulos:2022:EFH

[PSK^{+}22] Lazaros Papadopoulos, Dimitrios Soudris, Christoph Kessler, August Ernstsson, Johan Ahlqvist, Nikos Vasilas, Athanasios I. Papadopoulos, Panos Seferlis, Charles Prouveur, Matthieu Haefele, Samuel Thibault, Athanasios Salamanis, Theodoros Ioakimidis, and Dionysios Kehagias. EXA2PRO: a framework for high development productivity on heterogeneous computing systems. IEEE Transactions on Parallel and Distributed Systems, 33(4):792– 804, April 2022. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Pons:2020:PAC

[PSS^{+}20] L. Pons, J. Sahuquillo, V. Selfa, S. Petit, and J. Pons. Phaseaware cache partitioning to target both turnaround time and system performance. IEEE Transactions on Parallel and Distributed Systems, 31(11): 2556–2568, November 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Pei:2023:LNT

[PWX^{+}23] Shuyu Pei, Jigang Wen, Kun Xie, Gaogang Xie, and Kenli Li. On-line network traffic anomaly detection based on tensor sketch. IEEE Transactions on Parallel and Distributed Systems, 34(12):3028– 3045, December 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Pandey:2021:TTC

[PWZ^{+}21] Santosh Pandey, Zhibin Wang, Sheng Zhong, Chen Tian, Bolong Zheng, Xiaoye Li, Lingda Li, Adolfy Hoisie, Caiwen Ding, Dong Li, and Hang Liu. Trust: Triangle counting reloaded on GPUs. IEEE Transactions on Parallel and Distributed Systems, 32(11): 2646–2660, November 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Pan:2022:GSG

[PZL^{+}22] Zaifeng Pan, Feng Zhang, Hourun Li, Chenyang Zhang, Xiaoyong Du, and Dong Deng. G-SLIDE: a GPU-based sublinear deep learning engine via LSH sparsification. IEEE Transactions on Parallel and Distributed Systems, 33(11): 3015–3027, November 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Pan:2022:EDA

[PZZ^{+}22] Zaifeng Pan, Feng Zhang, Yanliang Zhou, Jidong Zhai, Xipeng Shen, Onur Mutlu, and Xiaoyong Du. Exploring data analytics without decompression on embedded GPU systems. IEEE Transactions on Parallel and Distributed Systems, 33(7):1553–1568, July 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Qin:2020:ECI

[QHC20] X. Qin, R. Hao, and J. Chang.

The existence of completely independent spanning trees for some compound graphs. IEEE Transactions on Parallel and Distributed Systems, 31(1):201– 210, January 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Qi:2024:PSM

[QJC^{+}24] Sheng Qi, Chao Jin, Mosharaf Chowdhury, Zhenming Liu, Xuanzhe Liu, and Xin Jin. Pyxis: Scheduling mixed tasks in disaggregated datacenters. IEEE Transactions on Parallel and Distributed Systems, 35(9): 1536–1550, September 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Qiu:2023:TSF

[QJZF23] Yudi Qiu, Jie Jiao, Xiaoyang Zeng, and Yibo Fan. Tagsharer-fusion directory: a scalable coherence directory with flexible entry formats. IEEE Transactions on Parallel and Distributed Systems, 34(1):262– 274, January 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Qu:2023:EES

[QLP^{+}23] Peng Qu, Hui Lin, Meng Pang, Xiaofei Liu, Weimin Zheng, and Youhui Zhang. ENLARGE: an efficient SNN simulation framework on GPU clusters. IEEE Transactions on Parallel and Distributed Systems, 34(9): 2529–2540, September 2023.

CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Qian:2024:EBP

[QQD^{+}24] Junyan Qian, Kunzhu Qiu, Hao Ding, Huimin Zhang, and Zhongyi Zhai. An efficient bottleneck planes exclusion method for reconfiguring 3D VLSI arrays. IEEE Transactions on Parallel and Distributed Systems, 35(2):250– 263, February 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Quraishi:2021:SSA

[QTR21] M. H. Quraishi, E. B. Tavakoli, and F. Ren. A survey of system architectures and techniques for FPGA virtualization. IEEE Transactions on Parallel and Distributed Systems, 32(9): 2216–2230, September 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Qu:2021:PFL

[QWHC21] X. Qu, S. Wang, Q. Hu, and X. Cheng. Proof of federated learning: a novel energyrecycling consensus algorithm. IEEE Transactions on Parallel and Distributed Systems, 32(8): 2074–2085, August 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Quan:2020:TSE

[QWYG20] Z. Quan, Z. Wang, T. Ye, and S. Guo. Task scheduling for energy consumption constrained parallel applications on heterogeneous computing systems. IEEE Transactions on Parallel and Distributed Systems, 31 (5):1165–1182, May 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Qiu:2020:FKF

[QXL^{+}20] Y. Qiu, J. Xie, H. Lv, W. Yin, W. Luk, L. Wang, B. Yu, H. Chen, X. Ge, Z. Liao, and X. Shi. FULL–KV: Flexible and ultra-low-latency inmemory key–value store system design on CPU-FPGA. IEEE Transactions on Parallel and Distributed Systems, 31(8): 1828–1444, August 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Qi:2024:ELA

[QXL^{+}24] Jiaxing Qi, Wencong Xiao, Mingzhen Li, Chaojie Yang, Yong Li, Wei Lin, Hailong Yang, Zhongzhi Luan, and Depei Qian. ElasticBatch: a learning-augmented elastic scheduling system for batch inference on MIG. IEEE Transactions on Parallel and Distributed Systems, 35(10):1708– 1720, October 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Qiu:2021:DCD

[QZCZ21] X. Qiu, W. Zhang, W. Chen, and Z. Zheng. Distributed and collective deep reinforcement learning for computation offloading: a practical perspective. IEEE Transactions on Parallel and Distributed Systems, 32(5):1085–1101, May 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Qu:2020:HPS

[QZFZ20] P. Qu, Y. Zhang, X. Fei, and W. Zheng. High performance simulation of spiking neural network on GPGPUs. IEEE Transactions on Parallel and Distributed Systems, 31(11): 2510–2523, November 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Rorabaugh:2022:BHT

[RCLJT22] Ariel Keller Rorabaugh, Silvina Caíno-Lores, Travis Johnston, and Michela Taufer. Building high-throughput neural architecture search workflows via a decoupled fitness prediction engine. IEEE Transactions on Parallel and Distributed Systems, 33(11):2913– 2926, November 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Rui:2023:SOM

[RCW^{+}23] Lanlan Rui, Shiyou Chen, Shuyun Wang, Zhipeng Gao, Xuesong Qiu, Wenjing Li, and Shaoyong Guo. SFC orchestration method for edge cloud and central cloud collaboration: QoS and energy consumption joint optimization combined with reputation assessment. IEEE Transactions on Parallel and Distributed Systems, 34(10):2735–2748, October 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Rabinia:2025:ADS

[RDBG25] Sanaz Rabinia, Niloofar Didar, Marco Brocanelli, and Daniel Grosu. Algorithms for data sharing-aware task allocation in edge computing systems. IEEE Transactions on Parallel and Distributed Systems, 36(1):15– 28, January 2025. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Raparti:2020:ANM

[RP20] V. Y. Raparti and S. Pasricha. Approximate NoC and memory controller architectures for GPGPU accelerators. IEEE Transactions on Parallel and Distributed Systems, 31(5):25– 39, May 2020. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Raffin:2025:DSB

[RT25] Guillaume Raffin and Denis Trystram. Dissecting the software-based measurement of CPU energy consumption: a comparative analysis. IEEE Transactions on Parallel and Distributed Systems, 36(1):96– 107, January 2025. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ren:2020:EAD

[RXL^{+}20] H. Ren, Z. Xu, W. Liang, Q. Xia, P. Zhou, O. F. Rana, A. Galis, and G. Wu. Efficient algorithms for delayaware NFV-Enabled multicasting in mobile edge clouds with resource sharing. IEEE Transactions on Parallel and Distributed Systems, 31(9):2050– 2066, September 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ren:2020:IJS

[RZLT20] R. Ren, Y. Zhu, C. Li, and X. Tang. Interval job scheduling with machine launch cost. IEEE Transactions on Parallel and Distributed Systems, 31(12):2776–2788, December 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Singh:2024:SFW

[SBM24] Ajay Singh, Trevor Alexander Brown, and Ali José Mashtizadeh. Simple, fast and widely applicable concurrent memory reclamation via neutralization. IEEE Transactions on Parallel and Distributed Systems, 35(2): 203–220, February 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Srivastava:2023:PFC

[SCA23] Ankit Srivastava, Sriram P.

Chockalingam, and Srinivas Aluru. A parallel framework for constraint-based Bayesian network learning via Markov blanket discovery. IEEE Transactions on Parallel and Distributed Systems, 34(6):1699– 1715, June 2023. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). See critiques [LCZ^{+}23, SGH^{+}23, CRZ^{+}23, GGL^{+}23].

Shi:2021:MWM

[SCL21a] S. Shi, X. Chu, and B. Li. MG-WFBP: Merging gradients wisely for efficient communication in distributed deep learning. IEEE Transactions on Parallel and Distributed Systems, 32(8):1903–1917, August 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sun:2021:CPN

[SCL^{+}21b] Wei-Fang Sun, Hung-Hsin Chen, Shao-Fu Lin, Yuan-Ching Lin, Jing-Wei Wu, En-Te Lin, and Jerry Chou. Critique of planetary normal mode computation: Parallel algorithms, performance, and reproducibility by SCC Team From National Tsing Hua University. IEEE Transactions on Parallel and Distributed Systems, 32(11):2623–2626, November 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). See [SLX^{+}21b].

Sun:2023:PDC

[SCS^{+}23] Hui Sun, Yiru Chen, Kewei Sha, Shaoyuan Huang, Xiaofei Wang, and Weisong Shi.

A proactive on-demand content placement strategy in edge intelligent gateways. IEEE Transactions on Parallel and Distributed Systems, 34(7): 2072–2090, July 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sahni:2021:MHM

[SCYJ21] Y. Sahni, J. Cao, L. Yang, and Y. Ji. Multi-hop multi-task partial computation offloading in collaborative edge computing. IEEE Transactions on Parallel and Distributed Systems, 32 (5):1133–1145, May 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Staffolani:2023:PRW

[SDBM23] Alessandro Staffolani, Victor-Alexandru Darvariu, Paolo Bellavista, and Mirco Musolesi. RLQ: Workload allocation with reinforcement learning in distributed queues. IEEE Transactions on Parallel and Distributed Systems, 34(3):856– 868, March 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sun:2021:CAW

[SDHQ21] H. Sun, S. Dai, J. Huang, and X. Qin. Co-Active: a workloadaware collaborative cache management scheme for NVMe SSDs. IEEE Transactions on Parallel and Distributed Systems, 32(6):1437–1451, June 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sunny:2024:CCR

[SDMC24] Chilankamol Sunny, Satyajit Das, Kevin J. M. Martin, and Philippe Coussy. CREPE: Concurrent reverse-modulo-scheduling and placement for CGRAs. IEEE Transactions on Parallel and Distributed Systems, 35(7):1293–1306, July 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Schwarzrock:2021:RNI

[SdR^{+}21] J. Schwarzrock, C. C. de Oliveira, M. Ritt, A. F. Lorenzon, and A. C. S. Beck. A runtime and non-intrusive approach to optimize EDP by tuning threads and CPU frequency for OpenMP applications. IEEE Transactions on Parallel and Distributed Systems, 32(7):1713–1724, July 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Spirovska:2021:OCC

[SDZ21] K. Spirovska, D. Didona, and W. Zwaenepoel. Optimistic causal consistency for georeplicated key–value stores. IEEE Transactions on Parallel and Distributed Systems, 32 (3):527–542, March 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Shi:2024:AQA

[SFW^{+}24] Jiuchen Shi, Kaihua Fu, Jiawen Wang, Quan Chen, Deze Zeng, and Minyi Guo. Adaptive QoS-Aware microservice deployment with excessive loads via intra- and inter-datacenter scheduling. IEEE Transactions on Parallel and Distributed Systems, 35(9):1565– 1582, September 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Shayan:2021:BBS

[SFYB21] M. Shayan, C. Fung, C. J. M.

Yoon, and I. Beschastnikh. Biscotti: a blockchain system for private and secure federated learning. IEEE Transactions on Parallel and Distributed Systems, 32(7):1513– 1525, July 2021. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Si:2023:CPF

[SGH^{+}23] Jiaqi Si, Junyi Guo, Zhewen Hao, Wenyang He, Ruihan Li, Yueyang Pan, Zhenxin Fu, and Chun Fan. Critique of A Parallel Framework for Constraint-Based Bayesian Network Learning via Markov Blanket Discovery by SCC Team From Peking University. IEEE Transactions on Parallel and Distributed Systems, 34 (6):1720–1722, June 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). See [SCA23].

Shah:2020:AAM

[SGJ^{+}20] A. Shah, R. Ganesan, S. Jajodia, P. Samarati, and H. Cam. Adaptive alert management for balancing optimal performance among distributed CSOCs using reinforcement learning. IEEE Transactions on Parallel and Distributed Systems, 31 (1):16–33, January 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Soumagne:2022:AHE

[SHC^{+}22] Jerome Soumagne, Jordan Henderson, Mohamad Chaarawi, Neil Fortner, Scot Breitenfeld, Songyu Lu, Dana Robinson, Elena Pourmal, and Johann Lombardi. Accelerating HDF5 I/O for exascale using DAOS. IEEE Transactions on Parallel and Distributed Systems, 33 (4):903–914, April 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Song:2023:FTL

[SHZ^{+}23] Mingyang Song, Zhongyun Hua, Yifeng Zheng, Tao Xiang, and Xiaohua Jia. FCDedup: a two-level deduplication system for encrypted data in fog computing. IEEE Transactions on Parallel and Distributed Systems, 34(10):2642–2656, October 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sun:2020:DEE

[SJLN20] Y. Sun, G. Jiang, S. Lam, and F. Ning. Designing energyefficient MPSoC with untrustworthy 3PIP cores. IEEE Transactions on Parallel and Distributed Systems, 31(1):51– 63, January 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sim:2020:IIS

[SKV^{+}20] H. Sim, A. Khan, S. S. Vazhkudai, S. Lim, A. R. Butt, and Y. Kim. An integrated indexing and search service for distributed file systems. IEEE Transactions on Parallel and Distributed Systems, 31(10): 2375–2391, October 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

ski:2022:MES

[sKW22] Maciej Kokoci ski, Tadeusz Kobus, and Pawe T. Wojciechowski. On mixing eventual and strong consistency: Acute cloud types. IEEE Transactions on Parallel and Distributed Systems, 33(6):1338– 1356, June 2022. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Saquib:2023:RVD

[SKW23] Nazmus Saquib, Chandra Krintz, and Rich Wolski. Replicated versioned data structures for wide-area distributed systems. IEEE Transactions on Parallel and Distributed Systems, 34(1):207–224, January 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Shen:2020:CRA

[SL20] Z. Shen and P. P. C. Lee. Crossrack-aware updates in erasurecoded data centers: Design and evaluation. IEEE Transactions on Parallel and Distributed Systems, 31(10):2315– 2328, October 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sun:2023:DTC

[SLG^{+}23] Wei Sun, Ang Li, Tong Geng, Sander Stuijk, and Henk Corporaal. Dissecting tensor cores via microbenchmarks: Latency, throughput and numeric behaviors. IEEE Transactions on Parallel and Distributed Systems, 34(1):246–261, January 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Song:2023:IFD

[SLHH23] Jiankang Song, Limei Lin, Yanze Huang, and Sun-Yuan Hsieh. Intermittent fault diagnosis of split-star networks and its applications. IEEE Transactions on Parallel and Distributed Systems, 34(4):1253– 1264, April 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Shivakumar:2023:SSF

[SLKA23] Shruti Shivakumar, Jiajia Li, Ramakrishnan Kannan, and Srinivas Aluru. Sparse symmetric format for Tucker decomposition. IEEE Transactions on Parallel and Distributed Systems, 34(6):1743– 1756, June 2023. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Shu:2020:TUW

[SLLL20] J. Shu, F. Li, S. Li, and Y. Lu. Towards unaligned writes optimization in cloud storage with high-performance SSDs. IEEE Transactions on Parallel and Distributed Systems, 31(12):2923–2937, December 2020. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Shen:2020:EFE

[SLX20] M. Shen, G. Luo, and N. Xiao.

EEPC: a framework for energyefficient parallel control of connected cars. IEEE Transactions on Parallel and Distributed Systems, 31(1):64–79, January 2020. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Shen:2021:CGP

[SLX21a] M. Shen, G. Luo, and N. Xiao.

Coarse-grained parallel routing with recursive partitioning for FPGAs. IEEE Transactions on Parallel and Distributed Systems, 32(4):884– 899, April 2021. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Shi:2021:PNM

[SLX^{+}21b] Jia Shi, Ruipeng Li, Yuanzhe Xi, Yousef Saad, and Maarten V. de Hoop. Planetary normal mode computation: Parallel algorithms, performance, and reproducibility. IEEE Transactions on Parallel and Distributed Systems, 32(11):2609– 2622, November 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). See critiques [SCL^{+}21b, BK21, ZZH^{+}21, MKKS21, LCG^{+}21, CFM^{+}21b].

Song:2023:PCV

[SLY^{+}23] Zhuoran Song, Wanzhen Liu, Tao Yang, Fangxin Liu, Naifeng Jing, and Xiaoyao Liang. A point cloud video recognition acceleration framework based on tempo-spatial information. IEEE Transactions on Parallel and Distributed Systems, 34(12):3224– 3237, December 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sun:2024:AAT

[SLY^{+}24] Qingxiao Sun, Yi Liu, Hailong Yang, Zhonghui Jiang, Zhongzhi Luan, and Depei Qian. Adaptive auto-tuning framework for global exploration of stencil optimization on GPUs. IEEE Transactions on Parallel and Distributed Systems, 35(1):20–33, January 2024. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Shi:2021:CSD

[SLZR21] Li Shi, Yang Liu, Junwei Zhang, and Thomas Robertazzi. Coflow scheduling in data centers: Routing and bandwidth allocation. IEEE Transactions on Parallel and Distributed Systems, 32(11):2661– 2675, November 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Shi:2020:LAB

[SMCH20] T. Shi, H. Ma, G. Chen, and S. Hartmann. Locationaware and budget-constrained service deployment for composite applications in multi-cloud environment. IEEE Transactions on Parallel and Distributed Systems, 31(8):1954– 1969, August 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Samani:2023:IMR

[SMK^{+}23] Zahra Najafabadi Samani, Narges Mehran, Dragi Kimovski, Shajulin Benedict, Nishant Saurabh, and Radu Prodan. Incremental multilayer resource partitioning for application placement in dynamic fog. IEEE Transactions on Parallel and Distributed Systems, 34 (6):1877–1896, June 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Suthakar:2021:OLA

[SMSK21] U. Suthakar, L. Magnoni, D. R.

Smith, and A. Khan. Optimised lambda architecture for monitoring scientific infrastructure. IEEE Transactions on Parallel and Distributed Systems, 32(6):1395–1408, June 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Singh:2020:WMW

[SNK20] V. K. Singh, B. Nathani, and M. Kumar. WEED-MC: Wavelet transform for energy efficient data gathering and matrix completion. IEEE Transactions on Parallel and Distributed Systems, 31(5):1066– 1073, May 2020. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Shekofteh:2020:CEC

[SNN^{+}20] S.-Kazen Shekofteh, Hamid Noori, Mahmoud Naghibzadeh, Holger Fröning, and Hadi Sadog Yazdi. cCUDA: Effective coscheduling of concurrent kernels on GPUs. IEEE Transactions on Parallel and Distributed Systems, 31(4):766– 778, April 2020. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sudo:2020:TOL

[SOI^{+}20] Y. Sudo, F. Ooshita, T. Izumi, H. Kakugawa, and T. Masuzawa. Time-optimal leader election in population protocols. IEEE Transactions on Parallel and Distributed Systems, 31(11):2620–2632, November 2020. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Shibata:2020:SPV

[SP20] Naoki Shibata and Francesco Petrogalli. SLEEF: a portable vectorized library of C standard mathematical functions. IEEE Transactions on Parallel and Distributed Systems, 31 (6):1316–1327, June 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Shaffer:2023:LCD

[SPCT23] Tim Shaffer, Thanh Son Phung, Kyle Chard, and Douglas Thain. Landlord: Coordinating dynamic software environments to reduce container sprawl. IEEE Transactions on Parallel and Distributed Systems, 34(5):1376–1389, May 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Siavoshani:2020:CLB

[SPPS20] M. J. Siavoshani, F. Parvaresh, A. Pourmiri, and S. P. Shariatpanahi. Coded load balancing in cache networks. IEEE Transactions on Parallel and Distributed Systems, 31(2):347– 358, February 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Song:2024:ILL

[SPS^{+}24] Enge Song, Tian Pan, Haoyu Song, Qiang Fu, Yingjiang Liu, Chenhao Jia, Chuanying Yuan, Minglan Gao, Jiao Zhang, Tao Huang, and Yunjie Liu. INT-Label: Lightweight in-band network-wide telemetry via distributed labeling. IEEE Transactions on Parallel and Distributed Systems, 35 (5):751–767, May 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sidik:2020:PTB

[SPZE20] B. Sidik, R. Puzis, P. Zilberman, and Y. Elovici. PALE: Time bounded practical agile leader election. IEEE Transactions on Parallel and Distributed Systems, 31(2):470– 485, February 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Saad:2021:PMP

[SQR^{+}21] M. Saad, Z. Qin, K. Ren, D. Nyang, and D. Mohaisen. e-PoS: Making proof-of-stake decentralized and fair. IEEE Transactions on Parallel and Distributed Systems, 32(8): 1961–1973, August 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Slota:2020:SMC

[SRD^{+}20] G. M. Slota, C. Root, K. Devine, K. Madduri, and S. Rajamanickam. Scalable, multiconstraint, complex-objective graph partitioning. IEEE Transactions on Parallel and Distributed Systems, 31(12): 2789–2801, December 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Shahrouz:2021:GGA

[SSH21] S. Shahrouz, S. Salehkaleybar, and M. Hashemi. gIM: GPU accelerated RIS-based influence maximization algorithm. IEEE Transactions on Parallel and Distributed Systems, 32(10): 2386–2399, October 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Singh:2021:QAT

[SSKG21] Ashutosh Kumar Singh, Deepika Saxena, Jitendra Kumar, and Vrinda Gupta. A quantum approach towards the adaptive prediction of cloud workloads. IEEE Transactions on Parallel and Distributed Systems, 32(12):2893–2905, December 2021. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Skrzypczak:2020:RFT

[SSS20] J. Skrzypczak, F. Schintke, and T. Schütt. RMWPaxos: Faulttolerant in-place consensus sequences. IEEE Transactions on Parallel and Distributed Systems, 31(10):2392–2405, October 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Schneider:2020:CPP

[ST20] T. Schneider and A. Treiber. A comment on privacy-preserving scalar product protocols as proposed in SPOC. IEEE Transactions on Parallel and Distributed Systems, 31(3):543– 546, March 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Shen:2023:IRR

[SWGB23] Haiying Shen, Haoyu Wang, Jiechao Gao, and Rajkumar Buyya. An instability-resilient renewable energy allocation system for a cloud datacenter. IEEE Transactions on Parallel and Distributed Systems, 34(3): 1020–1034, March 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Szustak:2020:CPO

[SWOM20] L. Szustak, R. Wyrzykowski, T. Olas, and V. Mele. Correlation of performance optimizations and energy consumption for stencil-based application on Intel Xeon scalable processors. IEEE Transactions on Parallel and Distributed Systems, 31(11):2582– 2593, November 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Shen:2024:HTG

[SYD^{+}24] Shiyu Shen, Hao Yang, Wangchen Dai, Hong Zhang, Zhe Liu, and Yunlei Zhao. Highthroughput GPU implementation of Dilithium post-quantum digital signature. IEEE Transactions on Parallel and Distributed Systems, 35(11):1964– 1976, November 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sun:2022:LCC

[SYS^{+}22] Jingwei Sun, Tao Yan, Hao Sun, Huancheng Lin, and Guangzhong Sun. Lossy compression of communication traces using recurrent neural networks. IEEE Transactions on Parallel and Distributed Systems, 33(11):3106– 3116, November 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Srinuan:2020:CME

[SYT20] P. Srinuan, X. Yuan, and N. Tzeng. Cooperative memory expansion via OS kernel support for networked computing systems. IEEE Transactions on Parallel and Distributed Systems, 31(11):2650– 2667, November 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Suyyagh:2020:ETA

[SZ20] A. Suyyagh and Z. Zilic. Energy and task-aware partitioning on single-ISA clustered heterogeneous processors. IEEE Transactions on Parallel and Distributed Systems, 31(2):306– 317, February 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Su:2023:MTS

[SZCL23] Xiaoxin Su, Yipeng Zhou, Laizhong Cui, and Jiangchuan Liu. On model transmission strategies in federated learning with lossy communications. IEEE Transactions on Parallel and Distributed Systems, 34 (4):1173–1185, April 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sun:2020:EPP

[SZM20] S. Sun, R. Zhang, and H. Ma. Efficient parallelism of postquantum signature scheme SPHINCS. IEEE Transactions on Parallel and Distributed Systems, 31(11):2542– 2555, November 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Scheffler:2023:SSS

[SZS^{+}23] Paul Scheffler, Florian Zaruba, Fabian Schuiki, Torsten Hoefler, and Luca Benini. Sparse stream semantic registers: a lightweight ISA extension accelerating general sparse linear algebra. IEEE Transactions on Parallel and Distributed Systems, 34(12):3147–3161, December 2023. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Song:2024:CSG

[SZZY24] Jie Song, Peimeng Zhu, Yanfeng Zhang, and Ge Yu. CloudSimPer: Simulating geodistributed datacenters powered by renewable energy mix. IEEE Transactions on Parallel and Distributed Systems, 35 (4):531–547, April 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tuli:2022:GTS

[TCJ22a] Shreshth Tuli, Giuliano Casale, and Nicholas R. Jennings. GOSH: Task scheduling using deep surrogate models in fog computing environments. IEEE Transactions on Parallel and Distributed Systems, 33(11): 2821–2833, November 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tuli:2022:MAA

[TCJ22b] Shreshth Tuli, Giuliano Casale, and Nicholas R. Jennings. MCDS: AI augmented workflow scheduling in mobile edge cloud computing systems. IEEE Transactions on Parallel and Distributed Systems, 33(11):2794–2807, November 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tang:2022:CEW

[TCT^{+}22] Xiaoyong Tang, Wenbiao Cao, Huiya Tang, Tan Deng, Jing Mei, Yi Liu, Cheng Shi, Meng Xia, and Zeng Zeng. Costefficient workflow scheduling algorithm for applications with deadline constraint on heterogeneous clouds. IEEE Transactions on Parallel and Distributed Systems, 33(9):2079– 2092, September 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Titos-Gil:2020:CIB

[TGFPRA20] R. Titos-Gil, R. Fernández-Pascual, A. Ros, and M. E. Acacio. Concurrent irrevocability in best-effort hardware transactional memory. IEEE Transactions on Parallel and Distributed Systems, 31 (6):1301–1315, June 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Titos-Gil:2022:DDS

[TGFPRA22] Rubén Titos-Gil, Ricardo Fernández-Pascual, Alberto Ros, and Manuel E. Acacio. DeTraS: Delaying stores for friendly-fire mitigation in hardware transactional memory. IEEE Transactions on Parallel and Distributed Systems, 33 (1):1–13, January 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tian:2024:SOC

[TJD^{+}24] Rui Tian, Jiazhi Jiang, Jiangsu Du, Dan Huang, and Yutong Lu. Sophisticated orchestrating concurrent DLRM training on CPU/GPU platform. IEEE Transactions on Parallel and Distributed Systems, 35(11):2177–2192, November 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tang:2022:TAP

[TKRB22] Houjun Tang, Quincey Koziol, John Ravi, and Suren Byna. Transparent asynchronous parallel I/O using background threads. IEEE Transactions on Parallel and Distributed Systems, 33(4):891–902, April 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Trott:2022:KPM

[TLGA^{+}22] Christian R. Trott, Damien Lebrun-Grandié, Daniel Arndt, Jan Ciesko, Vinh Dang, Nathan Ellingwood, Rahulkumar Gayatri, Evan Harvey, Daisy S. Hollman, Dan Ibanez, Nevin Liber, Jonathan Madsen, Jeff Miles, David Poliakoff, Amy Powell, Sivasankaran Rajamanickam, Mikael Simberg, Dan Sunderland, Bruno Turcksin, and Jeremiah Wilke. Kokkos 3: Programming model extensions for the exascale era. IEEE Transactions on Parallel and Distributed Systems, 33(4):805– 817, April 2022. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tauro:2022:MSM

[TLH22] Brian R. Tauro, Conghao Liu, and Kyle C. Hale. Modeling speedup in multi-OS environments. IEEE Transactions on Parallel and Distributed Systems, 33(6):1436–1450, June 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tian:2020:PPR

[TLQ^{+}20] C. Tian, B. Li, L. Qin, J. Zheng, J. Yang, W. Wang, G. Chen, and W. Dou. P-PFC:

Reducing tail latency with predictive PFC in lossless data center networks. IEEE Transactions on Parallel and Distributed Systems, 31(6):1447– 1459, June 2020. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tampouratzis:2023:NIS

[TMP23] Nikolaos Tampouratzis, Panagiotis Mousouliotis, and Ioannis Papaefstathiou. A novel integrated simulation framework for cyber-physical systems modelling. IEEE Transactions on Parallel and Distributed Systems, 34(10):2684– 2698, October 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Turchetto:2020:GDS

[TPV20] M. Turchetto, A. D. Palù, and R. Vacondio. A general design for a scalable MPI-GPU multiresolution 2D numerical solver. IEEE Transactions on Parallel and Distributed Systems, 31 (5):1036–1047, May 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Toha:2021:TGM

[TRN^{+}21] T. R. Toha, A. S. M. Rizvi, J. Noor, M. A. Adnan, and A. B. M. A. Al Islam. Towards greening MapReduce clusters considering both computation energy and cooling energy. IEEE Transactions on Parallel and Distributed Systems, 32 (4):931–942, April 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tarraf:2024:MMH

[TSC^{+}24] Ahmad Tarraf, Martin Schreiber, Alberto Cascajo, Jean-Baptiste Besnard, Marc-André Vef, Dominik Huber, Sonja Happ, André Brinkmann, David E. Singh, Hans-Christian Hoppe, Alberto Miranda, Antonio J. Peña, Rui Machado, Marta Garcia-Gasulla, Martin Schulz, Paul Carpenter, Simon Pickartz, Tiberiu Rotaru, Sergio Iserte, Victor Lopez, Jorge Ejarque, Heena Sirwani, Jesus Carretero, and Felix Wolf. Malleability in modern HPC systems: Current experiences, challenges, and future opportunities. IEEE Transactions on Parallel and Distributed Systems, 35(9):1551– 1564, September 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tang:2023:PGD

[TSLC23] Zhenheng Tang, Shaohuai Shi, Bo Li, and Xiaowen Chu. GossipFL: a decentralized federated learning framework with sparsified and adaptive communication. IEEE Transactions on Parallel and Distributed Systems, 34(3):909–922, March 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Thakkar:2021:RRN

[TSV21] Hiren Kumar Thakkar, Prasan Kumar Sahoo, and Bharadwaj Veeravalli. RENDA: Resource and network aware data placement algorithm for periodic workloads in cloud. IEEE Transactions on Parallel and Distributed Systems, 32(12): 2906–2920, December 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tan:2021:OLA

[TSW^{+}21] G. Tan, C. Shui, Y. Wang, X. Yu, and Y. Yan. Optimizing the LINPACK algorithm for large-scale PCIe-based CPU– GPU heterogeneous systems. IEEE Transactions on Parallel and Distributed Systems, 32(9): 2367–2380, September 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Taherpour:2024:HFA

[TW24] Amirhossein Taherpour and Xiaodong Wang. HybridChain: Fast, accurate, and secure transaction processing with distributed learning. IEEE Transactions on Parallel and Distributed Systems, 35(6):813– 827, June 2024. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tang:2022:PFE

[TWX22] Chenlei Tang, Jiguang Wan, and Changsheng Xie. FenceKV: Enabling efficient range query for key–value separation. IEEE Transactions on Parallel and Distributed Systems, 33(12): 3375–3386, December 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tan:2020:AIP

[TWY^{+}20] Y. Tan, B. Wang, Z. Yan, W. Srisa-an, X. Chen, and D. Liu. APMigration: Improving performance of hybrid memory performance via an adaptive page migration method. IEEE Transactions on Parallel and Distributed Systems, 31(2):266–278, February 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tang:2020:CBW

[TWYL20] Q. Tang, K. Wang, K. Yang, and Y. Luo. Congestionbalanced and welfare-maximized charging strategies for electric vehicles. IEEE Transactions on Parallel and Distributed Systems, 31(12):2882–2895, December 2020. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tan:2021:AAR

[TXG^{+}21] Cheng Tan, Chenhao Xie, Tong Geng, Andres Marquez, Antonino Tumeo, Kevin Barker, and Ang Li. ARENA: Asynchronous reconfigurable accelerator ring to enable datacentric parallel computing. IEEE Transactions on Parallel and Distributed Systems, 32(12):2880–2892, December 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tan:2021:IPD

[TXX^{+}21] Y. Tan, C. Xu, J. Xie, Z. Yan, H. Jiang, W. Srisaan, X. Chen, and D. Liu. Improving the performance of deduplication-based storage cache via content-driven cache management methods. IEEE Transactions on Parallel and Distributed Systems, 32(1):214– 228, January 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Uddin:2021:MID

[UXL^{+}21] M. P. Uddin, Y. Xiang, X. Lu, J. Yearwood, and L. Gao. Mutual information driven federated learning. IEEE Transactions on Parallel and Distributed Systems, 32(7):1526– 1538, July 2021. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

vanDaalen:2023:PPP

[vDIDB23] Florian van Daalen, Lianne Ippel, Andre Dekker, and Inigo Bermejo. Privacy preserving nparty scalar product protocol. IEEE Transactions on Parallel and Distributed Systems, 34(4): 1060–1066, April 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Versluis:2020:WTA

[VMT^{+}20] L. Versluis, R. Mathá, S. Talluri, T. Hegeman, R. Prodan, E. Deelman, and A. Iosup. The workflow trace archive: Openaccess data from public and private computing infrastructures. IEEE Transactions on Parallel and Distributed Systems, 31(9): 2170–2184, September 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

vanDongen:2020:ESP

[vV20] G. van Dongen and D. Van den Poel. Evaluation of stream processing frameworks. IEEE Transactions on Parallel and Distributed Systems, 31(8): 1845–1858, August 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2023:CMC

[WBS23] Chunyang Wang, Yuebin Bai, and Desen Sun. CD-MSA: Cooperative and deadline-aware scheduling for efficient multitenancy on DNN accelerators. IEEE Transactions on Parallel and Distributed Systems, 34 (7):2091–2106, July 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2020:GPE

[WC20] Q. Wang and X. Chu. GPGPU performance estimation with core and memory frequency scaling. IEEE Transactions on Parallel and Distributed Systems, 31(12):2865–2881, December 2020. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wen:2024:CJR

[WCD^{+}24] Zhaojie Wen, Qiong Chen, Quanfeng Deng, Yipei Niu, Zhen Song, and Fangming Liu. ComboFunc: Joint resource combination and container placement for serverless function scaling with heterogeneous container. IEEE Transactions on Parallel and Distributed Systems, 35(11):1989– 2005, November 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wen:2024:JOP

[WCN^{+}24] Zhaojie Wen, Qiong Chen, Yipei Niu, Zhen Song, Quanfeng Deng, and Fangming Liu. Joint optimization of parallelism and resource configuration for serverless function steps. IEEE Transactions on Parallel and Distributed Systems, 35(4):560–576, April 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wei:2021:ABK

[WCT21] Y.-W. Wei, W.-M. Chen, and H.-H. Tsai. Accelerating the Bron–Kerbosch algorithm for maximal clique enumeration using GPUs. IEEE Transactions on Parallel and Distributed Systems, 32(9):2352– 2366, September 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wei:2023:SFG

[WCW^{+}23] Jianghong Wei, Xiaofeng Chen, Jianfeng Wang, Xinyi Huang, and Willy Susilo. Securing finegrained data sharing and erasure in outsourced storage systems. IEEE Transactions on Parallel and Distributed Systems, 34(2):552–566, February 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2023:EGI

[WDCK23] Ziheng Wang, Xiaoshe Dong, Heng Chen, and Yan Kang. Efficient GPU implementations of post-quantum signature XMSS. IEEE Transactions on Parallel and Distributed Systems, 34 (3):938–954, March 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2021:ESM

[WDJ21] S. Wang, Z. Ding, and C. Jiang.

Elastic scheduling for microservice applications in clouds. IEEE Transactions on Parallel and Distributed Systems, 32 (1):98–115, January 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wan:2020:MAC

[WDL^{+}20] B. Wan, J. Dang, Z. Li, H. Gong, F. Zhang, and S. Oh. Modeling analysis and costperformance ratio optimization of virtual machine scheduling in cloud computing. IEEE Transactions on Parallel and Distributed Systems, 31(7):1518– 1532, July 2020. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:2023:HAS

[WDZ^{+}23] Zhaorui Wu, Yuhui Deng, Yi Zhou, Lin Cui, and Xiao Qin. HashCache: Accelerating serverless computing by skipping duplicated function execution. IEEE Transactions on Parallel and Distributed Systems, 34(12):3192–3206, December 2023. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2024:AAR

[WFY^{+}24] Siqi Wang, Tianyu Feng, Hailong Yang, Xin You, Bangduo Chen, Tongxuan Liu, Zhongzhi Luan, and Depei Qian. AtRec: Accelerating recommendation model training on CPUs. IEEE Transactions on Parallel and Distributed Systems, 35(6):750– 763, June 2024. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2023:NLM

[WGBS23] Chen Wang, Yanfei Guo, Pavan Balaji, and Marc Snir. Near-lossless MPI tracing and proxy application autogeneration. IEEE Transactions on Parallel and Distributed Systems, 34(1):123–140, January 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2020:UML

[WGLZ20] C. Wang, L. Gong, X. Li, and X. Zhou. A ubiquitous machine learning accelerator with automatic parallelization on FPGA. IEEE Transactions on Parallel and Distributed Systems, 31 (10):2346–2359, October 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2023:ADG

[WGN^{+}23] Yanhong Wang, Tianchan Guan, Dimin Niu, Qiaosha Zou, Hongzhong Zheng, C.- J. Richard Shi, and Yuan Xie. Accelerating distributed GNN training by codes. IEEE Transactions on Parallel and Distributed Systems, 34(9):2598– 2614, September 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2022:ECS

[WGQ^{+}22] Haozhao Wang, Song Guo, Zhihao Qu, Ruixuan Li, and Ziming Liu. Errorcompensated sparsification for communication-efficient decentralized training in edge environment. IEEE Transactions on Parallel and Distributed Systems, 33(1):14–25, January 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:2023:DAC

[WGW^{+}23] Feijie Wu, Song Guo, Haozhao Wang, Haobo Zhang, Zhihao Qu, Jie Zhang, and Ziming Liu. From deterioration to acceleration: a calibration approach to rehabilitating step asynchronism in federated optimization. IEEE Transactions on Parallel and Distributed Systems, 34 (5):1548–1559, May 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2021:TED

[WHC^{+}21] Zhaokang Wang, Weiwei Hu, Guowang Chen, Chunfeng Yuan, Rong Gu, and Yihua Huang. Towards efficient distributed subgraph enumeration via backtracking-based framework. IEEE Transactions on Parallel and Distributed Systems, 32(12):2953–2969, December 2021. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wan:2022:IPE

[WHG^{+}22] Lipeng Wan, Axel Huebl, Junmin Gu, Franz Poeschel, Ana Gainaru, Ruonan Wang, Jieyang Chen, Xin Liang, Dmitry Ganyushin, Todd Munson, Ian Foster, Jean-Luc Vay, Norbert Podhorszki, Kesheng Wu, and Scott Klasky. Improving I/O performance for exascale applications through online data layout reorganization. IEEE Transactions on Parallel and Distributed Systems, 33 (4):878–890, April 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2023:IMD

[WHL^{+}23] Zhilin Wang, Qin Hu, Ruinian Li, Minghui Xu, and Zehui Xiong. Incentive mechanism design for joint resource allocation in blockchain-based federated learning. IEEE Transactions on Parallel and Distributed Systems, 34(5):1536– 1547, May 2023. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:2021:AFL

[WHLM21] W. Wu, L. He, W. Lin, and R. Mao. Accelerating federated learning over reliabilityagnostic clients in mobile edge computing systems. IEEE Transactions on Parallel and Distributed Systems, 32(7): 1539–1551, July 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:2023:FSF

[WHLM23] Wentai Wu, Ligang He, Weiwei Lin, and Carsten Maple. FedProf: Selective federated learning based on distributional representation profiling. IEEE Transactions on Parallel and Distributed Systems, 34 (6):1942–1953, June 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2021:FAT

[WHM^{+}21] J. Wang, J. Hu, G. Min, A. Y.

Zomaya, and N. Georgalas. Fast adaptive task offloading in edge computing based on meta reinforcement learning. IEEE Transactions on Parallel and Distributed Systems, 32(1): 242–253, January 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2023:FEM

[WHM^{+}23] Jin Wang, Jia Hu, Jed Mills, Geyong Min, Ming Xia, and Nektarios Georgalas. Federated ensemble model-based reinforcement learning in edge computing. IEEE Transactions on Parallel and Distributed Systems, 34(6):1848– 1859, June 2023. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2024:ACR

[WHM^{+}24] Zhe Wang, Jia Hu, Geyong Min, Zhiwei Zhao, and Zi Wang. Agile cache replacement in edge computing via offline-online deep reinforcement learning. IEEE Transactions on Parallel and Distributed Systems, 35(4):663– 674, April 2024. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wong:2021:MMF

[WHRL21] A. Wong, E. Heymann, D. Rexachs, and E. Luque. Middleware to manage fault tolerance using semi-coordinated checkpoints. IEEE Transactions on Parallel and Distributed Systems, 32(2):254–268, February 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2022:RRV

[WIBD22] Yuan Wang, Hideaki Ishii, François Bonnet, and Xavier Défago. Resilient real-valued consensus in spite of mobile malicious agents on directed graphs. IEEE Transactions on Parallel and Distributed Systems, 33(3):586–603, March 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2021:PBS

[WJG^{+}21] Y. Wang, X. Jiang, N. Guan, Z. Guo, X. Liu, and W. Yi. Partitioning-based scheduling of OpenMP task systems with tied tasks. IEEE Transactions on Parallel and Distributed Systems, 32(6):1322– 1339, June 2021. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2024:BBA

[WJL^{+}24] Peng Wang, Hong Jiang, Yu Liu, Zhelong Zhao, Ke Zhou, and Zhihai Huang. Beyond Belady to attain a seemingly unattainable byte miss ratio for content delivery networks. IEEE Transactions on Parallel and Distributed Systems, 35(11):1949–1963, November 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wei:2024:IIA

[WJZ^{+}24] Cunyang Wei, Haipeng Jia, Yunquan Zhang, Jianyu Yao, Chendi Li, and Wenxuan Cao. IrGEMM: an inputaware tuning framework for irregular GEMM on ARM and X86 CPUs. IEEE Transactions on Parallel and Distributed Systems, 35(9):1672– 1689, September 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2020:CHP

[WL20] K. Wang and A. Louri. CURE: a high-performance, low-power, and reliable network-on-chip design using reinforcement learning. IEEE Transactions on Parallel and Distributed Systems, 31(9):2125– 2138, September 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2020:NLC

[WLF^{+}20] G. Wang, C. Lin, J. Fan, B. Cheng, and X. Jia. A novel low cost interconnection architecture based on the generalized hypercube. IEEE Transactions on Parallel and Distributed Systems, 31(3):647– 662, March 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2022:DDH

[WLF^{+}22] Junchang Wang, Dunwei Liu, Xiong Fu, Fu Xiao, and Chen Tian. DHash: Dynamic hash tables with non-blocking regular operations. IEEE Transactions on Parallel and Distributed Systems, 33(12):3274– 3290, December 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2020:AFG

[WLH^{+}20a] L. Wang, L. Liu, J. Han, X. Wang, S. Yin, and S. Wei. Achieving flexible global reconfiguration in NoCs using reconfigurable rings. IEEE Transactions on Parallel and Distributed Systems, 31(3):611– 622, March 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:2020:TDC

[WLH20b] X. Wu, P. Loiseau, and E. Hyytiä. Toward designing cost-optimal policies to utilize IaaS clouds with online learning. IEEE Transactions on Parallel and Distributed Systems, 31(3):501–514, March 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2023:PBE

[WLH^{+}23] Zhiwei Wang, Peinan Li, Rui Hou, Zhihao Li, Jiangfeng Cao, XiaoFeng Wang, and Dan Meng. HE-Booster: an efficient polynomial arithmetic acceleration on GPUs for fully homomorphic encryption. IEEE Transactions on Parallel and Distributed Systems, 34(4): 1067–1081, April 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2020:CRM

[WLL^{+}20] J. Wang, T. Liu, Q. Liu, X. He, H. Luo, and W. He. Compression ratio modeling and estimation across error bounds for lossy compression. IEEE Transactions on Parallel and Distributed Systems, 31 (7):1621–1635, July 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Witte:2020:EDA

[WLM^{+}20] P. A. Witte, M. Louboutin, H. Modzelewski, C. Jones, J. Selvage, and F. J. Herrmann. An event-driven approach to serverless seismic imaging in the cloud. IEEE Transactions on Parallel and Distributed Systems, 31(9):2032– 2049, September 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2023:LSM

[WLP^{+}23] Zhaohua Wang, Zhenyu Li, Heng Pan, Guangming Liu, Yunfei Chen, Qinghua Wu, Gareth Tyson, and Gang Cheng. Large-scale measurements and prediction of DC-WAN traffic. IEEE Transactions on Parallel and Distributed Systems, 34(5):1390– 1405, May 2023. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wen:2020:GPD

[WLY^{+}20] Z. Wen, T. Lin, R. Yang, S. Ji, R. Ranjan, A. Romanovsky, C. Lin, and J. Xu. GA-Par: Dependable microservice orchestration framework for geo-distributed clouds. IEEE Transactions on Parallel and Distributed Systems, 31(1):129– 143, January 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2022:DAD

[WLY22] Lipeng Wang, Qiong Luo, and Shengen Yan. DIESEL+: Accelerating distributed deep learning tasks on image datasets. IEEE Transactions on Parallel and Distributed Systems, 33 (5):1173–1184, May 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wei:2023:SDS

[WLZ^{+}23] Wenqi Wei, Ling Liu, Jingya Zhou, Ka-Ho Chow, and Yanzhao Wu. Securing distributed SGD against gradient leakage threats. IEEE Transactions on Parallel and Distributed Systems, 34(7):2040– 2054, July 2023. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2023:AAB

[WMG^{+}23] Mingyue Wang, Yinbin Miao, Yu Guo, Hejiao Huang, Cong Wang, and Xiaohua Jia. AESM2 attribute-based encrypted search for multi-owner and multi-user distributed systems. IEEE Transactions on Parallel and Distributed Systems, 34(1):92–107, January 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:2020:TDS

[WNA^{+}20] D. Wu, X. Nie, E. Asmare, D. I. Arkhipov, Z. Qin, R. Li, J. A. McCann, and K. Li. Towards distributed SDN: Mobility management and flow scheduling in software defined urban IoT. IEEE Transactions on Parallel and Distributed Systems, 31(6):1400– 1418, June 2020. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2020:TDD

[WNL20] H. Wang, D. Niu, and B. Li. Turbo: Dynamic and decentralized global analytics via machine learning. IEEE Transactions on Parallel and Distributed Systems, 31(6):1372– 1386, June 2020. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:2022:VEV

[WPG^{+}22] Haotian Wu, Zhe Peng, Songtao Guo, Yuanyuan Yang, and Bin Xiao. VQL: Efficient and verifiable cloud query services for blockchain systems. IEEE Transactions on Parallel and Distributed Systems, 33 (6):1393–1406, June 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2024:TOE

[WPG^{+}24] Daoce Wang, Jesus Pulido, Pascal Grosset, Sian Jin, Jiannan Tian, Kai Zhao, James Ahrens, and Dingwen Tao. TAC+: Optimizing errorbounded lossy compression for 3D AMR simulations. IEEE Transactions on Parallel and Distributed Systems, 35(3):421– 438, March 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2021:OCC

[WPZ^{+}21] S. Wang, A. Pi, X. Zhou, J. Wang, and C.-Z. Xu. Overlapping communication with computation in parameter server for scalable DL training. IEEE Transactions on Parallel and Distributed Systems, 32(9): 2144–2159, September 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2022:EPS

[WPZ22] Shaoqi Wang, Aidi Pi, and Xiaobo Zhou. Elastic parameter server: Accelerating ML training with scalable resource scheduling. IEEE Transactions on Parallel and Distributed Systems, 33(5):1128– 1143, May 2022. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2020:EPE

[WQKH20] X. Wang, X. Qian, A. Knoll, and K. Huang. Efficient performance estimation and workgroup size pruning for OpenCL kernels on GPUs. IEEE Transactions on Parallel and Distributed Systems, 31(5):1089– 1106, May 2020. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2025:CEL

[WQZD25] Pengwei Wang, Junye Qiao, Yuying Zhao, and Zhijun Ding. Cost-effective and low-latency data placement in edge environment based on PageRankinspired regional value. IEEE Transactions on Parallel and Distributed Systems, 36(2):185– 196, February 2025. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Weakley:2023:GER

[WR23] Le Mai Weakley and Tim Robinson. Guest editorial reproducibility initiative at the SC conference series: a preface to the special section. IEEE Transactions on Parallel and Distributed Systems, 34 (6):1697–1698, June 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2022:NTI

[WRLS22] Maolin Wang, Seyedramin Rasoulinezhad, Philip H. W. Leong, and Hayden K.-H. So. NITI: Training integer neural networks using integeronly arithmetic. IEEE Transactions on Parallel and Distributed Systems, 33(11):3249– 3261, November 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:2023:TIC

[WSHJ23] Wenchao Wu, Xuanhua Shi, Ligang He, and Hai Jin. TurboMGNN: Improving concurrent GNN training tasks on GPU with fine-grained kernel fusion. IEEE Transactions on Parallel and Distributed Systems, 34(6):1968–1981, June 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:2022:ORS

[WSLX22] Si Wu, Zhirong Shen, Patrick P. C. Lee, and Yinlong Xu. Optimal repair-scaling trade-off in locally repairable codes: Analysis and evaluation. IEEE Transactions on Parallel and Distributed Systems, 33(1):56– 69, January 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2020:QGA

[WSM^{+}20] S. Wang, W. Sun, L. Ma, W. Lv, and X. Cheng. Quantum game analysis on extrinsic incentive mechanisms for P2P services. IEEE Transactions on Parallel and Distributed Systems, 31(1):159–170, January 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2023:AST

[WSX^{+}23] Xun Wang, Ruibao Song, Junmin Xiao, Tong Li, and Xueqi Li. Accelerating k-shape time series clustering algorithm using GPU. IEEE Transactions on Parallel and Distributed Systems, 34(10):2718– 2734, October 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:2024:EEC

[WUR^{+}24] Di Wu, Rehmat Ullah, Philip Rodgers, Peter Kilpatrick, Ivor Spence, and Blesson Varghese. EcoFed: Efficient communication for DNN partitioningbased federated learning. IEEE Transactions on Parallel and Distributed Systems, 35(3):377– 390, March 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2023:DEO

[WVSL23] Xiaoli Wang, Bharadwaj Veeravalli, Jiaming Song, and Honghu Liu. On the design and evaluation of an optimal security-and-time cognizant data placement for dynamic fog environments. IEEE Transactions on Parallel and Distributed Systems, 34(2):489– 500, February 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:2024:GEI

[WWJL24] Jing Wu, Lin Wang, Qirui Jin, and Fangming Liu. Graft: Efficient inference serving for hybrid deep learning with SLO guarantees via DNN realignment. IEEE Transactions on Parallel and Distributed Systems, 35(2):280–296, February 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:2024:PCD

[WXH^{+}24] Changmao Wu, Zhengwei Xu, Xiaoming He, Qi Lou, Yuanyuan Xia, and Shuman Huang. Proactive caching with distributed deep reinforcement learning in 6G cloudedge collaboration computing. IEEE Transactions on Parallel and Distributed Systems, 35(8): 1387–1399, August 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2020:FAT

[WXHZ20] H. Wang, H. Xu, L. Huang, and Y. Zhai. Fast and accurate traffic measurement with hierarchical filtering. IEEE Transactions on Parallel and Distributed Systems, 31(10):2360– 2374, October 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2024:HHP

[WXT^{+}24] Cheng Wang, Kun Xie, Jiazheng Tian, Jigang Wen, Xiaocan Li, Gaogang Xie, and Kenli Li. HPETC: History priority enhanced tensor completion for network distance measurement. IEEE Transactions on Parallel and Distributed Systems, 35(6):857– 873, June 2024. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2024:FEF

[WXX^{+}24] Zhiyuan Wang, Hongli Xu, Yang Xu, Zhida Jiang, Jianchun Liu, and Suo Chen. FAST: Enhancing federated learning through adaptive data sampling and local training. IEEE Transactions on Parallel and Distributed Systems, 35(2):221– 236, February 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2024:TMG

[WXZ24] Ran Wang, Cheng Xu, and Xiaotong Zhang. Toward materials genome big-data: a blockchain-based secure storage and efficient retrieval method. IEEE Transactions on Parallel and Distributed Systems, 35(9): 1630–1643, September 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:2021:FSB

[WYW21] X. Wu, X. Yao, and C.-L. Wang. FedSCR: Structurebased communication reduction for federated learning. IEEE Transactions on Parallel and Distributed Systems, 32 (7):1565–1577, July 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2022:CPI

[WYW^{+}22] Zhonghua Wang, Ting Yao, Jiguang Wan, Hong Jiang, Qiu Cui, Liu Tang, Yiwen Zhang, and Qiuyang Zhang. ComboTree: a persistent indexing structure with universal operational efficiency and scalability. IEEE Transactions on Parallel and Distributed Systems, 33(10):2277–2290, October 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:2024:BDL

[WYZ^{+}24] Donglei Wu, Weihao Yang, Xiangyu Zou, Hao Feng, Dingwen Tao, Shiyi Li, Wen Xia, and Binxing Fang. BIRD+: Design of a lightweight communication compressor for resourceconstrained distribution learning platforms. IEEE Transactions on Parallel and Distributed Systems, 35(11):2193– 2207, November 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2024:OPT

[WYZY24] Liang Wang, Jinzhe Yang, Jidong Zhai, and Guangwen Yang. Optimizing I/O performance through effective vCPU scheduling interference management. IEEE Transactions on Parallel and Distributed Systems, 35(12):2315–2330, December 2024. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:2023:TTB

[WZGM23] Yulong Wu, Weizhe Zhang, Nan Guan, and Yehan Ma. TDTA: Topology-based realtime DAG task allocation on identical multiprocessor platforms. IEEE Transactions on Parallel and Distributed Systems, 34(11):2895– 2909, November 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2022:OPM

[WZHW22] Yiming Wang, Weizhe Zhang, Meng Hao, and Zheng Wang. Online power management for multi-cores: a reinforcement learning based approach. IEEE Transactions on Parallel and Distributed Systems, 33(4):751– 764, April 2022. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2022:DGE

[WZL^{+}22] Farui Wang, Weizhe Zhang, Shichao Lai, Meng Hao, and Zheng Wang. Dynamic GPU energy optimization for machine learning training workloads. IEEE Transactions on Parallel and Distributed Systems, 33(11):2943–2954, November 2022. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Weng:2023:DAB

[WZL^{+}23] Tongfeng Weng, Xu Zhou, Kenli Li, Kian-Lee Tan, and Keqin Li. Distributed approaches to butterfly analysis on large dynamic bipartite graphs. IEEE Transactions on Parallel and Distributed Systems, 34(2):431–445, February 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2022:HNH

[WZY^{+}22] Ziliang Wang, Xiaohong Zhang, Meng Yan, Ling Xu, and Dan Yang. HSA-Net: Hiddenstate-aware networks for highprecision QoS prediction. IEEE Transactions on Parallel and Distributed Systems, 33(6): 1421–1435, June 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2020:CPV

[WZZ^{+}20] T. Wang, J. Zhou, G. Zhang, T. Wei, and S. Hu. Customer perceived value- and risk-aware multiserver configuration for profit maximization. IEEE Transactions on Parallel and Distributed Systems, 31 (5):1074–1088, May 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xia:2021:OCD

[XCH^{+}21a] X. Xia, F. Chen, Q. He, J. Grundy, M. Abdelrazek, and H. Jin. Online collaborative data caching in edge computing. IEEE Transactions on Parallel and Distributed Systems, 32(2):281–294, February 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xia:2021:CEA

[XCH^{+}21b] X. Xia, F. Chen, Q. He, J. C. Grundy, M. Abdelrazek, and H. Jin. Cost-effective app data distribution in edge computing. IEEE Transactions on Parallel and Distributed Systems, 32 (1):31–44, January 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xia:2022:DUP

[XCH^{+}22] Xiaoyu Xia, Feifei Chen, Qiang He, Guangming Cui, John C. Grundy, Mohamed Abdelrazek, Xiaolong Xu, and Hai Jin. Data, user and power allocations for caching in multi-access edge computing. IEEE Transactions on Parallel and Distributed Systems, 33(5):1144– 1155, May 2022. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xia:2023:DMR

[XCL^{+}23] Junxu Xia, Geyao Cheng, Lailong Luo, Deke Guo, Pin Lv, and Bowen Sun. The doctrine of MEAN: Realizing deduplication storage at unreliable edge. IEEE Transactions on Parallel and Distributed Systems, 34 (10):2811–2826, October 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xu:2020:TAE

[XHQC20] B. Xu, J. Huang, X. Qin, and Q. Cao. Traffic-aware erasurecoded archival schemes for inmemory stores. IEEE Transactions on Parallel and Distributed Systems, 31(12):2938– 2953, December 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xue:2024:HAH

[XHY^{+}24] Runzhen Xue, Dengke Han, Mingyu Yan, Mo Zou, Xiaocheng Yang, Duo Wang, Wenming Li, Zhimin Tang, John Kim, Xiaochun Ye, and Dongrui Fan. HiHGNN: Accelerating HGNNs through parallelism and data reusability exploitation. IEEE Transactions on Parallel and Distributed Systems, 35(7):1122– 1138, July 2024. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xia:2024:OLA

[XJX24] Qiufen Xia, Zhiwei Jiao, and Zichuan Xu. Online learning algorithms for contextaware video caching in D2D edge networks. IEEE Transactions on Parallel and Distributed Systems, 35(1):1–19, January 2024. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xu:2020:MTD

[XLL^{+}20a] K. Xu, L. Lv, T. Li, M. Shen, H. Wang, and K. Yang. Minimizing tardiness for dataintensive applications in heterogeneous systems: a matching theory perspective. IEEE Transactions on Parallel and Distributed Systems, 31(1):144– 158, January 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xu:2020:DDD

[XLL^{+}20b] L. Xu, M. Lyu, Z. Li, Y. Li, and Y. Xu. Deterministic data distribution for efficient recovery in erasure-coded storage systems. IEEE Transactions on Parallel and Distributed Systems, 31(10):2248–2262, October 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xie:2024:ACE

[XLZD24] Renyou Xie, Chaojie Li, Xiaojun Zhou, and Zhaoyang Dong. Accelerating communicationefficient federated multi-task learning with personalization and fairness. IEEE Transactions on Parallel and Distributed Systems, 35(11):2239– 2253, November 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xu:2022:HSA

[XMW^{+}22] Rui Xu, Sheng Ma, Yaohua Wang, Yang Guo, Dongsheng Li, and Yuran Qiao. Heterogeneous systolic array architecture for compact CNNs hardware accelerators. IEEE Transactions on Parallel and Distributed Systems, 33(11):2860– 2871, November 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xu:2024:XSO

[XMW^{+}24] Jie Xu, Yulong Ming, Zihan Wu, Cong Wang, and Xiaohua Jia. X-Shard: Optimistic crossshard transaction processing for sharding-based blockchains. IEEE Transactions on Parallel and Distributed Systems, 35 (4):548–559, April 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xue:2023:HBN

[XRS^{+}23] Jie Xue, Liwen Ren, Bosheng Song, Yujie Guo, Jie Lu, Xiyu Liu, Guanzhong Gong, and Dengwang Li. Hypergraphbased numerical neural-like P systems for medical image segmentation. IEEE Transactions on Parallel and Distributed Systems, 34(4):1202– 1214, April 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xu:2023:ROM

[XSC^{+}23] Lei Xu, Honghui Shang, Xin Chen, Yunquan Zhang, Lifang Wang, Xingyu Gao, and Haifeng Song. Redesigning OpenKMC for multicomponent trillion-atom simulations on the new Sunway supercomputer. IEEE Transactions on Parallel and Distributed Systems, 34(7):1997– 2010, July 2023. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xie:2023:JCR

[XTH^{+}23] Tian Xie, Sanchal Thakkar, Ting He, Patrick McDaniel, and Quinn Burke. Joint caching and routing in cache networks with arbitrary topology. IEEE Transactions on Parallel and Distributed Systems, 34(8): 2237–2250, August 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xie:2023:CDC

[XWDC23] Li’an Xie, Ting Wang, Shuyi Du, and Haibin Cai. CERT-DF: a computing-efficient and robust distributed deep forest framework with low communication overhead. IEEE Transactions on Parallel and Distributed Systems, 34(12):3280– 3293, December 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xu:2020:HUS

[XWJ^{+}20] X. Xu, F. Wang, H. Jiang, Y. Cheng, D. Feng, and Y. Zhang. A hybrid update strategy for I/O-efficient out-of-core graph processing. IEEE Transactions on Parallel and Distributed Systems, 31(8): 1767–1782, August 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xu:2023:PII

[XXC^{+}23] Fei Xu, Jianian Xu, Jiabin Chen, Li Chen, Ruitao Shang, Zhi Zhou, and Fangming Liu. iGniter: Interference-aware GPU resource provisioning for predictable DNN inference in the cloud. IEEE Transactions on Parallel and Distributed Systems, 34(3):812–827, March 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xiao:2020:DGC

[XXM^{+}20] W. Xiao, J. Xue, Y. Miao, Z. Li, C. Chen, M. Wu, W. Li, and L. Zhou. Distributed graph computation meets machine learning. IEEE Transactions on Parallel and Distributed Systems, 31(7):1588– 1604, July 2020. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xu:2023:AAS

[XXP^{+}23] Jie Xu, Qingyuan Xie, Sen Peng, Cong Wang, and Xiaohua Jia. AdaptChain: Adaptive scaling blockchain with transaction deduplication. IEEE Transactions on Parallel and Distributed Systems, 34 (6):1909–1922, June 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xu:2024:ESA

[XXW^{+}24] Zichuan Xu, Guangyuan Xu, Hao Wang, Weifa Liang, Qiufen Xia, and Shangguang Wang.

Enabling streaming analytics in satellite edge computing via timely evaluation of big data queries. IEEE Transactions on Parallel and Distributed Systems, 35(1):105–122, January 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xu:2024:PPT

[XXWS24] Yin Xu, Mingjun Xiao, Jie Wu, and He Sun. Privacy preserving task push in spatial crowdsourcing with unknown popularity. IEEE Transactions on Parallel and Distributed Systems, 35(11):2039– 2053, November 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xie:2021:RCC

[XYL^{+}21] G. Xie, K. Yang, H. Luo, R. Li, and S. Hu. Reliability and confidentiality coverification for parallel applications in distributed systems. IEEE Transactions on Parallel and Distributed Systems, 32 (6):1353–1368, June 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xiao:2022:MAL

[XYW22] Danyang Xiao, Chengang Yang, and Weigang Wu. Mixing activations and labels in distributed training for split learning. IEEE Transactions on Parallel and Distributed Systems, 33(11):3165– 3177, November 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xia:2020:DFC

[XZJ^{+}20] W. Xia, X. Zou, H. Jiang, Y. Zhou, C. Liu, D. Feng, Y. Hua, Y. Hu, and Y. Zhang. The design of fast contentdefined chunking for data deduplication based storage systems. IEEE Transactions on Parallel and Distributed Systems, 31(9):2017– 2031, September 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xie:2020:SER

[XZL20] G. Xie, G. Zeng, and R. Li.

Safety enhancement for realtime parallel applications in distributed automotive embedded systems: a stable stopping approach. IEEE Transactions on Parallel and Distributed Systems, 31(9):2067– 2080, September 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xu:2021:EAI

[XZL^{+}21] Z. Xu, L. Zhao, W. Liang, O. F.

Rana, P. Zhou, Q. Xia, W. Xu, and G. Wu. Energy-aware inference offloading for DNNdriven applications in mobile edge clouds. IEEE Transactions on Parallel and Distributed Systems, 32(4):799– 814, April 2021. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xiang:2024:CER

[XZW^{+}24] Zhengzhe Xiang, Yuhang Zheng, Dongjing Wang, Javid Taheri, Zengwei Zheng, and Minyi Guo. Cost-effective and robust service provisioning in multiaccess edge computing. IEEE Transactions on Parallel and Distributed Systems, 35(10): 1765–1779, October 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xia:2024:RFL

[XZY^{+}24] Yaqi Xia, Zheng Zhang, Donglin Yang, Chuang Hu, Xiaobo Zhou, Hongyang Chen, Qianlong Sang, and Dazhao Cheng. Redundancy-free and load-balanced TGNN training with hierarchical pipeline parallelism. IEEE Transactions on Parallel and Distributed Systems, 35(11):1904– 1919, November 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yeung:2022:HIA

[YBY^{+}22] Gingfung Yeung, Damian Borowiec, Renyu Yang, Adrian Friday, Richard Harper, and Peter Garraghan. Horus: Interference-aware and predictionbased scheduling in deep learning systems. IEEE Transactions on Parallel and Distributed Systems, 33(1):88–100, January 2022. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yan:2020:IFS

[YCA^{+}20] H. Yan, H. R. Cherian, E. C.

Ahn, X. Qian, and L. Duan. iCELIA: a full-stack framework for STT-MRAM-based deep learning acceleration. IEEE Transactions on Parallel and Distributed Systems, 31(2):408– 422, February 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yao:2022:WFC

[YCZC22] Zihang Yao, Rong Chen, Binyu Zang, and Haibo Chen. Wukong+G: Fast and concurrent RDF query processing using RDMA-assisted GPU graph exploration. IEEE Transactions on Parallel and Distributed Systems, 33(7):1619– 1635, July 2022. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yu:2023:COA

[YDL23] Enda Yu, Dezun Dong, and Xiangke Liao. Communication optimization algorithms for distributed deep learning systems: a survey. IEEE Transactions on Parallel and Distributed Systems, 34(12):3294–3308, December 2023. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yang:2023:HFL

[YFB^{+}23] Zhengjie Yang, Sen Fu, Wei Bao, Dong Yuan, and Albert Y. Zomaya. Hierarchical federated learning with momentum acceleration in multi-tier networks.

IEEE Transactions on Parallel and Distributed Systems, 34 (10):2629–2641, October 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yang:2024:OFS

[YFD^{+}24] Weiling Yang, Jianbin Fang, Dezun Dong, Xing Su, and Zheng Wang. Optimizing full-spectrum matrix multiplications on ARMv8 multi-core CPUs. IEEE Transactions on Parallel and Distributed Systems, 35(3):439–454, March 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yang:2023:AAD

[YHD^{+}23] Siling Yang, Shuibing He, Hexiao Duan, Weijian Chen, Xuechen Zhang, Tong Wu, and Yanlong Yin. APQ: Automated DNN pruning and quantization for ReRAM-Based accelerators. IEEE Transactions on Parallel and Distributed Systems, 34(9): 2498–2511, September 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yang:2020:PAS

[YHS^{+}20] R. Yang, C. Hu, X. Sun, P. Garraghan, T. Wo, Z. Wen, H. Peng, J. Xu, and C. Li. Performance-aware speculative resource oversubscription for large-scale clusters. IEEE Transactions on Parallel and Distributed Systems, 31(7):

1499–1517, July 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yuan:2023:PCE

[YHT^{+}23] Liang Yuan, Qiang He, Siyu Tan, Bo Li, Jiangshan Yu, Feifei Chen, and Yun Yang. CoopEdge+: Enabling decentralized, secure and cooperative multi-access edge computing based on blockchain. IEEE Transactions on Parallel and Distributed Systems, 34 (3):894–908, March 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yang:2024:PDB

[YJWM24] Jian Yang, Jiantong Jiang, Zeyi Wen, and Ajmal Mian. Parallel and distributed Bayesian network structure learning. IEEE Transactions on Parallel and Distributed Systems, 35(4):517– 530, April 2024. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yang:2023:EED

[YLC^{+}23] Lei Yang, Yuwei Liao, Xin Cheng, Mengyuan Xia, and Guoqi Xie. Efficient edge data management framework for IIoT via prediction-based data reduction. IEEE Transactions on Parallel and Distributed Systems, 34(12):3309– 3322, December 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yeh:2020:SEP

[YLL^{+}20] E. Yeh, P. Lin, X. Lin, J. Jeng, and Y. Fang. System error prediction for business support systems in telecommunications networks. IEEE Transactions on Parallel and Distributed Systems, 31(11):2723– 2733, November 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yang:2021:EED

[YLL21] D. Yang, J. Liu, and J. Lai. EDGES: an efficient distributed graph embedding system on GPU clusters. IEEE Transactions on Parallel and Distributed Systems, 32(7):1892– 1902, July 2021. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yang:2023:SEI

[YLS^{+}23] Kaicheng Yang, Sheng Long, Qilong Shi, Yuanpeng Li, Zirui Liu, Yuhan Wu, Tong Yang, and Zhengyi Jia. SketchINT: Empowering INT with TowerSketch for per-flow per-switch measurement. IEEE Transactions on Parallel and Distributed Systems, 34(11):2876– 2894, November 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yang:2021:RAR

[YLT^{+}21] S. Yang, F. Li, S. Trajanovski, R. Yahyapour, and X. Fu. Recent advances of resource allocation in network function virtualization. IEEE Transactions on Parallel and Distributed Systems, 32(2):295–314, February 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yao:2022:WWA

[YLW^{+}22] Yuan Yao, Shuangyang Liu, Sikai Wu, Jinyu Wang, Jinting Ni, Gang Yang, and Yu Zhang. WAMP^{2}2S: Workload-aware GPU performance model based pseudo-preemptive real-time scheduling for the airborne embedded system. IEEE Transactions on Parallel and Distributed Systems, 33(11):2767– 2780, November 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Youness:2021:OWA

[YOM21] H. Youness, A. Omar, and M. Moness. An optimized weighted average makespan in fault-tolerant heterogeneous MPSoCs. IEEE Transactions on Parallel and Distributed Systems, 32(8):1933–1946, August 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yasar:2022:BBT

[YRBÇ22] Abdurrahman Yašar, Sivasankaran Rajamanickam, Jonathan W. Berry, and Ümit V. Çatalyürek. A block-based triangle counting algorithm on heterogeneous environments. IEEE Transactions on Parallel and Distributed Systems, 33(2):444– 458, February 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yu:2023:TLC

[YRQ23] Chenle Yu, Sara Royuela, and Eduardo Quiñones. Taskgraph: a low contention OpenMP tasking framework. IEEE Transactions on Parallel and Distributed Systems, 34(8):2325– 2336, August 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yang:2022:DRL

[YS22] Yuanhao Yang and Hong Shen. Deep reinforcement learning enhanced greedy optimization for online scheduling of batched tasks in cloud HPC systems. IEEE Transactions on Parallel and Distributed Systems, 33(11):3003–3014, November 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ye:2022:AFD

[YSG^{+}22] Zhisheng Ye, Peng Sun, Wei Gao, Tianwei Zhang, Xiaolin Wang, Shengen Yan, and Yingwei Luo. Astraea: a fair deep learning scheduler for multitenant GPU clusters. IEEE Transactions on Parallel and Distributed Systems, 33(11): 2781–2793, November 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ye:2021:DFE

[YSZL21] Q. Ye, Y. Sun, J. Zhang, and J. Lv. A distributed framework for EA-based NAS. IEEE Transactions on Parallel and Distributed Systems, 32 (7):1753–1764, July 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yu:2020:EPW

[YT20] C. Yu and S. Tsao. Efficient and portable workgroup size tuning. IEEE Transactions on Parallel and Distributed Systems, 31(2):455–469, February 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yu:2023:DNB

[YTL^{+}23] Bin Yu, Cong Tian, Xu Lu, Nan Zhang, and Zhenhua Duan. A distributed networkbased runtime verification of full regular temporal properties. IEEE Transactions on Parallel and Distributed Systems, 34(1):76–91, January 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yao:2024:TEG

[YTL^{+}24] Feng Yao, Qian Tao, Shengyuan Lin, Yanfeng Zhang, Wenyuan Yu, Shufeng Gong, Qiange Wang, Ge Yu, and Jingren Zhou. Towards efficient graph processing in geo-distributed data centers. IEEE Transactions on Parallel and Distributed Systems, 35(11):2147– 2160, November 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yao:2020:PCG

[YW20] X. Yao and C. Wang. Probabilistic consistency guarantee in partial quorum-based data store. IEEE Transactions on Parallel and Distributed Systems, 31(8):1815–1827, August 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yu:2020:ALB

[YWH^{+}20] Y. Yu, W. Wang, R. Huang, J. Zhang, and K. B. Letaief. Achieving load-balanced, redundancyfree cluster caching with selective partition. IEEE Transactions on Parallel and Distributed Systems, 31(2):439– 454, February 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ye:2021:SSD

[YWH^{+}21] Z. Ye, Y. Wang, S. He, C. Xu, and X. Sun. Sova: a softwaredefined autonomic framework for virtual network allocations. IEEE Transactions on Parallel and Distributed Systems, 32(1): 116–130, January 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yu:2024:FHI

[YWL^{+}24] Hanfei Yu, Hao Wang, Jian Li, Xu Yuan, and Seung-Jong Park. Freyr^{+}+: Harvesting idle resources in serverless computing via deep reinforcement learning. IEEE Transactions on Parallel and Distributed Systems, 35(11):2254– 2269, November 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yang:2023:BHC

[YWS^{+}23] Zhijie Yang, Lei Wang, Wei Shi, Yao Wang, Junbo Tie, Feng Wang, Xiang Yu, Linghui Peng, Chao Xiao, Xun Xiao, Yao Yao, Gan Zhou, Xuhu Yu, Rui Gong, Xia Zhao, Yuhua Tang, and Weixia Xu. Back to homogeneous computing: a tightly-coupled neuromorphic processor with neuromorphic ISA. IEEE Transactions on Parallel and Distributed Systems, 34(11):2910– 2927, November 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yin:2020:CHC

[YWZ^{+}20] Y. Yin, J. Wu, X. Zhou, L. Eeckhout, A. Qouneh, T. Li, and Z. Yu. COPA: Highly cost-effective power back-up for green datacenters. IEEE Transactions on Parallel and Distributed Systems, 31(4):967– 980, April 2020. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yang:2024:FFH

[YXDL24] Meilin Yang, Jian Xu, Wenbo Ding, and Yang Liu. Fed-HAP: Federated hashing with global prototypes for cross-silo retrieval. IEEE Transactions on Parallel and Distributed Systems, 35(4):592–603, April 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yuan:2024:OMG

[YYL^{+}24] Fan Yuan, Xiaojian Yang, Shengguo Li, Dezun Dong, Chun Huang, and Zheng Wang. Optimizing multi-grid preconditioned conjugate gradient method on multi-cores. IEEE Transactions on Parallel and Distributed Systems, 35(5):768– 779, May 2024. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yu:2020:SBS

[YYW^{+}20] J. Yu, W. Yang, F. Wang, D. Dong, J. Feng, and Y. Li. Spatially bursty I/O on supercomputers: Causes, impacts and solutions. IEEE Transactions on Parallel and Distributed Systems, 31(12):2908– 2922, December 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yuan:2020:EEB

[YYZ^{+}20] X. Yuan, X. Yuan, Y. Zhang, B. Li, and C. Wang. Enabling encrypted Boolean queries in geographically distributed databases. IEEE Transactions on Parallel and Distributed Systems, 31(3):634–646, March 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yu:2023:ATR

[YZC^{+}23] Se-Young Yu, Qingyang Zeng, Jim Chen, Yan Chen, and Joe Mambretti. AIDTN: Towards a real-time AI optimized DTN system with NVMeoF. IEEE Transactions on Parallel and Distributed Systems, 34 (6):1731–1742, June 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yu:2021:CHA

[YZJ^{+}21] T. Yu, R. Zhong, V. Janjic, P. Petoumenos, J. Zhai, H. Leather, and J. Thomson. Collaborative heterogeneityaware OS scheduler for asymmetric multicore processors. IEEE Transactions on Parallel and Distributed Systems, 32 (5):1224–1237, May 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yu:2020:LSA

[YZL^{+}20] T. Yu, W. Zhao, P. Liu, V. Janjic, X. Yan, S. Wang, H. Fu, G. Yang, and J. Thomson. Large-scale automatic K-means clustering for heterogeneous many-core supercomputer. IEEE Transactions on Parallel and Distributed Systems, 31(5):997–1008, May 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yang:2024:VDV

[YZL24] Jiaqi Yang, Hao Zheng, and Ahmed Louri. Versa-DNN: a versatile architecture enabling high-performance and energyefficient multi-DNN acceleration. IEEE Transactions on Parallel and Distributed Systems, 35(2):349–361, February 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yu:2021:OOP

[YZS^{+}21] H. Yu, Z. Zheng, J. Shen, C. Miao, C. Sun, H. Hu, J. Bi, J. Wu, and J. Wang. Octans: Optimal placement of service function chains in manycore systems. IEEE Transactions on Parallel and Distributed Systems, 32(9):2202– 2215, September 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yang:2023:ODF

[YZSX23] Lei Yang, Can Zheng, Xiaoyuan Shen, and Guoqi Xie. OfpCNN: On-demand finegrained partitioning for CNN inference acceleration in heterogeneous devices. IEEE Transactions on Parallel and Distributed Systems, 34(12):3090– 3103, December 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yi:2020:ECI

[YZWT20] D. Yi, X. Zhou, Y. Wen, and R. Tan. Efficient computeintensive job allocation in data centers via deep reinforcement learning. IEEE Transactions on Parallel and Distributed Systems, 31(6):1474– 1485, June 2020. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2022:PSC

[ZBB^{+}22] Junchao Zhang, Jed Brown, Satish Balay, Jacob Faibussowitsch, Matthew Knepley, Oana Marin, Richard Tran Mills, Todd Munson, Barry F. Smith, and Stefano Zampini. The PetscSF scalable communication layer. IEEE Transactions on Parallel and Distributed Systems, 33(4):842–853, April 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2023:NNI

[ZCHZ23] Zining Zhang, Yao Chen, Bingsheng He, and Zhenjie Zhang. NIOT: a novel inference optimization of transformers on modern CPUs. IEEE Transactions on Parallel and Distributed Systems, 34(6):1982– 1995, June 2023. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zheng:2022:NFF

[ZCJ^{+}22] Size Zheng, Renze Chen, Yicheng Jin, Anjiang Wei, Bingyang Wu, Xiuhong Li, Shengen Yan, and Yun Liang. NeoFlow: a flexible framework for enabling efficient compilation for high performance DNN training. IEEE Transactions on Parallel and Distributed Systems, 33(11):3220– 3232, November 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2020:IOP

[ZCJY20] W. Zhang, Q. Cao, H. Jiang, and J. Yao. Improving overall performance of TLC SSD by exploiting dissimilarity of flash pages. IEEE Transactions on Parallel and Distributed Systems, 31(2):332–346, February 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhou:2022:HMA

[ZCL^{+}22] Yuxuan Zhou, Wanzhong Chen, Linlin Li, Linlin Gong, and Tao Zhang. Heterogeneous multi-agent system for braincomputer interaction in routing and forwarding with memristive neuron networks. IEEE Transactions on Parallel and Distributed Systems, 33(11): 3233–3248, November 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhou:2020:HRM

[ZCW^{+}20] J. Zhou, Y. Chen, W. Wang, S. He, and D. Meng. A highly reliable metadata service for large-scale distributed file systems. IEEE Transactions on Parallel and Distributed Systems, 31(2):374–392, February 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2021:EPS

[ZCZ^{+}21] F. Zhang, Z. Chen, C. Zhang, A. C. Zhou, J. Zhai, and X. Du. An efficient parallel secure machine learning framework on GPUs. IEEE Transactions on Parallel and Distributed Systems, 32(9):2262– 2276, September 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhong:2022:CMM

[ZCZ^{+}22] Runxin Zhong, Jiajie Chen, Chen Zhang, Mingshu Zhai, Zeyu Song, Yutian Wang, Wentao Han, Lin Gan, and Jidong Zhai. Critique of MemXCT: Memory-centric X-ray CT reconstruction with massive parallelization by SCC Team from Tsinghua University. IEEE Transactions on Parallel and Distributed Systems, 33(9): 2050–2053, September 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhao:2023:LSM

[ZDC^{+}23] Hailiang Zhao, Shuiguang Deng, Feiyi Chen, Jianwei Yin, Schahram Dustdar, and Albert Y. Zomaya. Learning to schedule multi-server jobs with fluctuated processing speeds. IEEE Transactions on Parallel and Distributed Systems, 34(1): 234–245, January 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zeng:2022:ADB

[ZDK^{+}22] Jing Zeng, Ding Ding, Kaixuan Kang, HuaMao Xie, and Qian Yin. Adaptive DRLbased virtual machine consolidation in energy-efficient cloud data center. IEEE Transactions on Parallel and Distributed Systems, 33(11):2991– 3002, November 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhao:2021:FCA

[ZDL^{+}21] K. Zhao, S. Di, S. Li, X. Liang, Y. Zhai, J. Chen, K. Ouyang, F. Cappello, and Z. Chen. FT-CNN: Algorithm-based fault tolerance for convolutional neural networks. IEEE Transactions on Parallel and Distributed Systems, 32(7):1677– 1689, July 2021. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhou:2023:BEC

[ZF23] Hai Zhou and Dan Feng. Boosting erasure-coded multi-stripe repair in rack architecture and heterogeneous clusters: Design and analysis. IEEE Transactions on Parallel and Distributed Systems, 34(8):2251– 2264, August 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhou:2022:BAS

[ZFH22] Hai Zhou, Dan Feng, and Yuchong Hu. Bandwidth-aware scheduling repair techniques in erasure-coded clusters: Design and analysis. IEEE Transactions on Parallel and Distributed Systems, 33(12):3333– 3348, December 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2020:IRPa

[ZFW^{+}20] Y. Zhang, M. Fu, X. Wu, F. Wang, Q. Wang, C. Wang, X. Dong, and H. Han. Improving restore performance of packed datasets in deduplication systems via reducing persistent fragmented chunks. IEEE Transactions on Parallel and Distributed Systems, 31 (7):1651–1664, July 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2020:OSP

[ZFY^{+}20] P. Zhang, J. Fang, C. Yang, C. Huang, T. Tang, and Z. Wang. Optimizing streaming parallelism on heterogeneous many-core architectures. IEEE Transactions on Parallel and Distributed Systems, 31(8): 1878–1896, August 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhou:2021:FTD

[ZGG21] T. Zhou, L. Gao, and X. Guan.

A fault-tolerant distributed framework for asynchronous iterative computations. IEEE Transactions on Parallel and Distributed Systems, 32(8): 2062–2073, August 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2021:LDI

[ZGM21] H. Zhang, X. Geng, and H. Ma. Learning-driven interferenceaware workload parallelization for streaming applications in heterogeneous cluster. IEEE Transactions on Parallel and Distributed Systems, 32(1):1– 15, January 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhao:2022:SGM

[ZGNZ22] Chen Zhao, Wu Gao, Feiping Nie, and Huiyang Zhou. A survey of GPU multitasking methods supported by hardware architecture. IEEE Transactions on Parallel and Distributed Systems, 33(6):1451– 1463, June 2022. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhou:2021:PHA

[ZGQ^{+}21] Q. Zhou, S. Guo, Z. Qu, P. Li, L. Li, M. Guo, and K. Wang. Petrel: Heterogeneity-aware distributed deep learning via hybrid synchronization. IEEE Transactions on Parallel and Distributed Systems, 32(5): 1030–1043, May 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhai:2023:FBF

[ZGZ^{+}23] Yujia Zhai, Elisabeth Giem, Kai Zhao, Jinyang Liu, Jiajun Huang, Bryan M. Wong, Christian R. Shelton, and Zizhong Chen. FT-BLAS: a fault tolerant high performance BLAS implementation on x86 CPUs. IEEE Transactions on Parallel and Distributed Systems, 34(12):3207–3223, December 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2024:HOE

[ZGZ^{+}24] Bowen Zhang, Huaxi Gu, Grace Li Zhang, Yintang Yang, Ziteng Ma, and Ulf Schlichtmann. A 3D hybrid opticalelectrical NoC using novel mapping strategy based DCNN dataflow acceleration. IEEE Transactions on Parallel and Distributed Systems, 35(7): 1139–1154, July 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2024:AAD

[ZHH^{+}24] Kai Zhang, Jiahui Hong, Zhengying He, Yinan Jing, and X. Sean Wang. AdaptChain: Adaptive data sharing and synchronization for NFV systems on heterogeneous architectures. IEEE Transactions on Parallel and Distributed Systems, 35 (7):1281–1292, July 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2023:SHC

[ZHP^{+}23] Huaipeng Zhang, Nhut-Minh Ho, Dogukan Yigit Polat, Peng Chen, Mohamed Wahib, Truong Thao Nguyen, Jintao Meng, Rick Siow Mong Goh, Satoshi Matsuoka, Tao Luo, and Weng-Fai Wong. Simeuro: a hybrid CPU–GPU parallel simulator for neuromorphic computing chips. IEEE Transactions on Parallel and Distributed Systems, 34(10):2767– 2782, October 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2023:HFT

[ZHQ^{+}23] Hui Zhang, Rong-Xia Hao, Xiao-Wen Qin, Cheng-Kuan Lin, and Sun-Yuan Hsieh. The high faulty tolerant capability of the alternating group graphs. IEEE Transactions on Parallel and Distributed Systems, 34(1): 225–233, January 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2022:LAD

[ZHX^{+}22] Qinglong Zhang, Rui Han, Gaofeng Xin, Chi Harold Liu, Guoren Wang, and Lydia Y. Chen. Lightweight and accurate DNN-based anomaly detection at edge. IEEE Transactions on Parallel and Distributed Systems, 33(11):2927– 2942, November 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhu:2021:GAF

[ZJGD21] B. Zhu, Y. Jiang, M. Gu, and Y. Deng. A GPU acceleration framework for motif and discord based pattern mining. IEEE Transactions on Parallel and Distributed Systems, 32(8): 1987–2004, August 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2023:DED

[ZJH^{+}23] Shuai Zhang, Zite Jiang, Xingzhong Hou, Mingyu Li, Mengting Yuan, and Haihang You. DRONE: an efficient distributed subgraph-centric framework for processing largescale power-law graphs. IEEE Transactions on Parallel and Distributed Systems, 34(2):463– 474, February 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zarebavani:2020:CCB

[ZJHS20] B. Zarebavani, F. Jafarinejad, M. Hashemi, and S. Salehkaleybar. cuPC: CUDA-based parallel PC algorithm for causal structure learning on GPU. IEEE Transactions on Parallel and Distributed Systems, 31 (3):530–542, March 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zheng:2024:DTP

[ZJS^{+}24] Qiushi Zheng, Jiong Jin, Zhishu Shen, Libing Wu, Iftekhar Ahmad, and Yong Xiang. Distributed task processing platform for infrastructure-less IoT networks: a multi-dimensional optimization approach. IEEE Transactions on Parallel and Distributed Systems, 35(12): 2392–2404, December 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2024:FEF

[ZKL24] Mi Zhang, Qihan Kang, and Patrick P. C. Lee. FlexRaft: Exploiting flexible erasure coding for minimum-cost consensus and fast recovery. IEEE Transactions on Parallel and Distributed Systems, 35(10): 1826–1840, October 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhou:2020:ASG

[ZKP20] S. Zhou, R. Kannan, and V. K. Prasanna. Accelerating stochastic gradient descent based matrix factorization on FPGA. IEEE Transactions on Parallel and Distributed Systems, 31(8):1897–1911, August 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhao:2024:SNP

[ZL24] Yuzhen Zhao and Xiyu Liu. Spiking neural P systems with Microglia. IEEE Transactions on Parallel and Distributed Systems, 35(7):1239–1250, July 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhao:2022:VVA

[ZLC^{+}22] Shixiong Zhao, Fanxin Li, Xusheng Chen, Xiuxian Guan, Jianyu Jiang, Dong Huang, Yuhao Qing, Sen Wang, Peng Wang, Gong Zhang, Cheng Li, Ping Luo, and Heming Cui. vPipe: a virtualized acceleration system for achieving efficient and scalable pipeline parallel DNN training. IEEE Transactions on Parallel and Distributed Systems, 33(3):489– 506, March 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhou:2020:TEI

[ZLCL20] Z. Zhou, F. Liu, S. Chen, and Z. Li. A truthful and efficient incentive mechanism for demand response in green datacenters. IEEE Transactions on Parallel and Distributed Systems, 31(1):1–15, January 2020. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhuang:2023:EAH

[ZLCW23] Hongbin Zhuang, Xiao-Yan Li, Jou-Ming Chang, and Dajin Wang. An efficient algorithm for Hamiltonian path embedding of k-ary n-cubes under the partitioned edge fault model. IEEE Transactions on Parallel and Distributed Systems, 34 (6):1802–1815, June 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2023:PFE

[ZLD^{+}23] Yu Zhang, Duo Liu, Moming Duan, Li Li, Xianzhang Chen, Ao Ren, Yujuan Tan, and Chengliang Wang. FedMDS: an efficient model discrepancyaware semi-asynchronous clustered federated learning framework. IEEE Transactions on Parallel and Distributed Systems, 34(3):1007–1019, March 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zou:2023:RRT

[ZLGZ23] An Zou, Jing Li, Christopher D. Gill, and Xuan Zhang. RTGPU: Real-time GPU scheduling of hard deadline parallel tasks with finegrain utilization. IEEE Transactions on Parallel and Distributed Systems, 34(5):1450– 1465, May 2023. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2023:AMP

[ZLJ^{+}23] Yilian Zhang, Yao Liu, Penglong Jiao, Yiping Zhou, and Tongquan Wei. Automatic multi-parameter performance modeling of HPC applications on a new Sunway supercomputer. IEEE Transactions on Parallel and Distributed Systems, 34(11):2965– 2977, November 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhou:2022:FOM

[ZLK^{+}22] Amelie Chi Zhou, Jianming Lao, Zhoubin Ke, Yi Wang, and Rui Mao. FarSpot: Optimizing monetary cost for HPC applications in the cloud spot market. IEEE Transactions on Parallel and Distributed Systems, 33(11):2955–2967, November 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2022:ILA

[ZLL22a] Qi Zhang, Yi Liu, and Tao Liu. iBalancer: Load-aware in-server flow scheduling for sub-millisecond tail latency. IEEE Transactions on Parallel and Distributed Systems, 33(8): 1761–1774, August 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhao:2022:UUF

[ZLL^{+}22b] Minghao Zhao, Zhenhua Li, Wei Liu, Jian Chen, and Xingyao Li. UFC2: Userfriendly collaborative cloud. IEEE Transactions on Parallel and Distributed Systems, 33(9): 2163–2182, September 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2024:PPD

[ZLP^{+}24] Lan Zhang, Anran Li, Hongyi Peng, Feng Han, Fan Huang, and Xiang-Yang Li. Privacypreserving data selection for horizontal and vertical federated learning. IEEE Transactions on Parallel and Distributed Systems, 35(11):2054– 2068, November 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhu:2020:SPM

[ZLR^{+}20] J. Zhu, X. Li, R. Ruiz, W. Li, H. Huang, and A. Y. Zomaya. Scheduling periodical multistage jobs with fuzziness to elastic cloud resources. IEEE Transactions on Parallel and Distributed Systems, 31(12): 2819–2833, December 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhou:2022:TES

[ZLRY22] Zihao Zhou, Yanan Li, Xuebin Ren, and Shusen Yang. Towards efficient and stable kasynchronous federated learning with unbounded stale gradients on non-IID data. IEEE Transactions on Parallel and Distributed Systems, 33(12): 3291–3305, December 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zheng:2024:PDR

[ZLT^{+}24] Dongyu Zheng, Lei Liu, Guoming Tang, Yi Wang, and Weichao Li. Power demand reshaping using energy storage for distributed edge clouds. IEEE Transactions on Parallel and Distributed Systems, 35(2): 362–376, February 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2020:HPG

[ZLW20] T. Zhang, X. Liu, and X. Wang.

High performance GPU tensor completion with tubalsampling pattern. IEEE Transactions on Parallel and Distributed Systems, 31(7):1724– 1739, July 2020. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhao:2021:TSS

[ZLW^{+}21] Yunjian Zhao, Zhi Liu, Yidi Wu, Guanxian Jiang, James Cheng, Kunlong Liu, and Xiao Yan. Timestamped state sharing for stream analytics. IEEE Transactions on Parallel and Distributed Systems, 32(11): 2691–2704, November 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhou:2023:ADL

[ZLW^{+}23] Huan Zhou, Mingze Li, Ning Wang, Geyong Min, and Jie Wu. Accelerating deep learning inference via model parallelism and partial computation offloading. IEEE Transactions on Parallel and Distributed Systems, 34(2):475–488, February 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2020:CTE

[ZLWW20] T. Zhang, X. Liu, X. Wang, and A. Walid. cuTensor-Tubal: Efficient primitives for tubal-rank tensor learning operations on GPUs. IEEE Transactions on Parallel and Distributed Systems, 31(3):595–610, March 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zou:2020:POR

[ZLX^{+}20] X. Zou, T. Lu, W. Xia, X. Wang, W. Zhang, H. Zhang, S. Di, D. Tao, and F. Cappello. Performance optimization for relative-error-bounded lossy compression on scientific data. IEEE Transactions on Parallel and Distributed Systems, 31(7):1665–1680, July 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zeng:2022:DPF

[ZLYL22] Yiming Zeng, Yixuan Lin, Yuanyuan Yang, and Ji Liu. Differentially private federated temporal difference learning. IEEE Transactions on Parallel and Distributed Systems, 33(11):2714–2726, November 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhao:2023:AIA

[ZLZ^{+}23] Gongming Zhao, Jiawei Liu, Yutong Zhai, Hongli Xu, and Huang He. Alleviating the impact of abnormal events through multi-constrained VM placement. IEEE Transactions on Parallel and Distributed Systems, 34(5):1508–1523, May 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2023:FMD

[ZMP23] Chi Zhang, Yuan Meng, and Viktor Prasanna. A framework for mapping DRL algorithms with prioritized replay buffer onto heterogeneous platforms. IEEE Transactions on Parallel and Distributed Systems, 34 (6):1816–1829, June 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhou:2022:ATA

[ZMS^{+}22] Keren Zhou, Xiaozhu Meng, Ryuichi Sai, Dejan Grubisic, and John Mellor-Crummey. An automated tool for analysis and tuning of GPU-accelerated code in HPC applications. IEEE Transactions on Parallel and Distributed Systems, 33 (4):854–865, April 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2024:AEA

[ZMY^{+}24] Yanxi Zhang, Muyu Mei, Dongqi Yan, Xu Zhang, Qinghai Yang, and Mingwu Yao. Age-of-event aware: Sampling period optimization in a threestage wireless cyber-physical system with diverse parallelisms. IEEE Transactions on Parallel and Distributed Systems, 35(8):1360–1372, August 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2022:NNA

[ZPL^{+}22] Penghao Zhang, Heng Pan, Zhenyu Li, Penglai Cui, Ru Jia, Peng He, Zhibin Zhang, Gareth Tyson, and Gaogang Xie. NetSHa: In-network acceleration of LSH-based distributed search. IEEE Transactions on Parallel and Distributed Systems, 33(9):2213– 2229, September 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2022:CEC

[ZQM^{+}22] Xu Zhang, Zhengnan Qi, Geyong Min, Wang Miao, Qilin Fan, and Zhan Ma. Cooperative edge caching based on temporal convolutional networks. IEEE Transactions on Parallel and Distributed Systems, 33(9): 2093–2105, September 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhou:2024:DET

[ZRC^{+}24] Haoran Zhou, Wei Rang, Hongyang Chen, Xiaobo Zhou, and Dazhao Cheng. DeepTM: Efficient tensor management in heterogeneous memory for DNN training. IEEE Transactions on Parallel and Distributed Systems, 35(11):1920– 1935, November 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhou:2023:INI

[ZRXF23] Yu Zhou, Yanli Ren, Mengtian Xu, and Guorui Feng. An improved NSGA-III algorithm based on deep Q-networks for cloud storage optimization of blockchain. IEEE Transactions on Parallel and Distributed Systems, 34(5):1406– 1419, May 2023. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhu:2023:LMC

[ZSFX23] Weiguo Zhu, Yongqi Sun, Rongqiang Fang, and Baomin Xu. A low-memory community detection algorithm with hybrid sparse structure and structural information for largescale networks. IEEE Transactions on Parallel and Distributed Systems, 34(10):2671– 2683, October 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zheng:2021:FTS

[ZSH^{+}21] Z. Zheng, X. Shi, L. He, H. Jin, S. Wei, H. Dai, and X. Peng. Feluca: a twostage graph coloring algorithm with color-centric paradigm on GPU. IEEE Transactions on Parallel and Distributed Systems, 32(1):160–173, January 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2021:YTL

[ZSL^{+}21] F. Zhang, J. Su, W. Liu, B. He, R. Wu, X. Du, and R. Wang.

YuenyeungSpTRSV: a threadlevel and warp-level fusion synchronization-free sparse triangular solve. IEEE Transactions on Parallel and Distributed Systems, 32(9):2321– 2337, September 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2023:SMC

[ZSL^{+}23] Changwu Zhang, Hao Sun, Shuman Li, Yaohua Wang, Haiyan Chen, and Hengzhu Liu. A survey of memorycentric energy efficient computer architecture. IEEE Transactions on Parallel and Distributed Systems, 34(10): 2657–2670, October 2023. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhong:2024:SEF

[ZSL^{+}24] Yuchen Zhong, Guangming Sheng, Juncheng Liu, Jinhui Yuan, and Chuan Wu. Swift: Expedited failure recovery for large-scale DNN training. IEEE Transactions on Parallel and Distributed Systems, 35(9): 1644–1656, September 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhai:2022:GE

[ZSP22] Jidong Zhai, Min Si, and Antonio J. Peña. Guest editorial. IEEE Transactions on Parallel and Distributed Systems, 33(11):2644–2647, November 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhao:2020:THP

[ZSW^{+}20] Z. Zhao, W. Sheng, Q. Wang, W. Yin, P. Ye, J. Li, and Z. Mao. Towards higher performance and robust compilation for CGRA modulo scheduling. IEEE Transactions on Parallel and Distributed Systems, 31(9): 2201–2219, September 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhao:2022:EAF

[ZSW^{+}22] Zongyi Zhao, Xingang Shi, Zhiliang Wang, Qing Li, Han Zhang, and Xia Yin. Efficient and accurate flow record collection with HashFlow. IEEE Transactions on Parallel and Distributed Systems, 33(5): 1069–1083, May 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhou:2020:CAP

[ZSX^{+}20] A. C. Zhou, B. Shen, Y. Xiao, S. Ibrahim, and B. He. Costaware partitioning for efficient large graph processing in geo-distributed datacenters. IEEE Transactions on Parallel and Distributed Systems, 31 (7):1707–1723, July 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhao:2021:LSA

[ZTA^{+}21] N. Zhao, V. Tarasov, H. Albahar, A. Anwar, L. Rupprecht, D. Skourtis, A. K. Paul, K. Chen, and A. R. Butt. Large-scale analysis of Docker images and performance implications for container storage systems. IEEE Transactions on Parallel and Distributed Systems, 32(4):918– 930, April 2021. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2024:ANC

[ZTL24] Sai Zhang, Li Tang, and Yan-Jun Liu. Adaptive neural control for a network of parabolic PDEs with event-triggered mechanism. IEEE Transactions on Parallel and Distributed Systems, 35(7):1320–1330, July 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2022:MAG

[ZW22a] Zhaorui Zhang and Choli Wang. MIPD: an adaptive gradient sparsification framework for distributed DNNs training. IEEE Transactions on Parallel and Distributed Systems, 33(11):3053–3066, November 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2022:SSA

[ZW22b] Zhaorui Zhang and Choli Wang. SaPus: Self-adaptive parameter update strategy for DNN training on multi-GPU clusters. IEEE Transactions on Parallel and Distributed Systems, 33(7):1569–1580, July 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2020:IEP

[ZWK^{+}20] S. Zhang, Q. Wang, Y. Kanemasa, H. Shan, and L. Hu. The impact of event processing flow on asynchronous server efficiency. IEEE Transactions on Parallel and Distributed Systems, 31(3):565–579, March 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhou:2021:CDD

[ZWL^{+}21] Q. Zhou, K. Wang, H. Lu, W. Xu, Y. Sun, and S. Guo. Canary: Decentralized distributed deep learning via gradient sketch and partition in multi-interface networks. IEEE Transactions on Parallel and Distributed Systems, 32(4):900– 917, April 2021. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhou:2024:SAS

[ZWSF24] Yang Zhou, Fang Wang, Zhan Shi, and Dan Feng. The static allocation is not a static: Optimizing SSD address allocation through boosting static policy. IEEE Transactions on Parallel and Distributed Systems, 35(8): 1373–1386, August 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2022:DAN

[ZXG^{+}22] Feng Zhang, Erkang Xue, Ruixin Guo, Guangzhi Qu, Gansen Zhao, and Albert Y. Zomaya. DS-ADMM++: a novel distributed quantized ADMM to speed up differentially private matrix factorization. IEEE Transactions on Parallel and Distributed Systems, 33(6):1289–1302, June 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2021:DAC

[ZXGZ21] P. Zhang, H. Xue, S. Gao, and J. Zhang. Distributed adaptive consensus tracking control for multi-agent system with communication constraints. IEEE Transactions on Parallel and Distributed Systems, 32(6):1293–1306, June 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhao:2024:AAR

[ZXL^{+}24] Lu Zhao, Fu Xiao, Bo Li, Jian Zhou, Xiaolong Xu, and Yun Yang. Availability-aware revenue-effective application deployment in multi-access edge computing. IEEE Transactions on Parallel and Distributed Systems, 35(7):1268– 1280, July 2024. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2024:MME

[ZXW^{+}24] Zheng Zhang, Yaqi Xia, Hulin Wang, Donglin Yang, Chuang Hu, Xiaobo Zhou, and Dazhao Cheng. MPMoE: Memory efficient MoE for pre-trained models with adaptive pipeline parallelism. IEEE Transactions on Parallel and Distributed Systems, 35(6):843–856, June 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2023:EDD

[ZYD^{+}23] Shiwei Zhang, Xiaodong Yi, Lansong Diao, Chuan Wu, Siyu Wang, and Wei Lin. Expediting distributed DNN training with device topology-aware graph deployment. IEEE Transactions on Parallel and Distributed Systems, 34(4):1281– 1293, April 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2020:IRPb

[ZYF^{+}20] Y. Zhang, Y. Yuan, D. Feng, C. Wang, X. Wu, L. Yan, D. Pan, and S. Wang. Improving restore performance for in-line backup system combining deduplication and delta compression. IEEE Transactions on Parallel and Distributed Systems, 31(10):2302– 2314, October 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zheng:2022:ACI

[ZYK^{+}22] Quan Zheng, Tao Yang, Yuanzhi Kan, Xiaobin Tan, Jian Yang, and Xiaofeng Jiang. On the analysis of cache invalidation with LRU replacement. IEEE Transactions on Parallel and Distributed Systems, 33 (3):654–666, March 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2020:HTB

[ZYL^{+}20] W. Zhang, Z. Yan, Y. Lin, C. Zhao, and L. Peng. A high throughput B+tree for SIMD architectures. IEEE Transactions on Parallel and Distributed Systems, 31(3):707– 720, March 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhao:2020:OGD

[ZYM^{+}20] L. Zhao, Y. Yang, A. Munir, A. X. Liu, Y. Li, and W. Qu. Optimizing geodistributed data analytics with coordinated task scheduling and routing. IEEE Transactions on Parallel and Distributed Systems, 31(2):279– 293, February 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2023:PCE

[ZYW^{+}23] Ruixiao Zhang, Changpeng Yang, Xiaochan Wang, Tianchi Huang, Chenglei Wu, Jiangchuan Liu, and Lifeng Sun. Practical cloud-edge scheduling for largescale crowdsourced live streaming. IEEE Transactions on Parallel and Distributed Systems, 34(7):2055–2071, July 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhong:2022:EPR

[ZYX^{+}22] Kai Zhong, Zhibang Yang, Guoqing Xiao, Xingpei Li, Wangdong Yang, and Kenli Li. An efficient parallel reinforcement learning approach to cross-layer defense mechanism in industrial control systems. IEEE Transactions on Parallel and Distributed Systems, 33(11):2979–2990, November 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2023:EBB

[ZZC^{+}23] Qingyang Zhang, Zhiming Zhang, Jie Cui, Hong Zhong, Yang Li, Chengjie Gu, and Debiao He. Efficient blockchainbased data integrity auditing for multi-copy in decentralized storage. IEEE Transactions on Parallel and Distributed Systems, 34(12):3162–3173, December 2023. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zeng:2024:OTO

[ZZD^{+}24a] Tianyu Zeng, Xiaoxi Zhang, Jingpu Duan, Chao Yu, Chuan Wu, and Xu Chen. An offlinetransfer-online framework for cloud-edge collaborative distributed reinforcement learning. IEEE Transactions on Parallel and Distributed Systems, 35(5):720–731, May 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2024:FFL

[ZZD^{+}24b] Yuanhong Zhang, Weizhan Zhang, Haipeng Du, Caixia Yan, Li Liu, and Qinghua Zheng. FHVAC: Feature-level hybrid video adaptive configuration for machine-centric live streaming. IEEE Transactions on Parallel and Distributed Systems, 35(5):780–795, May 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2021:IML

[ZZG^{+}21a] C. Zhang, F. Zhang, X. Guo, B. He, X. Zhang, and X. Du. iMLBench: a machine learning benchmark suite for CPU-GPU integrated architectures. IEEE Transactions on Parallel and Distributed Systems, 32 (7):1740–1752, July 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2021:JTS

[ZZG^{+}21b] J. Zhang, X. Zhou, T. Ge, X. Wang, and T. Hwang. Joint task scheduling and containerizing for efficient edge computing. IEEE Transactions on Parallel and Distributed Systems, 32(8):2086–2100, August 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2020:MDF

[ZZH^{+}20a] J. Zhang, K. Zhou, P. Huang, X. He, M. Xie, B. Cheng, Y. Ji, and Y. Wang. Minority disk failure prediction based on transfer learning in large data centers of heterogeneous disk systems. IEEE Transactions on Parallel and Distributed Systems, 31(9):2155– 2169, September 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhou:2020:ESC

[ZZH^{+}20b] K. Zhou, Y. Zhang, P. Huang, H. Wang, Y. Ji, B. Cheng, and Y. Liu. Efficient SSD cache for cloud block storage via leveraging block reuse distances. IEEE Transactions on Parallel and Distributed Systems, 31(11):2496– 2509, November 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2021:CPN

[ZZH^{+}21] Chen Zhang, Chenggang Zhao, Jiaao He, Shengqi Chen, Liyan Zheng, Kezhao Huang, Wentao Han, and Jidong Zhai. Critique of planetary normal mode computation: Parallel algorithms, performance, and reproducibility by SCC Team From Tsinghua University. IEEE Transactions on Parallel and Distributed Systems, 32(11):2631– 2634, November 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). See [SLX^{+}21b].

Zhang:2023:PPB

[ZZM^{+}23] Yiwen Zhang, Jian Zhou, Xinhao Min, Song Ge, Jiguang Wan, Ting Yao, and Daohui Wang. PetaKV: Building efficient key–value store for file system metadata on persistent memory. IEEE Transactions on Parallel and Distributed Systems, 34(3):843–855, March 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2024:RPB

[ZZN^{+}24] Bowen Zhang, Shengan Zheng, Liangxu Nie, Zhenlin Qi, Hongyi Chen, Linpeng Huang, and Hong Mei. Revisiting PM-based B^{+}-tree with persistent CPU cache. IEEE Transactions on Parallel and Distributed Systems, 35(5):796– 813, May 2024. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2023:GFB

[ZZP23] Bingyi Zhang, Hanqing Zeng, and Viktor K. Prasanna. GraphAGILE: an FPGA-Based overlay accelerator for lowlatency GNN inference. IEEE Transactions on Parallel and Distributed Systems, 34(9): 2580–2597, September 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2021:DCA

[ZZQ^{+}21] S. Zhang, S. Zhang, Z. Qian, J. Wu, Y. Jin, and S. Lu. DeepSlicing: Collaborative and adaptive CNN inference with low latency. IEEE Transactions on Parallel and Distributed Systems, 32(9):2175– 2187, September 2021. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2022:PHP

[ZZS^{+}22] Feng Zhang, Jidong Zhai, Xipeng Shen, Onur Mutlu, and Xiaoyong Du. POCLib: a highperformance framework for enabling near orthogonal processing on compression. IEEE Transactions on Parallel and Distributed Systems, 33(2):459– 475, February 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhou:2020:ERS

[ZZSC20] W. Zhou, Y. Zhao, X. Shen, and W. Chen. Enabling runtime SpMV format selection through an overhead conscious method. IEEE Transactions on Parallel and Distributed Systems, 31(1):80–93, January 2020. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2021:FGM

[ZZY^{+}21] F. Zhang, C. Zhang, L. Yang, S. Zhang, B. He, W. Lu, and X. Du. Fine-grained multiquery stream processing on integrated architectures. IEEE Transactions on Parallel and Distributed Systems, 32(9): 2303–2320, September 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zeng:2024:HPG

[ZZZ^{+}24a] Gang Zeng, Jianfeng Zhu, Yichi Zhang, Ganhui Chen, Zhenhai Yuan, Shaojun Wei, and Leibo Liu. A high-performance genomic accelerator for accurate sequence-to-graph alignment using dynamic programming algorithm. IEEE Transactions on Parallel and Distributed Systems, 35(2):237– 249, February 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2024:BDA

[ZZZ^{+}24b] Liang Zhang, Wenli Zheng, Kuangyu Zheng, Hongzi Zhu, Chao Li, and Minyi Guo. Bayesian-driven automated scaling in stream computing with multiple QoS targets. IEEE Transactions on Parallel and Distributed Systems, 35 (7):1251–1267, July 2024. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).